A Complete Bibliography of Publications in

Conservation Biology: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

07 April 2025 Version 1.16

Title word cross-reference

30 × 30 [HXW^{+}23, MBLD^{+}25]. p [Hen22, IR22, Sta22a]. -values [Hen22, IR22, Sta22a]. 0 [Ano22i]. 19 [BBL^{+}20, GSS^{+}22]. 2 [Sár23]. 2000 [Con21, PSP^{+}20, RDS^{+}24, STA^{+}22b]. 2020 [Ano20b]. 2021 [Ano22b, Ano24a]. 2022 [Ano23a, Ano23b, Ano24a]. 2023 [Ano24b]. 2024 [Ano24b]. 2030 [FZG^{+}24]. 2nd [Ano22i]. 3 [Cli23, KWY^{+}21]. 3-D [KWY^{+}21]. 5 [Ano23n]. 6 [Ano22j].

978 [Ano22i, Ano22j, Ano23n, Cli23, Sár23]. 978-0-19-284346-3 [Ano22j]. 978-0-19-289551-6 [Ano22j]. 978-0-19-874625-6 [Ano22j]. 978-0-19-885324-4 [Ano22j]. 978-0-226-81594-7 [Ano23n]. 978-0-691-18047-2 [Sár23]. 978-0-691-20417-8 [Ano22j]. 978-0-691-20760-5 [Ano23n]. 978-0-8165-4148-5 [Ano22j]. 978-0-86154-017-4 [Ano22i]. 978-1-107-69430-9 [Cli23]. 978-1-107-69815-4 [Ano22i]. 978-1-324-00508-7 [Ano23n]. 978-1-4214-4457-4 [Ano23n]. 978-1-64283-202-0 [Ano22i]. 978-1-7372851-1-3 [Cli23]. 978-1-78924-206-5 [Ano22i]. 978-2-913688-38-4 [Sár23].

abatement [MAB^{+}23, RRH^{+}24]. Ability [YTW^{+}25, SWBD20]. abiotic [TFT^{+}24]. absence [DSF^{+}23]. Abundance [GCSZ21, BP21, DSZ21, FOYZ22, FJH21, GFF23, HMC23a]. abuse [CC24]. accepting [CCKT25]. access [RHGRSG21]. accessibility [LWC^{+}24]. Accounting [YVA22, MWH^{+}22, WDH^{+}21]. accumulation [LWD^{+}20]. achieve [MBLD^{+}25, SRCS23, WSC^{+}23]. Achieving [JCD^{+}23, NFD20, ODS^{+}24, RzEHB24]. Acid [Rap21]. acidification [HK21]. acoustic [VMP^{+}24, WKG^{+}21]. acquiring [ACD^{+}24]. across [BF21, DSC^{+}24, EHC^{+}21, ECVL25, GBRK^{+}20, HNC^{+}24, KTG^{+}22, MHP23, MCC^{+}23b, PMLM20, RDS^{+}24, RGD^{+}25, SCB^{+}25, SH21, TDC^{+}23, TTDdLM^{+}21, VRY^{+}23, VBK^{+}23, WFH^{+}21]. Act [FSC^{+}21, Ver23, NCE^{+}21, SENTI24, TENI23]. action [BDF^{+}22, Bou23, EGCS25, GBA23, HK21, ROK^{+}22, VM20, WMK^{+}21]. actionable [BE23, KBW^{+}23]. actions [DFS^{+}25, DS22, MCM^{+}21, RRH^{+}24, WRW^{+}20, WKT^{+}21, vESN^{+}20]. activities [CWMB23]. activity [CCL23, JZLP^{+}23, MHL^{+}22]. actors [STW22]. adapt [Ala25]. adaptation [HGCK22, IJK^{+}21, LdSB^{+}20, STSJ22]. adapted [CS21]. Adapting [BHM21]. Adaptive [CGM^{+}24, MBW^{+}22, STJ20, TGM^{+}22, XDA^{+}21, ZAS^{+}24]. added [YVA22]. address [DDM24, DFH^{+}21, GBA23, HSG^{+}21, JDB^{+}21, McC24]. Addressing [HSHA20, JSA^{+}20, MTS^{+}21, MJDMAS20, RzEHB24, SB21]. adequacy [KLVW20]. Adivasi [BNS24]. adjacent [HBJ^{+}20]. adjusting [GPJJ^{+}21]. adjustment [GGJ^{+}20]. administration [IDJ^{+}25]. adoption [KRB^{+}22, RW23]. adult [LFD^{+}25]. advance [KWY^{+}21, LKL^{+}21, RCRF20]. advances [CCPE^{+}25]. Advancing [DS25, VDO^{+}24]. advice [AFV20]. advocacy [BO21]. aeroconservation [HVA^{+}21]. affect [MEK21, WB23]. affected [VNC^{+}24]. affects [DFC21]. affirmation [HSHA20]. Africa [GR21, HtSR25, KGK^{+}24, MSMH22, PPN^{+}24]. African [CVT^{+}25, CTM^{+}22, DNCC22, GJC^{+}23, Imb22, LB22, SPA^{+}22, SHS^{+}22, YCH^{+}24]. after [AAvdW^{+}22, BFS24, BAGV24, MPM^{+}23, PLTC^{+}23, WCW^{+}22]. against [FJZ^{+}20, KvZS^{+}25, SWBD20]. Age [LXZ^{+}22, LWS^{+}24]. agencies [VBK^{+}23]. agency [CBB^{+}20, EKH23, Uts24]. Agenda [Cor21, SAB^{+}22, WMK^{+}21]. aggregating [EPL^{+}24]. aggregation [CFC21]. Agricultural [CCCB^{+}23, ANR23, BSRM22, FBG^{+}25, YMB^{+}21]. agriculture [KSF^{+}24]. agrienvironments [RW23]. agroforestry [MR22]. aimed [ZWC^{+}24c]. airspace [LS21]. al [CC20, CO21, CMAS20, EDA^{+}21, EHBT20, FJZ^{+}21, GN20, KOO^{+}21, LSÁH^{+}20, MCLM22, MBD^{+}21, NAS21b, NBB^{+}21, PG20, Ras22, Sag20b, SGB^{+}20, SÁT21, STSJ22, SP21, Wal24, WSW^{+}22, dG21]. Alarming [NSB^{+}21]. Aldo [Pic20]. algal [DNCC22]. Alien [KBB^{+}24, BAB^{+}23b]. Aligning [GBA23]. allegations [LSM24]. alliances [RCRF20]. allocating [MCM^{+}21]. allocation [WAA^{+}21]. along [GJC^{+}23, LWD^{+}20, LLYK24, SFC20]. alpha [JBE^{+}22a]. alpine [RMW23]. alternative [SdVD^{+}22]. alternatives [HBY^{+}23, NLD^{+}20]. Amazon [CVT^{+}21, JTD^{+}20, MEM^{+}22]. Amazonia [EMV^{+}20]. Amazonian [ECVL25, FGB^{+}23]. America [BPR24b, CRC^{+}21, GTAdPA24, HTC22, MMN^{+}20, PS20, Uts24, dlLBB^{+}23]. American [AP21, CFN^{+}24, EHC^{+}21, NRN^{+}24, SP23a, TRR^{+}20]. among [ADSG21, BGVK20, Ele20, GG24, HLL^{+}24, ITAAG^{+}24, MCM^{+}21, Moo21, PKC^{+}23b, WDHM21]. amount [CQS^{+}22]. amphibian [DSM24, FGL^{+}21, FW22, MWY^{+}21, MBV^{+}24, TAA^{+}24]. amphibians [Ala23, ABCH^{+}21, CST^{+}24, LFD^{+}25, LDC^{+}24, MHL^{+}22]. amplify [DBM^{+}25]. amusing [LSM^{+}20]. analyses [SNC^{+}22]. Analysis [Cli23, ABC^{+}23, BHG^{+}23, BAA^{+}25, CO20, CO21, DdLI^{+}25, FMGS21, GFF23, GGSRF20, JSGJ^{+}21, KFP^{+}21, LFT^{+}21, LVT^{+}20, LDC^{+}24, MHM^{+}24, MWS^{+}25, ML22, NLD^{+}20, NAG^{+}20, PHK^{+}25, RCRJ^{+}24, SMLL24, SCC23, SP21, SSKBL21, TGJ^{+}20, WJ21, WLA20, WWVC^{+}25]. Analyzing [FBG^{+}25]. ancestors [SQJ^{+}22]. ancestry [MPM^{+}23]. ancient [DBR^{+}23]. Andean [PKP20]. Andes [MHP23]. angiosperms [BOB^{+}20]. Animal [EKH23, CHW^{+}23, DFW^{+}24, LEF^{+}24, MS23, PS24, STJ20, WBB^{+}20, WKG^{+}21, ZLZ^{+}21]. Animals [Ano23n, CVT^{+}21, CC24, FLW^{+}22, FSBE22, GCSZ21, HQY^{+}24, SQJ^{+}22]. ant [ABCH^{+}21]. antagonisms [SMLL24]. Antagonistic [GGJ^{+}20]. Antarctic [GGLB25, ZWC^{+}24b]. Antarctica [BLW23, PLC22, ZSL^{+}24a]. Anthropocene [Gru20]. Anthropogenic [BOB^{+}20, QZX^{+}24, ACD^{+}24, ANR23, DLL22, GHR21, JM21, JZLP^{+}23, KFP^{+}21, KBB^{+}23, LHB^{+}22, WLR^{+}22]. Anticipating [ACD^{+}24]. ants [BDF^{+}22, GSB23]. Aotearoa [HCC^{+}21]. Apennine [MPM^{+}23]. apocalypse [Hal22]. Application [BHG^{+}23, NLD^{+}20, PWW^{+}22, SGA^{+}24, GGBD20, GBB^{+}20, Ano23n]. applications [MHC21]. applied [FSBE22, LGD^{+}25]. Applying [BBG^{+}20, Cli23, HCC^{+}21, TGM^{+}22, WBC^{+}20]. appraisal [CO20]. approach [BDQ^{+}25, Cas23, GWB^{+}24, HHM24, HM24, JAB^{+}21a, MSC^{+}24, MBD^{+}23, NCdlLZ24, RBLN21, SCW^{+}24, TF24, VFRV24, VG20, XST^{+}20]. Approaches [KSFP^{+}21, ADSG21, BHM21, CBD^{+}21, KBW^{+}23, KBCD20, PHB^{+}25, SGB^{+}20, YWH^{+}24]. aquaculture [CLM22, LMY^{+}24, ODS^{+}24]. aquarium [SS23]. aquariums [EAdC^{+}21, MWS^{+}25]. aquatic [BHSM^{+}22, KKE^{+}24, PHK^{+}25, SD22, VFKT24]. arboreal [BP21]. archipelagos [ECVL25]. archives [LMG^{+}23]. Arctic [Bus22, CRAN^{+}24]. area [And23, BKK^{+}21, BVH22, CTP^{+}24, CTM^{+}22, Coo24, FCL^{+}22, GJPJ^{+}22, GWB^{+}24, GCS^{+}21, HBJ^{+}20, HWP^{+}21, HTM25, LDD^{+}22, MOO^{+}24, MNMPGN23, MCF^{+}24, MFW^{+}24, MSMW21, MSJ20, NDS^{+}20, NFT^{+}20, PLC22, RCBT21, SCdM^{+}24, SBH^{+}23, SAB^{+}22, VFKT24, WCY^{+}25, PHB^{+}25]. area-based [BKK^{+}21, CTP^{+}24, Coo24, MOO^{+}24, SCdM^{+}24]. areas [ACD^{+}24, Ano23d, ARKA22, BBH^{+}25, BLW23, CAFM^{+}23, CMC^{+}23, COOH^{+}25, CFC21, GPJJ^{+}21, GGB20b, GTAdPA24, GCS^{+}24, HNC^{+}24, HGAB^{+}24, HC22, HMS^{+}23, IDJ^{+}25, KGK^{+}24, KBCD20, LdSB^{+}20, LZG^{+}25, LLYK24, MHJ^{+}24, MSMH22, NSB^{+}21, NLV20, OCC^{+}21, POH21, PMSC22, PSJG23, SDB^{+}24, SWBD20, TE20, TJC21, WBR^{+}24, WOF^{+}20, YTW^{+}25, YVA22, ZTCF21, ZYNC25, ZSL24b]. Argentine [ABCH^{+}21]. arising [NRSF21]. Arizona [Ano22j]. articles [HPM^{+}20, The24]. artificial [CO24, GWR23, KAD^{+}22]. Asia [CSCA^{+}23, CSGG25, GFK^{+}22, NGST21, QZX^{+}24]. Asian [Cli23, BNS24, TFT^{+}24]. Assam [BNS24]. assemblage [VPVdAMF20]. assemblages [FBDA22, MHS^{+}21]. assess [HBB^{+}24]. Assessing [ARK^{+}20, BRJ25, CTM^{+}22, DFB^{+}24, GMEV^{+}21, HHL^{+}24, HCZW24, IJ24, LMY^{+}24, OCC^{+}21, PWW^{+}22, RRR^{+}24, RW23, Ric21, RJV^{+}24, RCBT21, TBP^{+}22, Ano23d, ART^{+}25, CMC^{+}23, GAB^{+}21, KLVW20, MBD^{+}21, PHK^{+}25, RCRF20]. assessment [Ano24e, BAGK22, BPR24a, CND^{+}23, CVT^{+}25, CO21, DTV^{+}25, GWB^{+}24, LdSB^{+}20, MTH^{+}24, MCF^{+}21, MWH^{+}22, NSL^{+}23, OABDCM24, PLD^{+}21, RBC^{+}22, RML^{+}20, RDS^{+}24, SR20, SOW^{+}22, ZSVK21, dlLBB^{+}23]. assessments [And23, DS25, DBT21, GRP^{+}25, HOT^{+}25, KIA^{+}22, LS21, LFT^{+}21, MSC^{+}24, NCE^{+}21, RSRG^{+}20, RWC^{+}22, TGM^{+}22, VLVD^{+}23, WLB^{+}23]. assets [WGZ24]. assigned [JBE^{+}22b]. Assigning [WSW23]. Assisted [CS21, BB21, GBMD^{+}22, PB21]. associated [SCC23, TENI23, ZSL24b]. Association [MSHE23]. associations [ZWC^{+}24a]. assumptions [TSÁ20]. Atacama [VLPVM^{+}24]. Athabasca [PWW^{+}22]. Atlantic [MCP^{+}25, FGB^{+}23]. attempts [RBD^{+}21]. attention [BAGV24, Van23]. attitudes [BMM^{+}24, CVT^{+}25, MJDMAS20, MBL^{+}23, NFD20, STES23, The24, VG20, WTAD21, vESN^{+}20]. attributed [ARKA22]. Attributes [SENTI24]. attribution [CEM^{+}23]. AU$29.99 [Ano22i]. Audience [NCdlLZ24]. August [Ano23a, Ano24a]. Australia [vESN^{+}20]. Australian [CST^{+}24, TA22, TJB^{+}23]. author [YHW^{+}25]. authority [SFC20]. autocorrelation [NDS^{+}20]. Automated [LWRZ^{+}21, ZSVK21, EM21, WLB^{+}23]. availability [LWC^{+}24]. average [MWW^{+}24]. avian [AAvdW^{+}22, CCL23, GTC21, HRCA21, KSF^{+}24, LWG21]. Ávila [CBD^{+}21]. avoid [LGD^{+}25, POH^{+}22]. Avoiding [GG24, SDH^{+}25]. Awards [Ano20c, Ano21a, Ano22c, Ano24c]. awareness [MGJF21].

B [Cli23]. badges [JMB21]. balance [BBE^{+}25]. Balancing [PKC^{+}22]. Baltimore [Ano23n]. ban [MSRP21]. banks [NSL^{+}23]. bans [WDHR22]. Barcelona [Cli23]. Barrier [TCG^{+}21, WFH^{+}21]. barriers [HSHA20, INO23]. based [Ala23, BKK^{+}21, Bou23, CTP^{+}24, CP20b, CBD^{+}21, CHA^{+}24, Coo24, CCW24, DNCC22, DSF^{+}23, DBT21, FDP^{+}23, GW21, GCSZ21, GMEV^{+}21, GZBP24, HCC^{+}21, HAE^{+}24, HHB23, KGK^{+}24, KBCD20, LWC^{+}24, LKF^{+}23, MOO^{+}24, MHD^{+}22, MVC^{+}21, MR20b, MWH^{+}22, MCC^{+}23a, NJLC21, PAFB^{+}22, PSP^{+}20, PLC22, PLMF23, RSRG^{+}20, RWC^{+}22, SCdM^{+}24, SBG^{+}25, STW^{+}24, SCW^{+}24, THM^{+}21, TA22, WLB^{+}23, WKT^{+}21, WSW23, WJP^{+}20, WSEL^{+}24, WWVC^{+}25, YLY^{+}20, YWH^{+}24, Ano22j]. baseline [HRCA21]. baselines [FCL^{+}22, MR22]. Basin [JTD^{+}20, MWH^{+}22, SCB^{+}25]. bat [ADKM24, CSGG25, DSZ21, HOWM24, KxGG^{+}25, LB22, TSS^{+}22, WFBT^{+}25]. bats [CRC^{+}21, CLD^{+}21, CO24, MJPJ^{+}24, WWVC^{+}25]. Bayesian [PWW^{+}22, VG20]. bean [JPK^{+}21]. bear [HWG^{+}22, MPM^{+}23]. Bears [Ano23n]. Beauséjour [Cli23]. bee [CAFM^{+}23, TG21, Sár23]. bees [Sár23]. Beginner [Ano22i]. behavior [BSG^{+}23, CBG^{+}21, GHR21, JPK^{+}21, KSF^{+}24, LBBT24, LHB^{+}22, MTS^{+}21, MWV^{+}20, MWS^{+}25, NJLC21, Sár23, SBG^{+}22, TWMM^{+}23, WLA20, WSvRLK20, ZWC^{+}24c, dLMGK22]. behavior-change [SBG^{+}22]. behavioral [OVC^{+}24]. behaviors [CHA^{+}24, LTJ^{+}20, NFD20, RHGRSG21]. behind [FWZ^{+}24, Hul22]. being [MHGC21, TB25]. belief [LNYY20]. beliefs [MWS^{+}25]. belonging [CFN^{+}24]. benchmarks [FMGS21]. beneath [MEA^{+}22]. beneficial [ODS^{+}24]. benefit [DS22, PP25, SSKBL21]. Benefits [GPJJ^{+}21, KRB^{+}22, BGD25, GBMD^{+}22, HMC^{+}23b, SBG^{+}22, SSN^{+}23, SSKBL21]. Best [DPC^{+}24, CLD^{+}21]. beta [JBE^{+}22a]. better [FLW^{+}22, HRCA21, PKC^{+}23a, SD20]. between [CAFM^{+}23, CXJ^{+}22, CSHM^{+}20, DW23, GCS^{+}24, GJC^{+}23, HK21, HWG^{+}22, HPM^{+}20, KBB^{+}23, MCF^{+}24, MFW^{+}24, PNXY^{+}22, RRH^{+}24, TAA^{+}24, VM20]. Beyond [MBD^{+}21, CC20, DLL22]. bias [MBH^{+}21]. biased [CAM^{+}21]. biases [HLL^{+}24]. Bibliometric [CHW^{+}23]. big [COOH^{+}25]. bile [HWG^{+}22]. biobanking [HMC^{+}23b]. bioblitz [GFF23]. Biodiversity [Ano23n, PS24, PHB^{+}25, BBGL21, BBH^{+}25, BKK^{+}21, BHG^{+}23, BGAE25, BSSG21, dOCVJ^{+}23, CN22, CHR24, CP20b, CGA^{+}22, CAA^{+}24, CGL^{+}25, CCW24, DNCC22, DFS^{+}25, DM20, DAN^{+}21, DFC21, FGB^{+}23, FZG^{+}24, GBA23, GSS^{+}22, GFF23, GNH^{+}22, HRHB20, HK21, HBB^{+}24, HC22, HWLC22, HSG^{+}21, HMM^{+}24, HGCK22, ITAAG^{+}24, JGAH25, KB21, KCH^{+}20, KFG^{+}20, LC24, Lau20, MBLD^{+}25, MBH^{+}21, MBD^{+}21, MGJF21, MCG^{+}21, MK21, MWH^{+}22, Mol22, MTM^{+}22, NTS^{+}21, NSL^{+}23, NSB^{+}21, NRSF21, PNXY^{+}22, PGRDOGG23, PBB^{+}22, PSO^{+}22, RzEHB24, RJAR23, RLSH24, RF23, RCRJ^{+}24, RBJN21, RMD^{+}20, SP23a, SBH^{+}23, SBG^{+}22, SSN^{+}23, SDB^{+}24, SSCR^{+}23, SdVD^{+}22, SQJ^{+}22, SCW^{+}24, SBW^{+}22, SRCS23, SWBD20, SWC^{+}21, TCK^{+}20, TRR^{+}20, The24, TzEN^{+}24, VDO^{+}24, WGAB^{+}21, WSC^{+}23, WGZ24, ZEO^{+}22]. biodiversity-conservation [RMD^{+}20]. Biodiversity. [Ano22i]. bioenergy [NRSF21]. biogeographic [RDS^{+}24]. Biological [SAB^{+}22, BLF^{+}23, Hul22, MSB25, PPE^{+}20, RAG^{+}20]. Biology [Ano22j, Ano20v, CHP20, Dav20, HPM^{+}20, KTH^{+}22, MHC21, PG20, PSS20, Sag20a, Sag20b, Tay20, Ano24c, WSEL^{+}24, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano22c]. biomonitoring [STF^{+}20]. biophysical [BVH22, BMG^{+}22, PSJG23]. biosecurity [GKH^{+}24, OC24]. biota [IJ24]. biotechnology [AP21]. Biotic [KZM23, FBDA22, SMLL24, TFT^{+}24]. Bird [MSKM21, PKP20, Vin22, Ano24f, AOI22, BSRM22, BPR24b, EHC^{+}21, FMGS21, GGBD20, JPAR^{+}22, JAB^{+}21a, KFU^{+}24, MR22, NHSM^{+}24, ODG^{+}22, PSP^{+}20, RHS^{+}20, TFT^{+}24, VNC^{+}24]. bird-borne [NHSM^{+}24]. Bird-friendly [MSKM21]. Birds [Ano22i, Sco20, AL21, ANR23, BMA^{+}20, CLD^{+}21, DFB^{+}24, Eva21, FLH21, HVA^{+}21, HPG20, MNMPGN23, MBB^{+}23, MHP23, MSMW21, MSHE23, NRN^{+}24, PHB^{+}21, XST^{+}20, Ano22j]. Birdsong [Wal24]. Black [MCW^{+}24]. bleaching [TCG^{+}21]. bloom [DNCC22]. blue [SSKBL21]. boater [KKE^{+}24]. Bob [ÁRAB^{+}24]. body [BMG^{+}22, NFT^{+}20]. body-size [BMG^{+}22]. Bonsall [Ano22j]. Book [Ano20a, Ano22a, Ano23a, Ano24a, Sár23]. Books [Ano20b, Ano22b, Ano23b, Ano24b]. Boosting [INO23, TDC^{+}23, FBG^{+}25]. borders [WYB^{+}21]. Borne [Ano22j, NHSM^{+}24]. Borner [HL20]. both [CvH^{+}22]. Bottom [HBB^{+}24]. Bottom-up [HBB^{+}24]. boundaries [SH21]. boundary [TB25]. Brazil [BGMEHP20, LdSB^{+}20]. Brazilian [MCP^{+}25]. breeding [CCL23, DDM24, HMC^{+}23b, JAB^{+}21a, OAY^{+}24, WFH^{+}21]. bridge [BP21, MMS^{+}20]. Bridging [JSGJ^{+}21, KBM^{+}21, RAG^{+}20, Ric22]. Briggs [Ano22j]. British [Ano24g]. broad [FBDA22]. broad-scale [FBDA22]. broadleaved [QZX^{+}24]. brown [MPM^{+}23]. budget [Nov23]. budgets [WDHM21]. buffers [RCRJ^{+}24]. build [SACA^{+}20]. Building [HRCA21, MBV^{+}24, PWB^{+}21, Rob24, TFT^{+}24]. buildings [EHC^{+}21]. Buisson [Sár23]. Bumblebees [Sár23, MDG^{+}21]. burial [DBR^{+}23]. bushmeat [BF21, NJ20]. businesses [HBB^{+}24]. butterflies [OPF24, RKR^{+}21]. butterfly [PSP^{+}20, vBDD^{+}20]. bycatch [AAS21, CGM^{+}24, TGJ^{+}20].

C [Ano22i, Ano22j, Cli23]. CABI [Ano22i]. cacti [MMG^{+}23, VLPVM^{+}24]. call [BDF^{+}22, OLF24]. calls [LWRZ^{+}21]. Cambridge [Ano22i, Cli23]. camera [CGB^{+}24, GCSZ21, HHB23, VMP^{+}24]. camera-trap [GCSZ21]. Campesino [PS20]. can [BBL^{+}20, CN22]. Canada [KFN^{+}21, WBR^{+}24]. cancers [HMM^{+}21]. Capability [COOH^{+}25, COF^{+}25]. Capacity [COF^{+}25, GCA^{+}22, MO25, TGM^{+}22]. capitalism [MTM^{+}22]. Caption [Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g]. captions [FCH^{+}24]. captive [DDM24, HQY^{+}24, MSHE23]. captive- [MSHE23]. capture [BTH^{+}20, COOH^{+}25]. carbon [BBE^{+}25, MLIE25, Nov23]. caribou [CvH^{+}22, DSS^{+}23]. carnivore [DHE^{+}21, JDM^{+}21, WMM^{+}20]. carnivores [ARKA22, DAD23, RRR^{+}24, SHS^{+}22, TF24]. carrageenophyte [KCS20]. carrion [FBS^{+}23]. cascades [HBL20]. case [Cie24, CTM^{+}22, CGL^{+}25, HOWM24, SKS22, SSN^{+}23, WSEL^{+}24]. cat [FWZ^{+}24, LGD^{+}25, POH^{+}22]. Catalyzing [FDP^{+}23, MWV^{+}20]. categories [Huy23]. category [Van23]. Cats [Moo21, LGL20, LSÁH^{+}20, MMN^{+}20]. cave [MCLM22, WWVC^{+}25, Ano23n]. Celebrating [Pic20]. celebrities [BAGV24]. celebrity [OVC^{+}24]. celebrity-endorsed [OVC^{+}24]. cemeteries [IJ24]. center [MBK^{+}20]. centering [WAB^{+}24]. Central [PPN^{+}24, EMV^{+}20, GTAdPA24, MMN^{+}20, SPA^{+}22]. century [DPC^{+}24]. cetacean [LMG^{+}23]. cetaceans [Sco22, TGJ^{+}20]. challenge [BOAN23, CAM^{+}21]. Challenges [ARSL24, ART^{+}25, BDF^{+}22, FWZ^{+}24, NBB^{+}21, TFJ^{+}25, ZEO^{+}22, CND^{+}23, CCPE^{+}25, CCCB^{+}23, GBB^{+}20, HOT^{+}25, MSS^{+}24, NCE^{+}21, Sco22, SB21, VFKT24, WSW^{+}22]. champion [ÁRAB^{+}24]. Change [GTC21, Ala25, ADKM24, BCA^{+}21, CN21, CS21, CVT^{+}21, CHA^{+}24, CTM^{+}22, CCKT25, DAD23, EN24, EM21, GWR23, HRCA21, HBY^{+}23, HMS^{+}23, HCZW24, JZLP^{+}23, KMC^{+}23, KxGG^{+}25, KSFP^{+}21, LNYY20, LdSB^{+}20, LHB^{+}22, LWC^{+}24, LCEM24, LLYK24, LWS^{+}24, MHJD23, MSC^{+}24, MDB^{+}24, MWV^{+}20, MBH^{+}21, MSRP21, NRSF21, PKP20, RML^{+}20, SP23a, SBG^{+}22, SPK^{+}22, TWMM^{+}23, YCS^{+}23, ZWC^{+}24c, dLMGK22]. changes [BJP^{+}20, LNDM21, LMG^{+}23, MSMH22, WB23]. Changing [FLW^{+}22, Gon21, AMC^{+}24, BMM^{+}23, GW21, HSN^{+}21, KKKS^{+}20, LC24, MTC^{+}20, MLS^{+}22, TGM^{+}22]. Characteristics [MCB^{+}20, GJPJ^{+}22, ZYM^{+}24]. characterize [LH21]. Characterizing [BRW21, EGBJ23, HGM^{+}22]. Chaudhary [LFT^{+}21]. chestnut [AP21]. Chicago [Ano23n]. Child [GN20]. children [BZT^{+}22]. China [FLW^{+}22, HWG^{+}22, HZW^{+}21, KxGG^{+}25, LMY^{+}24, LLYK24, MHL^{+}22, WYB^{+}21, ZTCF21, ZWC^{+}24c, ZYM^{+}24, ZZ24, ZSL24b]. Chinese [GLL^{+}23, MMN^{+}20]. Chinese-led [MMN^{+}20]. Chittka [Sár23]. choice [NJ20]. choices [Des21, YHW^{+}25]. choosing [MVC^{+}21]. choruses [LWRZ^{+}21]. chytrid [FW22]. circuit [KWY^{+}21]. CITES [AFV20, RR24, VFRV24, WSSC22]. CITES-listed [RR24]. cities [VGW^{+}25]. Citizen [JJ21, FSBE22, GRP^{+}25, PSP^{+}20, VBK^{+}23]. citizenship [JJ21]. city [TFT^{+}24]. clam [LMY^{+}24]. Clark [SÁT21]. class [LDD^{+}22]. Classification [KBB^{+}24, Cas23, MCH^{+}20, SD22]. classifying [RHFV23]. Cliffs [MSLBS23]. Climate [BBM^{+}20, HGCK22, Ala23, Ala25, AMC^{+}24, Ano24f, BJP^{+}20, BPR24b, BCA^{+}21, BMM^{+}23, CN21, CS21, CTM^{+}22, EN24, GGJ^{+}20, GPJJ^{+}21, GJPJ^{+}22, GWR23, GW21, GBDM^{+}20, GMEV^{+}21, HSN^{+}21, HBY^{+}23, HMS^{+}23, HCZW24, JZLP^{+}23, KMC^{+}23, KFU^{+}24, KxGG^{+}25, KTG^{+}22, LdSB^{+}20, LWC^{+}24, LLYK24, MHJD23, MSC^{+}24, MSMH22, MDB^{+}24, MHL^{+}22, NRN^{+}24, PSV^{+}21, RJAR23, RML^{+}20, Ric21, SP23a, STSJ22, TGM^{+}22, YLY^{+}20, YCS^{+}23, ZBM^{+}24]. climate-change [EN24, LdSB^{+}20]. climate-driven [Ala23, BJP^{+}20, MSMH22]. climates [MCLM22]. climatic [DAD23, FBDA22]. close [BNS24]. Closing [MCF^{+}24, ROK^{+}22]. cloud [PKP20]. co [Bus22, SWCP^{+}22, VMP^{+}24]. co-management [SWCP^{+}22]. co-occurrence [VMP^{+}24]. co-productive [Bus22]. Coast [TW24]. coastal [CXJ^{+}22, CGL^{+}25, DYM25, GRS^{+}23, HMC23a, LNYY20, LEF^{+}24, MSMW21, WWB^{+}22, WSK^{+}22]. coastlines [LLYK24]. cobenefits [RCRF20]. cocoa [BSRM22, MR22]. Coexistence [Cor20, CBD^{+}21, GCS^{+}24, HHMO24, JDM^{+}21, JRKF21, JSKR22, JS25, JSA^{+}20, KKKS^{+}20, PBV21, Poo25, SÁT21, TSÁ20]. cofiltering [FBDA22]. cognitive [BTDS20, DRFL^{+}24]. cold [CS21, HTS^{+}22, MHJD23]. cold-adapted [CS21]. cold-water [HTS^{+}22]. collaboration [HSHA20, MW25]. Collaborative [CBD^{+}21]. collapse [AMC^{+}24, DTV^{+}25, KBB^{+}23]. collected [DSZ21]. collecting [DMR^{+}21, MTH^{+}24]. collection [FWP^{+}21]. collections [ADG^{+}20, DH21]. Collective [MWW^{+}24, VDO^{+}24]. collector [MMG^{+}23]. collision [CSGG25, SDH^{+}25]. collisions [EHC^{+}21, TFT^{+}24]. colonial [SBP^{+}23]. colonization [FGL^{+}21]. comanagement [HCC^{+}21]. combat [KTG^{+}22, MTS^{+}21]. Combating [Uts24]. Combined [GZBLS^{+}20, HLZ^{+}21, MSMH22]. Combining [CGB^{+}24, JDB^{+}21, KAD^{+}22, VMR21, KBL^{+}23]. comfort [HSN^{+}21]. commercial [ARSL24, DDM24]. common [DBC^{+}23, CCCB^{+}23]. commonly [ZBM^{+}24]. commons [McC24, MO25]. Communicating [WSW^{+}22, MR20a]. communication [PGRDOGG23]. communities [BJP^{+}20, BHSM^{+}22, FOYZ22, HCZW24, ICMO23, KZM23, PRS^{+}25, Ric21, ZSL^{+}24a]. Community [BGB^{+}24, GWR24, BSRM22, BAGK22, BAA^{+}25, CHA^{+}24, CFN^{+}24, DFS^{+}25, FDP^{+}23, GGJ^{+}20, KBCD20, LNYY20, MMG^{+}23, MCP^{+}25, PKP20, PLD^{+}21, SBW^{+}22, SOW^{+}22, TAA^{+}24, YWH^{+}24]. community-based [CHA^{+}24, FDP^{+}23, KBCD20, YWH^{+}24]. community-level [BAGK22]. community-sourced [SOW^{+}22]. Company [Ano23n]. comparative [LDC^{+}24]. Comparing [HHB23, MCC^{+}23a]. compassion [BA22a, CMAS20]. compassionate [BA22a, CC22, WBB^{+}20]. compensation [RzEHB24]. competition [BOB^{+}20, CHA^{+}24]. competitive [BB21]. complementarity [PHB^{+}25]. complementary [SGB^{+}20]. complements [RMW23]. Complex [BJP^{+}20, CQS^{+}22, AHM^{+}22, MB21]. complexity [HWK^{+}20]. Compliance [TRR^{+}20, IDJ^{+}25, TWS^{+}25]. component [LFT^{+}21, SQJ^{+}22]. components [TTDdLM^{+}21]. composition [BSRM22, HMS^{+}23, SBW^{+}22]. comprehensive [LHB^{+}24]. computational [HWK^{+}20]. computer [VHHT21]. concept [Coo24, Per20]. concepts [DFC21, MSPM20]. Conceptual [Ano23d, CMC^{+}23]. concern [LSM^{+}20]. concerns [DFH^{+}21]. concordance [MFW^{+}24]. conditions [BVH22, PKC^{+}23a]. Conducting [WDGM21, DBT21]. conferences [KRB^{+}22]. configuration [TG21]. conflict [ARKA22, FWZ^{+}24, GCS^{+}24, JDB^{+}21, LZG^{+}25, LTJ^{+}20, PMLM20, THSA24, TSÁ20, WLR^{+}24, WMM^{+}20]. conflicting [LGL20]. conflicts [AOI22, BHRBK20, CLB20, May22, McC24, SP22, TTDdLM^{+}21]. Congo [DAN^{+}21]. Connecting [Gon21]. connection [MBL^{+}23, STA^{+}22b, WLA20]. connections [GCS^{+}24, KKE^{+}24]. Connectivity [GJC^{+}23, CQS^{+}22, DSF^{+}23, INO23, MRS22, MCC^{+}23a, MCC^{+}23b, RCBT21, SNC^{+}22, SD20, SMSS^{+}21, WSB^{+}20]. Consensus [SVP^{+}22, CBD^{+}21]. consensus-based [CBD^{+}21]. consent [LC24]. Consequences [SMSS^{+}21, APE^{+}21, MPM^{+}23, PCH22]. Conservation [Ano24d, Bla24, Cor21, DSvE^{+}25, Far21, GBDM^{+}20, Gon21, IB20, Imb22, MSB25, POH21, Poo25, SPA^{+}22, Sch21, Vin22, VD20, WYB^{+}21, WH22, WAA^{+}21, ZXXL23, ABRCM21, Ala23, ADSG21, AL21, ÁRAB^{+}24, And25, Ano20v, Ano23d, Arm23, AP23, ANR23, ART^{+}25, BWN20, BDG^{+}23, BNW20, BNB^{+}21, BBH^{+}25, BBM^{+}20, BH23, BHM20, BGMEHP20, BGB^{+}24, BKK^{+}21, BHG^{+}23, BTDS20, BGAE25, BA22a, BBG^{+}20, Bou23, BBL22, BGKKC22, BIdL^{+}20, BDQ^{+}25, BHSM^{+}22, BHM21, BSSG21, BAA^{+}25, BLW23, Bus22, BBL^{+}20, dOCVJ^{+}23, CMP^{+}24, CC20, CACM23, CVT^{+}25, CBM20, CCPE^{+}25, CLB20, CEM^{+}23, CTP^{+}24, CP20b, CMC^{+}23, CXJ^{+}22, CGM^{+}25, CGA^{+}22, CFA^{+}24, CAM^{+}21, CSHM^{+}20, Cie24, CBG^{+}21, CBD22, CHA^{+}24, CC22, CC24, CHW^{+}23, CFN^{+}24, Coo24, CLJ^{+}21, CMAS20, CCW24, DNCC22, DS25, DTSE25, DBR^{+}23, DW23, DFS^{+}25, Des21, DdLI^{+}25]. conservation [DFH^{+}21, DBC^{+}23, DWC^{+}20, DEGL^{+}24, Dre24, DSC^{+}24, DFC21, EKH23, ECVL25, EM21, EGCS25, FDP^{+}23, FOYZ22, FFAS23, FR21, GCC^{+}22, GW21, GGB^{+}20a, GSS^{+}22, GRS^{+}23, GNH^{+}22, GAB^{+}22, GB22, GR21, GAB^{+}21, GN20, GZBP24, GSMW21, HSN^{+}21, HOT^{+}25, HRCA21, HVVA^{+}21, HSSM^{+}25, HCD^{+}20, HCA^{+}22, HtSR25, HWLC22, HPM^{+}20, HAE^{+}24, HCJH^{+}22, HL20, HMM^{+}24, HTM25, HMC^{+}23b, HLL^{+}24, HPG20, HZW^{+}21, HSHA20, Huy23, HGCK22, IDD^{+}22, IJ24, JBB23, JPAR^{+}22, JAB^{+}21a, Jen22, JTD^{+}20, JCD^{+}23, JP21, JBE^{+}22b, JAB^{+}21b, KBM^{+}21, KTH^{+}22, KBW^{+}23, KCH^{+}20, KOO^{+}21, KvZS^{+}25, KIA^{+}22, KRB^{+}22, KFG^{+}20, KLVW20, LNYY20, LKL^{+}21, LS21, Lau20, LB22, LVT^{+}20, LGL20, LMG^{+}23, LWG21, LHB^{+}24, LWC^{+}24, LLYK24, LKF^{+}23, LFD^{+}25, MS23, MOO^{+}24, MTC^{+}20, MHD^{+}22, MTS^{+}21, MGS^{+}23, MBD^{+}23, MBLD^{+}25, MWY^{+}21, MHGC21, MCF^{+}24, MWC^{+}20, MWS^{+}25]. conservation [MJPJ^{+}24, MHSL^{+}23, MBL^{+}23, MGJF21, MCG^{+}21, MBK^{+}20, MCM^{+}21, MCW^{+}24, MTM^{+}22, MW25, ME25, MMHR20, MRS22, MCC^{+}23a, MCC^{+}23b, MSN^{+}22, MEK21, MEA^{+}22, NCdlLZ24, NBB^{+}21, NAG^{+}20, NJLC21, NFD20, NLV20, NCE^{+}21, NBD^{+}24, OGF^{+}21, OVC^{+}24, OCC^{+}21, ODS^{+}24, PNXY^{+}22, PAFB^{+}22, PS24, PHB^{+}21, PCH22, PHMVL^{+}22, PMS^{+}20, PS20, PRS^{+}25, PSS20, PKC^{+}22, PKC^{+}23b, PMR^{+}22, PHB^{+}25, PWW^{+}22, PSO^{+}22, Ras22, RW23, Red23, RGBC20, Ric22, RF23, Rob24, ROK^{+}22, RMW23, RHS^{+}20, RMD^{+}20, SKS22, SCdM^{+}24, SBG^{+}20, SC23, SMB^{+}20, San20, SGB^{+}20, SS23, SEB^{+}20, SNC^{+}22, SBP^{+}23, Sco24, SBG^{+}22, SJRG23, SSG^{+}23, STW22, SDH^{+}25, SF23, SWMM24, SD20, SHA^{+}23, SGA^{+}24, SZH^{+}23, SOW^{+}22, Sra23, SFC20, SENTI24, SN22, SB21, STA^{+}22b, Sul24, SACA^{+}20, TDC^{+}23, Tay20, TCK^{+}20, TWMM^{+}23, TWGGB^{+}25, THM^{+}21]. conservation [TFJ^{+}25, TA22, TzEN^{+}24, TENI23, TB25, TB23, TFFM22, VG20, VFKT24, VBC^{+}25, VPS22, WLB^{+}23, WBB^{+}20, WLB^{+}20, WL20, WCY^{+}25, WRW^{+}20, WDGM21, WKT^{+}21, WJ21, WSW23, WSC^{+}23, WJP^{+}20, WSB^{+}20, WDH^{+}21, WSvRLK20, WSEL^{+}24, WBC^{+}20, WDHM21, XST^{+}20, XAAY^{+}21, XDA^{+}21, YCS^{+}23, YWH^{+}24, YMB^{+}21, YHW^{+}25, YVA22, ZBM^{+}24, ZWC^{+}24a, ZAS^{+}24, ZSVK21, dG21, dLMGK22, dlLBB^{+}23, vESN^{+}20, Ano24c, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano22c]. conservation-relevant [CBG^{+}21]. conservationist [PKC^{+}23a]. Conservationists [Moo21]. conserve [BZT^{+}22, CS21, MNMPGN23, SBH^{+}23]. conserved [KGK^{+}24, WBR^{+}24]. Conserving [GGB20b, Gru20, JBE^{+}22a, MBB^{+}23, Val20, GBRK^{+}20, KSZ^{+}23, MBD^{+}21]. consideration [FCL^{+}22, SKS22]. considerations [BIdL^{+}20, GBB^{+}20, HTM25, MHSL^{+}23]. Considering [BE23]. consolidating [And25]. constellations [GCS^{+}24]. constraints [TFFM22]. Consumer [ZYM^{+}24, ZWC^{+}24c]. consumers [HWG^{+}22]. consumption [CVT^{+}21, EMV^{+}20, LH21, TWHB^{+}21]. contagion [BO21]. contaminant [DSM24]. contamination [MHR^{+}22]. contemporary [BWN20, BPN^{+}24, HZW^{+}21]. content [LMG^{+}23]. Context [Gru21, ARK^{+}20, KxGG^{+}25, LFT^{+}21, MTC^{+}20, RvTW22, SMB^{+}20]. context-dependent [RvTW22]. contexts [CHW^{+}23]. contextualizing [WMM^{+}20]. continental [BOAN23, KFN^{+}21]. continuity [KSZ^{+}23]. contrast [WJ21]. Contrasting [LMG^{+}23, RCRJ^{+}24]. contribute [SP23b]. contributed [QOF^{+}21]. Contribution [DBR^{+}23, HTS^{+}22, Ano23d, CMC^{+}23]. Contributions [KGK^{+}24, FSBE22, VLVD^{+}23]. Control [Ano23n, BGD25, MJDMAS20, POH^{+}22, SJRG23, WJP^{+}20]. controversy [Hen22, SVP^{+}22]. Convention [Pit21, SAB^{+}22]. Convergence [TCG^{+}21]. conversion [BNI25, GGJ^{+}20, IMI^{+}21, MWC^{+}20]. Cooling [STSJ22, TSS^{+}22]. cooperation [BZT^{+}22]. coordination [McC24]. COP15 [CHR24]. COP16 [And25]. coproduction [BMM^{+}23, CACM23, VLVD^{+}23]. Copyright [Ano22e, Ano23i, Ano23j]. coral [DBM^{+}25, KTG^{+}22, MDB^{+}24, MHS^{+}21, MSRP21, QHS22, TCG^{+}21, TFFM22]. Correlates [EHC^{+}21, EN24, KxGG^{+}25]. Corridor [Ano22j]. corridors [BRW21, GH22, SFC20]. Cost [TGJ^{+}20, LKF^{+}23, MBI^{+}22, SAB^{+}22, SSKBL21, WRW^{+}20]. Cost-effective [TGJ^{+}20]. cost-effectively [WRW^{+}20]. cost-sharing [LKF^{+}23]. Costa [KSF^{+}24]. costs [Arm23, BAB^{+}23b, CBB^{+}20, GBRK^{+}20, GBMD^{+}22, HMC^{+}23b]. counteract [YTW^{+}25]. counterfactual [JAB^{+}21a]. counterfactuals [BSSG21, HAE^{+}24]. countries [EHC^{+}21, GJC^{+}23]. country [EOR^{+}24, MSKM21]. counts [KBL^{+}23, MBD^{+}21, WLB^{+}20]. Coupling [KWY^{+}21]. Courchamp [PG20]. Covelo [Ano22i]. Cover [Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, HRHB20, KAD^{+}22, LKA^{+}22, LDD^{+}22, SDB^{+}24, SUA22, YCS^{+}23]. coverage [GLL^{+}23, HWP^{+}21, MNMPGN23, MSJ20, WCY^{+}25]. COVID [BBL^{+}20, GSS^{+}22]. COVID-19 [BBL^{+}20, GSS^{+}22]. crayfish [OC24]. create [HRCA21]. created [MMHR20, SBP^{+}23]. Creating [GCS^{+}24, HHM24, MGS^{+}23, GKH^{+}24]. creative [LGD^{+}25]. Crespin [LSÁH^{+}20]. crime [ABRCM21, GB22, ME25, UEH^{+}23]. criminal [ME25]. criminology [ME25]. crises [MR20a]. crisis [GBA23, SB21]. criteria [RF23]. Critical [CWMB23, BRW21, Cas23, CO20, CC22, DWC^{+}20, QOF^{+}21, RGBC20]. critically [BSG^{+}23, FMGS21, HHL^{+}24]. crop [JDB^{+}21, NRSF21, SR20]. cropland [BNI25, Gok21]. cross [RVLD20]. cross-scale [RVLD20]. crossing [LEF^{+}24]. crossroads [Poo25]. crowding [CBG^{+}21, VPVdAMF20]. Cuban [DBT21]. Cultural [VFKT24, DBR^{+}23, GRS^{+}23, LGL20, ZZ24]. culture [Jen22, LVT^{+}20]. culturomic [VMR21]. Culturomics [JBC^{+}21, CLJ^{+}21, Huy23, JCD^{+}23, LH21, MGJF21]. cumulative [HTC22]. cumulative-effects [HTC22]. curation [CTL^{+}23]. current [MHL^{+}22, RCBT21, YHW^{+}25]. Cuthbert [Sag20b].

D [Ano22j, CCKT25, KWY^{+}21]. damage [RGJCN^{+}23, SR20]. damages [JDB^{+}21]. dangers [EDA^{+}21]. Darling [MWH^{+}22]. Darwin [HMM^{+}21]. Data [PSO^{+}22, APE^{+}21, BBH^{+}25, CSL^{+}23, CAA^{+}24, COOH^{+}25, CLJ^{+}21, DSF^{+}23, DSZ21, DS22, DFH^{+}21, DH21, DMR^{+}21, EOR^{+}24, FOYZ22, FSC^{+}21, GFF23, GCSZ21, GSB23, GWB^{+}24, HOT^{+}25, HSG^{+}21, HOWM24, KBL^{+}23, LH21, LFD^{+}25, MSN^{+}22, PSP^{+}20, POH^{+}22, PWW^{+}22, QOF^{+}21, RML^{+}20, RGBC20, STW^{+}24, VRBAB^{+}24, XDA^{+}21]. data-deficient [CSL^{+}23, FSC^{+}21]. data-driven [GWB^{+}24]. data-poor [DS22, POH^{+}22, RML^{+}20]. Deal [CCCB^{+}23]. Debate [PG20, CC22, CC24, EGBJ23]. debt [MWC^{+}20]. decade [HSG^{+}21]. December [Ano23a, Ano23b, Ano20a, Ano24a]. Decision [LFT^{+}21, WBJ^{+}20, BH23, BHG^{+}23, BHM21, DMR^{+}21, EHBT20, FSC^{+}21, GGLB25, GKH^{+}24, GBB^{+}20, HCC^{+}21, HCA^{+}22, KBB^{+}24, MHM^{+}24, MCC^{+}23b, MBI^{+}22, NJ20, RBC^{+}22, TF24, WGAB^{+}21]. decision-making [BH23, EHBT20, FSC^{+}21, GGLB25, GKH^{+}24, KBB^{+}24, MCC^{+}23b, RBC^{+}22]. Decision-support [WBJ^{+}20]. decision-tree [TF24]. decisions [BMM^{+}23, HSN^{+}21, MEK21]. decline [NSB^{+}21, QZX^{+}24, SPA^{+}22].

Decolonizing [dG21]. decrease [GG24]. deep [LNYY20, MVC^{+}22, MBLD^{+}25, Sra23, ZSVK21]. deep-belief [LNYY20]. deep-sea [MBLD^{+}25]. deepening [GB22]. Deepwater [GGBD20, SMT^{+}22]. Defaunation [BBE^{+}25, WB23]. defaunation-induced [WB23]. deficiency [HSG^{+}21]. deficient [CSL^{+}23, FSC^{+}21]. Defining [HC22, Kho22]. definitions [CHP20]. deforestation [GCJB^{+}22, MCP^{+}25, MBL^{+}23, SPK^{+}22, TWS^{+}25]. deforested [FPC^{+}24]. degazettement [YTW^{+}25]. degradation [GGSRF20, GCJB^{+}22, ZYNC25]. deliver [BE23, MK21]. DeLong [Ano22j]. demographic [JSGJ^{+}21, RMFP^{+}23]. demographically [JPAR^{+}22]. denialism [LSM24]. dense [LNDM21]. Dependence [Ala25]. dependency [GGB^{+}20a, MCC^{+}23a]. dependency-based [MCC^{+}23a]. dependent [Gok21, IMI^{+}21, NAG^{+}20, RvTW22]. depicting [FCH^{+}24]. depletion [EOR^{+}24]. depopulating [TE20]. depopulation [Ano24f, KFU^{+}24]. deposition [RKR^{+}21]. deriving [HAE^{+}24]. described [RR24]. desert [LB22, OAY^{+}24, ZWC^{+}24b, VLPVM^{+}24]. Design [HTM25, KGK^{+}24, NDS^{+}20, SJRG23]. designations [MFW^{+}24]. designed [SLM^{+}20]. Designing [OVC^{+}24, VGW^{+}25]. designs [Kho22]. despite [OAY^{+}24, WSW23]. destruction [MVC^{+}21]. detect [MVC^{+}22]. detection [EM21, FOYZ22, HGM^{+}22, KAD^{+}22, LWRZ^{+}21, RCL^{+}22, VMP^{+}24]. detections [HHB23]. Determinants [BP21]. determine [KBL^{+}23, MR22, RSRG^{+}20, WTAD21]. Determining [RR24, WSvRLK20, PMLM20]. Developing [MB21, Coo24, Ric22]. Development [SWCP^{+}22, AOI22, BNI25, KWY^{+}21, LEF^{+}24, MMN^{+}20, NSL^{+}23, PNXY^{+}22, VNC^{+}24, YLY^{+}20]. device [EPL^{+}24]. devices [CFC21]. devil [HMM^{+}21]. dichotomy [CO21, DLL22]. Dickie [Ano23n]. difference [PKC^{+}22]. differences [CSHM^{+}20]. different [FCL^{+}22, SP23b, TJB^{+}23]. differential [GGSRF20]. differentiating [ZWC^{+}24a]. differentiation [KZM23]. difficulty [Hen22]. diffusion [NJLC21]. Digital [CLJ^{+}21, RGJCN^{+}23]. dilemma [CXJ^{+}22]. dimensional [MBLD^{+}25]. dimensions [CCW24, DYM25, FWZ^{+}24]. directing [CCKT25]. directions [DWC^{+}20, VBC^{+}25]. disappearance [Cli23]. disciplinary [NWSL^{+}24, WSEL^{+}24]. discipline [SVP^{+}22, SB21]. disciplines [HWLC22, LGD^{+}25]. discourse [SP21]. discovered [KSC^{+}20]. Disease [Ano22j, BGD25, LKL^{+}21, MHM^{+}24, PMS^{+}20, RGD^{+}25, STSJ22, WFBT^{+}25]. Diseases [Ano22j]. Disentangling [TFT^{+}24]. dispersal [SMSS^{+}21, WB23]. Disrupted [PSV^{+}21, WH22]. distal [McC24]. distance [DEGL^{+}24]. distribution [BF21, HZW^{+}21, LMG^{+}23, LXZ^{+}22, MHRF20, MHL^{+}22, MEK21, RVLD20, SCB^{+}21, TDC^{+}23, ZBM^{+}24, ZSL24b]. distributions [CGB^{+}24, GPJJ^{+}21, GZBLS^{+}20]. disturbance [ANR23, LHB^{+}22, QZX^{+}24, SBW^{+}22, VPZ25]. disturbances [BRJ25, SCW^{+}24]. dive [Sra23]. divergence [ZAS^{+}24]. Divergent [MHJD23]. diverse [CC20, DFS^{+}25, WMK^{+}21]. Diversification [MDB^{+}24, ODG^{+}22]. diversified [MSKM21]. diversify [JAB^{+}21b].

Diversifying [Ano24d]. Diversity [Ano23n, Kut20, BSRM22, CAFM^{+}23, CSHM^{+}20, COOH^{+}25, CTL^{+}23, DFC21, GGSRF20, GR21, HVVA^{+}21, HMC23a, IMI^{+}21, JTD^{+}20, JBE^{+}22a, KMC^{+}23, LB22, LHB^{+}22, MCF^{+}21, MBW^{+}22, MNMPGN23, NWSL^{+}24, PMSC22, SHH^{+}23, WSEL^{+}24, SAB^{+}22]. DNA [KAD^{+}22, MCF^{+}21, STF^{+}20, YCH^{+}24]. documentary [ABC^{+}23]. Documenting [Cli23]. dog [SBG^{+}25]. dogmas [FFAS23]. dogs [BHM20]. Dolins [Cli23]. dolphins [BTM^{+}23]. domains [RAG^{+}20]. domestic [SQJ^{+}22]. dominance [VPZ25]. dominant [QZX^{+}24]. dominated [SNC^{+}22]. doses [Nov23]. downgrading [YTW^{+}25]. Downs [Ano22i]. downsizing [YTW^{+}25]. Drake [Ano22j]. drifting [CFC21, EPL^{+}24]. driven [Ala23, BJP^{+}20, CSS^{+}24, GWB^{+}24, MSMH22]. Drivers [VBK^{+}23, FGL^{+}21, HGAB^{+}24, ITAAG^{+}24, LTJ^{+}20, McC24, RMFP^{+}23, RHFV23, TFT^{+}24, TW24]. drives [San20]. Driving [Val20, Hul22, PSJG23, QZX^{+}24, SSKBL21]. drought [vBDD^{+}20]. Dryland [Ano22j]. due [LMY^{+}24, VPVdAMF20]. Dugatkin [Ano23n]. during [HGM^{+}22, JPAR^{+}22, LBBT24, OAY^{+}24]. Dynamic [Val20, CCKT25, HVA^{+}21, OGMMG22, SBE25, WBJ^{+}20, WLR^{+}24]. dynamics [ABM^{+}22, BGB^{+}24, DTSE25, FGL^{+}21, PKP20, RVLD20, Ano22j].

Earth [EM21, Red23, WMK^{+}21]. earthworms [ZBM^{+}24]. East [QZX^{+}24, TW24]. Eco [Ano22j]. Eco-Evolutionary [Ano22j]. ecocentrism [TCK^{+}20]. Ecological [BHSM^{+}22, SdVD^{+}22, SUA22, Veg24, ARK^{+}20, AOI22, BSSG21, CWMB23, CCKT25, DMR^{+}21, HWK^{+}20, HVA^{+}21, KWY^{+}21, LMG^{+}23, LHB^{+}24, MHM^{+}22, PS24, PHK^{+}25, RzEHB24, RCBT21, RAG^{+}20, SCB^{+}21, TFT^{+}24, TTDdLM^{+}21, TJC21, VRBAB^{+}24, WMM^{+}20, WKG^{+}21]. ecologically [CRAN^{+}24, KBCD20, ODS^{+}24]. Ecology [Ano22i, Ano22j, Gon20, BGMEHP20, HPM^{+}20, LKL^{+}21, PSS20, RMN^{+}21, SD22]. Economic [BAB^{+}23b, BBGL21, GW21, PNXY^{+}22, PP25, SdVD^{+}22, TBJ^{+}24]. ecosystem [AMC^{+}24, DL24, DTV^{+}25, DSC^{+}24, GTC21, GZBP24, HRHB20, ITAAG^{+}24, KBB^{+}23, LNDM21, LHB^{+}24, MMA^{+}22, ML22, MWH^{+}22, MSRP21, POH^{+}22, PSV^{+}21, RSRG^{+}20, RAG^{+}20, SH21, SBE25, SCC23, TRR^{+}20, TJE^{+}24, VFKT24, WSB^{+}20, XAAY^{+}21, Ano22j]. ecosystem-based [GZBP24, RSRG^{+}20, Ano22j]. ecosystem-services [RSRG^{+}20]. ecosystems [BMH^{+}24, HTC22, MSLBS23, PB21, SP23b, TDS^{+}23, ZWC^{+}24b, KGK^{+}24]. Edge [HBL20, GBDM^{+}20, HBJ^{+}20, VRY^{+}23]. Edicions [Cli23]. Editing [AR21, PSS20]. edition [Ano22i]. Editions [Sár23]. editor [Ano23n]. editors [Ano22i, Ano22j, Ano23n, Cli23]. education [BGKKC22, RLSH24, SS23]. Edwards [NAS21b]. Effect [CDSB23, NJ20, Van23, VPZ25, GGJ^{+}20, MWS^{+}25, vBDD^{+}20]. Effective [Ala23, MJPJ^{+}24, AAS21, CND^{+}23, CHR24, CTP^{+}24, Coo24, DS25, HC22, KFG^{+}20, Ric22, TGJ^{+}20]. effectively [WRW^{+}20]. Effectiveness [MMS^{+}20, RMD^{+}20, SLM^{+}20, FCL^{+}22, GCJB^{+}22, MSMW21, NDS^{+}20, RW23, RDS^{+}24, TJC21]. Effects [ABCH^{+}21, BDG^{+}23, BMG^{+}22, CLM22, CHA^{+}24, CCL23, FGB^{+}23, FCH^{+}24, GGBD20, HRHB20, HMC23a, IJK^{+}21, JZLP^{+}23, KFU^{+}24, KSF^{+}24, KBB^{+}23, LKA^{+}22, LEF^{+}24, LBBT24, LSM^{+}20, LZG^{+}25, MHC21, MCP^{+}25, MBH^{+}21, MHL^{+}22, MSMW21, NDS^{+}20, NFT^{+}20, NRSF21, PSP^{+}20, RHGRSG21, RGD^{+}25, SS24, STA^{+}22b, SSEN20, TJB^{+}23, TG21, VPVdAMF20, VLPVM^{+}24, WLR^{+}22, WCY^{+}25, YCS^{+}23, ZBM^{+}24, ZTCF21, ZWC^{+}24b, ZSL^{+}24a, dLMGK22, Ano23c, ART^{+}25, BGVK20, BHS^{+}22, CLGP23, CLD^{+}21, CQS^{+}22, GZBLS^{+}20, GWR23, GGSRF20, GCS^{+}21, HBL20, HBJ^{+}20, HLZ^{+}21, HTC22, ICMO23, JM21, JP21, LB22, MHJD23, MWW^{+}24, MDG^{+}21, MRTV24, NTS^{+}21, ODG^{+}22, OCC^{+}21, RKR^{+}21, SMT^{+}22, VRY^{+}23, Ano24f]. Efficacy [NJLC21, SBP^{+}23, SBG^{+}25]. Efficacy-based [NJLC21]. efficiency [LLYK24]. efficient [SGB^{+}20]. effort [BF21, RRH^{+}24]. efforts [LWC^{+}24, SENTI24, XST^{+}20]. Efrat [SGB^{+}20]. egg [MHJD23]. egg-laying [MHJD23]. eggs [GGBD20]. Eight [Ano23n]. elapid [ARSL24]. Eleanor [BAB^{+}23a]. elephant [CVT^{+}25, GLL^{+}23, SR20, VPS22]. elephant-induced [SR20]. elephants [BNS24]. elevations [MHP23]. Eliciting [DFS^{+}25]. Elinor [MOO^{+}24]. elusive [BTH^{+}20]. Embedding [YWH^{+}24]. embryonic [GGBD20]. embryos [MWW^{+}24]. emergence [MHSL^{+}23]. emergency [PMS^{+}20]. Emergent [CBM20]. Emerging [DYM25, CXJ^{+}22, RBLN21, WDHM21]. Emotion [BNB^{+}21, MHC21]. empirical [YLY^{+}20]. empirically [APE^{+}21]. Empowering [DM20]. empowerment [PRS^{+}25]. encourage [HHMO24]. end [SF23]. Endangered [FSC^{+}21, NCE^{+}21, SENTI24, TENI23, BSG^{+}23, CvH^{+}22, CS24, EPL^{+}24, FMGS21, GZBP24, HHL^{+}24, JP21, OAY^{+}24, PWW^{+}22, VRY^{+}23, ZXXL23]. endemic [HPG20, JZLP^{+}23, PPN^{+}24, RWC^{+}22]. endemic-rich [PPN^{+}24]. endorsed [OVC^{+}24]. energy [CLD^{+}21, VNC^{+}24]. enforcement [SGA^{+}24]. engagement [BGB^{+}24, CHR24, CWMB23, DW23, ITAAG^{+}24, SWMM24]. Engaging [RLSH24, STES23]. Engine [EM21]. engineering [MCLM22, San20]. English [CGA^{+}22]. enhance [CWMB23, GZBP24, HOWM24, MW25]. Enhancing [HK21, TB23, WSEL^{+}24]. ensemble [PB21]. Ensuring [SACA^{+}20]. entities [AHM^{+}22]. envelopes [GBDM^{+}20]. Environmental [KCS20, KBB^{+}24, MAC20, ABRCM21, ART^{+}25, BO21, BPH^{+}22, DH21, FBDA22, GWR24, GBB^{+}20, Hal22, Hyd24, JJ21, KSFP^{+}21, KAD^{+}22, MCF^{+}21, MR20a, MHS^{+}21, PHMVL^{+}22, RBLN21, RMFP^{+}23, STF^{+}20, SCW^{+}24, TCG^{+}21]. environments [BOB^{+}20, STSJ22]. epiphyte [Ric21]. epistemic [BAA^{+}25, NWSL^{+}24]. epistemological [KvZS^{+}25]. equitable [DS25, KBCD20, PP25, SAB^{+}22, VDO^{+}24]. equity [CSS^{+}24, HHMO24, Lau20, MW25, YWH^{+}24]. equivalency [PHK^{+}25]. eradication [BFS24, PLTC^{+}23]. eradications [PCH22]. erosion [JPAR^{+}22, OAY^{+}24]. Erratum [Ano23d, Ano23c, Ano24e, Ano24f, BPR24b, Ano22d]. error [VG20]. errors [DPC^{+}24, Sta22a]. Essential [TWGGB^{+}25]. established [GJC^{+}23]. Establishing [CTP^{+}24]. estimate [And23, APE^{+}21, FL20, HRCA21]. estimates [LKF^{+}23, MBH^{+}21, NDS^{+}20]. Estimating [LNDM21, LWD^{+}20, ZLZ^{+}21, MLS^{+}22, MBD^{+}23, MBI^{+}22]. Estimation [DSZ21, GCSZ21, NFT^{+}20]. ethic [Pic20]. Ethical [BIdL^{+}20, BHG^{+}23, MSS^{+}24]. Ethics [KB21, THM^{+}21, LSÁH^{+}20, NBB^{+}21, San20, Tay20]. eudaimonic [WSvRLK20]. EUR [Sár23]. Eurasia [DBR^{+}23]. Eurasian [GJC^{+}23]. Europe [Sár23, Ala25, BRJ25, Con21, SZH^{+}23]. European [BDF^{+}22, CCCB^{+}23, FZG^{+}24, MSJ20, RDS^{+}24, TRR^{+}20, ZEO^{+}22, ZBM^{+}24]. evaluate [DL24, NLD^{+}20, WGAB^{+}21]. Evaluating [AAvdW^{+}22, BHM20, BLW23, dOCVJ^{+}23, DSS^{+}23, DTV^{+}25, FJH21, GSMW21, HVVA^{+}21, KBCD20, MHJ^{+}24, MHGC21, MBL^{+}23, MSJ20, PCH22, RGJCN^{+}23, RBD^{+}21, SJRG23, SRCS23, TF24, TJE^{+}24, TW24, TJC21, BYHT^{+}20, BSSG21, NAS21b, SN22]. Evaluation [WTAD21, CEM^{+}23, DEGL^{+}24, GW21, GCJB^{+}22, JBE^{+}22b, ZXR24]. evaluations [Des21]. events [GTAdPA24, MWW^{+}24]. evergreen [QZX^{+}24]. everyday [VGW^{+}25]. Evidence [WLB^{+}23, CCPE^{+}25, CAM^{+}21, HAE^{+}24, MR20b, PAFB^{+}22, RMN^{+}21, RMW23, STW22, SN22, WJP^{+}20]. Evidence-based [WLB^{+}23, HAE^{+}24, MR20b, PAFB^{+}22]. evidence-review [RMN^{+}21]. Evident [RvTW22]. evolution [JP21]. Evolutionary [Ano22j, BOB^{+}20, DH21, MBB^{+}23, Sra23, Ano22j]. Ex [ADG^{+}20, FJZ^{+}20, DH21, FJZ^{+}21, KOO^{+}21, WJ21]. examine [NHSM^{+}24]. Examining [Cor20, TPP^{+}25]. exceptional [WBC^{+}20]. exemptions [TWS^{+}25]. existing [BHT^{+}24, MCF^{+}24, SACA^{+}20]. exotic [NCdlLZ24, PLKP24]. Expanding [Sal24, TFFM22, MSJ20]. expansion [CXJ^{+}22, LZG^{+}25]. experience [DRFL^{+}24, RHGRSG21]. experienced [TAA^{+}24]. experiencing [HBY^{+}23]. experiment [HGM^{+}22, MHD^{+}22]. Experimental [IDD^{+}22, STW22, ABM^{+}22, RGD^{+}25]. Experimentally [DFW^{+}24]. experiments [NJ20]. Expert [MTH^{+}24, DHNM22, DTV^{+}25, FSC^{+}21, HHB23, MHRF20, RHFV23, RMW23, STW^{+}24, TDS^{+}23]. expert-based [STW^{+}24]. expert-mediated [DHNM22]. expiring [WH22]. explain [vESN^{+}20]. explainable [DL24]. explicit [CVT^{+}25, NAG^{+}20, YLY^{+}20]. exploitation [FL20, HOWM24]. explore [SH21, SCW^{+}24]. Exploring [HBY^{+}23, HOWM24, ITAAG^{+}24, MHSL^{+}23, MMHR20, SSN^{+}23, VHHT21, vESN^{+}20]. exposed [KTG^{+}22]. Exposing [EOR^{+}24]. Exposure [GTAdPA24, VGB^{+}24]. extant [BFS24]. Extending [GRS^{+}23, Red23]. extensive [PWW^{+}22, YCH^{+}24]. extent [LNDM21, VNC^{+}24]. external [MMHR20]. extinct [ADG^{+}20, RHFV23]. Extinction [KSC^{+}20, RWC^{+}22, SWM^{+}24, BMA^{+}20, BTM^{+}23, BA22b, FJZ^{+}20, HLZ^{+}21, KCT22, KFN^{+}21, LDC^{+}24, LTH^{+}21, LWS^{+}24, MCH^{+}20, MHM^{+}22, RRH^{+}24, TBJ^{+}24, TBP^{+}22, VLPVM^{+}24, Ano22i]. extinction-risk [MCH^{+}20]. extinctions [vBDD^{+}20]. extirpated [BFS24]. extirpations [BPN^{+}24, PKP20]. extreme [BPH^{+}22, GTAdPA24, MDG^{+}21].

F [Cli23]. face [BPH^{+}22, CN21, TPP^{+}25]. Facebook [CAA^{+}24]. faced [ADKM24]. facilitate [HCC^{+}21, JRKF21, TWS^{+}25]. Facilitating [WSC^{+}23, NJLC21]. facing [ZXXL23]. Fact [Sag20a, Sag20b]. factor [CGM^{+}25]. Factors [TENI23, BMG^{+}22, CVT^{+}25, JDM^{+}21, MHS^{+}21, NAG^{+}20, PSJG23, SSKBL21, ZSL24b]. facts [MR20a]. failure [BBM^{+}20, DBC^{+}23]. fair [IDJ^{+}25, PP25]. fallacy [WAA^{+}21]. False [CO21]. family [ZSVK21]. Farhadinia [KOO^{+}21]. farm [Ogg20]. farm-program [Ogg20]. farmed [HWG^{+}22]. farming [KCS20]. farmland [ODG^{+}22]. Fauna [Ano23n]. feasibility [CCKT25]. fecundity [RMFP^{+}23]. federal [QGC^{+}25]. fences [MMHR20]. fencing [STJ20]. Field [Cli23, Sco21, AP23, HGM^{+}22, JM21]. fill [QOF^{+}21]. finance [MLIE25, NSL^{+}23]. Financial [ZZ24, HMC^{+}23b]. find [FMGS21]. fine [BHM21]. fine-tune [BHM21]. fingerprints [GTC21]. fins [QOF^{+}21]. Fire [Len21, DFW^{+}24, PLMF23, SCB^{+}25, SD20]. First [PSV^{+}21, GRS^{+}23]. Fish [Fur20, CFC21, EPL^{+}24, FJH21, HMC23a, HCZW24, IMI^{+}21, JTD^{+}20, LKA^{+}22, LLYK24, MCF^{+}21, MHJ^{+}24, MBLD^{+}25, MHS^{+}21, MHR^{+}22, OAY^{+}24, PMSC22, QGC^{+}25, RvTW22, TDC^{+}23]. fished [BMG^{+}22]. fisher [QOF^{+}21]. fisher-contributed [QOF^{+}21]. Fisheries [Ano22j, ABRCM21, CGM^{+}24, HTS^{+}22, Jen22, McC24, MWV^{+}20, MSRP21, NHSM^{+}24, QOF^{+}21, SWCP^{+}22, TBJ^{+}24]. fishers [RHGRSG21]. fishery [AHM^{+}22, BPH^{+}22]. fishes [TA22]. fishing [DAN^{+}21, HCZW24, MHS^{+}21, RMFP^{+}23, WOF^{+}20]. fit [GSMW21]. Fitness [MPM^{+}23, Ele20]. fits [LGD^{+}25]. fitting [MEK21]. Five [KBW^{+}23, KFG^{+}20]. flagship [SLM^{+}20]. Flames [Ano22i]. flats [CMP^{+}24, HWP^{+}21]. flawed [NAS21a]. fleets [WOF^{+}20]. Flexibility [Dre24, SSN^{+}23]. flora [RWC^{+}22]. flow [GBMD^{+}22]. fly [FBS^{+}23]. flying [FGB^{+}23]. flytraps [MTH^{+}24]. flyway [GJC^{+}23]. Focus [BO21, DSC^{+}24, NFD20]. focused [MGS^{+}23, NLV20]. following [TCG^{+}21]. food [DAN^{+}21, GMEV^{+}21]. footprint [BMG^{+}22]. forage [FJH21]. foraging [BF21]. Forces [Val20, Hul22, VPS22]. Forecast [Bou23]. Forecast-based [Bou23]. Forecasting [MRS22, HVA^{+}21]. forest [ADKM24, AP21, BP21, CP20a, CHA^{+}24, CCL23, DS22, FLH21, FPC^{+}24, GGB^{+}20a, GTC21, HBJ^{+}20, HGCK22, ICMO23, IMI^{+}21, LKA^{+}22, MLIE25, ML22, PKP20, PSJG23, PBB^{+}22, RJAR23, SMLL24, SBE25, MCP^{+}25]. forest-offsetting [GGB^{+}20a]. forests [BBGL21, BRJ25, BPN^{+}24, BZT^{+}22, BBE^{+}25, HXW^{+}23, KZM23, MSN^{+}22, NGST21, Ogg20, QZX^{+}24, RGJCN^{+}23, RCRJ^{+}24, YLY^{+}20]. form [SSEN20]. formal [BBL22]. forms [PHMVL^{+}22]. fossil [Gok21]. fossil-fuel-dependent [Gok21]. foundations [HSHA20]. Four [CND^{+}23]. Fragmentation [VRY^{+}23, BHS^{+}22, CCL23, CQS^{+}22, GGSRF20, SMLL24, TJE^{+}24]. fragmented [ABM^{+}22]. fragments [HBJ^{+}20]. Frames [MR20a]. framework [BYHT^{+}20, DRFL^{+}24, GW21, HK21, MCM^{+}21, MBI^{+}22, NWSL^{+}24, RJC^{+}22, Sal24, SD22, WMM^{+}20]. frameworks [Ano23d, CMC^{+}23, GN20, LS21]. Framing [Bus22, KFG^{+}20, Kus20, MR20b, STW22, SBE25]. France [Sár23, SBP^{+}23]. Free [LC24, JP21]. frequentist [PWW^{+}22]. Freshwater [JTD^{+}20, BJP^{+}20, DNCC22, GNH^{+}22, HLZ^{+}21, HMM^{+}24, NSB^{+}21, RCRJ^{+}24, SWM^{+}24, SZH^{+}23, TDC^{+}23, TA22]. freshwaters [TRR^{+}20]. friendly [MSKM21]. frog [FGB^{+}23, LWRZ^{+}21]. front [HL20]. frontier [MBL^{+}23, SPK^{+}22]. frontline [PSV^{+}21]. fuel [Gok21]. function [ARK^{+}20, TRR^{+}20]. Functional [Ano22j, LHB^{+}22, AAvdW^{+}22, CQS^{+}22, GTC21, MCF^{+}21, MNMPGN23, SD20]. functionally [MSB25]. functioning [DBM^{+}25]. functions [FPC^{+}24]. Fundamental [RGBC20]. Funding [PAFB^{+}22]. fungal [RGD^{+}25, WFBT^{+}25, WWVC^{+}25]. Fungi [Kut20, OLF24]. further [LNYY20]. Future [Len21, Ano24e, APE^{+}21, BNI25, BKK^{+}21, BPR24a, CP20a, CGL^{+}25, Far21, GMEV^{+}21, LWC^{+}24, MSMH22, MDB^{+}24, SBH^{+}23, VBC^{+}25]. futures [WMK^{+}21].

G [Ano23n, Sár23]. gain [RzEHB24, SRCS23]. Galindo [Sal24]. games [DMR^{+}21]. Gap [DdLI^{+}25, MTS^{+}21, MCF^{+}24, ROK^{+}22, VM20]. gaps [JSGJ^{+}21, JM21, QOF^{+}21, RzEHB24]. gardens [MA20]. gastropod [GNH^{+}22]. gender [Lau20]. gene [GBMD^{+}22, San20]. Generalization [ARKA22]. generating [GFF23]. Generation [BMA^{+}20]. generational [CVT^{+}21]. genes [CvH^{+}22]. Genetic [OAY^{+}24, SHH^{+}23, CSHM^{+}20, CTL^{+}23, DSF^{+}23, GGSRF20, HVVA^{+}21, HMC^{+}23b, MPM^{+}23, NW22, RBD^{+}21, San20, VH22, WJ21]. genome [PSS20, ZAS^{+}24]. genome-wide [ZAS^{+}24]. Genomic [JPAR^{+}22, MBW^{+}22, XDA^{+}21]. genomics [Sra23]. genoscape [RHS^{+}20]. genoscape-network [RHS^{+}20]. genotyping [YCH^{+}24]. Genus [Kut20]. geodiversity [TB23]. geographic [DH21, Hul22, VNC^{+}24]. Geology [IMI^{+}21]. Geology-dependent [IMI^{+}21]. Georgia [CGL^{+}25]. Ghisbain [Sár23]. giant [CQS^{+}22, MCB^{+}20, PPN^{+}24, LHB^{+}24]. gibbons [YCS^{+}23, ZTCF21]. gill [ZYM^{+}24]. Gillnet [AAS21]. GIS [Cli23]. Global [Ano22i, Arm23, BMH^{+}24, EAdC^{+}21, HGAB^{+}24, HWP^{+}21, ICMO23, KCH^{+}20, MDG^{+}21, NSL^{+}23, OC24, RRH^{+}24, WWB^{+}22, ZWC^{+}24a, ADKM24, BSP22, BVH22, dOCVJ^{+}23, CFS^{+}21, CP20a, CTL^{+}23, CCW24, DFB^{+}24, Eva21, FW22, GSB23, Gok21, GAB^{+}21, GCJB^{+}22, HK21, HBB^{+}24, HBY^{+}23, HXW^{+}23, HOWM24, LS21, MBS^{+}22, MHM^{+}22, PWB^{+}21, PHMVL^{+}22, RMW23, SMLL24, SWMM24, SOW^{+}22, TB23, WGAB^{+}21, WWVC^{+}25, ZXR24]. globally [GSS^{+}22, KBL^{+}23, SNC^{+}22, SSKBL21, WCY^{+}25]. goals [HXW^{+}23, PHB^{+}21, SWCP^{+}22]. Going [CC20]. golden [QHS22]. gondii [RRR^{+}24]. González [Sal24]. Google [EM21, Huy23]. govern [MO25]. governance [ART^{+}25, GAB^{+}22, LC24, MOO^{+}24, McC24, Mol22, RzEHB24, RBLN21, SWMM24, THM^{+}21, TFFM22, WAB^{+}24]. governments [AFV20]. Gracia [JAB^{+}21b]. gradient [VRY^{+}23]. gradients [MMHR20]. grant [BGB^{+}24]. graph [DSF^{+}23]. graph-based [DSF^{+}23]. grasshopper [BTH^{+}20]. Grassland [BPR24b, AL21, BNI25, JAB^{+}21a, NRN^{+}24, PHB^{+}21, RVLD20]. gray [HHL^{+}24]. Great [Imb22, TCG^{+}21, WFH^{+}21, MHR^{+}22]. greater [Hul22, ZAS^{+}24]. Green [CCCB^{+}23, LKF^{+}23, CLS20, NTS^{+}21]. greenspace [TG21]. grief [FR21]. groups [AHM^{+}22, GBRK^{+}20]. grouse [ZAS^{+}24]. Growth [CGA^{+}22, MTM^{+}22]. guardians [RBLN21]. Guide [Ano23n, CACM23, CEM^{+}23, KFG^{+}20, RBJN21, STK^{+}21, Ano22i]. Guidelines [VH22, WLB^{+}23]. Guiding [CCH^{+}21]. Guinea [BHT^{+}24]. guitarfish [GZBP24]. Gulf [THSA24]. gull [THSA24]. Gut [HQY^{+}24]. Gwich’in [PSV^{+}21].

Habitat [MWH^{+}22, MSPM20, SHS^{+}22, And23, ABM^{+}22, BRW21, BPN^{+}24, BHS^{+}22, BHT^{+}24, BMM^{+}23, CMP^{+}24, CvH^{+}22, CP20b, CQS^{+}22, DSS^{+}23, DBT21, EM21, FBG^{+}25, GGJ^{+}20, Gok21, GGSRF20, HTS^{+}22, HHB23, LS21, LZG^{+}25, LWC^{+}24, LDD^{+}22, MVC^{+}21, SNC^{+}22, VPVdAMF20, ZYNC25]. Habitat-based [MWH^{+}22]. habitat-masked [And23]. Habitat-tree [MSPM20]. Habitats [Vin22, DS22, STF^{+}20]. half [DPC^{+}24]. hardback [Sár23]. hardcover [Ano22i, Ano22j, Ano23n, Sár23]. harm [HM24]. harness [GCA^{+}22]. Hart [Ano22i]. harvest [BPH^{+}22, GGB20b]. harvesting [CAA^{+}24]. havens [JP21]. Hawaii [RWC^{+}22]. Hawley [Ano22j]. Healers [PS24]. health [FLW^{+}22, KIA^{+}22, PKC^{+}23a, STW^{+}24]. heart [GGBD20]. Heat [Len21, MWW^{+}24]. helmeted [HHL^{+}24]. help [GJPJ^{+}22]. Hemisphere [WWVC^{+}25]. Hendry [Ano22j]. herbaceous [CSHM^{+}20]. Heritage [GR21, BOB^{+}20, DBR^{+}23, MBK^{+}20, ZZ24]. herpetofauna [ICMO23, KMC^{+}23]. heterogeneity [CMP^{+}24, YVA22]. hibernacula [TSS^{+}22]. hibernating [CRC^{+}21]. Hickey [Cli23]. Hidden [RMFP^{+}23]. hierarchical [VG20]. high [ARKA22, BOB^{+}20, DBMS24, MSN^{+}22, TDC^{+}23, TzEN^{+}24, WSW23, XAAY^{+}21, ZLZ^{+}21]. high-competition [BOB^{+}20]. high-conservation-value [MSN^{+}22]. high-level [TzEN^{+}24]. high-resolution [TDC^{+}23, ZLZ^{+}21]. high-seas [XAAY^{+}21]. Highlighting [BGAE25]. Historical [Ano24g, BWN20, BPN^{+}24, HZW^{+}21, LMG^{+}23, MSN^{+}22]. historically [HRCA21]. history [HLZ^{+}21, MR22, MAC20, MBB^{+}23, SSG^{+}23]. Holy [BGAE25]. homogenization [KZM23]. Honoring [ÁRAB^{+}24]. hope [DW23, FR21]. Hopkins [Ano23n]. Horizon [GGBD20, SMT^{+}22]. hornbill [HHL^{+}24]. horticulture [RR24]. host [Ele20, WWVC^{+}25]. hotspot [SBW^{+}22]. hotspots [CAFM^{+}23, JTD^{+}20]. Human [Cor20, KKKS^{+}20, SBW^{+}22, TB25, Vin22, Ano24f, AOI22, ARKA22, BMG^{+}22, BF21, CBM20, CBD^{+}21, CBB^{+}20, DYM25, DRFL^{+}24, DFC21, EGCS25, FLW^{+}22, FCL^{+}22, FCH^{+}24, FWZ^{+}24, GCS^{+}24, GZBP24, HLZ^{+}21, HWP^{+}21, JDM^{+}21, JRKF21, JSKR22, JS25, JSA^{+}20, KFU^{+}24, LZG^{+}25, LTJ^{+}20, MHL^{+}22, MBK^{+}20, NFD20, PS24, PMLM20, PBV21, PSV^{+}21, Sal24, SÁT21, SNC^{+}22, TTDdLM^{+}21, TSÁ20, VHHT21, VPZ25, WLR^{+}24, WLA20, WTAD21, ZWC^{+}24a, ZXR24]. human-dominated [SNC^{+}22]. Human-Made [Vin22]. human-pressure [FCL^{+}22]. human-wildlife [CBM20, TSÁ20]. humaneness [HFW20]. humanities [HCJH^{+}22]. humanity [SSCR^{+}23]. hunt [DHE^{+}21]. Hunting [Pri21b, Sul24, BF21, DHE^{+}21, EOR^{+}24, EGBJ23, GZBLS^{+}20, MCB^{+}20, NJ20, PS20, WDHR22]. Huyvaert [Ano22j]. hybrid [TF24]. hybridization [CS21, DLL22, SBG^{+}25]. hybrids [CS24]. hydroelectric [RvTW22]. hydroperiod [CS24]. hyperdispersal [BM24]. hypotheses [BF21].

Iberian [Ala23]. iconic [HtSR25]. Identification [Kut20, ZSL24b, LMG^{+}23, Sco22]. identified [TGJ^{+}20]. identify [Ano23c, BRW21, CLGP23, EPL^{+}24, LFD^{+}25, MSN^{+}22, RJC^{+}22]. Identifying [BB21, BOPM^{+}24, CCW24, DHNM22, DBT21, GH22, KTG^{+}22, PSJG23, PLKP24, Ric22, SP23a, TDS^{+}23]. identity [BMM^{+}24, HSHA20, vESN^{+}20]. identity-related [HSHA20]. ideology [JGAH25]. iDNA [FBS^{+}23]. iEcology [JBC^{+}21]. ignoring [SMSS^{+}21]. III [Sta22a]. Illegal [ABRCM21, MMN^{+}20, SP21, DAN^{+}21, EOR^{+}24, HHL^{+}24, HLL^{+}24, HM24, MMG^{+}23, MTH^{+}24, NHSM^{+}24, OGMMG22, OABDCM24, ZWC^{+}24c]. illumination [AAS21]. illustrated [BTH^{+}20]. illustration [Hen22]. imagery [ZLZ^{+}21]. images [FCH^{+}24, Won24]. imaging [MCF^{+}24]. immediate [GBMD^{+}22]. Impact [BSRM22, BAGV24, KCT22, KBB^{+}24, PMSC22, BYHT^{+}20, BSSG21, BLW23, BMM^{+}23, Cas23, CGM^{+}25, DS25, Des21, DSS^{+}23, DEGL^{+}24, GAB^{+}21, GWB^{+}24, HOT^{+}25, HBB^{+}24, HLZ^{+}21, HtSR25, JAB^{+}21a, JBE^{+}22b, MVC^{+}21, MK21, NFD20, RRR^{+}24, SJRG23, STW22]. Impacts [Ano22j, BHT^{+}24, CVT^{+}21, KMC^{+}23, SBE25, BMH^{+}24, BBE^{+}25, CGL^{+}25, CBB^{+}20, FJH21, GH22, GHR21, GSS^{+}22, HNC^{+}24, HWK^{+}20, IMI^{+}21, KFP^{+}21, KCS20, MTH^{+}24, MSMH22, PB21, PLD^{+}21, RMFP^{+}23, SBG^{+}20, SMLL24, SCW^{+}24, Veg24, WJP^{+}20, ZXR24, Ano22i]. impasse [LGL20]. impending [SDH^{+}25]. imperfect [TPP^{+}25]. implement [LGD^{+}25]. implementation [CGM^{+}25, HTC22, MTS^{+}21]. Implications [BLF^{+}23, VMP^{+}24, BMA^{+}20, HLL^{+}24, KCS20, LWG21, NCE^{+}21, PHMVL^{+}22, SdVD^{+}22, ZWC^{+}24a]. Importance [Ano22j, BCA^{+}21, CMP^{+}24, CTL^{+}23, DSM24, GB22, Gru21, HAE^{+}24, Hul22, KIA^{+}22, KSZ^{+}23, SGB^{+}20, Sár20, SAB^{+}22, BGAE25, FBG^{+}25, GSS^{+}22, HHMO24, JDM^{+}21, OGF^{+}21, PPN^{+}24, PG20, TDS^{+}23]. important [SNC^{+}22, VGB^{+}24]. imported [WSSC22]. improve [CGB^{+}24, JBE^{+}22b, KAD^{+}22, MCH^{+}20, RAG^{+}20, TFFM22, VLVD^{+}23, WKG^{+}21]. improved [PCH22]. Improving [CO24, FCL^{+}22, IDJ^{+}25, MW25, NCE^{+}21, PLC22, RJAR23, WDH^{+}21, BDQ^{+}25, CBD22]. imputation [JSGJ^{+}21]. inbreeding [APE^{+}21, CQS^{+}22, FMGS21]. incentive [ML22, WJP^{+}20]. incentive-based [WJP^{+}20]. incentives [Ogg20]. include [LS21, OLF24, Red23]. Including [SCB^{+}21, MSC^{+}24]. inclusion [CFN^{+}24, NWSL^{+}24, WSEL^{+}24, YWH^{+}24]. inclusive [And25, HHM24, VDO^{+}24]. incomplete [WDH^{+}21]. inconceivable [MCB^{+}20]. inconsistencies [PLKP24]. Incorporating [BSG^{+}23, CSHM^{+}20, GRP^{+}25, WSB^{+}20, XDA^{+}21]. Incorporation [GZBP24]. increase [BOB^{+}20]. increased [MHSL^{+}23, SCB^{+}21]. increases [Gok21]. Increasing [JMB21, ZWC^{+}24b]. independent [DSF^{+}23]. index [MGJF21]. India [BNS24, JSKR22]. Indian [MFW^{+}24]. indicator [GTC21, HQY^{+}24]. Indicators [GCD^{+}23, MHGC21, RBJN21, SWC^{+}21, TB25]. indices [WGAB^{+}21]. Indigenous [OGF^{+}21, RBLN21, BWN20, SCB^{+}21, Uts24, Bus22, ECVL25, GKH^{+}24, JSKR22, JS25, SDB^{+}24, VFKT24, WAB^{+}24]. Indirect [LB22]. individual [GHR21, SBG^{+}25]. individual-based [SBG^{+}25]. individualism [SB21]. Indo [ARSL24]. Indo-Pacific [ARSL24]. Indonesia [LCEM24]. Indonesian [PS24]. induced [SR20, WB23]. industrial [WOF^{+}20]. inequality [JSA^{+}20]. Infectious [PMS^{+}20, Ano22j]. Inference [Ano23n, WKG^{+}21]. inferences [WDH^{+}21]. inferred [MHB^{+}22]. inferring [VMP^{+}24]. Influence [ADKM24, MRTV24, BMM^{+}24, MHJ^{+}24, MHS^{+}21, MBL^{+}23, MRS22, NAG^{+}20, ZAS^{+}24]. influences [Ric21]. inform [AL21, COF^{+}25, CFA^{+}24, DTV^{+}25, HNC^{+}24, HZW^{+}21, KBB^{+}24, PMR^{+}22, RBC^{+}22, TGM^{+}22, ZWC^{+}24c]. Information [Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano22e, Ano23i, Ano23j, Ano23e, Ano23f, Ano23g, Ano23h, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano25a, Ano25b, DBT21, KWY^{+}21, LMG^{+}23, RAG^{+}20, WKT^{+}21]. informed [HRCA21, LC24]. infrastructure [NSL^{+}23]. inquiry [GBA23]. insect [BJP^{+}20, MA20, OPF24, RBC^{+}22]. insect-resistant [OPF24]. Insects [Car20]. insight [Ano24g, RMW23]. Insights [JS25, DBT21, JSKR22, RGBC20, SMB^{+}20, SCB^{+}21, ZWC^{+}24c]. institutional [GSMW21]. instruments [GW21]. insular [PLTC^{+}23, TBP^{+}22]. insurance [FJZ^{+}20]. intact [WWB^{+}22]. intangible [JDM^{+}21]. integrated [ADSG21, SGA^{+}24]. Integrating [And23, LFD^{+}25, ME25, MCC^{+}23b, RVLD20, SD20, VRBAB^{+}24, HMC^{+}23b, TFJ^{+}25, TB23]. Integration [SPK^{+}22, Cie24, CHW^{+}23, GB22, QGC^{+}25, TB25]. integrity [GTC21, RJAR23]. intellectual [May22]. intelligence [KTH^{+}22, Sár23]. intensification [BSRM22, LCEM24]. intentional [PCH22]. intentions [STES23]. interaction [FCH^{+}24, Sal24, WTAD21]. interactions [CLS20, EPL^{+}24, GG24, GZBP24, KBB^{+}23, NHSM^{+}24, PS24, TPP^{+}25, VHHT21, VPZ25]. Interactive [ODG^{+}22, MRTV24]. intercontinental [MMS^{+}20]. interdisciplinarity [Hul22, SH21]. interdisciplinary [BDQ^{+}25, DTSE25, KSFP^{+}21, Mol22]. Interest [Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano22j, Ano22g, Ano23l, Ano23m, Ano24q, Ano24r, Ano22i, Ano22f, Ano22h, Ano23n, Ano23k, Ano24n, Ano24o, Ano24p, Ano25c, Ano25d, dOCVJ^{+}23, May22, MCG^{+}21, SP22, VMR21]. interests [NBD^{+}24]. interface [BGMEHP20]. intergroup [CHA^{+}24]. interior [VRY^{+}23]. internal [MMHR20]. international [CST^{+}24, SDH^{+}25, SF23]. internet [STK^{+}21, THM^{+}21]. Internet-based [THM^{+}21]. intersection [NBB^{+}21]. Interspecies [THSA24]. interspecific [SD22]. intervention [BHRBK20, DDM24, MPM^{+}23, RW23]. interventions [JBE^{+}22b, Kho22, MWV^{+}20, MBI^{+}22, NJLC21, OVC^{+}24, OABDCM24, PMS^{+}20, SACA^{+}20, TWMM^{+}23, ZWC^{+}24c, dLMGK22]. interview [PMR^{+}22]. Intimidation [KvZS^{+}25]. intolerance [JSA^{+}20]. intraseasonal [RVLD20]. introduced [FLH21, WRBL^{+}23]. Introducing [DBC^{+}23, MOO^{+}24]. Introduction [CLR21, DWC^{+}20, AP21, HCA^{+}22, Ano22i]. introductions [PCH22]. invader [HBY^{+}23]. Invasion [CHP20, CBB^{+}20, JBC^{+}21, BLF^{+}23, Dav20, KKE^{+}24, OC24, PG20, Sag20a, Sag20b, SVP^{+}22]. invasions [Hul22, MSB25, PPE^{+}20, RAG^{+}20]. invasive [BOAN23, BAB^{+}23b, COF^{+}25, CS24, FBS^{+}23, GWR24, GCA^{+}22, LSM24, ODMG^{+}20, PPN^{+}24, Won24, Ano22i]. investigating [JM21]. investigation [CHW^{+}23]. investment [TGJ^{+}20]. investments [DEGL^{+}24]. involves [BIdL^{+}20]. Irish [Ano24g]. irreplaceability [BSP22]. irrigation [GWR23]. ISBN [Ano22i, Ano22j, Ano23n, Cli23, Sár23]. Island [Ano22i, DBM^{+}25, AP23, GWR24, PNXY^{+}22, PPN^{+}24, POH^{+}22]. islands [BP21, BAB^{+}23b, Eva21, COF^{+}25]. isolated [ZAS^{+}24]. Issue [Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano22e, Ano23i, Ano23j, Ano23e, Ano23f, Ano23g, Ano23h, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano25a, Ano25b]. Italy [DPC^{+}24]. IUCN [And23, BYHT^{+}20, Cas23, CSL^{+}23, CGB^{+}24, DBT21, GRP^{+}25, KGK^{+}24, KBB^{+}24, LKF^{+}23, LDC^{+}24, MSC^{+}24, MHB^{+}22, RWC^{+}22, SHH^{+}23, Van23]. ivory [GLL^{+}23].

J [Ano22i, Ano22j, Ano23n, Cli23]. J.P [Ano22j]. Jane [BAB^{+}23a]. January [Ano20a, Ano24b]. Japan [Ano24f, KFU^{+}24, LKA^{+}22, STES23, TE20]. jeopardizes [PGRDOGG23]. Johns [Ano23n]. Joining [VPS22]. journals [CGM^{+}25, CBD22]. Jucker [dG21]. Judas [RCL^{+}22]. judgment [DTV^{+}25]. June [Ano22b]. just [WMK^{+}21]. justice [Arm23, ME25, NWSL^{+}24, PSO^{+}22, SKS22, SÁT21, STA^{+}22b]. Justus [Ano22i].

Kauai [RWC^{+}22]. Keeping [WDHR22]. Kenya [HTM25, PMLM20]. Key [GBB^{+}20, Ano23d, BBH^{+}25, CMC^{+}23, HQY^{+}24, LMY^{+}24, NSB^{+}21, SDB^{+}24, PHB^{+}25]. Khalatbari [FJZ^{+}21]. Killing [CMAS20, CC24]. Kingdom [Kut20, Ano24e, BPR24a]. Knowledge [BMM^{+}23, VLVD^{+}23, CGB^{+}24, DM20, FSC^{+}21, JM21, JS25, MHC21, MWS^{+}25, PS24, ROK^{+}22, SH21, SCB^{+}21, TPP^{+}25, VM20, WAB^{+}24]. knowledge-action [ROK^{+}22]. Knowles [Ano23n]. known [DBT21]. Kohler [CC20]. Kubatko [Ano23n]. Kusmanoff [MR20b].

L [Ano22i, Ano23n, Cli23, Sár23]. L.S [Ano23n]. laboratories [DMR^{+}21]. lack [CLD^{+}21, Hyd24]. ladybird [SHA^{+}23]. lagged [BHS^{+}22]. lakes [DNCC22, MHR^{+}22]. Land [LCEM24, Poz20, BP21, EN24, GZBLS^{+}20, GW21, HRHB20, HBJ^{+}20, KSFP^{+}21, KTG^{+}22, LDD^{+}22, MR22, MMS^{+}20, NRSF21, PKP20, PHB^{+}21, Pic20, Ric21, SJRG23, SPK^{+}22, SSEN20, WSK^{+}22, WDH^{+}21, YCS^{+}23, YVA22, ZXR24, dlLBB^{+}23]. land-bridge [BP21]. land-change [KSFP^{+}21]. land-cover [YCS^{+}23]. land-sea [WSK^{+}22]. Land-use [LCEM24, EN24, MR22, NRSF21, PKP20, Ric21]. land-use-based [GW21]. land-use-change [SPK^{+}22]. landbirds [GJC^{+}23]. landowner [ML22]. lands [OGF^{+}21, SDB^{+}24]. Landscape [BNI25, BRW21, Cie24, MRS22, PBB^{+}22, RJC^{+}22, SBG^{+}20, SNC^{+}22, SSG^{+}23, SHS^{+}22, SMSS^{+}21, YHW^{+}25]. Landscape-scale [BNI25, PBB^{+}22, RJC^{+}22]. Landscapes [Gon21, Val20, ABM^{+}22, ANR23, BF21, FBG^{+}25, HGCK22, JBE^{+}22a, TTDdLM^{+}21, WLR^{+}22, YMB^{+}21]. language [CCPE^{+}25, CGA^{+}22]. Laos [LC24, WYB^{+}21]. Large [BAGK22, DHE^{+}21, ARKA22, CFC21, EHBT20, GBMD^{+}22, HHMO24, INO23, KSZ^{+}23, MCF^{+}24, MMS^{+}20, RCBT21, SHS^{+}22, TDC^{+}23, TW24, WOF^{+}20]. large-area [MCF^{+}24]. Large-sample-size [BAGK22]. large-scale [EHBT20, INO23]. largest [BLF^{+}23, BFS24, EOR^{+}24]. Laridae [SBP^{+}23]. larval [MCC^{+}23b]. last [FJZ^{+}21, MEM^{+}22, QZX^{+}24]. latent [KLVW20]. Latin [PS20, Uts24]. law [SGA^{+}24]. laws [ODMG^{+}20]. Lawson [CO21]. lay [GCA^{+}22]. laying [MHJD23]. leaders [Ric22]. leadership [GKH^{+}24]. Leakage [FZG^{+}24]. learn [BBL^{+}20]. Learning [SS23, UEH^{+}23, DL24, DBC^{+}23, IJK^{+}21, MVC^{+}22, ZSVK21]. led [MMN^{+}20]. legacy [ÁRAB^{+}24]. Legal [PP25, HHL^{+}24, NHSM^{+}24, OGMMG22, SP21, SFC20, dlLBB^{+}23]. legality [RR24]. legislative [FLW^{+}22]. lemur [TPP^{+}25]. lengths [BMA^{+}20]. Lens [Poz20]. Leopards [Pri21a]. Leopold [Pic20]. less [KTG^{+}22]. Lessons [CHR24, HWLC22, MHD^{+}22, MEA^{+}22, SF23, BBL^{+}20, KFG^{+}20, Lau20, Sch21]. Let [Dav20]. lethal [MJDMAS20]. level [BAGK22, EN24, OCC^{+}21, TzEN^{+}24, VGB^{+}24]. Leveraging [MVC^{+}22]. lexicon [SCdM^{+}24]. LGBTQIA [CFN^{+}24]. license [DHE^{+}21]. life [GCD^{+}23, HLZ^{+}21, HL20, SSG^{+}23, WLB^{+}20, WMK^{+}21]. life-history [HLZ^{+}21]. light [WH22]. limit [LTJ^{+}20, RRR^{+}24]. Limitations [CLD^{+}21]. limits [GGB20b, MLS^{+}22, SW25]. line [HL20]. Linguistic [GR21]. Link [Ano22j, MMN^{+}20]. Linking [SSG^{+}23, SWCP^{+}22, TTDdLM^{+}21, TBJ^{+}24]. links [WSK^{+}22]. List [BYHT^{+}20, DBT21, KGK^{+}24, MHB^{+}22, BTM^{+}23, CSL^{+}23, GRP^{+}25, LDC^{+}24, MSC^{+}24, SHH^{+}23, Van23]. listed [RR24, WSSC22]. listing [SENTI24]. lists [SP23b]. literature [CGA^{+}22, LVT^{+}20, NAS21a, SACA^{+}20]. lived [SMT^{+}22]. livelihood [LNYY20]. livelihoods [PLD^{+}21, SBG^{+}20]. livestock [WMM^{+}20]. living [BNS24, PKC^{+}22]. local [EHBT20, ITAAG^{+}24, LC24, PLD^{+}21, Ver23, vBDD^{+}20]. local-scale [EHBT20]. location [SSN^{+}23]. logics [GBA23]. London [Ano22i]. Long [FGL^{+}21, APE^{+}21, DEGL^{+}24, GBMD^{+}22, MBD^{+}23, MRTV24, OAY^{+}24, SPA^{+}22, SMT^{+}22, TJB^{+}23]. long-distance [DEGL^{+}24]. long-lived [SMT^{+}22]. Long-term [FGL^{+}21, APE^{+}21, GBMD^{+}22, MBD^{+}23, MRTV24, OAY^{+}24, SPA^{+}22, TJB^{+}23]. Longitudinal [MBW^{+}22]. Loss [FPC^{+}24, NGST21, BPN^{+}24, Gok21, HGAB^{+}24, LHB^{+}22, MPM^{+}23, PSJG23, SDB^{+}24, VPVdAMF20]. losses [YTW^{+}25]. lost [AAvdW^{+}22, MMA^{+}22]. low [CGM^{+}25, DBMS24, PWW^{+}22]. low-impact-factor [CGM^{+}25]. low-resolution [PWW^{+}22]. lower [MHP23]. lowland [NGST21]. Lynn [CMAS20]. Lynx [Cli23].

M [Ano22j, Cli23, Sár23]. machine [DL24, HPM^{+}20]. Madagascar [BHS^{+}22, Nut25, RRR^{+}24]. Made [Vin22]. magnitude [YCH^{+}24]. Maine [THSA24]. mainland [NGST21]. Mainstreaming [HCJH^{+}22]. maintenance [CLM22, WLR^{+}22]. maize [OPF24]. major [HMM^{+}24, RGD^{+}25, TJE^{+}24]. make [MCB^{+}20]. Making [SKS22, BH23, BGB^{+}24, DMR^{+}21, EHBT20, FSC^{+}21, GGLB25, GKH^{+}24, GBB^{+}20, KBB^{+}24, MCC^{+}23b, PKC^{+}22, RBC^{+}22]. mammal [BPN^{+}24, BHT^{+}24, PLMF23]. mammalian [FJZ^{+}20, Kho22, NFT^{+}20]. mammals [Ala25, BHS^{+}22, DBT21, FLH21, FBS^{+}23, GZBLS^{+}20, HHMO24, HZW^{+}21, LHB^{+}22, MMS^{+}20, OGF^{+}21, RSRG^{+}20, SWM^{+}24, TJB^{+}23, Van23]. manage [HBB^{+}24]. managed [HMC^{+}23b, HGCK22]. Management [Ano22j, Fur20, AHM^{+}22, ADKM24, ANR23, BDG^{+}23, BSG^{+}23, BMH^{+}24, BPH^{+}22, BGD25, BTC^{+}22, BAGK22, COF^{+}25, CXJ^{+}22, CGM^{+}24, DS22, EKH23, FJZ^{+}20, FJZ^{+}21, FLH21, GWR24, GCA^{+}22, GKH^{+}24, GZBP24, HMM^{+}21, HC22, HWLC22, HHM24, HTC22, HTM25, KGK^{+}24, KOO^{+}21, KSFP^{+}21, KTG^{+}22, LGD^{+}25, MOO^{+}24, MWV^{+}20, MHM^{+}24, Mol22, ODMG^{+}20, PCH22, RCRF20, RML^{+}20, RSRG^{+}20, SBG^{+}25, STSJ22, SLM^{+}20, SD20, SWCP^{+}22, TGM^{+}22, TE20, TJB^{+}23, VBK^{+}23, WBJ^{+}20, WLR^{+}24]. Managing [BH23, CS24, DLL22, BOAN23]. mangrove [CXJ^{+}22, DSC^{+}24, GAB^{+}22, HGAB^{+}24, HPG20]. Manipulating [MWY^{+}21]. manipulation [CS24]. map [CP20a, LDD^{+}22, WL20]. Mapping [AL21, SFC20, BSP22, BTDS20, DRFL^{+}24, HOWM24, MAB^{+}23, RRH^{+}24, TDC^{+}23, VM20]. maps [And23, CGB^{+}24, MHRF20]. March [Ano20b]. Marine [MSRP21, ABRCM21, Ano23d, ARSL24, BWN20, BDG^{+}23, BMG^{+}22, CMC^{+}23, CSCA^{+}23, CFC21, DS22, FJH21, GWB^{+}24, HTC22, HTM25, HCZW24, Jen22, KBCD20, KTG^{+}22, LEF^{+}24, LLYK24, MHJ^{+}24, MCF^{+}24, MFW^{+}24, MWC^{+}20, MCC^{+}23b, MBI^{+}22, OCC^{+}21, PMSC22, RSRG^{+}20, RCBT21, RAG^{+}20, STF^{+}20, SS24, SWBD20, TBP^{+}22, TJC21, WOF^{+}20, WSK^{+}22, XDA^{+}21]. marine-protected [KBCD20]. market [BLF^{+}23, CP20b, LCEM24, OABDCM24, The24, WSSC22]. market-based [CP20b]. marketing [RMD^{+}20]. Markets [CBG^{+}21, BSH24, OGMMG22, SdVD^{+}22, ZZ24]. Markus [HL20]. Marshak [Ano22j]. marsupials [HMC^{+}23b]. Martell [Kus20]. masked [And23]. mass [TCG^{+}21, TAA^{+}24, Won24]. masterclass [Sco24]. Matching [MHRF20, SWC^{+}21, Ras22, SEB^{+}20]. May [Ano20b, Ano24b]. Mayo [Hen22, IR22, SP22, Sta22a]. MD [Ano23n]. me [Wal24]. meaning [BKK^{+}21, PG20]. meaningful [CLS20]. means [SMB^{+}20]. measure [AAS21, JAB^{+}21a, JPK^{+}21, LVT^{+}20, MCG^{+}21]. measured [SS23]. measures [CTP^{+}24, Coo24, HtSR25, LFD^{+}25, RRR^{+}24]. Measuring [GGLB25, HWK^{+}20, LHB^{+}24]. meat [EMV^{+}20, MEM^{+}22]. Mechanisms [WJP^{+}20, HNC^{+}24]. Media [MR20b, CFA^{+}24, DHNM22, DFH^{+}21, EOR^{+}24, GSB23, HCD^{+}20, LH21, MVC^{+}22, The24, VHHT21, Won24]. mediated [DHNM22]. Mediation [HBJ^{+}20]. Medical [AP21]. Mediterranean [NSB^{+}21, SCB^{+}25, VGB^{+}24]. meet [PHB^{+}21]. meeting [VFRV24]. meets [ZZ24]. megadiverse [FPC^{+}24]. megadrought [OAY^{+}24]. megafauna [FJZ^{+}20, HLZ^{+}21, MVC^{+}22]. megafire [WCW^{+}22]. megaherbivore [LB22]. Meierhofer [STSJ22]. members [CFN^{+}24]. memes [LSM^{+}20]. memoriam [BAB^{+}23a]. mental [PKC^{+}23a]. mesophotic [BMH^{+}24]. Message [Kus20, KFG^{+}20, SBE25]. messaging [MHD^{+}22, NCdlLZ24]. Meta [GGSRF20, ML22, RCRJ^{+}24, SCC23, WJ21, WLA20, KFP^{+}21, MWS^{+}25, SMLL24]. Meta-analysis [GGSRF20, ML22, RCRJ^{+}24, SCC23, WJ21, WLA20, KFP^{+}21, MWS^{+}25, SMLL24]. metabarcoding [FBS^{+}23]. metabolic [GGBD20]. metacommunity [BB21]. metadata [CTL^{+}23]. metapopulation [ABM^{+}22, HBY^{+}23, HPG20, vBDD^{+}20]. metapopulation-scale [HBY^{+}23]. metapopulations [MBV^{+}24]. method [FBDA22, FL20, IR22, JPK^{+}21]. methodological [HAE^{+}24]. methodology [SN22]. Methods [Ano23n, CEM^{+}23, CLJ^{+}21, DL24, KSFP^{+}21, MCH^{+}20, MCC^{+}23a, PHK^{+}25, RMN^{+}21, SPK^{+}22]. metric [MGJF21]. Metrics [SP23b, DNCC22, HWK^{+}20, HBB^{+}24, MVC^{+}21, MBLD^{+}25, SdVD^{+}22]. Mexico [ADSG21, GNH^{+}22, ITAAG^{+}24, WBR^{+}24]. Meyer [NBB^{+}21]. Michael [Ano20v, Tay20]. microbial [ZWC^{+}24b]. microbiome [HQY^{+}24, Red23, RGD^{+}25]. microbiota [Ele20]. microclimate [ADKM24]. microhabitats [KSZ^{+}23]. Microplastic [MHR^{+}22]. migrating [GJC^{+}23]. migration [BB21, CS21, CVT^{+}21, LWS^{+}24, PB21]. migratory [BSG^{+}23, CMP^{+}24, DEGL^{+}24, HVA^{+}21, LMY^{+}24, RHS^{+}20, SSN^{+}23, XST^{+}20]. Mind [Sár23]. minimum [NW22, RF23]. Mining [TzEN^{+}24]. misinformation [HCD^{+}20]. mismatch [CAFM^{+}23]. mismatches [RRH^{+}24]. Misrepresentation [Won24]. misrepresenting [EDA^{+}21]. Misunderstandings [LSÁH^{+}20]. mitigate [TSS^{+}22]. Mitigating [GWR23, GH22]. Mitigation [BTC^{+}22, SR20, TGJ^{+}20]. Mittermeier [Cli23]. mixed [KSFP^{+}21, SN22, SHS^{+}22]. mixed-use [SHS^{+}22]. Mnúxv [WAB^{+}24]. mobulid [ZYM^{+}24]. model [GFF23, LWD^{+}20, MHRF20, MLS^{+}22, OGMMG22, RHS^{+}20, WAB^{+}24, WBC^{+}20, YLY^{+}20]. Modeling [GHR21, HMC^{+}23b, KKE^{+}24, NW22, POH^{+}22, PLMF23, SMT^{+}22, Des21, FBDA22, GLL^{+}23, PB21, RVLD20, SBG^{+}25, SPK^{+}22, SCB^{+}21, TFT^{+}24]. models [DSF^{+}23, DTV^{+}25, EHBT20, HWK^{+}20, HHB23, LGL20, LTH^{+}21, MB21, MEK21, PMR^{+}22, PWW^{+}22, SMSS^{+}21]. modern [LMG^{+}23]. modification [CGL^{+}25, ICMO23, ZXR24]. moist [GCJB^{+}22]. moisture [ZWC^{+}24b]. monitor [FBS^{+}23, GCD^{+}23, STK^{+}21]. monitoring [BTH^{+}20, CRAN^{+}24, DHNM22, GSB23, HQY^{+}24, ITAAG^{+}24, JCD^{+}23, MOO^{+}24, MBW^{+}22, MBS^{+}22, RGJCN^{+}23, VH22, WKG^{+}21]. montane [ANR23, GTC21, Ric21]. moral [BNW20, BNB^{+}21, CC24, HSHA20, LSÁH^{+}20]. mortalities [LMY^{+}24, TAA^{+}24]. mortality [DSZ21, DBMS24, MLS^{+}22, RvTW22, STJ20]. most [Ala25, RGBC20]. moths [OPF24]. Motivating [BZT^{+}22]. motivational [YWH^{+}24]. Motivations [TWHB^{+}21]. mounds [DBR^{+}23]. mountain [DAD23, TJE^{+}24]. movement [Ala25, BRW21, DSS^{+}23]. movements [KKE^{+}24, LEF^{+}24]. Moving [FMGS21]. much [SP23b]. multiactor [BHM21]. multidata [HLL^{+}24]. multidimensional [BSG^{+}23, MHGC21]. multidisciplinary [MBD^{+}23]. multilateral [WSSC22]. multiplayer [DMR^{+}21]. multiple [BSSG21, CCW24, GWB^{+}24, KCT22, MR22, MCM^{+}21, NW22, NLD^{+}20, RJC^{+}22, SW25, WSB^{+}20, ZXXL23]. multiple-population [NLD^{+}20]. multiple-stressor [GWB^{+}24]. multipredator [BFS24]. multiscale [ODG^{+}22]. multispecies [AHM^{+}22, BBH^{+}25, GFF23, MRTV24, NCE^{+}21, PMR^{+}22, PLD^{+}21, SÁT21]. multitaxa [MCP^{+}25]. multitaxonomic [RDS^{+}24]. multiuse [SSG^{+}23]. Murray [MWH^{+}22]. museomics [Bla24]. Mussel [BA22b]. mussels [SZH^{+}23]. mutualisms [SMLL24]. Myanmar [WYB^{+}21]. Myths [TSÁ20].

N [Cli23]. Nagoya [PP25]. Nam [LC24]. named [BAGV24]. nanopore [YCH^{+}24]. NAP [Sár23]. narrative [RLSH24]. narratives [STW22, TFFM22]. Nation [PSV^{+}21]. National [LHB^{+}24, CND^{+}23, CTM^{+}22, JGAH25, MAB^{+}23, MJDMAS20, MMA^{+}22, NBD^{+}24, VFRV24, VHHT21]. nationality [vESN^{+}20]. Nations [GRS^{+}23]. native [BLF^{+}23, Cas23, Eva21, FBS^{+}23, GCS^{+}21, LWD^{+}20, Per20, TG21, CHP20, PG20]. Natura [Con21, PSP^{+}20, RDS^{+}24, STA^{+}22b]. natural [CCPE^{+}25, DLL22, GGJ^{+}20, PBB^{+}22, SSCR^{+}23, TGM^{+}22, YLY^{+}20]. natural-anthropogenic [DLL22]. natural-resource [TGM^{+}22]. Naturalness [SC23, CP20a]. nature [BSH24, CND^{+}23, CLB20, CLS20, DRFL^{+}24, DFC21, KSF^{+}24, MBL^{+}23, Nut25, Rob24, STA^{+}22b, TB23, VHHT21, VLVD^{+}23, VGW^{+}25, WLA20]. Natusch [EDA^{+}21, SP21]. Navigating [DTSE25, HOT^{+}25, MSS^{+}24]. Near [HVA^{+}21, TJB^{+}23]. Near-term [HVA^{+}21]. near-threatened [TJB^{+}23]. nearshore [DYM25, PHK^{+}25]. necessarily [FJZ^{+}21]. necessary [LFT^{+}21]. Need [CP20a, GN20, LS21, PPE^{+}20, BBL22, HM24, LEF^{+}24, MLIE25, TCK^{+}20]. needed [HBB^{+}24, MDB^{+}24]. needs [BE23, BBG^{+}20, MCF^{+}24, SWC^{+}21]. Negative [RKR^{+}21, NTS^{+}21]. neglected [HSG^{+}21, SQJ^{+}22]. negotiations [WH22]. Neighbouring [Sár23]. nematode [ZSL^{+}24a]. neoliberal [BAA^{+}25]. Neotropical [BPN^{+}24, GTC21]. Nepenthes [RR24]. nesting [SBP^{+}23]. nests [GWR23]. net [MK21, RzEHB24, SRCS23, SSKBL21]. Netflix [ABC^{+}23]. Netherlands [BOB^{+}20]. network [BVH22, BAA^{+}25, CTP^{+}24, CRAN^{+}24, Con21, GBB^{+}20, HBY^{+}23, MSJ20, RGBC20, RHS^{+}20, SS24, SMSS^{+}21, XST^{+}20]. networks [KBCD20, KWY^{+}21, LNYY20, NJLC21, dLMGK22]. neutral [MBW^{+}22, XDA^{+}21]. newly [RR24]. NewSouth [Ano22i]. newspaper [GLL^{+}23]. next [HSG^{+}21]. Nibbelink [Cli23]. Niche [SCW^{+}24, DAD23, HMS^{+}23, TFT^{+}24]. Niche-based [SCW^{+}24]. nitrogen [RKR^{+}21]. NJ [Ano22j, Ano23n, Sár23]. No [LC24, GGB20b, LLYK24]. no-take [GGB20b, LLYK24]. nocturnal [LHB^{+}22]. noise [JM21, LBBT24]. non [BLF^{+}23, BHSM^{+}22, Cas23, CGA^{+}22, Eva21, GCS^{+}21, LWD^{+}20]. non-English-language [CGA^{+}22]. non-native [BLF^{+}23, Cas23, Eva21, GCS^{+}21, LWD^{+}20]. non-perennial [BHSM^{+}22]. Nonadditive [BGVK20]. nonbreeding [GGJ^{+}20, GPJJ^{+}21, LBBT24, VGB^{+}24]. noncaptive [KIA^{+}22]. nondetection [FOYZ22]. nongovernmental [TFJ^{+}25]. nonlinear [BJP^{+}20]. nontarget [PSP^{+}20]. normative [NJLC21]. norms [SBE25]. North [BWN20, BPR24b, CRC^{+}21, CFN^{+}24, EHC^{+}21, HTC22, NRN^{+}24, TRR^{+}20]. Northern [MHP23]. northwestern [KSF^{+}24]. Norton [Ano23n]. nose [CRC^{+}21, TSS^{+}22, WWVC^{+}25]. Not-So-Subtle [Ano23n]. Noted [Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano22i, Ano22j, Ano22f, Ano22g, Ano22h, Ano23n, Ano23k, Ano23l, Ano23m, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano25c, Ano25d]. Novel [FW22, GG24, MHC21]. November [Ano22b, Ano23b]. nutritional [TBJ^{+}24]. NY [Ano23n].

object [TB25]. objectives [BDG^{+}23]. objects [KAD^{+}22]. obligations [VFRV24]. observed [ZBM^{+}24]. Obstruction [RF23]. occupancy [And23, ABM^{+}22, PMR^{+}22]. occurred [Cli23]. occurrence [BAGK22, LFD^{+}25, VMP^{+}24]. ocean [BBH^{+}25, HK21, BBH^{+}25, MFW^{+}24, WOF^{+}20]. Oceania [DYM25]. Oceans [Sco20]. odyssey [AP23]. offenses [PWB^{+}21]. offering [AFV20]. officials [LC24]. offs [ADSG21, Dre24, PNXY^{+}22, RSRG^{+}20, SD22]. offset [SSN^{+}23, SdVD^{+}22]. offsets [Dre24, MVC^{+}21, MK21, SRCS23, TRR^{+}20]. offsetting [GGB^{+}20a, KB21, The24]. oil [GGBD20, LCEM24, RCRJ^{+}24, SMT^{+}22]. old [KSZ^{+}23]. oldest [LB22, LXZ^{+}22]. Oli [LFT^{+}21]. once [Cli23]. One [Ano22i, LGD^{+}25]. one-size-fits-all [LGD^{+}25]. ongoing [TWS^{+}25, WDHM21]. Online [DMR^{+}21, BLF^{+}23, BAGV24, HOWM24, JCD^{+}23, MSS^{+}24, OC24, The24, WDGM21]. Open [KTH^{+}22, JMB21, ROK^{+}22, SW25]. Open-source [KTH^{+}22]. ophidiomycosis [RGD^{+}25]. opinion [HHB23, RHFV23]. Opportunities [CCCB^{+}23, CCPE^{+}25, DHNM22, FWZ^{+}24, GR21, KTG^{+}22, MAB^{+}23, MWC^{+}20, TFJ^{+}25, UEH^{+}23, ZEO^{+}22]. opportunity [Arm23, MHJD23]. optimal [WSB^{+}20]. optimism [DW23, Far21]. Optimization [BTH^{+}20]. optimize [DH21]. Optimizing [LLYK24, YMB^{+}21]. Options [Cor20, RML^{+}20]. Orangutans [Poz20]. orchid [ZSVK21]. organisms [Won24]. organization [BO21, TFJ^{+}25]. organizational [HNC^{+}24]. organizations [QGC^{+}25]. ornamental [HHL^{+}24, OC24]. Osterkamp [Ano22j]. other [CTP^{+}24, Coo24, HWLC22, Pit21, WGZ24]. Our [Ano22j, ABC^{+}23]. outcome [ZAS^{+}24]. outcomes [ART^{+}25, BDG^{+}23, BDQ^{+}25, CN22, CBM20, GGB^{+}20a, JGAH25, MHGC21, MEK21, NAG^{+}20, ODS^{+}24, POH21, POH^{+}22, RJAR23, RBD^{+}21, SS23, SD20, WSC^{+}23]. outdoor [LGL20, LGD^{+}25]. output [EAdC^{+}21]. overcome [LGL20]. overexploitation [MLIE25, RJV^{+}24]. overstretch [SPA^{+}22]. Overton [JBB23]. oviparous [MWW^{+}24]. Owen [Ano22j]. Oxford [Ano22j]. oyster [GRS^{+}23, SCC23].

P [Ano22j, Cli23, Sár23]. Pacific [ARSL24, BWN20, COF^{+}25, WOF^{+}20]. package [HSSM^{+}25]. page [Ano22e, Ano23i, Ano23j]. paired [VMP^{+}24]. Pakistan [Pri21a]. Palau [SF23]. Palearctic [WFBT^{+}25]. palm [LCEM24, RCRJ^{+}24]. pan [CRAN^{+}24]. pan-Arctic [CRAN^{+}24]. Panama [MMS^{+}20]. Panda [LHB^{+}24, CQS^{+}22]. pandas [MCB^{+}20]. pandemic [BBL^{+}20, GSS^{+}22]. pangolin [YCH^{+}24]. panic [LSÁH^{+}20]. paperback [Ano22i, Ano22j, Ano23n, Cli23]. parachute [SF23, Uts24]. paradigm [And25, AP21]. paradigms [GKH^{+}24]. parasite [BA22b, Ele20]. Parasites [GCC^{+}22, BDQ^{+}25]. Park [LHB^{+}24]. parks [CTM^{+}22, MJDMAS20, VHHT21]. parrots [DDM24]. partially [TJC21]. participants [IB20, ITAAG^{+}24]. participation [ML22]. Participatory [JRKF21, BHM21, CRAN^{+}24, RGJCN^{+}23]. partnership [Bus22, CSS^{+}24]. partnership-driven [CSS^{+}24]. partnerships [WSC^{+}23]. passing [RvTW22]. Past [Len21, MFW^{+}24]. patch [ABM^{+}22, RF23]. pathogen [FW22]. pathogens [TAA^{+}24]. Patterns [BFS24, KxGG^{+}25, VPVdAMF20, VMP^{+}24, ZWC^{+}24a]. pavement [MEA^{+}22]. payment [JDB^{+}21]. payments [HHMO24, ITAAG^{+}24]. peatland [RMW23]. pedagogy [CWMB23]. peer [CBD22]. penguins [BBG^{+}20]. People [OGF^{+}21, BIdL^{+}20, GCA^{+}22, GGA24, GCS^{+}24, VLVD^{+}23, VFKT24]. peoples [Bus22, SDB^{+}24]. percentage [CN22]. percentage-protected [CN22]. perception [CBM20]. perceptions [CHA^{+}24, DSC^{+}24, FCH^{+}24, GWR24, ITAAG^{+}24, STA^{+}22b, TCG^{+}21]. perennial [BHSM^{+}22]. Pereyra [CHP20]. performance [GGLB25, HC22, LTH^{+}21]. perils [MCLM22, NAS21a]. permanent [KRB^{+}22]. perpetrator [UEH^{+}23]. perpetuate [PLKP24]. persistence [BP21, FGL^{+}21, MVC^{+}21]. persistent [KCT22]. personality [CHW^{+}23, MS23]. personhood [WBB^{+}20]. Perspectives [AMC^{+}24, BKK^{+}21, CFN^{+}24, Cor21, DPC^{+}24, DFS^{+}25, MMG^{+}23, PP25]. perturbations [KBB^{+}23]. Peruvian [MEM^{+}22]. perverse [POH^{+}22]. pet [FW22, GSB23, NCdlLZ24, PLKP24, RJV^{+}24, ZWC^{+}24c]. Petrossian [NAS21b]. phenology [PSV^{+}21]. philanthropy [BGB^{+}24]. Philosophy [Ano22i]. photographs [MSN^{+}22, VHHT21]. phylogenetic [JSGJ^{+}21, MCF^{+}21]. Phylogeny [TA22]. Phylogeny-based [TA22]. physical [Cie24]. physiological [KFP^{+}21]. Pickrell [Ano22i]. pioneer [AP23]. pitcher [RR24]. Pitfalls [NAS21b, DBT21]. Place [Sco21, CLB20, CWMB23, GGA24]. plan [STJ20]. Planet [ABC^{+}23]. planetary [LWS^{+}24]. planned [BHT^{+}24]. planning [Ala23, AL21, BTDS20, CFA^{+}24, HVVA^{+}21, INO23, JRKF21, LMG^{+}23, LWG21, MBLD^{+}25, PHB^{+}25, RGBC20, SBH^{+}23, SN22, WYB^{+}21, WLR^{+}24, WSB^{+}20, XDA^{+}21]. plans [RVLD20]. Plant [Rap21, BOPM^{+}24, BGVK20, COF^{+}25, DPC^{+}24, FBDA22, GCA^{+}22, GGSRF20, HGM^{+}22, KFN^{+}21, KZM23, LWD^{+}20, TPP^{+}25, VH22, WLB^{+}23, WJ21, WBC^{+}20, ZSL^{+}24a]. plantations [RCRJ^{+}24]. Planting [MA20]. plants [ADG^{+}20, BLF^{+}23, CS21, CSHM^{+}20, KSC^{+}20, MMG^{+}23, RR24, VMR21]. plastic [MBI^{+}22]. Plateau [JZLP^{+}23, KMC^{+}23]. plates [ZYM^{+}24]. plea [CFS^{+}21]. pluralism [HAE^{+}24, SC23]. poaching [UEH^{+}23, VLPVM^{+}24, YCH^{+}24]. point [MHSL^{+}23]. policies [LGD^{+}25, MTS^{+}21, RzEHB24, TzEN^{+}24]. policy [BKK^{+}21, CHR24, GW21, OLF24, PHMVL^{+}22, SWC^{+}21, STA^{+}22b, TB23, CCCB^{+}23]. Political [Cor21]. politics [Jen22]. pollination [GHR21, RBC^{+}22]. pollinator [AAvdW^{+}22, FBG^{+}25, GHR21, MTS^{+}21]. Pollinators [Sár20, MA20]. pollution [MBI^{+}22, RGJCN^{+}23]. pollution-related [RGJCN^{+}23]. pond [CS24]. pondscapes [MBV^{+}24]. Poor [PGRDOGG23, TWS^{+}25, DS22, POH^{+}22, RML^{+}20]. poorly [DBT21]. popularity [VMR21]. population [APE^{+}21, BP21, CO20, CO21, CQS^{+}22, FL20, HRCA21, IJK^{+}21, KBL^{+}23, LFT^{+}21, MBD^{+}23, MRTV24, MBS^{+}22, NW22, NLD^{+}20, OCC^{+}21, QZX^{+}24, RMFP^{+}23, SMT^{+}22, STW^{+}24, WLR^{+}22, WAA^{+}21, ZLZ^{+}21, ZAS^{+}24, Ano22j]. population-level [OCC^{+}21]. populations [Ano24f, CLM22, DBM^{+}25, FGL^{+}21, FGB^{+}23, FBG^{+}25, GBMD^{+}22, JAB^{+}21a, KFU^{+}24, NW22, PLTC^{+}23, VH22, VNC^{+}24, WJ21, WB23, WFH^{+}21]. porpoises [BTM^{+}23]. Porter [Cli23]. posed [CFC21, DBMS24, HCD^{+}20]. position [ZSL^{+}24a]. positionality [CSS^{+}24]. Positioning [MBK^{+}20]. positive [BSH24, CN22, CLS20, MK21, MSRP21]. possibilities [DBT21]. post [HK21, MSJ20, Ric22]. post-2020 [HK21, MSJ20, Ric22]. postrelease [Ano23c, CLGP23]. Potential [NTS^{+}21, ADG^{+}20, Ano23c, Ano24e, BPR24a, CLGP23, DL24, DH21, EPL^{+}24, HOWM24, KxGG^{+}25, LKF^{+}23, MSJ20, OABDCM24, RRH^{+}24, SH21, SRCS23, ZAS^{+}24]. Power [Jen22, BGB^{+}24, SKS22, Ano23n]. pp [Ano22i, Ano22j, Ano23n, Cli23, Sár23]. Practical [CACM23, HTC22, HPG20, Kho22, VFRV24]. practice [BGMEHP20, Cie24, CO24, DdLI^{+}25, KBM^{+}21, MBK^{+}20, ME25, RW23, TB23, VFKT24, VRBAB^{+}24, WSEL^{+}24]. practices [CLD^{+}21, DPC^{+}24, HNC^{+}24, ODG^{+}22, VDO^{+}24]. Practitioner [SMB^{+}20, BGKKC22]. practitioners [PKC^{+}23b]. Pray [Pri25]. pre [FL20]. pre-exploitation [FL20]. precarious [THSA24]. precaution [SB21]. precautionary [Cas23, Kus20]. precluded [SENTI24]. predation [Eva21, POH^{+}22, WB23]. predator [DSS^{+}23, JP21, PMLM20, Ano22j]. predator-free [JP21]. predators [FJH21, Kho22, VPZ25, WRBL^{+}23]. predict [BPN^{+}24, PB21, STW^{+}24]. predictability [vBDD^{+}20]. predicted [MFW^{+}24]. Predicting [BHRBK20, CSGG25, PBB^{+}22, WB23, JGAH25]. predictions [BNI25, MHRF20]. predictor [BBM^{+}20]. predictors [AOI22, BOPM^{+}24, BAGK22]. preemptive [SENTI24, TENI23]. preferences [ZYM^{+}24]. prescribed [DFW^{+}24]. presence [DSF^{+}23, INO23, SWBD20, VG20]. Presenting [Kut20]. preserve [DH21, Ogg20]. Press [Ano22i, Ano22j, Ano23n, Cli23, Sár23]. Pressey [ÁRAB^{+}24]. pressing [BBG^{+}20]. Pressure [OPF24, FCL^{+}22, HWP^{+}21, VLPVM^{+}24]. pressures [MSMH22, SWBD20]. Prevalence [MMG^{+}23, MHB^{+}22, And23, DFB^{+}24]. prevent [UEH^{+}23]. preventative [RRR^{+}24]. Preventing [LWS^{+}24]. prey [DSS^{+}23, Pri25, WRBL^{+}23]. primary [BRJ25]. primate [FCH^{+}24]. Primates [Gru20, Cli23, GH22, KFP^{+}21, WCY^{+}25, Cli23]. Primatology [Cli23]. Princeton [Ano22j, Ano23n, Sár23]. principle [Kus20, SC23]. principles [CCH^{+}21, GAB^{+}22, McC24]. prior [LC24]. priorities [BHRBK20, CCW24, JTD^{+}20, MTH^{+}24, MSB25, TA22, ZXXL23]. prioritization [EHBT20, HSSM^{+}25, MFW^{+}24, MCC^{+}23a, RHS^{+}20]. prioritize [DFS^{+}25, LDC^{+}24, WRW^{+}20, WCW^{+}22, XST^{+}20]. Prioritizing [CSL^{+}23, DNCC22, DS22, HXW^{+}23, LWC^{+}24, LKF^{+}23, MWC^{+}20, STJ20, WKT^{+}21, YVA22]. prioritizr [HSSM^{+}25]. Priority [NLV20, JAB^{+}21b, MSLBS23]. priority-setting [JAB^{+}21b]. privacy [DFH^{+}21]. private [HNC^{+}24, POH21, PHB^{+}21, SJRG23, WDH^{+}21, dlLBB^{+}23]. privilege [SB21]. proactive [BGD25, VPS22]. probabilistic [WKT^{+}21]. probability [RCL^{+}22]. problem [THSA24]. problems [MHM^{+}24, NW22, Sch21]. process [And25, HCC^{+}21, MCC^{+}23b]. processes [BHM21, CHR24, Cie24, YLY^{+}20]. processing [CCPE^{+}25]. producing [KBW^{+}23]. production [JBE^{+}22a, SH21]. productive [Bus22]. productivity [Gok21, JDB^{+}21]. products [HWG^{+}22, TWHB^{+}21]. proenvironmental [WLA20]. program [MMA^{+}22, Ogg20, WJP^{+}20]. programs [DSM24, LKF^{+}23, ML22, SBG^{+}22, WKG^{+}21]. Progress [Ano22j, Coo24, Hyd24]. Projecting [SBG^{+}22]. projection [FBDA22]. projects [GCJB^{+}22, WDHM21]. promise [MCLM22]. Promote [Cor20, TWMM^{+}23]. Promoting [CLS20, KOO^{+}21, NWSL^{+}24, RW23, RML^{+}20]. propensity [LNYY20]. protect [FLW^{+}22, IB20, Pit21]. Protected [GJPJ^{+}22, MNMPGN23, SBH^{+}23, ACD^{+}24, Ano23d, BVH22, BLW23, CN22, CAFM^{+}23, CTP^{+}24, CMC^{+}23, CTM^{+}22, COOH^{+}25, CFC21, FCL^{+}22, GPJJ^{+}21, GGB20b, GTAdPA24, GWB^{+}24, HNC^{+}24, HGAB^{+}24, HWP^{+}21, HC22, HMS^{+}23, IDJ^{+}25, KGK^{+}24, KBCD20, LdSB^{+}20, LZG^{+}25, LLYK24, MHJ^{+}24, MSMH22, MFW^{+}24, MSMW21, MSJ20, NDS^{+}20, OPF24, OCC^{+}21, POH21, PMSC22, PSJG23, RCBT21, SAB^{+}22, SWBD20, TJC21, VFKT24, WCY^{+}25, WOF^{+}20, YTW^{+}25, YVA22, ZTCF21, ZYNC25, ZSL24b]. protected-area [HWP^{+}21, MSJ20]. Protecting [BBGL21, BDF^{+}22, BAB^{+}23b, SLM^{+}20, ZEO^{+}22]. protection [CFS^{+}21, HNC^{+}24, HRHB20, MSPM20, PLC22, SS24, TJE^{+}24, YVA22]. protocol [CAA^{+}24, PP25]. provision [ML22]. provisions [IB20]. proximate [McC24]. Psychological [LTJ^{+}20]. psychology [DWC^{+}20, PMLM20, WL20]. psychometric [KLVW20, WTAD21]. Public [STES23, MGJF21, MCG^{+}21, SBE25, Van23]. Publications [Ano22i, IB20]. Publishing [Ano22i, YHW^{+}25]. pulse [LWRZ^{+}21]. purposes [FWP^{+}21]. putatively [XDA^{+}21].

Qinghai [JZLP^{+}23, KMC^{+}23]. Qinghai-Tibetan [KMC^{+}23]. qualitative [CEM^{+}23, CBD22, GCA^{+}22, GAB^{+}22, MB21, SPK^{+}22]. quality [ABM^{+}22, CO21, FBG^{+}25, MW25, ZYNC25]. quantification [CP20b]. quantifications [And23]. quantify [STK^{+}21]. Quantifying [CGL^{+}25, Eva21, FOYZ22, HLL^{+}24, JBE^{+}22b, LTH^{+}21, MAB^{+}23, MVC^{+}21, RCL^{+}22, SW25, GAB^{+}21, SP23b, VG20]. Quantitative [BGD25, ZXR24, MBS^{+}22, SPK^{+}22]. quantity [FBG^{+}25]. Queer [MCW^{+}24]. questioning [IDD^{+}22]. questions [Ano23d, CMC^{+}23]. Qur’an [BGAE25].

R [Ano22j, Cli23, CCKT25, HSSM^{+}25]. radar [NHSM^{+}24]. rainbow [PWW^{+}22]. rainforest [FGB^{+}23]. rainforests [HBL20]. Ramsar [Pit21]. randomized [WJP^{+}20]. range [Ala23, And23, CGB^{+}24, KxGG^{+}25, MHRF20, MHP23, Per20, VRY^{+}23, CHP20, PG20]. range-restricted [MHP23]. ranges [JZLP^{+}23]. Rapid [GFF23, JP21, RBC^{+}22, RML^{+}20, BCA^{+}21]. Rapidly [RCRF20]. raptors [GJC^{+}23, ZXR24]. rare [BOPM^{+}24, BHM20, BGVK20, FSBE22, MVC^{+}22]. rarity [SSEN20, WWB^{+}22]. Rasmont [Sár23]. Rasolofoson [JBE^{+}22b]. rate [GGBD20, KSC^{+}20, LWRZ^{+}21, STW^{+}24]. Rates [SDB^{+}24, HGM^{+}22]. ray [Ano23d, CMC^{+}23]. rays [CSCA^{+}23]. razor [LMY^{+}24]. reading [HPM^{+}20]. real [JCD^{+}23]. real-time [JCD^{+}23]. realism [Nov23]. Realities [AFV20]. realized [DAD23]. Reallocating [WDHM21]. reasoning [THSA24]. reassessment [CSL^{+}23]. reassessments [LDC^{+}24]. rebuild [DBM^{+}25]. rebuilding [RML^{+}20]. recapture [BTH^{+}20]. Received [Ano20b, Ano22b, Ano23b, Ano24b]. recipient [PB21]. reckoning [Far21]. recognition [BBH^{+}25, MHSL^{+}23]. Recognizing [PKC^{+}23b, WBB^{+}20]. Recommendations [CBD22]. recommended [CLD^{+}21]. Reconciling [BSSG21]. Reconsidering [HFW20]. recorders [VMP^{+}24]. records [CFA^{+}24]. recovered [JPAR^{+}22]. Recovery [PLTC^{+}23, ARK^{+}20, BFS24, GAB^{+}21, JP21, NLD^{+}20, PHB^{+}21, PBB^{+}22, RVLD20, SHA^{+}23, WCW^{+}22]. recurring [GFF23, KBB^{+}23]. Red [BYHT^{+}20, BTM^{+}23, CSL^{+}23, DBT21, GRP^{+}25, KGK^{+}24, LDC^{+}24, MSC^{+}24, MHB^{+}22, SHH^{+}23, Van23, BDF^{+}22, SP23b]. Red-list [BTM^{+}23]. REDD [GCJB^{+}22]. Redefining [BSP22]. Redesigning [BPH^{+}22]. reduce [AAS21, CGM^{+}24, HM24, OABDCM24, RRH^{+}24, STJ20, SWBD20, TGJ^{+}20, WLR^{+}24]. reducing [Ano23c, CLGP23, GCJB^{+}22, LTJ^{+}20]. Reduction [Gok21, HM24, OGMMG22]. Reef [TCG^{+}21, WFH^{+}21, AHM^{+}22, DBM^{+}25, GRS^{+}23, MMA^{+}22, MHS^{+}21, MSRP21, QHS22, TFFM22]. reefs [KTG^{+}22, MDB^{+}24]. references [Rob24]. Reflections [Ano20v, JAB^{+}21b]. reflexivity [BBL22, CSS^{+}24, PKC^{+}23b]. reforestation [KCH^{+}20, QHS22]. reform [CCCB^{+}23]. Reframing [MHM^{+}24]. refugia [MDB^{+}24, RJC^{+}22, SP23a]. regime [JGAH25]. regimes [HOT^{+}25, RHGRSG21]. Region [MHP23, COF^{+}25, SSG^{+}23]. Region-wide [MHP23]. regionalization [Ala23]. regions [DS22, HNC^{+}24, MHJD23, MNMPGN23, RDS^{+}24, TJE^{+}24, WWB^{+}22, Ano22j, Sár23]. regression [FMGS21]. regrowth [PBB^{+}22]. regulating [ARSL24]. regulations [PLKP24, WH22]. regulatory [SW25]. reintroduced [APE^{+}21]. reintroduction [AAvdW^{+}22, DSM24, GMEV^{+}21, MBW^{+}22, STES23]. reintroductions [ABM^{+}22, HQY^{+}24, MRTV24]. related [HSHA20, KSZ^{+}23, RGJCN^{+}23]. Relationship [HPM^{+}20, TAA^{+}24]. Relationships [DW23, Ele20, GCS^{+}21, SKS22]. Relative [GBRK^{+}20, MHS^{+}21, FBG^{+}25, GSS^{+}22, LTH^{+}21, WGZ24]. relevance [BHSM^{+}22]. relevant [CBG^{+}21]. Reliability [GSB23, RMN^{+}21]. remote [COF^{+}25, LNDM21, MSN^{+}22]. remotely [DBT21]. remove [UEH^{+}23]. renewable [CLD^{+}21, VNC^{+}24]. renewable-energy [VNC^{+}24]. Renewing [Ano22j]. repeatable [GN20]. repetition [LWRZ^{+}21]. replacement [UEH^{+}23]. Reply [NBB^{+}21, CO21, FJZ^{+}21, GN20, JBE^{+}22b, LSÁH^{+}20, MR20b, NAS21b, PG20, Sag20b, SGB^{+}20, SÁT21, STSJ22]. reporting [VG20, WDH^{+}21]. Representation [BVH22, WSB^{+}20]. representative [KBCD20]. representativeness [WJ21]. representing [RBJN21]. Reproducibility [Sta22a]. reproduction [LFD^{+}25, TG21]. reproductive [MSHE23]. Reptile [SCB^{+}25, KAD^{+}22]. reptiles [CST^{+}24]. Republic [DAN^{+}21]. requirements [NFT^{+}20]. rescue [JPAR^{+}22, MPM^{+}23, RBD^{+}21]. research [BIdL^{+}20, CSS^{+}24, CBD22, DYM25, EAdC^{+}21, EGCS25, GCA^{+}22, GBB^{+}20, IB20, JM21, JAB^{+}21b, KBM^{+}21, KSFP^{+}21, KvZS^{+}25, KLVW20, LKL^{+}21, MTH^{+}24, MCF^{+}24, MSS^{+}24, MW25, NLV20, OLF24, RRH^{+}24, SF23, THM^{+}21, VDO^{+}24, WCY^{+}25, WTAD21, WMK^{+}21]. reserve [SS24]. reserves [KSF^{+}24, MSRP21, VD20]. resident [STA^{+}22b]. residue [BNW20]. resilience [KBB^{+}23, LLYK24]. resiliency [WAA^{+}21]. resistance [Ele20]. resistant [OPF24]. resisting [CCKT25]. resolution [PWW^{+}22, TDC^{+}23, ZLZ^{+}21]. resolve [ABM^{+}22]. resort [FJZ^{+}21]. Resource [WSK^{+}22, BBGL21, BAGK22, TGM^{+}22, WAA^{+}21]. resource-rich [BBGL21]. resources [ABRCM21, DdLI^{+}25, GMEV^{+}21, MCM^{+}21]. respond [GJPJ^{+}22]. Response [SP22, Wal24, ABC^{+}23, CC20, CBD^{+}21, CHP20, CMAS20, CBB^{+}20, EDA^{+}21, EHBT20, Hen22, HCZW24, IR22, KOO^{+}21, Kus20, LFT^{+}21, MCLM22, MBD^{+}21, PLMF23, Ras22, SP21, Sta22a, WSW^{+}22, WOF^{+}20, dG21, Ano22j]. Responses [FLH21, DFW^{+}24, EN24, LCEM24, MSC^{+}24, MR22, SCB^{+}25, SUA22, SSEN20]. responsibility [VDO^{+}24]. Restoration [Ano22j, ADG^{+}20, AAvdW^{+}22, AP21, DL24, DSS^{+}23, DBM^{+}25, GRS^{+}23, HMC23a, INO23, MMA^{+}22, ODS^{+}24, QHS22, RJAR23, RGJCN^{+}23, RMN^{+}21, SBE25, SCC23, SSKBL21, VRBAB^{+}24]. restored [SBP^{+}23]. restricted [MHP23]. restrictive [HtSR25]. Retaining [SSCR^{+}23]. rethink [VM20]. Rethinking [Gon20, Per20, PBV21, SR20]. retreats [MHP23]. return [TGJ^{+}20]. return-on-investment [TGJ^{+}20]. returns [BBGL21]. Review [Ano22a, BM24, PHK^{+}25, WBR^{+}24, Ano23c, Cas23, CLGP23, CBD22, CC22, DDM24, GAB^{+}22, HtSR25, MS23, MBS^{+}22, RMN^{+}21, TWMM^{+}23]. Reviewers [Ano20a, Ano23a, Ano24a]. Reviews [Sár23]. revisited [DSvE^{+}25, Ras22]. Revisiting [FFAS23]. revitalization [GRS^{+}23]. reward [JDB^{+}21]. Rewilding [CN21, Ano24e, BPR24a, CCH^{+}21, GGA24]. rhinoceros [VPS22]. Rica [KSF^{+}24]. rich [BBGL21, PPN^{+}24]. richness [BA22b, LKA^{+}22, LWD^{+}20, MFW^{+}24]. Rigid [ODMG^{+}20]. riparian [HRHB20, HMM^{+}24, RCRJ^{+}24]. rise [VGB^{+}24]. rising [MWW^{+}24]. risk [BSG^{+}23, BMA^{+}20, BTM^{+}23, BA22b, CBM20, CSGG25, DBMS24, GG24, HLZ^{+}21, KGK^{+}24, KCT22, KKE^{+}24, LTJ^{+}20, LDC^{+}24, MCH^{+}20, MHM^{+}22, OC24, RRH^{+}24, RJV^{+}24, RWC^{+}22, SWM^{+}24, TBJ^{+}24, TBP^{+}22, VLPVM^{+}24, WGZ24]. risk-based [KGK^{+}24]. risk-reducing [LTJ^{+}20]. Risks [CFC21, KOO^{+}21, FZG^{+}24, SSN^{+}23]. risky [PLKP24]. river [HBY^{+}23, INO23, KKE^{+}24, PWW^{+}22]. Rivers [Ano22j, BHSM^{+}22, FGB^{+}23]. road [STJ20]. roadmap [SHA^{+}23, SZH^{+}23]. roads [BHT^{+}24, DBMS24, LWD^{+}20]. robust [BB21, Kho22]. robustness [Des21]. rocky [BMH^{+}24]. rodent [LKL^{+}21, PLTC^{+}23]. Rodewald [Kus20]. Role [CGM^{+}25, JGAH25, BGKKC22, PMLM20, WSvRLK20]. roles [ABM^{+}22]. roosting [MJPJ^{+}24]. roosts [CO24]. routine [PPE^{+}20]. rules [BDG^{+}23, IDJ^{+}25, NJ20, QHS22]. rural [CVT^{+}21, LNYY20, TE20]. Rylands [Cli23].

S [Ano22j, Cli23]. safeguard [SSCR^{+}23]. safeguards [NSL^{+}23]. safety [CFN^{+}24]. sage [ZAS^{+}24]. sage-grouse [ZAS^{+}24]. Sagoff [CBB^{+}20]. Saharan [DNCC22]. Sahel [NTS^{+}21]. sake [CMAS20]. salamander [CS24]. sale [BLF^{+}23]. sample [BAGK22]. sampling [NDS^{+}20]. Santangeli [EHBT20]. Santiago [CBD^{+}21]. Santiago-Ávila [CBD^{+}21]. São [PPN^{+}24]. satellite [LNDM21, MSN^{+}22, ZLZ^{+}21]. savannas [SPA^{+}22]. saving [WSW^{+}22]. Savviness [WRBL^{+}23]. sawfish [LEF^{+}24]. Scale [CST^{+}24, GGB^{+}20a, BOAN23, BNI25, DTSE25, EHBT20, FBDA22, GHR21, HBY^{+}23, INO23, PBB^{+}22, RVLD20, RHGRSG21, RJC^{+}22, SWCP^{+}22, TBJ^{+}24, VMP^{+}24]. scales [DSC^{+}24, MCC^{+}23b, RGD^{+}25, WTAD21, YCH^{+}24, Cli23]. Scaling [BBH^{+}25, GCA^{+}22, PHB^{+}21, KBCD20]. scaling-up [KBCD20]. scenario [JRKF21]. Scenarios [DAD23]. scheme [GGB^{+}20a]. schemes [ITAAG^{+}24, MLIE25]. Schleicher [Ras22]. Science [Imb22, Poz20, BH23, BGMEHP20, BE23, BBG^{+}20, BBL22, BBL^{+}20, CACM23, Cie24, CO24, DTSE25, Des21, DdLI^{+}25, DFH^{+}21, FFAS23, FSBE22, GRP^{+}25, HCA^{+}22, JMB21, JS25, JJ21, KBW^{+}23, KSFP^{+}21, KvZS^{+}25, KRB^{+}22, LGD^{+}25, LSÁH^{+}20, MS23, MR20a, MW25, ME25, NAS21a, PSP^{+}20, QGC^{+}25, Ras22, ROK^{+}22, SEB^{+}20, SVP^{+}22, SF23, Sra23, SB21, THM^{+}21, TFJ^{+}25, TB25, Uts24, VBK^{+}23, VM20, WDGM21, WGAB^{+}21, WSEL^{+}24, dG21]. Sciences [Sco21, DSvE^{+}25, GB22, VBC^{+}25]. Scientific [FSBE22, CHR24, FWP^{+}21, IR22, LVT^{+}20, RMW23]. scientists [TWGGB^{+}25]. scope [CRC^{+}21, Hul22, PRS^{+}25]. Scotia [WH22]. sea [AAS21, EPL^{+}24, GWR23, MBLD^{+}25, RMFP^{+}23, VGB^{+}24, WSK^{+}22, WH22]. sea-level [VGB^{+}24]. seabird [DBM^{+}25, PLTC^{+}23, WFH^{+}21]. seabirds [BFS24, CRAN^{+}24, Sco22]. seagrass [HMC23a]. search [BHM20]. seas [XAAY^{+}21, LLYK24]. season [LBBT24]. seasonal [SUA22, VMR21]. seawater [STF^{+}20]. seaweed [KCS20]. secondary [QOF^{+}21]. sections [STJ20]. sector [PKC^{+}22]. secure [MDB^{+}24]. security [DAN^{+}21, NBD^{+}24]. seed [WB23]. segmentation [NCdlLZ24]. seized [YCH^{+}24]. Selection [CvH^{+}22, WLR^{+}24, MBH^{+}21, TDS^{+}23]. self [HSHA20]. self-affirmation [HSHA20]. semantics [CDSB23]. sensed [DBT21]. sensing [COF^{+}25, LNDM21, MSN^{+}22]. sensitive [JPK^{+}21]. Sensitivity [ANR23, NRN^{+}24, BPR24b, EN24, LHB^{+}24]. Sentiment [ABC^{+}23, LVT^{+}20, The24]. series [LNDM21]. Service [Imb22, GH22, SBE25]. services [DSC^{+}24, HRHB20, ITAAG^{+}24, LHB^{+}24, ML22, RBC^{+}22, RSRG^{+}20, SH21, SCC23, VFKT24, WSB^{+}20]. sets [DSF^{+}23]. setting [HWLC22, JAB^{+}21b, MAB^{+}23, SMB^{+}20]. settings [BHM21]. severity [CRC^{+}21, DFW^{+}24]. Shaffer [Cli23]. shaping [CVT^{+}25]. Shared [BGMEHP20, CBM20, MB21]. sharing [LKF^{+}23, PP25, SAB^{+}22]. shark [Ano23d, CMC^{+}23, CGM^{+}24, QOF^{+}21]. shark-fisheries [QOF^{+}21]. sharks [CSCA^{+}23, TBJ^{+}24]. shearwater [NHSM^{+}24]. Shifting [GKH^{+}24, JBB23, NFD20]. shifts [Ala23, AL21, KxGG^{+}25, SBW^{+}22]. shorebird [CXJ^{+}22, LMY^{+}24]. shorebirds [CMP^{+}24, SSN^{+}23]. shoreline [CGL^{+}25]. short [KBL^{+}23]. short-term [KBL^{+}23]. Show [Wal24]. shown [HPM^{+}20]. side [DDM24]. signal [MHM^{+}22]. significance [SP22, WKG^{+}21]. similarity [MCP^{+}25]. simulated [MPM^{+}23]. Simulating [EPL^{+}24, SBG^{+}25]. simulation [OGMMG22]. Singapore [KSC^{+}20]. site [MBH^{+}21, MCC^{+}23a]. site- [MCC^{+}23a]. site-selection [MBH^{+}21]. sites [GMEV^{+}21, LMY^{+}24, SBP^{+}23, GR21]. situ [ADG^{+}20, DH21, FJZ^{+}20, FJZ^{+}21, KOO^{+}21, WJ21]. size [BMG^{+}22, BAGK22, DFW^{+}24, FL20, LGD^{+}25, MBD^{+}23, NW22, NFT^{+}20, RF23, WLR^{+}24, ZLZ^{+}21]. skills [TWGGB^{+}25]. skin [RGD^{+}25]. slot [GGB20b]. small [Eva21, KCT22, NW22, RHGRSG21, SWCP^{+}22, TBJ^{+}24, VD20, WAA^{+}21]. small-scale [RHGRSG21, SWCP^{+}22, TBJ^{+}24]. snail [PPN^{+}24]. snake [Ano23c, CLGP23, RGD^{+}25]. snakes [ARSL24]. Snow [Pri21a, SUA22]. Social [HSN^{+}21, QGC^{+}25, WGZ24, BO21, BMM^{+}24, BSSG21, BAA^{+}25, CFA^{+}24, DHE^{+}21, DHNM22, DS25, DTSE25, DdLI^{+}25, DFH^{+}21, DSvE^{+}25, DMR^{+}21, EOR^{+}24, GSB23, GBB^{+}20, JS25, KvZS^{+}25, LGD^{+}25, LH21, MVC^{+}22, MJDMAS20, MHSL^{+}23, MW25, NAG^{+}20, NJLC21, POH21, RGBC20, SBE25, TWGGB^{+}25, TFJ^{+}25, TB25, TJC21, VHHT21, VBC^{+}25, WDGM21, dLMGK22, vESN^{+}20]. social-ecological [BSSG21, DMR^{+}21]. social-network [GBB^{+}20]. socially [KBCD20]. Society [WSEL^{+}24]. sociocultural [MTC^{+}20]. socioecological [HM24, WSK^{+}22]. Socioeconomic [AOI22, BAGK22, HOT^{+}25, MSMH22, PPN^{+}24, PSJG23, VRBAB^{+}24]. Software [WKT^{+}21]. soil [KAD^{+}22, ZEO^{+}22, ZWC^{+}24b, ZSL^{+}24a]. Solution [Pri21b]. solutions [LEF^{+}24, LGD^{+}25, MAC20]. Solving [Sco22]. songbird [LBBT24]. Sosnowski [NAS21b]. Soulé [Ano20v, Tay20]. sound [CRAN^{+}24]. source [BNB^{+}21, KTH^{+}22, WJ21]. sourced [MSHE23, SOW^{+}22]. sources [CLJ^{+}21, HLL^{+}24, HOWM24, VMR21]. South [MMN^{+}20, CTM^{+}22, SP23a, SBP^{+}23, dlLBB^{+}23]. Southeast [CSCA^{+}23, CSGG25, NGST21]. southeastern [ITAAG^{+}24, LMY^{+}24]. southern [HtSR25, JSKR22, PMLM20, BBH^{+}25, WWVC^{+}25]. space [GKH^{+}24]. spaces [Cli23, CLS20]. Spain [Cli23]. sparked [MAC20]. Spatial [CAFM^{+}23, SSKBL21, AL21, LXZ^{+}22, MCC^{+}23a, MCC^{+}23b, NDS^{+}20, PHB^{+}25, YVA22, ZWC^{+}24a, ZSL24b, Cli23]. spatially [CVT^{+}25, FGL^{+}21, NAG^{+}20, YLY^{+}20]. spatiotemporal [CGM^{+}24, DSZ21, LZG^{+}25]. spawning [GGB20b]. Speaking [SH21]. specialized [CvH^{+}22, IDD^{+}22, MSB25]. Speciation [Ano23n, TF24]. specie [BAGV24, PLMF23, SSEN20]. Species [EN24, FSC^{+}21, TENI23, ARK^{+}20, ADKM24, BOAN23, BSG^{+}23, BOPM^{+}24, BHM20, BGVK20, BYHT^{+}20, BAB^{+}23b, BMG^{+}22, BTH^{+}20, BCA^{+}21, CSL^{+}23, COF^{+}25, CS21, CGB^{+}24, CSGG25, DEGL^{+}24, FPC^{+}24, GFK^{+}22, GG24, GWR24, GMEV^{+}21, GBRK^{+}20, GAB^{+}21, GSMW21, GCS^{+}21, HWK^{+}20, HCC^{+}21, HtSR25, HLL^{+}24, HOWM24, JPAR^{+}22, JZLP^{+}23, JP21, KxGG^{+}25, KBL^{+}23, KOO^{+}21, KKE^{+}24, LKA^{+}22, LSM^{+}20, LWD^{+}20, LSM24, LKF^{+}23, LWS^{+}24, MHJD23, MHRF20, MAB^{+}23, MHB^{+}22, MVC^{+}21, MFW^{+}24, MBD^{+}21, MGJF21, MBS^{+}22, MEK21, NLD^{+}20, NSB^{+}21, NCE^{+}21, ODMG^{+}20, PCH22, PSP^{+}20, PMSC22, PLC22, PLKP24, RML^{+}20, RJV^{+}24, RHFV23, RR24, SP23b, SMT^{+}22, SSG^{+}23, SLM^{+}20, SCB^{+}21, SENTI24, TTDdLM^{+}21, TBP^{+}22, VRY^{+}23, VPVdAMF20, WLB^{+}23, WRW^{+}20, WCW^{+}22, WSW^{+}22, WSB^{+}20, Won24, WBC^{+}20, YMB^{+}21, Ano23n]. species-accumulation [LWD^{+}20]. species-based [PLC22]. Species-level [EN24]. species-specific [YMB^{+}21]. specific [YMB^{+}21]. specificity [WWVC^{+}25]. speculation [ZZ24]. Spicer [Ano22i]. spill [SMT^{+}22]. spirituality [Tay20]. sprawl [Con21]. spread [KKE^{+}24, RRR^{+}24]. springs [CFS^{+}21, Pit21]. spy [MSN^{+}22]. stability [PMSC22, SS24]. Stakeholder [DSC^{+}24, TCG^{+}21, VG20]. standard [GAB^{+}21, MSC^{+}24, SCdM^{+}24]. standardization [CLD^{+}21]. standardized [MCH^{+}20]. state [DL24, Nut25, QGC^{+}25, RCBT21, SOW^{+}22, VBK^{+}23]. States [Ano22j, CP20b, CGL^{+}25, GNH^{+}22, KFN^{+}21, QGC^{+}25, TW24, VBK^{+}23, WBR^{+}24, vESN^{+}20]. Statistical [Ras22, SEB^{+}20, SP22]. statistics [Hen22, May22]. Status [CSCA^{+}23, BMH^{+}24, BTM^{+}23, FOYZ22, GCD^{+}23, NCE^{+}21, PWW^{+}22, SHH^{+}23, WCY^{+}25, WSW23, YCS^{+}23, LKF^{+}23]. steppe [DBR^{+}23]. Steps [JAB^{+}21b, HMM^{+}24]. Sterling [BAB^{+}23a]. stewardship [MSMW21, RHGRSG21]. Stimulating [MO25]. stochastic [BB21, MLS^{+}22]. stock [RML^{+}20]. Stocks [Fur20, GGB20b]. stopover [LMY^{+}24]. Strand [Ano22j]. stranded [BP21]. strandings [TW24]. Strategies [VDO^{+}24, ANR23, BB21, BWN20, BPH^{+}22, CSHM^{+}20, MGS^{+}23, MTM^{+}22, MRS22, SR20, TGJ^{+}20, TJB^{+}23, WLR^{+}24, ZWC^{+}24a]. Strategy [FZG^{+}24, HSG^{+}21, TE20]. Stream [Ano22j, HRHB20, IMI^{+}21]. streams [SFC20]. strength [Ric21]. Strengthening [SB21]. Strengths [PHB^{+}25]. stress [KFP^{+}21, ODG^{+}22]. stressor [GWB^{+}24]. stressors [KBB^{+}23, RJC^{+}22, SW25]. Strive [Ano23n]. stronger [Cie24]. structural [YLY^{+}20]. structure [BMG^{+}22, TPP^{+}25]. structured [FGL^{+}21, GFF23]. studies [CLD^{+}21]. study [CDSB23, CTM^{+}22, CGL^{+}25, HOWM24, Kho22, PBV21, SSN^{+}23, WSEL^{+}24]. studying [BAA^{+}25]. sub [DNCC22]. sub-Saharan [DNCC22]. subject [MDB^{+}24]. Subjective [Des21]. sublethal [GGBD20]. submerged [PHK^{+}25]. Subterranean [Ano23n, GNH^{+}22, MJPJ^{+}24, STSJ22]. subterranean-roosting [MJPJ^{+}24]. Subtle [Ano23n]. subtropical [GFK^{+}22, KZM23]. success [AAvdW^{+}22, BOPM^{+}24, FDP^{+}23, RGJCN^{+}23, SR20]. Succession [Rap21]. succulent [MMG^{+}23]. suitability [BBM^{+}20, GMEV^{+}21, HHB23]. summer [vBDD^{+}20]. Sundarbans [UEH^{+}23]. supply [DDM24]. supply-side [DDM24]. Support [LKL^{+}21, BGD25, CSS^{+}24, CN22, CCCB^{+}23, DBC^{+}23, FSC^{+}21, MAB^{+}23, MA20, MHC21, MHD^{+}22, OGMMG22, RVLD20, SBE25, VBK^{+}23, WBJ^{+}20]. Supporting [EM21, PKC^{+}23a]. supportive [OAY^{+}24]. suppression [HBY^{+}23]. surf [MHJ^{+}24]. surface [KAD^{+}22]. surrogate [WRW^{+}20]. surrogates [HVVA^{+}21]. surveillance [CTL^{+}23]. survey [GFF23, KAD^{+}22]. surveys [CGB^{+}24, HGM^{+}22, LMG^{+}23, MVC^{+}22, PMR^{+}22, WDGM21]. survival [STW^{+}24, ZTCF21]. surviving [BRJ25]. sustainability [NAS21b]. Sustainable [Fur20, SWCP^{+}22, BPH^{+}22, CP20a, MLS^{+}22, MHB^{+}22, MWV^{+}20, VFRV24]. swamping [TF24]. Swiss [RKR^{+}21]. switch [HWG^{+}22]. sympatric [SLM^{+}20]. synchrony [AL21]. syndrome [CRC^{+}21, TSS^{+}22, WWVC^{+}25]. Synergies [ADSG21, HK21]. synthesis [CCPE^{+}25]. Synthesizing [SNC^{+}22]. synthetic [HWG^{+}22, MHC21, SJRG23]. system [CS24, JCD^{+}23, RCBT21, STW^{+}24]. Systematic [BSH24, HSSM^{+}25, HtSR25, TWMM^{+}23, WL20, Ano23c, CLGP23, GAB^{+}22, MS23, PHB^{+}25]. systems [ART^{+}25, BSSG21, CBM20, GHR21, MB21, WSK^{+}22].

T [Ano22i]. taboos [Sch21]. tackle [MLIE25]. tackling [Hul22]. tactics [Ano23c, CLGP23]. take [GGB20b, LLYK24]. tall [HXW^{+}23]. tangible [JDM^{+}21]. Tanzania [HTM25]. tarantula [RJV^{+}24]. target [MAB^{+}23, MBLD^{+}25, WDHR22]. targeted [Kho22]. targeting [DEGL^{+}24]. targets [CN22, HRCA21, MSJ20]. taxa [Cas23, KBB^{+}24]. taxonomic [GBRK^{+}20, ZSL^{+}24a]. taxonomy [DBC^{+}23, PWB^{+}21, Sár23]. Tea [BNS24]. technique [RCL^{+}22]. techniques [IDD^{+}22]. technology [Jen22, SOW^{+}22]. telemetry [KBL^{+}23]. temperate [BMH^{+}24, PHK^{+}25]. Temperature [HMS^{+}23, SS24]. temperatures [MWW^{+}24, MDG^{+}21, WWVC^{+}25]. temporal [PMSC22, SS24]. temporarily [WBR^{+}24]. term [APE^{+}21, FGL^{+}21, GBMD^{+}22, HVA^{+}21, KBL^{+}23, MBD^{+}23, MRTV24, OAY^{+}24, SPA^{+}22, TJB^{+}23]. terms [SCdM^{+}24]. terrestrial [ABCH^{+}21, BP21, BHS^{+}22, BHT^{+}24, BHSM^{+}22, CCW24, EN24, Kho22, KAD^{+}22, LDD^{+}22, MSB25, MHM^{+}22, NLV20, OGF^{+}21, SFC20, WBR^{+}24, ZYNC25]. territories [ECVL25]. territory [GWR24, PSV^{+}21, TDC^{+}23]. Terzo [Sár23]. Testing [GAB^{+}21, MFW^{+}24, Rap21, SWBD20, BF21, DSM24, DFW^{+}24]. tests [SACA^{+}20]. Their [Sár20, ADG^{+}20, ANR23, BMA^{+}20, GPJJ^{+}21, HBB^{+}24, IB20, PHMVL^{+}22, SWBD20]. theories [CC24, EGCS25]. theory [HPG20, HSHA20, KWY^{+}21, SGA^{+}24, Wal24, YWH^{+}24]. Think [Ver23]. thinking [BGMEHP20, FLW^{+}22, MGS^{+}23]. threat [Eva21, GFK^{+}22, HOWM24, MAB^{+}23, RCRF20, RRH^{+}24, SGA^{+}24, TCG^{+}21]. threat-management [RCRF20]. Threatened [MHB^{+}22, Ala25, BOAN23, BYHT^{+}20, BTH^{+}20, COOH^{+}25, FSBE22, GBRK^{+}20, GSMW21, HCC^{+}21, KBL^{+}23, LKA^{+}22, MSB25, SSG^{+}23, SQJ^{+}22, TDS^{+}23, TJB^{+}23, WBC^{+}20, ZTCF21]. Threats [DAN^{+}21, HCD^{+}20, NBD^{+}24, ACD^{+}24, ARKA22, BOB^{+}20, BGVK20, KCT22, KTG^{+}22, MCM^{+}21, SP23b, SHS^{+}22, TzEN^{+}24, TGJ^{+}20, ZXXL23]. Three [BBL^{+}20, Hal22, HMM^{+}24, Lau20, MBLD^{+}25, MK21, EHC^{+}21]. Three-dimensional [MBLD^{+}25]. thresholds [FMGS21, HC22, HWLC22].

Tibet [JZLP^{+}23]. Tibetan [KMC^{+}23]. tidal [CMP^{+}24, HWP^{+}21]. tiger [Pri25, SNC^{+}22, UEH^{+}23]. tiger-conservation [SNC^{+}22]. Time [BHS^{+}22, Car20, DWC^{+}20, JCD^{+}23, Kus20, LNDM21, NAG^{+}20]. time-dependent [NAG^{+}20]. Time-lagged [BHS^{+}22]. time-series [LNDM21]. timely [CTL^{+}23]. timing [AMC^{+}24]. tipping [MHSL^{+}23]. TOC [Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m]. tolerance [CDSB23, Ele20, VPZ25, WTAD21]. Tomé [PPN^{+}24]. tool [BSG^{+}23, BTC^{+}22, JPK^{+}21]. Tools [KLVW20, BHG^{+}23, CP20b, GCC^{+}22, RGJCN^{+}23, WBJ^{+}20]. top [VPZ25]. topic [GLL^{+}23]. topography [YLY^{+}20]. total [POH^{+}22]. totoaba [OABDCM24]. tourism [SPA^{+}22]. Towns [Gon20]. Toxoplasma [RRR^{+}24]. Tracking [GLL^{+}23, WOF^{+}20, BBH^{+}25, PPE^{+}20]. Trade [PNXY^{+}22, PPN^{+}24, ADSG21, CST^{+}24, DHNM22, DDM24, DFB^{+}24, Dre24, EDA^{+}21, EMV^{+}20, FW22, GSB23, HHL^{+}24, HtSR25, HHM24, HLL^{+}24, HM24, MMG^{+}23, MTH^{+}24, MEM^{+}22, MMN^{+}20, MSS^{+}24, MSHE23, NCdlLZ24, NAS21b, NAS21a, NJ20, OC24, OGMMG22, OABDCM24, OLF24, PLKP24, RSRG^{+}20, RJV^{+}24, RR24, SD22, SP21, STK^{+}21, VFRV24, VLPVM^{+}24, ZWC^{+}24c]. Trade-offs [PNXY^{+}22, ADSG21, Dre24, RSRG^{+}20, SD22]. traditional [JS25]. traffic [DBMS24]. traffication [Veg24]. trailing [GBDM^{+}20]. training [TWGGB^{+}25]. Trait [AHM^{+}22]. traits [AOI22, HLZ^{+}21, MSHE23, PLMF23]. trajectories [EPL^{+}24, MRTV24]. transboundary [HTM25]. Transcending [MTM^{+}22]. transect [KBL^{+}23]. transformation [LZG^{+}25]. transformative [HSN^{+}21, WSC^{+}23]. Transforming [Ogg20, PSS20, NCdlLZ24]. transfrontier [SBG^{+}20]. transgenic [AP21]. transitions [DL24]. Translating [EHBT20, LDD^{+}22, QHS22]. translocated [VH22]. translocation [BBM^{+}20, BOPM^{+}24, BTC^{+}22, DPC^{+}24, ZAS^{+}24]. translocations [Ano23c, BM24, BCA^{+}21, CLGP23]. transmissible [HMM^{+}21]. transparency [JMB21, MR20b, MW25]. transparent [GN20, SW25]. transportation [GH22]. trap [CGB^{+}24, GCSZ21, HHB23, LGD^{+}25]. traps [VMP^{+}24]. treaty [WSSC22, GGLB25]. Tree [Ano23n, FMGS21, GFK^{+}22, GCD^{+}23, KSZ^{+}23, MSPM20, QZX^{+}24, SDB^{+}24, TF24, WB23]. tree-related [KSZ^{+}23]. trees [CSHM^{+}20, KSZ^{+}23, LXZ^{+}22, ZXXL23]. trend [WSW23]. trend-based [WSW23]. Trends [Ano22i, CP20b, JM21, VBC^{+}25, WFH^{+}21, ZYNC25, DSZ21, EAdC^{+}21, FOYZ22, FMGS21, KBL^{+}23, LNDM21, TJB^{+}23, VMR21, Huy23]. Treves [CBD^{+}21]. triage [GBDM^{+}20, WAA^{+}21]. trial [WJP^{+}20]. triangulation [DHNM22]. Tribe [BNS24]. trigger [NSB^{+}21]. trolley [Sch21]. trophic [HBL20]. trophy [EGBJ23, MCB^{+}20]. Tropical [GFK^{+}22, ANR23, BBGL21, BP21, BBE^{+}25, EOR^{+}24, FPC^{+}24, GZBLS^{+}20, HBL20, KCS20, KZM23, MBL^{+}23, MLIE25, Ogg20, PNXY^{+}22, PBB^{+}22, RR24, SBE25, TFT^{+}24, WB23]. tropics [GCJB^{+}22]. trout [PWW^{+}22]. true [WSW^{+}22]. trust [SKS22]. Tucson [Ano22j]. tuna [CGM^{+}24]. tune [BHM21]. turbine [CSGG25, LBBT24]. turbines [DSZ21, RvTW22]. Turner [MCLM22]. Turning [MMA^{+}22]. turtle [AAS21, RMFP^{+}23, ZWC^{+}24c]. turtles [EPL^{+}24, GWR23]. Twitter [ABC^{+}23, BO21, EGBJ23]. Two [Nov23, FFAS23, LTH^{+}21, RGD^{+}25, TAA^{+}24]. type [NRSF21, Sta22a]. types [MDB^{+}24, WGZ24]. Typifying [BGKKC22].

U [Ano22j]. U.K [Ano22i, Ano22j, Cli23]. U.K. [JAB^{+}21a]. U.N. [SWCP^{+}22]. U.S. [NCE^{+}21, SENTI24, TENI23, WSSC22]. U.S.A [Ano22i, Ano22j, Ano23n, Sár23]. UK [Ano22j]. unappealing [LSM^{+}20]. uncertain [FJZ^{+}20, SBH^{+}23]. uncertainty [BH23, CHP20, JBE^{+}22b, RBJN21, WSW23]. uncover [VMR21]. Uncovering [YCH^{+}24]. underpinning [RzEHB24]. understand [BTDS20, LNYY20, NAG^{+}20, TTDdLM^{+}21]. Understanding [BF21, HNC^{+}24, HWG^{+}22, LGL20, RHFV23, YHW^{+}25, BNB^{+}21, DRFL^{+}24, RAG^{+}20]. understudied [TBP^{+}22]. undetected [LTH^{+}21]. undetected-extinction [LTH^{+}21]. undiscovered [KSC^{+}20]. UNESCO [GR21]. uneven [VPZ25]. ungulates [DAD23, RCL^{+}22]. unintended [PCH22]. Union [FZG^{+}24, MSJ20]. unique [MSB25]. unit [WLR^{+}24]. United [Ano22j, CGL^{+}25, Ano24e, BPR24a, CP20b, GNH^{+}22, KFN^{+}21, QGC^{+}25, TW24, VBK^{+}23, WBR^{+}24, vESN^{+}20]. universities [YWH^{+}24]. University [Ano22i, Ano22j, Ano23n, Cli23, Sár23]. unmarked [GCSZ21]. unrecorded [INO23]. unregulated [CST^{+}24, DAN^{+}21]. unrelated [Won24]. unreported [DAN^{+}21]. unsustainable [DDM24, MHB^{+}22]. unwanted [GG24]. update [LSM24]. Urban [Con21, EMV^{+}20, CVT^{+}21, IJ24, KWY^{+}21, RGJCN^{+}23, TG21, ZSL24b]. urbanization [CVT^{+}21]. Urgent [CFS^{+}21]. US$125.00 [Ano22j]. US$25.00 [Ano23n]. US$29.95 [Sár23]. US$30.00 [Ano23n]. US$45.00 [Ano22j]. US$49.95 [Ano23n]. US$50.00 [Ano22j]. US$95.00 [Ano23n]. Usage [McC24]. Use [EGCS25, FJZ^{+}21, FBS^{+}23, HZW^{+}21, Huy23, MCF^{+}21, MCH^{+}20, NHSM^{+}24, Pit21, WRW^{+}20, ZWC^{+}24c, ARSL24, CC24, EN24, GZBLS^{+}20, GW21, HRHB20, HBJ^{+}20, LCEM24, MHB^{+}22, MR22, NRSF21, PKP20, Ric21, SPK^{+}22, Sra23, SN22, SHS^{+}22, SSEN20, TWHB^{+}21, WTAD21]. used [CP20b]. users [WSK^{+}22]. uses [LWG21]. Using [APE^{+}21, ABM^{+}22, BTDS20, BPN^{+}24, CSS^{+}24, CFA^{+}24, DL24, DH21, FSC^{+}21, HPG20, KBB^{+}24, LNYY20, LGD^{+}25, LH21, LDC^{+}24, MR22, MCG^{+}21, MSN^{+}22, NAG^{+}20, PB21, PMR^{+}22, QOF^{+}21, RSRG^{+}20, STF^{+}20, VM20, WGAB^{+}21, WKG^{+}21, CO24, DFH^{+}21, FMGS21, STK^{+}21, TGJ^{+}20].

Validating [DSF^{+}23]. validation [IDD^{+}22]. valuation [PHK^{+}25, WGZ24]. Value [KBL^{+}23, IJ24, MGS^{+}23, MSN^{+}22, Sag20a, Sag20b, TBJ^{+}24, VD20, YVA22]. value-focused [MGS^{+}23]. valued [GFK^{+}22]. values [DBR^{+}23, HCC^{+}21, Hen22, IR22, MHD^{+}22, MBL^{+}23, Sta22a, WSvRLK20]. values-based [HCC^{+}21, MHD^{+}22]. Valuing [XAAY^{+}21, MBD^{+}21]. variability [BPH^{+}22, BPR24b, NRN^{+}24, RBJN21]. variable [HGM^{+}22]. variables [KLVW20]. variants [FW22]. variation [HNC^{+}24, LZG^{+}25, SUA22, SS24, SMSS^{+}21]. variety [Dav20]. Varying [Cli23, HWK^{+}20]. Vascular [KFN^{+}21, RWC^{+}22]. Vector [Ano22j]. Vector-Borne [Ano22j]. vegetation [MVC^{+}21, MSMH22, PHK^{+}25, SSCR^{+}23, TG21]. vegetation-based [MVC^{+}21]. venom [ABCH^{+}21]. Venus [MTH^{+}24]. Verrieres [Sár23]. version [SCdM^{+}24]. versus [GFF23, MSHE23, MCC^{+}23a]. vertebrate [COOH^{+}25, JZLP^{+}23, PMR^{+}22, STF^{+}20, SBW^{+}22]. vertebrates [BP21, CCW24, EN24, LDD^{+}22, MSB25, MHM^{+}22, NLV20]. very [ZLZ^{+}21]. viability [CO20, CO21, DTV^{+}25, IJK^{+}21, LFT^{+}21, NLD^{+}20]. viable [NW22]. victims [STW22]. video [MVC^{+}22]. Vietnam [WYB^{+}21]. view [WFBT^{+}25]. viewer [FCH^{+}24]. views [BGKKC22]. viii [Ano22i, Ano22j]. Villages [Gon20]. vineyards [MSKM21]. violence [KvZS^{+}25]. virtual [DMR^{+}21, KRB^{+}22]. virtue [BA22a]. vision [VHHT21]. visionary [Ano20v]. visions [Hal22]. visiting [MWS^{+}25]. visitor [MWS^{+}25]. visitors [CLS20]. visual [RLSH24]. vocalizations [LWG21, WKG^{+}21]. Voices [Ano24d, Dav20, MCW^{+}24]. voluntary [GCJB^{+}22, ML22, RW23, TWMM^{+}23]. Volunteers [Sco21]. Vulnerability [WWVC^{+}25, CTM^{+}22, DL24, LdSB^{+}20, NGST21]. vulnerable [MNMPGN23, SBG^{+}20]. vulture [SGB^{+}20]. vultures [EHBT20].

W [Ano22j, Ano23n]. WA [Ano22i]. Wall [NTS^{+}21]. Wallach [MBD^{+}21]. Wallingford [Ano22i]. warm [HTS^{+}22, MHJD23]. warming [Ano24f, GGJ^{+}20, GPJJ^{+}21, GJPJ^{+}22, HMS^{+}23, KFU^{+}24]. wars [May22]. wasp [TG21]. water [DSM24, HTS^{+}22, MWY^{+}21]. waterbird [BMM^{+}23, CLM22, RVLD20]. waterbirds [GGJ^{+}20, GPJJ^{+}21, GJPJ^{+}22, VGB^{+}24, WLR^{+}22]. watershed [MHD^{+}22]. Waterways [Sco20]. way [SW25]. Ways [Ano23n, BGMEHP20, MK21]. weather [BPR24b, GTAdPA24, NRN^{+}24]. welcome [Dav20]. well [MHGC21, TB25]. well-being [MHGC21, TB25]. western [HTC22, MFW^{+}24, NSB^{+}21, SCB^{+}25]. wet [JAB^{+}21a]. wetland [CXJ^{+}22]. wetlands [Pit21, VGB^{+}24]. whale [TW24]. whales [BTM^{+}23]. whaling [LMG^{+}23]. Wheels [Kut20]. where [Cli23]. white [CRC^{+}21, TSS^{+}22, WWVC^{+}25]. white-nose [CRC^{+}21, TSS^{+}22, WWVC^{+}25]. wide [MHP23, ZAS^{+}24]. Wiedenfeld [WSW^{+}22]. Wikipedia [MCG^{+}21]. Wild [Ano22j, Ano23n, MEM^{+}22, AR21, BNS24, CAFM^{+}23, CVT^{+}21, DFB^{+}24, EMV^{+}20, FLW^{+}22, HWG^{+}22, KFP^{+}21, KOO^{+}21, MHB^{+}22, MDG^{+}21, MMN^{+}20, MSHE23, RCL^{+}22, RRR^{+}24, SQJ^{+}22, TF24, WJ21]. wild-sourced [MSHE23]. Wilderness [PHMVL^{+}22]. wildfire [WGZ24].

Wildlife [Cor20, DBMS24, FWP^{+}21, SBG^{+}20, Ano24g, ARKA22, BRW21, BHRBK20, BM24, BMM^{+}24, BGD25, CBM20, CBD^{+}21, DHNM22, EKH23, EDA^{+}21, EOR^{+}24, FOYZ22, FWZ^{+}24, GCS^{+}24, GZBP24, HtSR25, HHM24, HLL^{+}24, HM24, HHB23, JRKF21, JSKR22, JS25, JSA^{+}20, JDB^{+}21, KKKS^{+}20, KIA^{+}22, LH21, LZG^{+}25, LTJ^{+}20, MTC^{+}20, MLS^{+}22, MJDMAS20, MHM^{+}24, MLIE25, MSS^{+}24, MMHR20, NAS21b, NAS21a, OGMMG22, OLF24, PWB^{+}21, PBV21, PLD^{+}21, QGC^{+}25, Sal24, SÁT21, SPA^{+}22, SP21, STK^{+}21, TTDdLM^{+}21, TWHB^{+}21, TSÁ20, TE20, VBK^{+}23, WSSC22, WLR^{+}24, WTAD21, ZWC^{+}24a, ZZ24]. will [WB23]. wind [CSGG25, DSZ21, LBBT24]. window [JBB23]. wine [MSKM21]. within [HC22, OGMMG22]. without [RCRJ^{+}24]. wolf [MBD^{+}23, SBG^{+}25]. wolves [CDSB23, STES23]. wood [BDF^{+}22, JBE^{+}22a]. wood-production [JBE^{+}22a]. working [PKC^{+}23a]. World [Ano22i, Gon21, GR21, BLF^{+}23, BVH22, BMA^{+}20, BFS24, BTM^{+}23, EOR^{+}24, FJZ^{+}20, KKKS^{+}20, LB22, LXZ^{+}22, MLS^{+}22, SWM^{+}24, TJE^{+}24, ZXXL23]. worldviews [BNS24, CC20]. Worldwide [GCS^{+}21, IJ24, NCE^{+}21]. would [JBE^{+}22b]. Wynne [Ano23n].

Xinlong [CTP^{+}24]. xv [Ano22i]. xvi [Ano22j, Ano23n]. xx [Ano22i]. xxii [Ano23n].

years [BJP^{+}20, GTC21, HRCA21, JZLP^{+}23, MEM^{+}22, MSPM20, MSRP21, PKP20, PLTC^{+}23, QZX^{+}24]. York [Ano23n]. Youth [SWMM24, RLSH24].

Zealand [FLH21, HCC^{+}21]. zone [MHJ^{+}24]. zones [HSN^{+}21]. zoning [SLM^{+}20]. zoo [SS23, WBC^{+}20]. zoonoses [GG24]. zoonotic [LKL^{+}21]. zoos [EAdC^{+}21, MWS^{+}25]. zootherapy [PS24].

References

Allman:2021:GIE

[AAS21] Phil Allman, Andrews Agyekumhene, and Leyna Stemle. Gillnet illumination as an effective measure to reduce sea turtle bycatch. Conservation Biology, 35(3):967–975, June 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andrews:2022:ESF

[AAvdW^{+}22] Caitlin E. Andrews, Sandra H. Anderson, Karin van der Walt, Rose Thorogood, and John G. Ewen. Evaluating the success of functional restoration after reintroduction of a lost avian pollinator. Conservation Biology, 36(4):e13892:1–e13892:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Acerbi:2023:SAT

[ABC^{+}23] Alberto Acerbi, John Burns, Unal Cabuk, Jakub Kryczka, Bethany Trapp, John Joseph Valletta, and Alex Mesoudi. Sentiment analysis of the Twitter response to Netflix’s Our Planet documentary. Conservation Biology, 37(4):e14060:1–e14060:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alvarez-Blanco:2021:EAA

[ABCH^{+}21] Paloma Alvarez-Blanco, Xim Cerdá, Abraham Hefetz, Raphaël Boulay, Alejandro Bertó-Moran, Carmen Díaz-Paniagua, Alain Lenoir, Johan Billen, H. Christoph Liedtke, Kamlesh R. Chauhan, Ganga Bhagavathy, and Elena Angulo. Effects of the Argentine ant venom on terrestrial amphibians. Conservation Biology, 35(1):216–226, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2022:UER

[ABM^{+}22] Doug P. Armstrong, Rebecca L. Boulton, Nikki McArthur, Susanne Govella, Nic Gorman, Rhonda Pike, and Yvan Richard. Using experimental reintroductions to resolve the roles of habitat quality and metapopulation dynamics on patch occupancy in fragmented landscapes. Conservation Biology, 36(3): e13843:1–e13843:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aceves-Bueno:2021:IFE

[ABRCM21] Eréndira Aceves-Bueno, Andrew J. Read, and Miguel A. Cisneros-Mata. Illegal fisheries, environmental crime, and the conservation of marine resources. Conservation Biology, 35(4): 1120–1129, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Albers:2024:AAT

[ACD^{+}24] Heidi J. Albers, Charlotte H. Chang, Sahan T. M. Dissanayake, Kate J. Helmstedt, Kailin Kroetz, Bistra Dilkina, Irene Zapata-Morán, Christoph Nolte, Leticia M. Ochoa-Ochoa, and Gwen Spencer. Anticipating anthropogenic threats in acquiring new protected areas. Conservation Biology, 38(2):e14176:1– e14176:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Abeli:2020:ESC

[ADG^{+}20] Thomas Abeli, Sarah Dalrymple, Sandrine Godefroid, Andrea Mondoni, Jonas V. Müller, Graziano Rossi, and Simone Orsenigo. Ex situ collections and their potential for the restoration of extinct plants. Conservation Biology, 34(2):303–313, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Andreozzi:2024:IMF

[ADKM24] Chelsea L. Andreozzi, Todd E. Dawson, Justin Kitzes, and Adina M. Merenlender. Influence of microclimate and forest management on bat species faced with global change. Conservation Biology, 38(4):e14246:1–e14246:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alatorre:2021:STO

[ADSG21] Andrea Alatorre, Johanna Depenthal, and Elizabeth Shapiro-Garza. Synergies and trade-offs among integrated conservation approaches in Mexico. Conservation Biology, 35(5):1451–1462, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aylesworth:2020:ROA

[AFV20] Lindsay Aylesworth, Sarah J. Foster, and Amanda C. J. Vincent. Realities of offering advice to governments on CITES. Conservation Biology, 34(3):644–653, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:2022:TGM

[AHM^{+}22] Louise Anderson, Peter Houk, Mark G. R. Miller, Javier Cuetos-Bueno, Curtis Graham, Kriskitina Kanemoto, Elizabeth Terk, Elizabeth McLeod, and Maria Beger. Trait groups as management entities in a complex, multispecies reef fishery. Conservation Biology, 36(3):e13866:1–e13866:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Allen:2021:MSS

[AL21] Michael C. Allen and Julie L. Lockwood. Mapping shifts in spatial synchrony in grassland birds to inform conservation planning. Conservation Biology, 35(3):1029–1038, June 2021.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alagador:2023:ECP

[Ala23] Diogo Alagador. Effective conservation planning of Iberian amphibians based on a regionalization of climate-driven range shifts. Conservation Biology, 37(2):e14026:1–e14026:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alagador:2025:DEM

[Ala25] Diogo Alagador. Dependence of Europe’s most threatened mammals on movement to adapt to climate change. Conservation Biology, 39(1):e14315:1–e14315:??, February 2025. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alaniz:2024:PTE

[AMC^{+}24] Alberto J. Alaniz, Pablo A. Marquet, Mario A. Carvajal, Pablo M. Vergara, Darío Moreira-Arce, Miguel A. Muzzio, and David A. Keith. Perspectives on the timing of ecosystem collapse in a changing climate. Conservation Biology, 38 (4):e14247:1–e14247:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:2023:IHM

[And23] Robert P. Anderson. Integrating habitat-masked range maps with quantifications of prevalence to estimate area of occupancy in IUCN assessments. Conservation Biology, 37(1): e14019:1–e14019:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anderson:2025:CPC

[And25] Christopher B. Anderson. COP16 and the process of consolidating an inclusive conservation paradigm. Conservation Biology, 39(1):e14438:1–e14438:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:BRJ

[Ano20a] Anonymous. Book reviewers, January–December 2019. Conservation Biology, 34(5):1319, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:BRM

[Ano20b] Anonymous. Books received (March 2019–May 2020). Conservation Biology, 34(5):1320–1322, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:CBA

[Ano20c] Anonymous. Conservation Biology awards. Conservation Biology, 34(3):537, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:CCCa

[Ano20d] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 34(1):i, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:CCCb

[Ano20e] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 34(2):i, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:CCCc

[Ano20f] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 34(3):i, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:CCCd

[Ano20g] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 34(4):i, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:CCCe

[Ano20h] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 34(5):i, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:CCCf

[Ano20i] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 34(6):i, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:IITa

[Ano20j] Anonymous. Issue information — TOC. Conservation Biology, 34(1):1–3, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:IITb

[Ano20k] Anonymous. Issue information — TOC. Conservation Biology, 34(2):299–301, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:IITc

[Ano20l] Anonymous. Issue information — TOC. Conservation Biology, 34(3):534–535, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:IITd

[Ano20m] Anonymous. Issue information — TOC. Conservation Biology, 34(4):777–779, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:IITe

[Ano20n] Anonymous. Issue information — TOC. Conservation Biology, 34(5):1057–1060, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:IITf

[Ano20o] Anonymous. Issue information — TOC. Conservation Biology, 34(6):1323–1326, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:NIa

[Ano20p] Anonymous. Noted with interest. Conservation Biology, 34(1): 295–297, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:NIb

[Ano20q] Anonymous. Noted with interest. Conservation Biology, 34 (2):531–532, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:NIc

[Ano20r] Anonymous. Noted with interest. Conservation Biology, 34 (3):775–776, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:NId

[Ano20s] Anonymous. Noted with interest. Conservation Biology, 34(4): 1055–1056, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:NIe

[Ano20t] Anonymous. Noted with interest. Conservation Biology, 34(5): 1318–1319, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:NIf

[Ano20u] Anonymous. Noted with interest. Conservation Biology, 34(6): 1602–1603, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2020:RMS

[Ano20v] Anonymous. Reflections on Michael Soulé, a visionary for conservation biology. Conservation Biology, 34(6):1592–1597, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:CBA

[Ano21a] Anonymous. Conservation Biology awards. Conservation Biology, 35(3):763, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:CCCa

[Ano21b] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 35(1):i, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:CCCb

[Ano21c] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 35(2):i, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:CCCc

[Ano21d] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 35(3):i, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:CCCd

[Ano21e] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 35(4):i, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:CCCe

[Ano21f] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 35(5):i, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:CCCf

[Ano21g] Anonymous. Cover caption: Conservation Biology. Conservation Biology, 35(6):i, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:IITa

[Ano21h] Anonymous. Issue information — TOC. Conservation Biology, 35(1):1–4, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:IITb

[Ano21i] Anonymous. Issue information — TOC. Conservation Biology, 35(2):391–394, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:IITc

[Ano21j] Anonymous. Issue information — TOC. Conservation Biology, 35(3):759–762, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:IITd

[Ano21k] Anonymous. Issue information — TOC. Conservation Biology, 35(4):1057–1060, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:IITe

[Ano21l] Anonymous. Issue information — TOC. Conservation Biology, 35(5):1361–1366, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:IITf

[Ano21m] Anonymous. Issue information — TOC. Conservation Biology, 35(6):1711–1714, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:NIa

[Ano21n] Anonymous. Noted with interest. Conservation Biology, 35(1): 388–390, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:NIb

[Ano21o] Anonymous. Noted with interest. Conservation Biology, 35 (2):754–757, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:NIc

[Ano21p] Anonymous. Noted with interest. Conservation Biology, 35 (3):1055–1056, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:NId

[Ano21q] Anonymous. Noted with interest. Conservation Biology, 35(4): 1357–1359, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:NIe

[Ano21r] Anonymous. Noted with interest. Conservation Biology, 35(5): 1707–1709, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2021:NIf

[Ano21s] Anonymous. Noted with interest. Conservation Biology, 35(6): 1974–1975, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:BR

[Ano22a] Anonymous. Book review. Conservation Biology, 36(1): e13882:1–e13882:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:BRJ

[Ano22b] Anonymous. Books received, June 2020–November 2021. Conservation Biology, 36(1):e13880:1–e13880:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:CBA

[Ano22c] Anonymous. Conservation Biology awards. Conservation Biology, 36(3):e13926:1–e13926:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:E

[Ano22d] Anonymous. Erratum. Conservation Biology, 36(4):e13975:1– e13975:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:IIC

[Ano22e] Anonymous. Issue information — copyright page. Conservation Biology, 36(6):e13775:1–e13775:??, December 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:NIc

[Ano22f] Anonymous. Noted with interest. Conservation Biology, 36(4): e13969:1–e13969:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:NId

[Ano22g] Anonymous. Noted with interest. Conservation Biology, 36(5): e13990:1–e13990:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:NIe

[Ano22h] Anonymous. Noted with interest. Conservation Biology, 36(6): e14027:1–e14027:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:NII

[Ano22i] Anonymous. Noted with interest: Invasive Birds. Global Trends and Impacts. Downs, C. T., and L. A. Hart, editors. 2020. CABI Publishing, Wallingford, U.K. xv + 381 pp. £135.00 (hardcover). ISBN 978-1-78924-206-5. Philosophy of Ecology. An Introduction. Justus, J. 2021. Cambridge University Press, Cambridge, U.K. viii + 215 pp. £22.99 (paperback). ISBN 978-1-107-69815-4. Biodiversity. A Beginner’s Guide. 2nd edition. Spicer, J. 2021. One World Publications, London, U.K. xx + 246 pp. £9.99 (paperback). ISBN 978-0-86154-017-4. Flames of Extinction. Pickrell, J. 2021. New-South, Island Press, Covelo, WA, U.S.A. 280 pp. AU$29.99 (hardcover). ISBN 978-1-64283-202-0. Conservation Biology, 36(1):e13881:1–e13881:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2022:NIb

[Ano22j] Anonymous. Noted with interest: Renewing Our Rivers: Stream Corridor Restoration in Dryland Regions. Briggs, M. K., and W. R. Osterkamp. 2020. University of Arizona Press, Tucson, AZ, U.S.A. 468 pp. US$50.00 (paperback). ISBN 978-0-8165-4148-5. Predator Ecology: Evolutionary Ecology of the Functional Response. DeLong, J.P. 2021. Oxford University Press, Oxford, U.K. viii + 168 pp. £34.99 (paperback). ISBN 978-0-19-289551-6. Ecosystem-Based Fisheries Management — Progress, Importance, and Impacts in the United States. Link, J. S., and A. R. Marshak. 2021. Oxford University Press, Oxford, UK. 712 pp. US$125.00 (hardcover). ISBN 978-0-19-284346-3. Eco-Evolutionary Dynamics. Hendry, A. P. 2020. Princeton University Press, Princeton, NJ, U.S.A. 416 pp. US$45.00 (paperback). ISBN 978-0-691-20417-8. Population Biology of Vector-Borne Diseases. Drake, J. M., M. Bonsall, and M. R. Strand, editors. 2021. Oxford University Press, Oxford, U.K. viii + 293 pp. £74.99 (hardcover). ISBN 978-0-19-885324-4. Infectious Disease Ecology of Wild Birds. Owen, J. C., D. M. Hawley, and K. P. Huyvaert, editors. 2021. Oxford University Press, Oxford, U. K. xvi + 266 pp. £37.99 (paperback). ISBN 978-0-19-874625-6. Conservation Biology, 36 (4):e13921:1–e13921:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:BRA

[Ano23a] Anonymous. Book reviewers (August 2021–December 2022).

Conservation Biology, 37(2):e14077:1–e14077:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:BRN

[Ano23b] Anonymous. Books received (November 2021–December 2022). Conservation Biology, 37(2):e14078:1–e14078:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:ESR

[Ano23c] Anonymous. Erratum to “A systematic review of snake translocations to identify potential tactics for reducing postrelease effects”. Conservation Biology, 37(3):e14129:1–e14129:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [CLGP23].

Anonymous:2023:ECF

[Ano23d] Anonymous. Erratum to “Conceptual frameworks and key questions for assessing the contribution of marine protected areas to shark and ray conservation”. Conservation Biology, 37(6):e14017:1–e14017:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [CMC^{+}23].

Anonymous:2023:IIa

[Ano23e] Anonymous. Issue information. Conservation Biology, 37(3): e13951:1–e13951:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:IIb

[Ano23f] Anonymous. Issue information. Conservation Biology, 37(4): e13953:1–e13953:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:IIc

[Ano23g] Anonymous. Issue information. Conservation Biology, 37(5): e13955:1–e13955:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:IId

[Ano23h] Anonymous. Issue information. Conservation Biology, 37(6): e13957:1–e13957:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:IICa

[Ano23i] Anonymous. Issue information — copyright page. Conservation Biology, 37(1):e13947:1–e13947:??, February 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:IICb

[Ano23j] Anonymous. Issue information — copyright page. Conservation Biology, 37(2):e13949:1–e13949:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:NIa

[Ano23k] Anonymous. Noted with interest. Conservation Biology, 37(1): e14053:1–e14053:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:NIb

[Ano23l] Anonymous. Noted with interest. Conservation Biology, 37 (2):e14079:1–e14079:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:NIc

[Ano23m] Anonymous. Noted with interest. Conservation Biology, 37 (3):e14107:1–e14107:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2023:NIP

[Ano23n] Anonymous. Noted with interest: Power in the Wild The Subtle and Not-So-Subtle Ways Animals Strive for Control over Others. Dugatkin, L. A. 2022. University of Chicago Press, Chicago, IL, U.S.A. xvi+187 pp. US$25.00 (hardcover). ISBN 978-0-226-81594-7. Cave Biodiversity Speciation and Diversity of Subterranean Fauna. Wynne, J. J., editor. 2022. Johns Hopkins University Press, Baltimore, MD, U.S.A. 311 pp. US$95.00 (hardcover). ISBN 978-1-4214-4457-4. Eight Bears. Dickie, G. 2023. W. W. Norton & Company, New York, NY, U.S.A. 272 pp. US$30.00 (hardcover). ISBN 978-1-324-00508-7. Species Tree Inference: a Guide to Methods and Application. Kubatko, L.S., and L. L. Knowles, editors. 2023. Princeton University Press, Princeton, NJ, U.S.A. xxii+324 pp. US$49.95 (paperback). ISBN 978-0-691-20760-5. Conservation Biology, 37(5): e14171:1–e14171:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:BRA

[Ano24a] Anonymous. Book reviewers (August 2021 to December 2022). Conservation Biology, 38(6):e14365:1–e14365:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:BRJ

[Ano24b] Anonymous. Books received (January 2023 to May 2024). Conservation Biology, 38(6):e14366:1–e14366:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:CBA

[Ano24c] Anonymous. Conservation Biology awards. Conservation Biology, 38(3):e14289:1–e14289:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:DVC

[Ano24d] Anonymous. Diversifying voices in conservation. Conservation Biology, 38(6):e14418:1–e14418:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:EAF

[Ano24e] Anonymous. Erratum to An assessment of future rewilding potential in the United Kingdom. Conservation Biology, 38(6): e14383:1–e14383:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [BPR24a].

Anonymous:2024:EEH

[Ano24f] Anonymous. Erratum to Effects of human depopulation and warming climate on bird populations in Japan. Conservation Biology, 38(2):e14252:1–e14252:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [KFU^{+}24].

Anonymous:2024:HIB

[Ano24g] Anonymous. Historical insight into British and Irish wildlife. Conservation Biology, 38(2):e14238:1–e14238:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:IIa

[Ano24h] Anonymous. Issue information. Conservation Biology, 38(1): e14116:1–e14116:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:IIb

[Ano24i] Anonymous. Issue information. Conservation Biology, 38(2): e14118:1–e14118:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:IIc

[Ano24j] Anonymous. Issue information. Conservation Biology, 38(3): e14120:1–e14120:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:IId

[Ano24k] Anonymous. Issue information. Conservation Biology, 38(4): e14122:1–e14122:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:IIe

[Ano24l] Anonymous. Issue information. Conservation Biology, 38(5): e14124:1–e14124:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:IIf

[Ano24m] Anonymous. Issue information. Conservation Biology, 38(6): e14126:1–e14126:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:NIa

[Ano24n] Anonymous. Noted with interest. Conservation Biology, 38(1): e14211:1–e14211:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:NIb

[Ano24o] Anonymous. Noted with interest. Conservation Biology, 38 (2):e14236:1–e14236:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:NIc

[Ano24p] Anonymous. Noted with interest. Conservation Biology, 38 (3):e14294:1–e14294:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:NId

[Ano24q] Anonymous. Noted with interest. Conservation Biology, 38(6): e14319:1–e14319:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2024:NIe

[Ano24r] Anonymous. Noted with interest. Conservation Biology, 38(6): e14330:1–e14330:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2025:IIa

[Ano25a] Anonymous. Issue information. Conservation Biology, 39(1): e14303:1–e14303:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2025:IIb

[Ano25b] Anonymous. Issue information. Conservation Biology, 39(2): e14304:1–e14304:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2025:NIa

[Ano25c] Anonymous. Noted with interest. Conservation Biology, 39(1): e14367:1–e14367:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Anonymous:2025:NIb

[Ano25d] Anonymous. Noted with interest. Conservation Biology, 39 (2):e14443:1–e14443:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ausprey:2023:STM

[ANR23] Ian J. Ausprey, Felicity L. Newell, and Scott K. Robinson. Sensitivity of tropical montane birds to anthropogenic disturbance and management strategies for their conservation in agricultural landscapes. Conservation Biology, 37(6):e14136:1– e14136:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Araneda:2022:SDE

[AOI22] Paola Araneda, Omar Ohrens, and José Tomás Ibarra. Socioeconomic development and ecological traits as predictors of human–bird conflicts. Conservation Biology, 36(1):e13859:1– e13859:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Aucott:2021:MBP

[AP21] Michael Aucott and Rex A. Parker. Medical biotechnology as a paradigm for forest restoration and introduction of the transgenic American chestnut. Conservation Biology, 35(1): 190–196, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2023:ICO

[AP23] Doug P. Armstrong and Kevin A. Parker. An island conservation odyssey from a pioneer in the field. Conservation Biology, 37(6):e14179:1–e14179:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2021:ULT

[APE^{+}21] Doug P. Armstrong, Elizabeth H. Parlato, Barbara Egli, Wendy J. Dimond, Renske Kwikkel, Åsa Berggren, Mhairi Mc-Cready, Kevin A. Parker, and John G. Ewen. Using longterm data for a reintroduced population to empirically estimate future consequences of inbreeding. Conservation Biology, 35(3):859–869, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Adams:2021:EW

[AR21] William M. Adams and Kent H. Redford. Editing the wild. Conservation Biology, 35(5):1701–1703, October 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Alvarez-Romero:2024:HLC

[ÁRAB^{+}24] Jorge G. Álvarez-Romero, Vanessa M. Adams, Natalie C. Ban, Morena Mills, and Piero Visconti. Honoring the legacy of a conservation champion: Bob Pressey (1953–2023). Conservation Biology, 38(1):e14168:1–e14168:??, February 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Akcakaya:2020:AEF

[ARK^{+}20] H. Resit Akçakaya, Ana S. L. Rodrigues, David A. Keith, E. J. Milner-Gulland, Eric W. Sanderson, Simon Hedges, David P. Mallon, Molly K. Grace, Barney Long, Erik Meijaard, and P. J. Stephenson. Assessing ecological function in the context of species recovery. Conservation Biology, 34(3):561–571, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ashish:2022:GTA

[ARKA22] Kumar Ashish, Tharmalingam Ramesh, Riddhika Kalle, and Rathinasamy Arumugam. Generalization of threats attributed to large carnivores in areas of high human–wildlife conflict. Conservation Biology, 36(5):e13974:1–e13974:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Armstrong:2023:GJO

[Arm23] Chris Armstrong. Global justice and the opportunity costs of conservation. Conservation Biology, 37(2):e14018:1–e14018:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Auliya:2024:CRC

[ARSL24] Mark Auliya, Arne Redsted Rasmussen, Kate L. Sanders, and Aaron Savio Lobo. Challenges of regulating commercial use of marine elapid snakes in the Indo-Pacific. Conservation Biology, 38(5):e14336:1–e14336:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ayambire:2025:CAE

[ART^{+}25] Raphael A. Ayambire, Trina Rytwinski, Jessica J. Taylor, Matthew W. Luizza, Matthew J. Muir, Cynthia Cadet, Derek Armitage, Nathan J. Bennett, Jeremy Brooks, Samantha H. Cheng, Jenny Martinez, Meenakshi Nagendran, Siri Öckerman, Shannon N. Rivera, Anne Savage, David S. Wilkie, Steven J. Cooke, and Joseph R. Bennett. Challenges in assessing the effects of environmental governance systems on conservation outcomes. Conservation Biology, 39(1):e14392:1–e14392:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bobier:2022:VCC

[BA22a] Christopher Bobier and Benjamin Allen. The virtue of compassion in compassionate conservation. Conservation Biology, 36(1):e13776:1–e13776:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brian:2022:MPR

[BA22b] Joshua I. Brian and David C. Aldridge. Mussel parasite richness and risk of extinction. Conservation Biology, 36(6): e13979:1–e13979:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bunce:2025:SNA

[BAA^{+}25] Brittany Bunce, Elia Apostolopoulou, Sara Maestre Andres, Alejandra Pizarro Choy, Marina Requena i Mora, and Dan Brockington. A social network analysis of an epistemic community studying neoliberal conservation. Conservation Biology, 39(2):e70001:1–e70001:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blair:2023:MEJ

[BAB^{+}23a] Mary E. Blair, Felicity Arengo, Nora Bynum, Georgina Cullman, Leo R. Douglas, Martha J. Groom, Christian J. Rivera, and Ana L. Porzecanski. In memoriam: Eleanor Jane Sterling (1960–2023). Conservation Biology, 37(3):e14099:1–e14099:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bodey:2023:ECP

[BAB^{+}23b] Thomas W. Bodey, Elena Angulo, Alok Bang, Céline Bellard, Jean Fantle-Lepczyk, Bernd Lenzner, Anna Turbelin, Yuya Watari, and Franck Courchamp. Economic costs of protecting islands from invasive alien species. Conservation Biology, 37(2):e14034:1–e14034:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brewer:2022:LSS

[BAGK22] Tom David Brewer, Neil Andrew, Bernd Gruber, and Johnathan Kool. Large-sample-size assessment of socioeconomic predictors of community-level resource management occurrence. Conservation Biology, 36(2):e13800:1–e13800:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blake:2024:ISO

[BAGV24] Katie Blake, Sean C. Anderson, Adam Gleave, and Diogo Veríssimo. Impact on species’ online attention when named after celebrities. Conservation Biology, 38(2):e14184:1– e14184:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Backus:2021:IRS

[BB21] Gregory A. Backus and Marissa L. Baskett. Identifying robust strategies for assisted migration in a competitive stochastic metacommunity. Conservation Biology, 35(6):1809–1820, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brodie:2025:DIC

[BBE^{+}25] Jedediah F. Brodie, Carolina Bello, Carine Emer, Mauro Galetti, Matthew S. Luskin, Anand Osuri, Carlos A. Peres, Annina Stoll, Nacho Villar, and Ana-Benítez López. Defaunation impacts on the carbon balance of tropical forests. Conservation Biology, 39(1):e14414:1–e14414:??, February 2025. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boersma:2020:ASP

[BBG^{+}20] P. D. Boersma, P. García Borboroglu, N. J. Gownaris, C. A. Bost, A. Chiaradia, S. Ellis, T. Schneider, P. J. Seddon, A. Simeone, P. N. Trathan, L. J. Waller, and B. Wienecke. Applying science to pressing conservation needs for penguins. Conservation Biology, 34(1):103–112, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ball:2021:PBE

[BBGL21] James G. C. Ball, Mark A. Burgman, Elizabeth D. Goldman, and Janeth Lessmann. Protecting biodiversity and economic returns in resource-rich tropical forests. Conservation Biology, 35(1):263–273, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Becker:2025:SOC

[BBH^{+}25] Sarah L. Becker, Charlotte Boyd, Jonathan M. Handley, Ben Raymond, Ryan Reisinger, Yan Ropert-Coudert, Nora Apelgren, Tammy E. Davies, Mary-Anne Lea, Mercedes Santos, Philip N. Trathan, Anton P. Van de Putte, Luis A. Huckstadt, Jean-Benoit Charrassin, and Cassandra M. Brooks. Scaling up ocean conservation through recognition of key biodiversity areas in the Southern Ocean from multispecies tracking data. Conservation Biology, 39(1):e14345:1–e14345:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buxton:2020:TLC

[BBL^{+}20] Rachel T. Buxton, Jordanna N. Bergman, Hsien-Yung Lin, Allison D. Binley, Stephanie Avery-Gomm, Richard Schuster, Dominique G. Roche, and Joseph R. Bennett. Three lessons conservation science can learn from the COVID-19 pandemic. Conservation Biology, 34(6):1331–1332, December 2020. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boyce:2022:NFR

[BBL22] Paul Boyce, Jonaki Bhattacharyya, and Wayne Linklater. The need for formal reflexivity in conservation science. Conservation Biology, 36(2):e13840:1–e13840:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bellis:2020:CSP

[BBM^{+}20] Joe Bellis, David Bourke, Joyce Maschinski, Katie Heineman, and Sarah Dalrymple. Climate suitability as a predictor of conservation translocation failure. Conservation Biology, 34 (6):1473–1481, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Butt:2021:IST

[BCA^{+}21] Nathalie Butt, Alienor L. M. Chauvenet, Vanessa M. Adams, Maria Beger, Rachael V. Gallagher, Danielle F. Shanahan, Michelle Ward, James E. M. Watson, and Hugh P. Possingham. Importance of species translocations under rapid climate change. Conservation Biology, 35(3):775–783, June 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Balzani:2022:CCA

[BDF^{+}22] Paride Balzani, Wouter Dekoninck, Heike Feldhaar, Anne Freitag, Filippo Frizzi, Jan Frouz, Alberto Masoni, Elva Robinson, Jouni Sorvari, and Giacomo Santini. Challenges and a call to action for protecting European red wood ants. Conservation Biology, 36(6):e13959:1–e13959:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ban:2023:EMO

[BDG^{+}23] Natalie C. Ban, Emily S. Darling, Georgina G. Gurney, Whitney Friedman, Stacy D. Jupiter, W. Peni Lestari, Irfan Yulianto, Sinta Pardede, Sukma A. R. Tarigan, Puji Prihatiningsih, Sangeeta Mangubhai, Waisea Naisilisili, Sirilo Dulunaqio, Josheena Naggea, Ravaka Ranaivoson, Vera N. Agostini, Gabby Ahmadia, Jessica Blythe, Stuart J. Campbell, Joachim Claudet, Courtney Cox, Graham Epstein, Estradivari, Margaret Fox, David Gill, Amber Himes-Cornell, Harry Jonas, Elizabeth Mcleod, Nyawira A. Muthiga, and Tim McClanahan. Effects of management objectives and rules on marine conservation outcomes. Conservation Biology, 37(6):e14156:1– e14156:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2025:IAI

[BDQ^{+}25] Timothy M. Brown, Alison M. Dunn, Rupert J. Quinnell, Ellen Clarke, Andrew A. Cunningham, and Simon J. Goodman. An interdisciplinary approach to improving conservation outcomes for parasites. Conservation Biology, 39(1):e14431:1–e14431:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bisbal:2023:CSN

[BE23] Gustavo A. Bisbal and Mitchell J. Eaton. Considering science needs to deliver actionable science. Conservation Biology, 37(1):e14013:1–e14013:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brodie:2021:UDB

[BF21] Jedediah F. Brodie and Jose M. V. Fragoso. Understanding the distribution of bushmeat hunting effort across landscapes by testing hypotheses about human foraging. Conservation Biology, 35(3):1009–1018, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bird:2024:PRE

[BFS24] Jeremy P. Bird, Richard A. Fuller, and Justine D. Shaw. Patterns of recovery in extant and extirpated seabirds after the world’s largest multipredator eradication. Conservation Biology, 38(4):e14239:1–e14239:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blankinship:2025:HIB

[BGAE25] Lisa A. Blankinship, Sarah Gillaspie, and Basil H. Aboul-Enein. Highlighting the importance of biodiversity conservation through the Holy Qur’an. Conservation Biology, 39(1): e14309:1–e14309:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Betsill:2024:CEP

[BGB^{+}24] Michele M. Betsill, Rebecca L. Gruby, Jeffrey Blackwatters, Ash Enrici, Elodie Le Cornu, Xavier Basurto, Chad English, Charlotte Hudson, Leah Meth, Imani Fairweather-Morrison, Dana Okano, and David Secord. Community engagement and power dynamics in conservation philanthropy grant making. Conservation Biology, 38(6):e14272:1–e14272:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bletz:2025:QSB

[BGD25] Molly C. Bletz, Evan H. Campbell Grant, and Graziella Di-Renzo. Quantitative support for the benefits of proactive management for wildlife disease control. Conservation Biology, 39(1):e14363:1–e14363:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brias-Guinart:2022:TCP

[BGKKC22] Aina Brias-Guinart, Kaisa Korhonen-Kurki, and Mar Cabeza. Typifying conservation practitioners’ views on the role of education. Conservation Biology, 36(4):e13893:1–e13893:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bertuol-Garcia:2020:SWT

[BGMEHP20] Diana Bertuol-Garcia, Carla Morsello, Charbel N. El-Hani, and Renata Pardini. Shared ways of thinking in Brazil about the science–practice interface in ecology and conservation. Conservation Biology, 34(2):449–461, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bernardo:2020:NEA

[BGVK20] Holly L. Bernardo, Rachel Goad, Pati Vitt, and Tiffany M. Knight. Nonadditive effects among threats on rare plant species. Conservation Biology, 34(4):1029–1034, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ben-Haim:2023:MUD

[BH23] Yakov Ben-Haim. Managing uncertainty in decision-making for conservation science. Conservation Biology, 37(6):e14164:1– e14164:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Biasetti:2023:ADT

[BHG^{+}23] Pierfrancesco Biasetti, Thomas B. Hildebrandt, Frank Göritz, Robert Hermes, Susanne Holtze, Jan Stejskal, Cesare Galli, Ilaria Pollastri, Alessia Muzzo, Isaac Lekolool, David Ndereeh, Patrick Omondi, Linus Kariuki, Domnic Mijele, Samuel Mutisya, Stephen Ngulu, and Barbara de Mori. Application of decision tools to ethical analysis in biodiversity conservation. Conservation Biology, 37(2):e14029:1–e14029:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2020:ECD

[BHM20] Emma M. Bennett, Cindy E. Hauser, and Joslin L. Moore. Evaluating conservation dogs in the search for rare species. Conservation Biology, 34(2):314–325, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buchs:2021:APP

[BHM21] Arnaud Buchs, Emeline Hassenforder, and Yves Meinard. Adapting participatory processes to fine-tune conservation approaches in multiactor decision settings. Conservation Biology, 35(3):804–815, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baynham-Herd:2020:PIP

[BHRBK20] Zachary Baynham-Herd, Steve Redpath, Nils Bunnefeld, and Aidan Keane. Predicting intervention priorities for wildlife conflicts. Conservation Biology, 34(1):232–243, February 2020.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Broekman:2022:TLE

[BHS^{+}22] Maarten J. E. Broekman, Jelle P. Hilbers, Aafke M. Schipper, Ana Benítez-López, Luca Santini, and Mark A. J. Huijbregts. Time-lagged effects of habitat fragmentation on terrestrial mammals in Madagascar. Conservation Biology, 36(5): e13942:1–e13942:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bruno:2022:ERN

[BHSM^{+}22] Daniel Bruno, Virgilio Hermoso, María Mar Sánchez-Montoya, Oscar Belmar, Cayetano Gutiérrez-Cánovas, and Miguel Cañedo-Argüelles. Ecological relevance of non-perennial rivers for the conservation of terrestrial and aquatic communities. Conservation Biology, 36(6):e13982:1–e13982:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Broekman:2024:IEP

[BHT^{+}24] Maarten J. E. Broekman, Jelle P. Hilbers, Marlee A. Tucker, Mark A. J. Huijbregts, and Aafke M. Schipper. Impacts of existing and planned roads on terrestrial mammal habitat in New Guinea. Conservation Biology, 38(1):e14152:1–e14152:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brittain:2020:ECW

[BIdL^{+}20] Stephanie Brittain, Harriet Ibbett, Emiel de Lange, Leejiah Dorward, Simon Hoyte, Agnese Marino, E. J. Milner-Gulland, Julia Newth, Sarobidy Rakotonarivo, Diogo Veríssimo, and Jerome Lewis. Ethical considerations when conservation research involves people. Conservation Biology, 34(4):925–933, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baranov:2020:CNC

[BJP^{+}20] Viktor Baranov, Jonas Jourdan, Francesca Pilotto, Rüdiger Wagner, and Peter Haase. Complex and nonlinear climatedriven changes in freshwater insect communities over 42 years. Conservation Biology, 34(5):1241–1251, October 2020. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bhola:2021:PAB

[BKK^{+}21] Nina Bhola, Helen Klimmek, Naomi Kingston, Neil D. Burgess, Arnout van Soesbergen, Colleen Corrigan, Jerry Harrison, and Marcel T. J. Kok. Perspectives on area-based conservation and its meaning for future biodiversity policy. Conservation Biology, 35(1):168–178, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blair:2024:CM

[Bla24] Mary E. Blair. Conservation museomics. Conservation Biology, 38(3):e14234:1–e14234:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Banerjee:2023:IBI

[BLF^{+}23] Achyut Kumar Banerjee, Tien Ming Lee, Hui Feng, Xinru Liang, Yuting Lin, Jiakai Wang, Minghui Yin, Hao Peng, and Yelin Huang. Implications for biological invasion of non-native plants for sale in the world’s largest online market. Conservation Biology, 37(4):e14055:1–e14055:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Burrows:2023:ECI

[BLW23] Joanna L. Burrows, Jasmine R. Lee, and Kerrie A. Wilson. Evaluating the conservation impact of Antarctica’s protected areas. Conservation Biology, 37(3):e14059:1–e14059:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bilby:2024:RHW

[BM24] Jack Bilby and Katherine Moseby. Review of hyperdispersal in wildlife translocations. Conservation Biology, 38(1):e14083:1– e14083:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bird:2020:GLW

[BMA^{+}20] Jeremy P. Bird, Robert Martin, H. Reşit Akçakaya, James Gilroy, Ian J. Burfield, Stephen T. Garnett, Andy Symes, Joseph Taylor, Çağan H. Şekercioğlu, and Stuart H. M. Butchart. Generation lengths of the world’s birds and their implications for extinction risk. Conservation Biology, 34(5): 1252–1261, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bosch:2022:EHF

[BMG^{+}22] Nestor E. Bosch, Jacquomo Monk, Jordan Goetze, Shaun Wilson, Russell C. Babcock, Neville Barrett, Jock Clough, Leanne M. Currey-Randall, David V. Fairclough, Rebecca Fisher, Brooke A. Gibbons, David Harasti, Euan S. Harvey, Michelle R. Heupel, Jamie L. Hicks, Thomas H. Holmes, Charlie Huveneers, Daniel Ierodiaconou, Alan Jordan, Nathan A. Knott, Hamish A. Malcolm, Dianne McLean, Mark Meekan, Stephen J. Newman, Ben Radford, Matthew J. Rees, Benjamin J. Saunders, Conrad W. Speed, Michael J. Travers, Corey B. Wakefield, Thomas Wernberg, and Tim J. Langlois. Effects of human footprint and biophysical factors on the bodysize structure of fished marine species. Conservation Biology, 36(2):e13807:1–e13807:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bell:2024:GSI

[BMH^{+}24] James J. Bell, Valerio Micaroni, Benjamin Harris, Francesca Strano, Manon Broadribb, and Alice Rogers. Global status, impacts, and management of rocky temperate mesophotic ecosystems. Conservation Biology, 38(1):e13945:1–e13945:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Byrd:2023:KCI

[BMM^{+}23] Kristin B. Byrd, Elliott Matchett, Claudia Mengelt, Tamara S. Wilson, Deanne DiPietro, Monica Moritsch, Erin Conlisk, Sam Veloz, Michael L. Casazza, and Matthew E. Reiter. Knowledge coproduction on the impact of decisions for waterbird habitat in a changing climate. Conservation Biology, 37(5):e14089:1– e14089:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Birdsong:2024:ISI

[BMM^{+}24] Max H. Birdsong, Alexander L. Metcalf, Elizabeth Covelli Metcalf, Holly Kathleen Nesbitt, and Justin A. Gude. The influence of social identity on attitudes toward wildlife. Conservation Biology, 38(6):e14243:1–e14243:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [Wal24].

Batavia:2021:ESM

[BNB^{+}21] Chelsea Batavia, Michael Paul Nelson, Jeremy T. Bruskotter, Megan S. Jones, Esty Yanco, Daniel Ramp, Marc Bekoff, and Arian D. Wallach. Emotion as a source of moral understanding in conservation. Conservation Biology, 35(5):1380–1387, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barnes:2025:LSP

[BNI25] Kevin W. Barnes, Neal D. Niemuth, and Rich Iovanna. Landscape-scale predictions of future grassland conversion to cropland or development. Conservation Biology, 39(1): e14346:1–e14346:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Banerjee:2024:ATT

[BNS24] Sayan Banerjee, Dibakar Nayak, and Anindya Sinha. Adivasi (Tea Tribe) worldviews of living close to wild Asian elephants in Assam, India. Conservation Biology, 38(6):e14397:1–e14397:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Batavia:2020:MRC

[BNW20] Chelsea Batavia, Michael Paul Nelson, and Arian D. Wallach. The moral residue of conservation. Conservation Biology, 34 (5):1114–1121, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barrios-ONeill:2021:FSC

[BO21] Daniel Barrios-O’Neill. Focus and social contagion of environmental organization advocacy on Twitter. Conservation Biology, 35(1):307–315, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baquero:2023:CMT

[BOAN23] Rocío A. Baquero, Francisco J. Oficialdegui, Daniel Ayllón, and Graciela G. Nicola. The challenge of managing threatened invasive species at a continental scale. Conservation Biology, 37(5):e14165:1–e14165:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bartish:2020:ATE

[BOB^{+}20] Igor V. Bartish, Wim A. Ozinga, Mark I. Bartish, G. W. Wieger Wamelink, Stephan M. Hennekens, Benjamin Yguel, and Andreas Prinzing. Anthropogenic threats to evolutionary heritage of angiosperms in the Netherlands through an increase in highcompetition environments. Conservation Biology, 34(6):1536– 1548, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bellis:2024:IPT

[BOPM^{+}24] Joe Bellis, Oyomoare Osazuwa-Peters, Joyce Maschinski, Matthew J. Keir, Elliott W. Parsons, Thomas N. Kaye, Michael Kunz, Jennifer Possley, Eric Menges, Stacy A. Smith, Daniela Roth, Debbie Brewer, William Brumback, James J. Lange, Christal Niederer, Jessica B. Turner-Skoff, Megan Bontrager, Richard Braham, Michelle Coppoletta, Karen D. Holl, Paula Williamson, Timothy Bell, Jayne L. Jonas, Kathryn McEachern, Kathy L. Robertson, Sandra J. Birnbaum, Adam Dattilo, John J. Dollard, Jr., Jeremie Fant, Wendy Kishida, Peter Lesica, Steven O. Link, Noel B. Pavlovic, Jackie Poole, Charlotte M. Reemts, Peter Stiling, David D. Taylor, Jonathan H. Titus, Priscilla J. Titus, Edith D. Adkins, Timothy Chambers, Mark W. Paschke, Katherine D. Heineman, and Matthew A. Albrecht. Identifying predictors of translocation success in rare plant species. Conservation Biology, 38(2):e14190:1–e14190:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Boult:2023:FBA

[Bou23] Victoria L. Boult. Forecast-based action for conservation. Conservation Biology, 37(3):e14054:1–e14054:??, June 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Benchimol:2021:DPP

[BP21] Maíra Benchimol and Carlos A. Peres. Determinants of population persistence and abundance of terrestrial and arboreal vertebrates stranded in tropical forest land-bridge islands. Conservation Biology, 35(3):870–883, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Blamey:2022:RHS

[BPH^{+}22] Laura K. Blamey, Éva E. Plagányi, Trevor Hutton, Roy A. Deng, Judy Upston, and Annie Jarrett. Redesigning harvest strategies for sustainable fishery management in the face of extreme environmental variability. Conservation Biology, 36 (3):e13864:1–e13864:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bogoni:2024:UHH

[BPN^{+}24] Juliano A. Bogoni, Carlos A. Peres, Ana B. Navarro, Vitor Carvalho-Rocha, and Mauro Galetti. Using historical habitat loss to predict contemporary mammal extirpations in neotropical forests. Conservation Biology, 38(4):e14245:1–e14245:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Brown:2024:AFR

[BPR24a] C. Brown, R. Prestele, and M. Rounsevell. An assessment of future rewilding potential in the United Kingdom. Conservation Biology, 38(4):e14276:1–e14276:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See erratum [Ano24e].

Brown:2024:EGB

[BPR24b] C. Brown, R. Prestele, and M. Rounsevell. Erratum to grassland bird sensitivity to weather and climate variability in North America. Conservation Biology, 38(4):e14417:1–e14417:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barredo:2025:ADS

[BRJ25] José I. Barredo, Inés Marí Rivero, and Klára Janoušková. Assessing disturbances in surviving primary forests of Europe. Conservation Biology, 39(2):e14404:1–e14404:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bastille-Rousseau:2021:CLM

[BRW21] Guillaume Bastille-Rousseau and George Wittemyer. Characterizing the landscape of movement to identify critical wildlife habitat and corridors. Conservation Biology, 35(1):346–359, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Barbour:2023:IMB

[BSG^{+}23] Nicole Barbour, George L. Shillinger, Eliezer Gurarie, Aimee L. Hoover, Philippe Gaspar, Julien Temple-Boyer, Tony Candela, William F. Fagan, and Helen Bailey. Incorporating multidimensional behavior into a risk management tool for a critically endangered and migratory species. Conservation Biology, 37(5):e14114:1–e14114:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bush:2024:SNP

[BSH24] Alex Bush, Katherine Hannah Simpson, and Nick Hanley. Systematic nature positive markets. Conservation Biology, 38(3): e14216:1–e14216:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Baisero:2022:RMG

[BSP22] Daniele Baisero, Richard Schuster, and Andrew J. Plumptre. Redefining and mapping global irreplaceability. Conservation Biology, 36(2):e13806:1–e13806:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bennett:2022:ICA

[BSRM22] Ruth E. Bennett, T. Scott Sillett, Robert A. Rice, and Peter P. Marra. Impact of cocoa agricultural intensification on bird diversity and community composition. Conservation Biology, 36(1):e13779:1–e13779:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bull:2021:RMC

[BSSG21] Joseph W. Bull, Niels Strange, Robert J. Smith, and Ascelin Gordon. Reconciling multiple counterfactuals when evaluating biodiversity conservation impact in social-ecological systems. Conservation Biology, 35(2):510–521, April 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bradley:2022:MTM

[BTC^{+}22] Holly S. Bradley, Sean Tomlinson, Michael D. Craig, Adam T. Cross, and Philip W. Bateman. Mitigation translocation as a management tool. Conservation Biology, 36(1):e13667:1– e13667:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Biedenweg:2020:UCM

[BTDS20] Kelly Biedenweg, David Trimbach, Jackie Delie, and Bessie Schwarz. Using cognitive mapping to understand conservation planning. Conservation Biology, 34(6):1364–1372, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Broder:2020:OCR

[BTH^{+}20] Linda Bröder, Laurent Tatin, Axel Hochkirch, Andreas Schuld, Lucas Pabst, and Aurélien Besnard. Optimization of capture– recapture monitoring of elusive species illustrated with a threatened grasshopper. Conservation Biology, 34(3):743–753, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Braulik:2023:RLS

[BTM^{+}23] Gill T. Braulik, Barbara L. Taylor, Gianna Minton, Giuseppe Notarbartolo di Sciara, Tim Collins, Lorenzo Rojas-Bracho, Enrique A. Crespo, Louisa S. Ponnampalam, Michael C. Double, and Randall R. Reeves. Red-list status and extinction risk of the world’s whales, dolphins, and porpoises. Conservation Biology, 37(5):e14090:1–e14090:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Buschman:2022:FCP

[Bus22] Victoria Qutuuq Buschman. Framing co-productive conservation in partnership with Arctic Indigenous peoples. Conservation Biology, 36(6):e13972:1–e13972:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Biber:2022:RWB

[BVH22] Matthias F. Biber, Alke Voskamp, and Christian Hof. Representation of the world’s biophysical conditions by the global protected area network. Conservation Biology, 36(2):e13822:1– e13822:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ban:2020:HCI

[BWN20] Natalie C. Ban, Emma Wilson, and Doug Neasloss. Historical and contemporary indigenous marine conservation strategies in the North Pacific. Conservation Biology, 34(1):5–14, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Betts:2020:FEI

[BYHT^{+}20] Jessica Betts, Richard P. Young, Craig Hilton-Taylor, Michael Hoffmann, Jon Paul Rodríguez, Simon N. Stuart, and E. J. Milner-Gulland. A framework for evaluating the impact of the IUCN red list of threatened species. Conservation Biology, 34 (3):632–643, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Bowie:2022:MCC

[BZT^{+}22] Aleah Bowie, Wen Zhou, Jingzhi Tan, Philip White, Tara Stoinski, Yanjie Su, and Brian Hare. Motivating children’s cooperation to conserve forests. Conservation Biology, 36(4): e13922:1–e13922:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chowdhury:2024:PHB

[CAA^{+}24] Shawan Chowdhury, Sultan Ahmed, Shofiul Alam, Corey T. Callaghan, Priyanka Das, Moreno Di Marco, Enrico Di Minin, Ivan Jarić, Mahzabin Muzahid Labi, Md. Rokonuzzaman, Uri Roll, Valerio Sbragaglia, Asma Siddika, and Aletta Bonn. A protocol for harvesting biodiversity data from Facebook. Conservation Biology, 38(6):e14257:1–e14257:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caro:2023:PGC

[CACM23] Tim Caro, Jeffrey Andrews, Matthew Clark, and Monique Borgerhoff Mulder. Practical guide to coproduction in conservation science. Conservation Biology, 37(1):e14011:1–e14011:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Casanelles-Abella:2023:SMB

[CAFM^{+}23] Joan Casanelles-Abella, Simone Fontana, Eliane Meier, Marco Moretti, and Betrand Fournier. Spatial mismatch between wild bee diversity hotspots and protected areas. Conservation Biology, 37(4):e14082:1–e14082:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Christie:2021:CBE

[CAM^{+}21] Alec P. Christie, Tatsuya Amano, Philip A. Martin, Silviu O. Petrovan, Gorm E. Shackelford, Benno I. Simmons, Rebecca K. Smith, David R. Williams, Claire F. R. Wordley, and William J. Sutherland. The challenge of biased evidence in conservation. Conservation Biology, 35(1):249–262, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cardoso:2020:TIN

[Car20] Pedro Cardoso. The time for insects is now. Conservation Biology, 34(6):1598–1600, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cassini:2023:CRP

[Cas23] Marcelo H. Cassini. A critical review of the precautionary approach of the IUCN impact classification for non-native taxa. Conservation Biology, 37(2):e14037:1–e14037:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cuthbert:2020:ICI

[CBB^{+}20] Ross N. Cuthbert, Sven Bacher, Tim M. Blackburn, Elizabeta Briski, Christophe Diagne, Jaimie T. A. Dick, Franz Essl, Piero Genovesi, Phillip J. Haubrock, Guillaume Latombe, Bernd Lenzner, Yves Meinard, Aníbal Pauchard, Petr Pyšek, Anthony Ricciardi, David M. Richardson, James C. Russell, Daniel Simberloff, and Franck Courchamp. Invasion costs, impacts, and human agency: response to Sagoff 2020. Conservation Biology, 34(6):1579–1582, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Sag20a, Sag20b].

Clark:2021:CCB

[CBD^{+}21] Douglas A. Clark, Ryan K. Brook, Ethan D. Doney, Beatrice Frank, Thomas S. Jung, David S. Lee, Simon Lambert, and Brenda Parlee. Collaborative and consensus-based approaches for human–wildlife coexistence: response to Treves and Santiago-Ávila 2020. Conservation Biology, 35(4):1334– 1336, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [TSÁ20].

Clark:2022:RIP

[CBD22] Douglas A. Clark, Ryan K. Brook, and Ethan Doney. Recommendations for improving peer review of qualitative research in conservation journals. Conservation Biology, 36(1):e13823:1– e13823:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cinner:2021:MCC

[CBG^{+}21] Joshua E. Cinner, Michele L. Barnes, Georgina G. Gurney, Stewart Lockie, and Cristian Rojas. Markets and the crowding out of conservation-relevant behavior. Conservation Biology, 35(3):816–823, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carter:2020:ECO

[CBM20] Neil H. Carter, Andres Baeza, and Nicholas R. Magliocca. Emergent conservation outcomes of shared risk perception in human-wildlife systems. Conservation Biology, 34(4):903–914, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carman:2020:GBD

[CC20] María Carman and Victoria González Carman. Going beyond diverse worldviews for conservation: response to Kohler et al. Conservation Biology, 34(1):286–288, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coghlan:2022:CRC

[CC22] Simon Coghlan and Adam P. A. Cardilini. A critical review of the compassionate conservation debate. Conservation Biology, 36(1):e13760:1–e13760:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coghlan:2024:UAM

[CC24] Simon Coghlan and Adam Cardilini. The use and abuse of moral theories in conservation debate about killing animals. Conservation Biology, 38(4):e14280:1–e14280:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cuadros-Casanova:2023:OCC

[CCCB^{+}23] Ivon Cuadros-Casanova, Andrea Cristiano, Dino Biancolini, Marta Cimatti, Andrea Antonio Sessa, Valeria Yeraldin Mendez Angarita, Chiara Dragonetti, Michela Pacifici, Carlo Rondinini, and Moreno Di Marco. Opportunities and challenges for Common Agricultural Policy reform to support the European Green Deal. Conservation Biology, 37(3):e14052:1– e14052:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carver:2021:GPR

[CCH^{+}21] Steve Carver, Ian Convery, Sally Hawkins, Rene Beyers, Adam Eagle, Zoltan Kun, Erwin Van Maanen, Yue Cao, Mark Fisher, Stephen R. Edwards, Cara Nelson, George D. Gann, Steve Shurter, Karina Aguilar, Angela Andrade, William J. Ripple, John Davis, Anthony Sinclair, Marc Bekoff, Reed Noss, Dave Foreman, Hanna Pettersson, Meredith Root-Bernstein, Jens-Christian Svenning, Peter Taylor, Sophie Wynne-Jones, Alan Watson Featherstone, Camilla Fløjgaard, Mark Stanley-Price, Laetitia M. Navarro, Toby Aykroyd, Alison Parfitt, and Michael Soulé. Guiding principles for rewilding. Conservation Biology, 35(6):1882–1893, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cravens:2025:DFR

[CCKT25] Amanda E. Cravens, Katherine R. Clifford, Corrine Knapp, and William R. Travis. The dynamic feasibility of resisting (R), accepting (A), or directing (D) ecological change. Conservation Biology, 39(2):e14331:1–e14331:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coddington:2023:EFF

[CCL23] Charles P. J. Coddington, W. Justin Cooper, and David A. Luther. Effects of forest fragmentation on avian breeding activity. Conservation Biology, 37(4):e14063:1–e14063:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chang:2025:NOC

[CCPE^{+}25] Charlotte H. Chang, Susan C. Cook-Patton, James T. Erbaugh, Luci Lu, Yuta J. Masuda, István Molnár, Dávid Papp, and Brian E. Robinson. New opportunities and challenges for conservation evidence synthesis from advances in natural language processing. Conservation Biology, 39(2):e14464:1– e14464:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cui:2024:IGC

[CCW24] Yu Cui, Carlos Pérez Carmona, and Zhiheng Wang. Identifying global conservation priorities for terrestrial vertebrates based on multiple dimensions of biodiversity. Conservation Biology, 38(3):e14205:1–e14205:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carlson:2023:ESS

[CDSB23] Shelby C. Carlson, Alia M. Dietsch, Kristina M. Slagle, and Jeremy T. Bruskotter. Effect of semantics in the study of tolerance for wolves. Conservation Biology, 37(2):e14003:1– e14003:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cheek:2023:GQA

[CEM^{+}23] Jennifer Zavaleta Cheek, Johanna Eklund, Nikolas Merten, Jeremy Brooks, and Daniel C. Miller. A guide to qualitative attribution methods for evaluation in conservation. Conservation Biology, 37(4):e14071:1–e14071:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chowdhury:2024:USM

[CFA^{+}24] Shawan Chowdhury, Richard A. Fuller, Sultan Ahmed, Shofiul Alam, Corey T. Callaghan, Priyanka Das, Ricardo A. Correia, Moreno Di Marco, Enrico Di Minin, Ivan Jarić, Mahzabin Muzahid Labi, Richard J. Ladle, M. Rokonuzzaman, Uri Roll, Valerio Sbragaglia, Asma Siddika, and Aletta Bonn. Using social media records to inform conservation planning. Conservation Biology, 38(1):e14161:1–e14161:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Curnick:2021:RLM

[CFC21] David J. Curnick, David A. Feary, and Geórgenes H. Cavalcante. Risks to large marine protected areas posed by drifting fish aggregation devices. Conservation Biology, 35(4):1222– 1232, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Collins:2024:PIS

[CFN^{+}24] Amy C. Collins, Abigail B. Feuka, Jasmine Nelson, Anahita K. Verahrami, and Sara Bombaci. Perspectives on inclusion, safety, and belonging from members of the North American LGBTQIA+ conservation community. Conservation Biology, 38(6):e14389:1–e14389:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cantonati:2021:UPG

[CFS^{+}21] Marco Cantonati, Roderick J. Fensham, Lawrence E. Stevens, Reinhard Gerecke, Douglas S. Glazier, Nico Goldscheider, Robert L. Knight, John S. Richardson, Abraham E. Springer, and Klement Tockner. Urgent plea for global protection of springs. Conservation Biology, 35(1):378–382, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chowdhury:2022:GNE

[CGA^{+}22] Shawan Chowdhury, Kristofer Gonzalez, M. Çisel Kemahlı Aytekin, Seung-Yun Baek, Michał Bełcik, Sandro Bertolino, Sjoerd Duijns, Yuqing Han, Kerstin Jantke, Ryosuke Katayose, Mu-Ming Lin, Elham Nourani, Danielle Leal Ramos, Marie-Morgane Rouyer, William Sidemo-Holm, Svetlana Vozykova, Veronica Zamora-Gutierrez, and Tatsuya Amano. Growth of non-English-language literature on biodiversity conservation. Conservation Biology, 36(4):e13883:1–e13883:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chen:2024:CCT

[CGB^{+}24] Cheng Chen, Alys Granados, Jedediah F. Brodie, Roland Kays, T. Jonathan Davies, Runzhe Liu, Jason T. Fisher, Jorge Ahumada, William McShea, Douglas Sheil, Jayasilan Mohd-Azlan, Bernard Agwanda, Mahandry H. Andrianarisoa, Robyn D. Appleton, Robert Bitariho, Santiago Espinosa, Melissa M. Grigione, Kristofer M. Helgen, Andy Hubbard, Cindy M. Hurtado, Patrick A. Jansen, Xuelong Jiang, Alex Jones, Elizabeth L. Kalies, Cisquet Kiebou-Opepa, Xueyou Li, Marcela Guimarães Moreira Lima, Erik Meyer, Anna B. Miller, Thomas Murphy, Renzo Piana, Rui-Chang Quan, Christopher T. Rota, Francesco Rovero, Fernanda Santos, Stephanie Schuttler, Aisha Uduman, Joanna Klees van Bommel, Hilary Young, and A. Cole Burton. Combining camera trap surveys and IUCN range maps to improve knowledge of species distributions. Conservation Biology, 38(3):e14221:1–e14221:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coleman:2025:QIF

[CGL^{+}25] Daniel J. Coleman, Rachel K. Gittman, Craig E. Landry, James E. Byers, Clark R. Alexander, G. Paul Coughlin, and C. Brock Woodson. Quantifying the impacts of future shoreline modification on biodiversity in a case study of coastal Georgia, United States. Conservation Biology, 39(1):e14301:1– e14301:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crespo:2024:ASM

[CGM^{+}24] Guillermo Ortuño Crespo, Shane Griffiths, Hilario Murua, Henrik Österblom, and Jon Lopez. Adaptive spatiotemporal management to reduce shark bycatch in tuna fisheries. Conservation Biology, 38(4):e14324:1–e14324:??, August 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Choi:2025:RLI

[CGM^{+}25] Jonathan J. Choi, Leo C. Gaskins, Joseph P. Morton, Julia A. Bingham, Ashley M. Blawas, Christine Hayes, Carmen Hoyt, Patrick N. Halpin, and Brian Silliman. Role of low-impact-factor journals in conservation implementation. Conservation Biology, 39(2):e14391:1–e14391:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark:2024:EPF

[CHA^{+}24] Matt Clark, Haji Masoud Hamad, Jeffrey Andrews, Vicken Hillis, and Monique Borgerhoff Mulder. Effects of perceptions of forest change and intergroup competition on communitybased conservation behaviors. Conservation Biology, 38(6): e14259:1–e14259:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Courchamp:2020:IBU

[CHP20] Franck Courchamp, Philip E. Hulme, and Petr Pyšek. Invasion biology and uncertainty in native range definitions: response to Pereyra 2019. Conservation Biology, 34(4):1041–1043, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [PG20].

Carroll:2024:LCE

[CHR24] Carlos Carroll, Sean Hoban, and Justina C. Ray. Lessons from COP15 on effective scientific engagement in biodiversity policy processes. Conservation Biology, 38(2):e14192:1–e14192:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Collins:2023:BII

[CHW^{+}23] Sydney M. Collins, Jack G. Hendrix, Quinn M. R. Webber, Sean P. Boyle, Katrien A. Kingdon, Robert J. Blackmore, Kyle J. N. d’Entremont, Jennifer Hogg, Juan P. Ibáñez, Joanie L. Kennah, Jessika Lamarre, Miguel Mejías, Levi Newediuk, Cerren Richards, Katrina Schwedak, Chirathi Wijekulathilake, and Julie W. Turner. Bibliometric investigation of the integration of animal personality in conservation contexts. Conservation Biology, 37(2):e14021:1–e14021:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cienciala:2024:CSI

[Cie24] Piotr Cienciala. A case for stronger integration of physical landscape processes in conservation science and practice. Conservation Biology, 38(3):e14229:1–e14229:??, June 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chapron:2020:PNC

[CLB20] Guillaume Chapron and José Vicente López-Bao. The place of nature in conservation conflicts. Conservation Biology, 34 (4):795–802, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Conkling:2021:LLS

[CLD^{+}21] Tara J. Conkling, Scott R. Loss, Jay E. Diffendorfer, Adam E. Duerr, and Todd E. Katzner. Limitations, lack of standardization, and recommended best practices in studies of renewable energy effects on birds and bats. Conservation Biology, 35(1): 64–76, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Choquette:2023:SRS

[CLGP23] Jonathan D. Choquette, Jacqueline D. Litzgus, Joanne X. Y. Gui, and Trevor E. Pitcher. A systematic review of snake translocations to identify potential tactics for reducing postrelease effects. Conservation Biology, 37(1):e14016:1–e14016:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See erratum [Ano23c].

Clink:2023:DDP

[Cli23] Dena J. Clink. Documenting the disappearance of primates from spaces where they once occurred: Spatial Analysis in Field Primatology: Applying GIS at Varying Scales. Dolins, F. L., C. A. Shaffer, L. M. Porter, J. R. Hickey, and N. P. Nibbelink, editors. 2021. Cambridge University Press, Cambridge, U.K. 445 pp. £36.96 (paperback). ISBN 978-1-107-69430-9. All Asian Primates. Beauséjour, S., A. B. Rylands, and R. A. Mittermeier. 2021. Lynx Edicions, Barcelona, Spain. 536 pp. £47.56 (paperback). ISBN 978-1-7372851-1-3. Conservation Biology, 37(5):e14140:1–e14140:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Correia:2021:DDS

[CLJ^{+}21] Ricardo A. Correia, Richard Ladle, Ivan Jarić, Ana C. M. Malhado, John C. Mittermeier, Uri Roll, Andrea Soriano-Redondo, Diogo Veríssimo, Christoph Fink, Anna Hausmann, Jhonatan Guedes-Santos, Reut Vardi, and Enrico Di Minin. Digital data sources and methods for conservation culturomics. Conservation Biology, 35(2):398–411, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cheng:2022:EAM

[CLM22] Chuyu Cheng, Jiajia Liu, and Zhijun Ma. Effects of aquaculture on the maintenance of waterbird populations. Conservation Biology, 36(5):e13913:1–e13913:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Correia:2021:I

[CLR21] Ricardo A. Correia, Richard Ladle, and Uri Roll. Introduction. Conservation Biology, 35(2):395–397, April 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Colleony:2020:PMP

[CLS20] Agathe Colléony, Liat Levontin, and Assaf Shwartz. Promoting meaningful and positive nature interactions for visitors to green spaces. Conservation Biology, 34(6):1373–1382, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crespin:2020:KCS

[CMAS20] Silvio J. Crespin, Dario Moreira-Arce, and Javier A. Simonetti. Killing with compassion for the sake of conservation: response to Lynn et al. 2019. Conservation Biology, 34(4):1035–1037, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [LSÁH^{+}20].

Chin:2023:CFK

[CMC^{+}23] Andrew Chin, Molloy John Molloy, Darren Cameron, Jon C. Day, Jessica Cramp, Karin G. Gerhardt, Michelle R. Heupel, Mark Read, and Colin A. Simpfendorfer. Conceptual frameworks and key questions for assessing the contribution of marine protected areas to shark and ray conservation. Conservation Biology, 37(1):e13917:1–e13917:??, February 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See erratum [Ano23d].

Cai:2024:IHH

[CMP^{+}24] Shangxiao Cai, Tong Mu, He-Bo Peng, Zhijun Ma, and David S. Wilcove. Importance of habitat heterogeneity in tidal flats to the conservation of migratory shorebirds. Conservation Biology, 38(1):e14153:1–e14153:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2021:RFC

[CN21] Carlos Carroll and Reed F. Noss. Rewilding in the face of climate change. Conservation Biology, 35(1):155–167, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2022:HPP

[CN22] Carlos Carroll and Reed F. Noss. How percentage-protected targets can support positive biodiversity outcomes. Conservation Biology, 36(4):e13869:1–e13869:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2023:FCE

[CND^{+}23] Carlos Carroll, Reed F. Noss, Lindsay M. Dreiss, Healy Hamilton, and Bruce A. Stein. Four challenges to an effective national nature assessment. Conservation Biology, 37(5):e14075:1– e14075:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chaudhary:2020:CAP

[CO20] Vratika Chaudhary and Madan K. Oli. A critical appraisal of population viability analysis. Conservation Biology, 34(1):26– 40, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [LFT^{+}21].

Chaudhary:2021:FDP

[CO21] Vratika Chaudhary and Madan K. Oli. False dichotomy in population viability analysis quality assessment: reply to Lawson et al. Conservation Biology, 35(5):1686–1688, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crawford:2024:ISP

[CO24] Reed D. Crawford and Joy M. O’Keefe. Improving the science and practice of using artificial roosts for bats. Conservation Biology, 38(1):e14170:1–e14170:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chan:2025:CCR

[COF^{+}25] Carrol M. H. Chan, Christopher J. Owers, Sascha Fuller, Matt W. Hayward, David Moverley, and Andrea S. Griffin. Capacity and capability of remote sensing to inform invasive plant species management in the Pacific Islands region. Conservation Biology, 39(1):e14344:1–e14344:??, February 2025. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Concepcion:2021:USN

[Con21] Elena D. Concepción. Urban sprawl into Natura 2000 network over Europe. Conservation Biology, 35(4):1063–1072, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cook:2024:PDC

[Coo24] Carly N. Cook. Progress developing the concept of other effective area-based conservation measures. Conservation Biology, 38(1):e14106:1–e14106:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cordier:2025:CBD

[COOH^{+}25] Javier M. Cordier, Luis Osorio-Olvera, Pablo Y. Huais, Ana N. Tomba, Fabricio Villalobos, and Javier Nori. Capability of big data to capture threatened vertebrate diversity in protected areas. Conservation Biology, 39(1):e14371:1–e14371:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cork:2020:EOP

[Cor20] Susan Catherine Cork. Examining options to promote human– wildlife coexistence. Conservation Biology, 34(4):1051–1056, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cork:2021:PCP

[Cor21] Susan Catherine Cork. Perspectives on conservation and the political agenda. Conservation Biology, 35(6):1973–1974, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chiarucci:2020:NGM

[CP20a] Alessandro Chiarucci and Gianluca Piovesan. Need for a global map of forest naturalness for a sustainable future. Conservation Biology, 34(2):368–372, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chiavacci:2020:TBH

[CP20b] Scott J. Chiavacci and Emily J. Pindilli. Trends in biodiversity and habitat quantification tools used for market-based conservation in the United States. Conservation Biology, 34(1): 125–136, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Connor:2022:CEH

[CQS^{+}22] Thomas Connor, Maiju Qiao, Kim Scribner, Jindong Zhang, Vanessa Hull, Wenke Bai, Ashton Shortridge, Rengui Li, Hemin Zhang, and Jianguo Liu. Complex effects of habitat amount and fragmentation on functional connectivity and inbreeding in a giant panda population. Conservation Biology, 36 (2):e13828:1–e13828:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clairbaux:2024:ESP

[CRAN^{+}24] Manon Clairbaux, Mia Rönkä, Tycho Anker-Nilssen, Yuri Artukhin, Jóhannis Danielsen, Maria Gavrilo, Grant Gilchrist, Erpur Snær Hansen, April Hedd, Robert Kaler, Kathy Kuletz, Bergur Olsen, Mark L. Mallory, Flemming Ravn Merkel, Hallvard Strøm, Jérôme Fort, and David Grémillet. An ecologically sound and participatory monitoring network for pan-Arctic seabirds. Conservation Biology, 38(6):e14287:1–e14287:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cheng:2021:SSW

[CRC^{+}21] Tina L. Cheng, Jonathan D. Reichard, Jeremy T. H. Coleman, Theodore J. Weller, Wayne E. Thogmartin, Brian E. Reichert, Alyssa B. Bennett, Hugh G. Broders, Joshua Campbell, Katherine Etchison, Daniel J. Feller, Richard Geboy, Traci Hemberger, Carl Herzog, Alan C. Hicks, Sandra Houghton, Jessica Humber, Joseph A. Kath, R. Andrew King, Susan C. Loeb, Ariane Massé, Katrina M. Morris, Holly Niederriter, Gerda Nordquist, Roger W. Perry, Richard J. Reynolds, D. Blake Sasse, Michael R. Scafini, Richard C. Stark, Craig W. Stihler, Steven C. Thomas, Gregory G. Turner, Shevenell Webb, Bradford J. Westrich, and Winifred F. Frick. The scope and severity of white-nose syndrome on hibernating bats in North America. Conservation Biology, 35(5):1586–1597, October 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Charles:2021:ASM

[CS21] Kimberly M. Charles and Ivana Stehlik. Assisted species migration and hybridization to conserve cold-adapted plants under climate change. Conservation Biology, 35(2):559–566, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cooper:2024:MIH

[CS24] Robert D. Cooper and H. Bradley Shaffer. Managing invasive hybrids with pond hydroperiod manipulation in an endangered salamander system. Conservation Biology, 38(2): e14167:1–e14167:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Clark-Shen:2023:SSA

[CSCA^{+}23] Naomi Clark-Shen, Andrew Chin, Sirachai Arunrugstichai, Jessica Labaja, Meira Mizrahi, Benaya Simeon, and Neil Hutchinson. Status of Southeast Asia’s marine sharks and rays. Conservation Biology, 37(1):e13962:1–e13962:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crane:2025:PRB

[CSGG25] Matt Crane, Inês Silva, Matthew J. Grainger, and George A. Gale. Predicting risk to bat species from wind turbine collision in Southeast Asia. Conservation Biology, 39(2):e14452:1– e14452:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chung:2020:IDB

[CSHM^{+}20] Mi Yoon Chung, Sungwon Son, Sonia Herrando-Moraira, Cindy Q. Tang, Masayuki Maki, Young-Dong Kim, Jordi López-Pujol, James L. Hamrick, and Myong Gi Chung. Incorporating differences between genetic diversity of trees and herbaceous plants in conservation strategies. Conservation Biology, 34(5):1142–1151, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cazalis:2023:PRD

[CSL^{+}23] Victor Cazalis, Luca Santini, Pablo M. Lucas, Manuela González-Suárez, Michael Hoffmann, Ana Benítez-López, Michela Pacifici, Aafke M. Schipper, Monika Böhm, Alexander Zizka, Viola Clausnitzer, Carsten Meyer, Martin Jung, Stuart H. M. Butchart, Pedro Cardoso, Giordano Mancini, H. Reşit Akçakaya, Bruce E. Young, Guillaume Patoine, and Moreno Di Marco. Prioritizing the reassessment of data-deficient species on the IUCN Red List. Conservation Biology, 37(6):e14139:1– e14139:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cadman:2024:UPR

[CSS^{+}24] Rachael Cadman, Alanna Syliboy, Michelle Saunders, Shelley Denny, Mary Denniston, Eleanor Barry, Breanna Bishop, Shannon Landovskis, and Megan Bailey. Using positionality and reflexivity to support equity in partnership-driven research. Conservation Biology, 38(6):e14396:1–e14396:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chekunov:2024:SUI

[CST^{+}24] Sebastian Chekunov, Oliver Stringham, Adam Toomes, Thomas Prowse, and Phillip Cassey. Scale of unregulated international trade in Australian reptiles and amphibians. Conservation Biology, 38(5):e14355:1–e14355:??, October 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Crandall:2023:ITM

[CTL^{+}23] Eric D. Crandall, Rachel H. Toczydlowski, Libby Liggins, Ann E. Holmes, Maryam Ghoojaei, Michelle R. Gaither, Briana E. Wham, Andrea L. Pritt, Cory Noble, Tanner J. Anderson, Randi L. Barton, Justin T. Berg, Sofia G. Beskid, Alonso Delgado, Emily Farrell, Nan Himmelsbach, Samantha R. Queeno, Thienthanh Trinh, Courtney Weyand, Andrew Bentley, John Deck, Cynthia Riginos, Gideon S. Bradburd, and Robert J. Toonen. Importance of timely metadata curation to the global surveillance of genetic diversity. Conservation Biology, 37(4):e14061:1–e14061:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Coldrey:2022:APA

[CTM^{+}22] Kevin M. Coldrey, Jane K. Turpie, Guy Midgley, Simon Scheiter, Lee Hannah, Patrick R. Roehrdanz, and Wendy B. Foden. Assessing protected area vulnerability to climate change in a case study of South African national parks. Conservation Biology, 36(5):e13941:1–e13941:??, October 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chen:2024:EPA

[CTP^{+}24] Xing Chen, Tengteng Tian, Han Pan, Yuyi Jin, Xiaodian Zhang, Biao Yang, and Li Zhang. Establishing a protected area network in Xinlong with other effective area-based conservation measures. Conservation Biology, 38(4):e14297:1– e14297:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cavedon:2022:SBH

[CvH^{+}22] Maria Cavedon, Bridgett vonHoldt, Mark Hebblewhite, Troy Hegel, Elizabeth Heppenheimer, Dave Hervieux, Stefano Mariani, Helen Schwantje, Robin Steenweg, Megan Watters, and Marco Musiani. Selection of both habitat and genes in specialized and endangered caribou. Conservation Biology, 36(4): e13900:1–e13900:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Chaves:2021:IRU

[CVT^{+}21] Willandia A. Chaves, Denis Valle, Aline S. Tavares, Thais Q. Morcatty, and David S. Wilcove. Impacts of rural to urban migration, urbanization, and generational change on consumption of wild animals in the Amazon. Conservation Biology, 35(4): 1186–1197, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Carroll:2025:SEA

[CVT^{+}25] Sarah L. Carroll, Susanne M. Vogel, Purity Nititi Taek, Clevers Tumuti, Divya Vasudev, Varun R. Goswami, Jake Wall, Stephen Mwiu, Robin S. Reid, and Jonathan Salerno. A spatially explicit assessment of factors shaping attitudes toward African elephant conservation. Conservation Biology, 39(1):e14408:1–e14408:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Cicchino:2023:CPP

[CWMB23] Amanda S. Cicchino, Andrea E. Weinberg, Laura B. Sample McMeeking, and Meena M. Balgopal. Critical pedagogy of place to enhance ecological engagement activities. Conservation Biology, 37(2):e14023:1–e14023:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Choi:2022:ECW

[CXJ^{+}22] Chi-Yeung Choi, Hui Xiao, Mingming Jia, Micha V. Jackson, Yi-Chien Lai, Nicholas J. Murray, Luke Gibson, and Richard A. Fuller. An emerging coastal wetland management dilemma between mangrove expansion and shorebird conservation. Conservation Biology, 36(5):e13905:1–e13905:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dragonetti:2023:SCR

[DAD23] Chiara Dragonetti, Valeria Y. Mendez Angarita, and Moreno Di Marco. Scenarios of change in the realized climatic niche of mountain carnivores and ungulates. Conservation Biology, 37 (3):e14035:1–e14035:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doherty:2021:TIU

[DAN^{+}21] Philip D. Doherty, Benoit C. Atsango, Gaston Ngassiki, Appolinaire Ngouembe, Nathalie Bréheret, Eva Chauvet, Brendan J. Godley, Lucie Machin, Baudelaire Dissondet Moundzoho, Richard J. Parnell, and Kristian Metcalfe. Threats of illegal, unregulated, and unreported fishing to biodiversity and food security in the Republic of the Congo. Conservation Biology, 35(5):1463–1472, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davis:2020:LWV

[Dav20] Mark A. Davis. Let’s welcome a variety of voices to invasion biology. Conservation Biology, 34(6):1329–1330, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dickson:2023:ICT

[DBC^{+}23] Iain Dickson, Stuart H. M. Butchart, Allison Catalano, David Gibbons, Julia P. G. Jones, Katie Lee-Brooks, Thomasina Oldfield, David Noble, Stuart Paterson, Sugoto Roy, Julien Semelin, Paul Tinsley-Marshall, Rosie Trevelyan, Hannah Wauchope, Sylvia Wicander, and William J. Sutherland. Introducing a common taxonomy to support learning from failure in conservation. Conservation Biology, 37(1):e13967:1–e13967:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dunn:2025:IRR

[DBM^{+}25] Ruth E. Dunn, Cassandra E. Benkwitt, Olivier Maury, Nicolas Barrier, Peter Carr, and Nicholas A. J. Graham. Island restoration to rebuild seabird populations and amplify coral reef functioning. Conservation Biology, 39(1):e14313:1– e14313:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Denneboom:2024:WMR

[DBMS24] Dror Denneboom, Avi Bar-Massada, and Assaf Shwartz. Wildlife mortality risk posed by high and low traffic roads. Conservation Biology, 38(2):e14159:1–e14159:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Deak:2023:CCH

[DBR^{+}23] Balázs Deák, Ádám Bede, Zoltán Rádai, Iwona Dembicz, Iva Apostolova, Péter Batáry, Róbert Gallé, Csaba Albert Tóth, József Dózsai, Ivan I. Moysiyenko, Barbara Sudnik-Wójcikowska, Maria Zachwatowicz, Georgi Nekhrizov, Fedor N. Lisetskii, Zhanna A. Buryak, Szabolcs Kis, Sándor Borza, Laura Godó, Tatyana M. Bragina, Ilya Smelansky, Ábel Molnár, Miklós Bán, Ferenc Báthori, Zoltán Árgay, János Dani, Réka Kiss, and Orsolya Valkó. Contribution of cultural heritage values to steppe conservation on ancient burial mounds of Eurasia. Conservation Biology, 37(6):e14148:1– e14148:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duncan:2021:IPP

[DBT21] Clare Duncan, Monika Böhm, and Samuel T. Turvey. Identifying the possibilities and pitfalls of conducting IUCN red list assessments from remotely sensed habitat information based on insights from poorly known Cuban mammals. Conservation Biology, 35(5):1598–1614, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Detoeuf:2025:GAS

[DdLI^{+}25] Diane Detoeuf, Emiel de Lange, Harriet Ibbett, Trisha Gupta, Constanza Monterrubio Solís, Krossy Mavakala, Mariana Labão Catapani, Heidi E. Kretser, Eleanor J. Milner-Gulland, Stephanie Brittain, Helen Newing, Brandie Fariss, Charlotte Spira, Harold N. Eyster, Nicole DeMello, Kenneth E. Wallen, Sara A. Thornton, Nathan J. Bennett, and Li Ling Choo. Gap analysis of social science resources for conservation practice. Conservation Biology, 39(2):e14463:1–e14463:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davies:2024:RCC

[DDM24] Alisa Davies, Neil D’Cruze, and Rowan Martin. A review of commercial captive breeding of parrots as a supply-side intervention to address unsustainable trade. Conservation Biology, 38(5):e14338:1–e14338:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donlan:2024:IEC

[DEGL^{+}24] C. Josh Donlan, Diana Eusse-González, Gloria M. Luque, Matthew E. Reiter, Viviana Ruiz-Gutierrez, Michael C. Allen, Richard Johnston-González, Orin J. Robinson, Guillermo Fernández, Eduardo Palacios, and Jorge Valenzuela. An impact evaluation of conservation investments targeting longdistance migratory species. Conservation Biology, 38(3): e14194:1–e14194:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Desbureaux:2021:SMC

[Des21] Sébastien Desbureaux. Subjective modeling choices and the robustness of impact evaluations in conservation science. Conservation Biology, 35(5):1615–1626, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donald:2024:AGP

[DFB^{+}24] Paul F. Donald, Eresha Fernando, Lauren Brown, Michela Busana, Stuart H. M. Butchart, Serene Chng, Alicia de la Colina, Juliana Machado Ferreira, Anuj Jain, Victoria R. Jones, Rocio Lapido, Kelly Malsch, Amy McDougall, Colum Muccio, Dao Nguyen, Willow Outhwaite, Silviu O. Petrovan, Ciara Stafford, William J. Sutherland, Oliver Tallowin, and Roger Safford. Assessing the global prevalence of wild birds in trade. Conservation Biology, 38(5):e14350:1–e14350:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ducarme:2021:HDH

[DFC21] Frédéric Ducarme, Fabrice Flipo, and Denis Couvet. How the diversity of human concepts of nature affects conservation of biodiversity. Conservation Biology, 35(3):1019–1028, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DiMinin:2021:HAD

[DFH^{+}21] Enrico Di Minin, Christoph Fink, Anna Hausmann, Jens Kremer, and Ritwik Kulkarni. How to address data privacy concerns when using social media data in conservation science. Conservation Biology, 35(2):437–446, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dean:2025:EDP

[DFS^{+}25] Angela J. Dean, Kelly S. Fielding, Liam D. G. Smith, Emma K. Church, and Kerrie A. Wilson. Eliciting diverse perspectives to prioritize community actions for biodiversity conservation. Conservation Biology, 39(2):e14372:1–e14372:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Doherty:2024:ETA

[DFW^{+}24] Tim S. Doherty, Daniel F. Bohórquez Fandiño, Darcy J. Watchorn, Sarah M. Legge, and Chris R. Dickman. Experimentally testing animal responses to prescribed fire size and severity. Conservation Biology, 38(3):e14231:1–e14231:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

deGracia:2021:DCS

[dG21] Naira de Gracia. Decolonizing conservation science: response to Jucker et al. 2018. Conservation Biology, 35(4):1321–1323, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DiSanto:2021:UEG

[DH21] Lionel N. Di Santo and Jill A. Hamilton. Using environmental and geographic data to optimize ex situ collections and preserve evolutionary potential. Conservation Biology, 35(2):733–744, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Darimont:2021:LCH

[DHE^{+}21] Chris T. Darimont, Hannah Hall, Lauren Eckert, Ilona Mihalik, Kyle Artelle, Adrian Treves, and Paul C. Paquet. Large carnivore hunting and the social license to hunt. Conservation Biology, 35(4):1111–1119, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davies:2022:IOE

[DHNM22] Alisa Davies, Amy Hinsley, Ana Nuno, and Rowan O. Martin. Identifying opportunities for expert-mediated triangulation in monitoring wildlife trade on social media. Conservation Biology, 36(2):e13858:1–e13858:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Delaney:2024:UEM

[DL24] John T. Delaney and Danelle M. Larson. Using explainable machine learning methods to evaluate vulnerability and restoration potential of ecosystem state transitions. Conservation Biology, 38(3):e14203:1–e14203:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Donfrancesco:2022:MHB

[DLL22] Valerio Donfrancesco and Rogelio Luque-Lora. Managing hybridization beyond the natural-anthropogenic dichotomy. Conservation Biology, 36(2):e13816:1–e13816:??, April 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

delaLama:2023:LAP

[dlLBB^{+}23] Rocío López de la Lama, Nathan Bennett, Janette Bulkan, David Boyd, and Kai M. A. Chan. A legal assessment of private land conservation in South America. Conservation Biology, 37 (4):e14068:1–e14068:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

deLange:2022:ESN

[dLMGK22] Emiel de Lange, E. J. Milner-Gulland, and Aidan Keane. Effects of social networks on interventions to change conservation behavior. Conservation Biology, 36(3):e13833:1–e13833:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Devictor:2020:EBK

[DM20] Vincent Devictor and Yves Meinard. Empowering biodiversity knowledge. Conservation Biology, 34(2):527–529, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Duthie:2021:OMG

[DMR^{+}21] A. Bradley Duthie, Jeroen Minderman, O. Sarobidy Rakotonarivo, Gabriela Ochoa, and Nils Bunnefeld. Online multiplayer games as virtual laboratories for collecting data on social-ecological decision making. Conservation Biology, 35(3): 1051–1053, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Danaher:2022:PCS

[DNCC22] Cody Danaher, Tim Newbold, Jeffrey Cardille, and Abbie S. A. Chapman. Prioritizing conservation in sub-Saharan African lakes based on freshwater biodiversity and algal bloom metrics. Conservation Biology, 36(5):e13914:1–e13914:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Caetano:2023:EGI

[dOCVJ^{+}23] Gabriel Henrique de Oliveira Caetano, Reut Vardi, Ivan Jarić, Ricardo A. Correia, Uri Roll, and Diogo Veríssimo. Evaluating global interest in biodiversity and conservation. Conservation Biology, 37(5):e14100:1–e14100:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DAgostino:2024:BPE

[DPC^{+}24] Martina D’Agostino, Luigi Cao Pinna, Marta Carboni, Silvia Assini, Gianluigi Bacchetta, Fabrizio Bartolucci, Lisa Brancaleoni, Fabrizio Buldrini, Angelino Carta, Bruno Cerabolini, Roberta Maria Ceriani, Umberto Clementi, Donatella Cogoni, Fabio Conti, Roberto Crosti, Alba Cuena-Lombraña, Marcello De Vitis, Attilio Di Giustino, Giuseppe Fabrini, Emanuele Farris, Giuseppe Fenu, Roberto Fiorentin, Bruno Foggi, Luigi Forte, Giuseppe Garfì, Rodolfo Gentili, Gian Pietro Giusso Del Galdo, Valentino Martinelli, Pietro Medagli, Domitilla Nonis, Simone Orsenigo, Luca Paoli, Simon Pierce, Maria Silvia Pinna, Franco Rainini, Sonia Ravera, Graziano Rossi, Aldo Schettino, Rosario Schicchi, Angelo Troìa, Laura Varone, Elena Zappa, and Thomas Abeli. Best practices, errors, and perspectives of half a century of plant translocation in Italy. Conservation Biology, 38(4):e14233:1–e14233:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drechsler:2024:FTO

[Dre24] Martin Drechsler. Flexibility trade-offs in conservation offsets. Conservation Biology, 38(1):e14144:1–e14144:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dan-Rakedzon:2024:FUH

[DRFL^{+}24] Nitzan Dan-Rakedzon, Whitney Fleming, Nurit Lissovsky, Susan Clayton, and Assaf Shwartz. A framework for understanding the human experience of nature through cognitive mapping. Conservation Biology, 38(6):e14283:1–e14283:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Delevaux:2022:PFM

[DS22] Jade M. S. Delevaux and Kostantinos A. Stamoulis. Prioritizing forest management actions to benefit marine habitats in data-poor regions. Conservation Biology, 36(2):e13792:1– e13792:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dawson:2025:ASI

[DS25] Neil M. Dawson and Helen Suich. Advancing social impact assessments for more effective and equitable conservation. Conservation Biology, 39(2):e14453:1–e14453:??, April 2025. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Drew:2024:SPM

[DSC^{+}24] Joshua Drew, Sevanaia Sakai, Akanisi Caginitoba, Lauren C. Warr, Jessica I. Espinosa, and Kelly H. Dunning. Stakeholder perceptions of mangrove ecosystem services across scales of conservation focus. Conservation Biology, 38(6):e14405:1– e14405:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Daniel:2023:VGB

[DSF^{+}23] Alexandrine Daniel, Paul Savary, Jean-Christophe Foltête, Aurélie Khimoun, Bruno Faivre, Anthony Ollivier, Cyril Éraud, Hervé Moal, Gilles Vuidel, and Stéphane Garnier. Validating graph-based connectivity models with independent presence–absence and genetic data sets. Conservation Biology, 37(3):e14047:1–e14047:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Danis:2024:IWC

[DSM24] Blake E. G. Danis, Pourya Sardari, and Vicki L. Marlatt. Importance of water contaminant testing in amphibian reintroduction programs. Conservation Biology, 38(6):e14347:1– e14347:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dickie:2023:EIC

[DSS^{+}23] Melanie Dickie, Geoff G. Sherman, Glenn D. Sutherland, Robert S. McNay, and Michael Cody. Evaluating the impact of caribou habitat restoration on predator and prey movement. Conservation Biology, 37(2):e14004:1–e14004:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dietsch:2025:CSS

[DSvE^{+}25] Alia M. Dietsch, Matthew J. Selinske, Lily M. van Eeden, Kwan-Lamar Blount-Hill, Rina S. Hauptfeld, Sayan Banerjee, Stephanie Brittain, Francisco Gelves-Gómez, Amit Kaushik, Patricia Manzano Fischer, M. Kutub Uddin, and Kenneth E. Wallen. Conservation and the social sciences revisited. Conservation Biology, 39(2):e14462:1–e14462:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Davy:2021:EST

[DSZ21] Christina M. Davy, Kelly Squires, and J. Ryan Zimmerling. Estimation of spatiotemporal trends in bat abundance from mortality data collected at wind turbines. Conservation Biology, 35(1):227–238, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

DePuy:2025:NSI

[DTSE25] Walker DePuy, Paul Thung, Viola Schreer, and Wendy M. Erb. Navigating scale and interdisciplinary dynamics in conservation social science. Conservation Biology, 39(2):e70005:1– e70005:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dorrough:2025:EME

[DTV^{+}25] Josh Dorrough, Samantha K. Travers, James Val, Mitchell L. Scott, Claudine J. Moutou, and Ian Oliver. Evaluating models of expert judgment to inform assessment of ecosystem viability and collapse. Conservation Biology, 39(2):e14370:1–e14370:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dean:2023:RBH

[DW23] Angela J. Dean and Kerrie A. Wilson. Relationships between hope, optimism, and conservation engagement. Conservation Biology, 37(2):e14009:1–e14009:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dietsch:2020:IND

[DWC^{+}20] A. M. Dietsch, K. E. Wallen, S. Clayton, H. E. Kretser, G. T. Kyle, Z. Ma, and A. Vercammen. Introduction: New directions in conservation psychology at a critical time. Conservation Biology, 34(6):1335–1338, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Dacks:2025:EHD

[DYM25] Rachel Dacks, Shreya Yadav, and Alexander Mawyer. Emerging human dimensions research in coastal and nearshore Oceania. Conservation Biology, 39(2):e14455:1–e14455:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Escribano:2021:GTR

[EAdC^{+}21] Nora Escribano, Arturo H. Ariño, Andrea Pino del Carpio, David Galicia, and Rafael Miranda. Global trends in research output by zoos and aquariums. Conservation Biology, 35(6): 1894–1902, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Esbach:2025:ACA

[ECVL25] Michael S. Esbach, Joel E. Correia, Gabriela Valdivia, and Flora Lu. Amazonian conservation across archipelagos of Indigenous territories. Conservation Biology, 39(2):e14407:1– e14407:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edwards:2021:DMW

[EDA^{+}21] David P. Edwards, Neil D’Cruze, Sandra Altherr, Alice Hughes, Jordi Janssen, Vincent Nijman, Stesha A. Pasachnik, Brett R. Scheffers, Chris R. Shepherd, Emerson Sy, and Mark Auliya. The dangers of misrepresenting wildlife trade: response to Natusch et al. 2021. Conservation Biology, 35(5): 1692–1694, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Evans:2023:CTH

[EGBJ23] Luke Christopher Evans, Matthew P. Greenwell, Victoria L. Boult, and Thomas Frederick Johnson. Characterizing the trophy hunting debate on Twitter. Conservation Biology, 37(4): e14070:1–e14070:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eyster:2025:UTH

[EGCS25] Harold N. Eyster, Rachelle K. Gould, Kai M. A. Chan, and Terre Satterfield. Use of theories of human action in recent conservation research. Conservation Biology, 39(2):e14461:1– e14461:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Efrat:2020:TLS

[EHBT20] Ron Efrat, Ohad Hatzofe, and Oded Berger-Tal. Translating large-scale prioritization models for vultures to local-scale decision-making: response to Santangeli et al. 2019. Conservation Biology, 34(5):1305–1307, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Elmore:2021:CBC

[EHC^{+}21] Jared A. Elmore, Stephen B. Hager, Bradley J. Cosentino, Timothy J. O’Connell, Corey S. Riding, Michelle L. Anderson, Marja H. Bakermans, Than J. Boves, David Brandes, Eric M. Butler, Michael W. Butler, Nicolette L. Cagle, Rafael Calderón-Parra, Angelo P. Capparella, Anqi Chen, Kendra Cipollini, April A. T. Conkey, Thomas A. Contreras, Rebecca I. Cooper, Clay E. Corbin, Robert L. Curry, Jerald J. Dosch, Karen L. Dyson, Erin E. Fraser, Ross A. Furbush, Natasha D. G. Hagemeyer, Kristine N. Hopfensperger, Daniel Klem, Jr., Elizabeth A. Lago, Ally S. Lahey, Craig S. Machtans, Jessa M. Madosky, Terri J. Maness, Kelly J. McKay, Sean B. Menke, Natalia Ocampo-Peñuela, Rubén Ortega-Álvarez, Amber L. Pitt, Aura Puga-Caballero, John E. Quinn, Amber M. Roth, Ryan T. Schmitz, Jaclyn L. Schnurr, Matthew E. Simmons, Alexis D. Smith, Claire W. Varian-Ramos, Eric L. Walters, Lindsey A. Walters, Jason T. Weir, Kathy Winnett-Murray, Iriana Zuria, Jesse Vigliotti, and Scott R. Loss. Correlates of bird collisions with buildings across three North American countries. Conservation Biology, 35(2):654–665, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Edelblutte:2023:AAW

[EKH23] Émilie Edelblutte, Roopa Krithivasan, and Matthew Nassif Hayek. Animal agency in wildlife conservation and management. Conservation Biology, 37(1):e13853:1–e13853:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Eleftheriou:2020:RAH

[Ele20] Andreas Eleftheriou. Relationships among host microbiota, parasite resistance or tolerance, and host fitness. Conservation Biology, 34(6):1327–1328, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Evans:2021:SHC

[EM21] Michael J. Evans and Jacob W. Malcom. Supporting habitat conservation with automated change detection in Google Earth Engine. Conservation Biology, 35(4):1151–1161, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

ElBizri:2020:UWM

[EMV^{+}20] Hani R. El Bizri, Thaís Q. Morcatty, João Valsecchi, Pedro Mayor, Jéssica E. S. Ribeiro, Carlos F. A. Vasconcelos Neto, Jéssica S. Oliveira, Keilla M. Furtado, Urânia C. Ferreira, Carlos F. S. Miranda, Ciclene H. Silva, Valdinei L. Lopes, Gerson P. Lopes, Caio C. F. Florindo, Romerson C. Chagas, Vincent Nijman, and Julia E. Fa. Urban wild meat consumption and trade in central Amazonia. Conservation Biology, 34 (2):438–448, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Etard:2024:SLC

[EN24] Adrienne Etard and Tim Newbold. Species-level correlates of land-use responses and climate-change sensitivity in terrestrial vertebrates. Conservation Biology, 38(3):e14208:1–e14208:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

ElBizri:2024:EIH

[EOR^{+}24] Hani R. El Bizri, Marcela A. Oliveira, Aline Pessutti Rampini, Simon Knoop, Julia E. Fa, Lauren Coad, Thais Queiroz Morcatty, Gabriel Favero Massocato, Arnaud L. J. Desbiez, João Vitor Campos-Silva, Daniel Zani La Laina, José Maurício Barbanti Duarte, Rafael Sá Leitão Barboza, Zilca Campos, Marcélia Basto da Silva, Sarah Mângia, Daniel J. Ingram, and Juliano A. Bogoni. Exposing illegal hunting and wildlife depletion in the world’s largest tropical country through social media data. Conservation Biology, 38(5): e14334:1–e14334:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Escalle:2024:SDF

[EPL^{+}24] Lauriane Escalle, J. Scutt Phillips, J. Lopez, J. M. Lynch, H. Murua, S. J. Royer, Y. Swimmer, J. Murua, Alex Sen Gupta, V. Restrepo, and G. Moreno. Simulating drifting fish aggregating device trajectories to identify potential interactions with endangered sea turtles. Conservation Biology, 38 (6):e14295:1–e14295:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Evans:2021:QGT

[Eva21] Thomas Evans. Quantifying the global threat to native birds from predation by non-native birds on small islands. Conservation Biology, 35(4):1268–1277, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Faria:2021:COR

[Far21] Luiz R. R. Faria. Conservation optimism and reckoning with the future. Conservation Biology, 35(2):745–747, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferrarini:2022:NMB

[FBDA22] Alessandro Ferrarini, Yang Bai, Junhu Dai, and Juha M. Alatalo. A new method for broad-scale modeling and projection of plant assemblages under climatic, biotic, and environmental cofiltering. Conservation Biology, 36(2):e13797:1–e13797:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fijen:2025:ARI

[FBG^{+}25] Thijs P. M. Fijen, Gabriella A. Bishop, Cristina Ganuza, Jeroen Scheper, and David Kleijn. Analyzing the relative importance of habitat quantity and quality for boosting pollinator populations in agricultural landscapes. Conservation Biology, 39(1):e14317:1–e14317:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fernandes:2023:UCF

[FBS^{+}23] Kristen Fernandes, Philip W. Bateman, Benjamin J. Saunders, Michael Bunce, Kristine Bohmann, and Paul Nevill. Use of carrion fly iDNA metabarcoding to monitor invasive and native mammals. Conservation Biology, 37(5):e14098:1–e14098:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Freund:2024:ECV

[FCH^{+}24] Cathryn A. Freund, Katherine A. Cronin, Michelle Huang, Natalie J. Robinson, Boaz Yoo, and Andrea L. DiGiorgio. Effects of captions on viewers’ perceptions of images depicting human– primate interaction. Conservation Biology, 38(3):e14199:1– e14199:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Feng:2022:IPA

[FCL^{+}22] Chun-Ting Feng, Ming Cao, Fang-Zheng Liu, Yue Zhou, Jin-Hong Du, Li-Bo Zhang, Wen-Jie Huang, Jian-Wu Luo, Jun-Sheng Li, and Wei Wang. Improving protected area effectiveness through consideration of different human-pressure baselines. Conservation Biology, 36(4):e13887:1–e13887:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fariss:2023:CSC

[FDP^{+}23] Brandie Fariss, Nicole DeMello, Kathryn A. Powlen, Christopher E. Latimer, Yuta Masuda, and Christina M. Kennedy. Catalyzing success in community-based conservation. Conservation Biology, 37(1):e13973:1–e13973:??, February 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferraro:2023:RTD

[FFAS23] Kristy M. Ferraro, Anthony L. Ferraro, A. Z. Andis Arietta, and Nathalie R. Sommer. Revisiting two dogmas of conservation science. Conservation Biology, 37(4):e14101:1–e14101:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ferrante:2023:EAF

[FGB^{+}23] Lucas Ferrante, Augusto Getirana, Fabricio Beggiato Baccaro, Jochen Schöngart, Ana Cristina Monteiro Leonel, Renato Gaiga, Michel Varajão Garey, and Philip Martin Fearnside. Effects of Amazonian flying rivers on frog biodiversity and populations in the Atlantic rainforest. Conservation Biology, 37 (3):e14033:1–e14033:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Falaschi:2021:LTD

[FGL^{+}21] Mattia Falaschi, Simone Giachello, Elia Lo Parrino, Martina Muraro, Raoul Manenti, and Gentile Francesco Ficetola. Longterm drivers of persistence and colonization dynamics in spatially structured amphibian populations. Conservation Biology, 35(5):1530–1539, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Free:2021:EIF

[FJH21] Christopher M. Free, Olaf P. Jensen, and Ray Hilborn. Evaluating impacts of forage fish abundance on marine predators. Conservation Biology, 35(5):1540–1551, October 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Farhadinia:2020:ESM

[FJZ^{+}20] Mohammad S. Farhadinia, Paul J. Johnson, Alexandra Zimmermann, Philip J. K. McGowan, Erik Meijaard, Mark Stanley-Price, and David W. Macdonald. Ex situ management as insurance against extinction of mammalian megafauna in an uncertain world. Conservation Biology, 34(4):988–996, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [KOO^{+}21].

Farhadinia:2021:UES

[FJZ^{+}21] Mohammad S. Farhadinia, Paul J. Johnson, Alexandra Zimmermann, Philip J. K. McGowan, Erik Meijaard, Mark Stanley-Price, and David W. Macdonald. Use of ex situ management not necessarily a last resort: reply to Khalatbari et al. 2021. Conservation Biology, 35(4):1331–1333, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [KOO^{+}21].

Foley:2020:MEP

[FL20] Catherine M. Foley and Heather J. Lynch. A method to estimate pre-exploitation population size. Conservation Biology, 34(1):256–265, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fea:2021:RNZ

[FLH21] Nyree Fea, Wayne Linklater, and Stephen Hartley. Responses of New Zealand forest birds to management of introduced mammals. Conservation Biology, 35(1):35–49, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fang:2022:CLT

[FLW^{+}22] Guirong Fang, Hong Liu, Ka Wu, Tongqing Wei, and Qing Wang. Changing legislative thinking in China to better protect wild animals and human health. Conservation Biology, 36(1): e13837:1–e13837:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Flanagan:2021:MTB

[FMGS21] Alison M. Flanagan, Bryce Masuda, Catherine E. Grueber, and Jolene T. Sutton. Moving from trends to benchmarks by using regression tree analysis to find inbreeding thresholds in a critically endangered bird. Conservation Biology, 35(4):1278– 1287, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Farr:2022:QCS

[FOYZ22] Matthew T. Farr, Timothy O’Brien, Charles B. Yackulic, and Elise F. Zipkin. Quantifying the conservation status and abundance trends of wildlife communities with detection– nondetection data. Conservation Biology, 36(6):e13934:1– e13934:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fuzessy:2024:LSF

[FPC^{+}24] Lisieux Fuzessy, Sandrine Pavoine, Laura Cardador, Joan Maspons, and Daniel Sol. Loss of species and functions in a deforested megadiverse tropical forest. Conservation Biology, 38 (4):e14250:1–e14250:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fischer:2021:GHC

[FR21] Joern Fischer and Maraja Riechers. From grief to hope in conservation. Conservation Biology, 35(5):1698–1700, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fontaine:2022:SCC

[FSBE22] Amélie Fontaine, Anouk Simard, Nicolas Brunet, and Kyle H. Elliott. Scientific contributions of citizen science applied to rare or threatened animals. Conservation Biology, 36(6):e13976:1– e13976:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fitzgerald:2021:UEK

[FSC^{+}21] Daniel B. Fitzgerald, David R. Smith, David C. Culver, Daniel Feller, Daniel W. Fong, Jeff Hajenga, Matthew L. Niemiller, Daniel C. Nolfi, Wil D. Orndorff, Barbara Douglas, Kelly O. Maloney, and John A. Young. Using expert knowledge to support endangered species act decision-making for data-deficient species. Conservation Biology, 35(5):1627–1638, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Furtado:2020:TSM

[Fur20] Miguel Furtado. Toward sustainable management of fish stocks. Conservation Biology, 34(6):1600–1602, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fu:2022:NCP

[FW22] Minjie Fu and Bruce Waldman. Novel chytrid pathogen variants and the global amphibian pet trade. Conservation Biology, 36(5):e13938:1–e13938:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fukushima:2021:WCS

[FWP^{+}21] Caroline Fukushima, Rick West, Thomas Pape, Lyubomir Penev, Leif Schulman, and Pedro Cardoso. Wildlife collection for scientific purposes. Conservation Biology, 35(1):5–11, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fu:2024:COH

[FWZ^{+}24] Changjian Fu, Fang Wang, Yumeng Zhao, Qin Zhu, Yunchao Luo, Yuhang Li, Ziye Zhang, Xueting Yan, Taozhu Sun, Yang Liu, and Zhongqiu Li. Challenges and opportunities in human dimensions behind cat–wildlife conflict. Conservation Biology, 38(6):e14253:1–e14253:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Fischer:2024:LBR

[FZG^{+}24] Richard Fischer, Eliza Zhunusova, Sven Günter, Susanne Iost, Franziska Schier, Jörg Schweinle, Holger Weimar, and Matthias Dieter. Leakage of biodiversity risks under the European Union Biodiversity Strategy 2030. Conservation Biology, 38 (3):e14235:1–e14235:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grace:2021:TGS

[GAB^{+}21] Molly K. Grace, H. Resit Akçakaya, Elizabeth L. Bennett, Thomas M. Brooks, Anna Heath, Simon Hedges, Craig Hilton-Taylor, Michael Hoffmann, Axel Hochkirch, Richard Jenkins, David A. Keith, Barney Long, David P. Mallon, Erik Meijaard, E. J. Milner-Gulland, Jon Paul Rodriguez, P. J. Stephenson, Simon N. Stuart, Richard P. Young, Pablo Acebes, Joanna Alfaro-Shigueto, Silvia Alvarez-Clare, Raphali Rodlis Andriantsimanarilafy, Marina Arbetman, Claudio Azat, Gianluigi Bacchetta, Ruchi Badola, Luís M. D. Barcelos, Joao Pedro Barreiros, Sayanti Basak, Danielle J. Berger, Sabuj Bhattacharyya, Gilad Bino, Paulo A. V. Borges, Raoul K. Boughton, H. Jane Brockmann, Hannah L. Buckley, Ian J. Burfield, James Burton, Teresa Camacho-Badani, Luis Santiago Cano-Alonso, Ruth H. Carmichael, Christina Carrero, John P. Carroll, Giorgos Catsadorakis, David G. Chapple, Guillaume Chapron, Gawsia Wahidunnessa Chowdhury, Louw Claassens, Donatella Cogoni, Rochelle Constantine, Christie Anne Craig, Andrew A. Cunningham, Nishma Dahal, Jennifer C. Daltry, Goura Chandra Das, Niladri Dasgupta, Alexandra Davey, Katharine Davies, Pedro Develey, Vanitha Elangovan, David Fairclough, Mirko Di Febbraro, Giuseppe Fenu, Fernando Moreira Fernandes, Eduardo Pinheiro Fernandez, Brittany Finucci, Rita Földesi, Catherine M. Foley, Matthew Ford, Michael R. J. Forstner, Néstor García, Ricardo Garcia-Sandoval, Penny C. Gardner, Roberto Garibay-Orijel, Marites Gatan-Balbas, Irene Gauto, Mirza Ghazanfar Ullah Ghazi, Stephanie S. Godfrey, Matthew Gollock, Benito A. González, Tandora D. Grant, Thomas Gray, Andrew J. Gregory, Roy H. A. van Grunsven, Marieka Gryzenhout, Noelle C. Guernsey, Garima Gupta, Christina Hagen, Christian A. Hagen, Madison B. Hall, Eric Hallerman, Kelly Hare, Tom Hart, Ruston Hartdegen, Yvette Harvey-Brown, Richard Hatfield, Tahneal Hawke, Claudia Hermes, Rod Hitchmough, Pablo Melo Hoffmann, Charlie Howarth, Michael A. Hudson, Syed Ainul Hussain, Charlie Huveneers, Hélène Jacques, Dennis Jorgensen, Suyash Katdare, Lydia K. D. Katsis, Rahul Kaul, Boaz Kaunda-Arara, Lucy Keith-Diagne, Daniel T. Kraus, Thales Moreira de Lima, Ken Lindeman, Jean Linsky, Edward Louis, Jr., Anna Loy, Eimear Nic Lughadha, Jeffrey C. Mangel, Paul E. Marinari, Gabriel M. Martin, Gustavo Martinelli, Philip J. K. McGowan, Alistair McInnes, Eduardo Teles Barbosa Mendes, Michael J. Millard, Claire Mirande, Daniel Money, Joanne M. Monks, Carolina Laura Morales, Nazia Naoreen Mumu, Raquel Negrao, Anh Ha Nguyen, Md. Nazmul Hasan Niloy, Grant Leslie Norbury, Cale Nordmeyer, Darren Norris, Mark O’Brien, Gabriela Akemi Oda, Simone Orsenigo, Mark Evan Outerbridge, Stesha Pasachnik, Juan Carlos Pérez-Jiménez, Charlotte Pike, Fred Pilkington, Glenn Plumb, Rita de Cassia Quitete Portela, Ana Prohaska, Manuel G. Quintana, Eddie Fanantenana Rakotondrasoa, Dustin H. Ranglack, Hassan Rankou, Ajay Prakash Rawat, James Thomas Reardon, Marcelo Lopes Rheingantz, Stephen C. Richter, Malin C. Rivers, Luke Rollie Rogers, Patrícia da Rosa, Paul Rose, Emily Royer, Catherine Ryan, Yvonne J. Sadovy de Mitcheson, Lily Salmon, Carlos Henrique Salvador, Michael J. Samways, Tatiana Sanjuan, Amanda Souza dos Santos, Hiroshi Sasaki, Emmanuel Schutz, Heather Ann Scott, Robert Michael Scott, Fabrizio Serena, Surya P. Sharma, John A. Shuey, Carlos Julio Polo Silva, John P. Simaika, David R. Smith, Julia L. Y. Spaet, Shanjida Sultana, Bibhab Kumar Talukdar, Vikash Tatayah, Philip Thomas, Angela Tringali, Hoang Trinh-Dinh, Chongpi Tuboi, Aftab Alam Usmani, Aída M. Vasco-Palacios, Jean-Christophe Vié, Evelyn Virens, Alan Walker, Bryan Wallace, Lauren J. Waller, Hongfeng Wang, Oliver R. Wearn, Merlijn van Weerd, Simon Weigmann, Daniel Willcox, John Woinarski, Jean W. H. Yong, and Stuart Young. Testing a global standard for quantifying species recovery and assessing conservation impact. Conservation Biology, 35(6):1833–1849, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Golebie:2022:QSR

[GAB^{+}22] Elizabeth J. Golebie, Miriam Aczel, Jacob J. Bukoski, Sophia Chau, Natali Ramirez-Bullon, Mimi Gong, and Noah Teller. A qualitative systematic review of governance principles for mangrove conservation. Conservation Biology, 36(1):e13850:1– e13850:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gore:2022:IDI

[GB22] Meredith L. Gore and Abigail Bennett. Importance of deepening integration of crime and conservation sciences. Conservation Biology, 36(1):e13710:1–e13710:??, February 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerber:2023:ALI

[GBA23] Leah R. Gerber, Chris J. Barton, and Derrick M. Anderson. Aligning the logics of inquiry and action to address the biodiversity crisis. Conservation Biology, 37(5):e14128:1–e14128:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guerrero:2020:KCC

[GBB^{+}20] A. M. Guerrero, M. Barnes, Ö. Bodin, I. Chadès, K. J. Davis, M. S. Iftekhar, C. Morgans, and K. A. Wilson. Key considerations and challenges in the application of social-network research for environmental decision making. Conservation Biology, 34(3):733–742, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilbert:2020:CTT

[GBDM^{+}20] Sophie L. Gilbert, Kate Broadley, Darcy Doran-Myers, Amanda Droghini, Jessica A. Haines, Anni Hämäläinen, Clayton T. Lamb, Eric W. Neilson, and Stan Boutin. Conservation triage at the trailing edge of climate envelopes. Conservation Biology, 34(1):289–292, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grummer:2022:ICL

[GBMD^{+}22] Jared A. Grummer, Tom R. Booker, Remi Matthey-Doret, Pirmin Nietlisbach, Andréa T. Thomaz, and Michael C. Whitlock. The immediate costs and long-term benefits of assisted gene flow in large populations. Conservation Biology, 36(4): e13911:1–e13911:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gordon:2020:RCC

[GBRK^{+}20] Elizabeth Robson Gordon, Nathalie Butt, Hanna Rosner-Katz, Allison D. Binley, and Joseph R. Bennett. Relative costs of conserving threatened species across taxonomic groups. Conservation Biology, 34(1):276–281, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gill:2022:SQR

[GCA^{+}22] Nicholas Gill, Laurie Chisholm, Jennifer Atchison, Sonia Graham, Gina Hawkes, Lesley Head, and Shaun McKiernan. Scaling up qualitative research to harness the capacity of lay people in invasive plant management. Conservation Biology, 36(6): e13929:1–e13929:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gagne:2022:PCT

[GCC^{+}22] Roderick B. Gagne, Kevin R. Crooks, Meggan E. Craft, Elliott S. Chiu, Nicholas M. Fountain-Jones, Jennifer L. Malmberg, Scott Carver, W. Chris Funk, and Sue VandeWoude. Parasites as conservation tools. Conservation Biology, 36(1): e13719:1–e13719:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gumbs:2023:IMS

[GCD^{+}23] Rikki Gumbs, Abhishek Chaudhary, Barnabas H. Daru, Daniel P. Faith, Félix Forest, Claudia L. Gray, Aida Kowalska, Who-Seung Lee, Roseli Pellens, Sebastian Pipins, Laura J. Pollock, James Rosindell, Rosa A. Scherson, and Nisha R. Owen. Indicators to monitor the status of the tree of life. Conservation Biology, 37(6):e14138:1–e14138:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guizar-Coutino:2022:GEE

[GCJB^{+}22] Alejandro Guizar-Coutiño, Julia P. G. Jones, Andrew Balmford, Rachel Carmenta, and David A. Coomes. A global evaluation of the effectiveness of voluntary REDD+ projects at reducing deforestation and degradation in the moist tropics. Conservation Biology, 36(6):e13970:1–e13970:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guo:2021:WEN

[GCS^{+}21] Qinfeng Guo, Xiaoyu Cen, Ruiyan Song, Michael L. McKinney, and Deli Wang. Worldwide effects of non-native species on species–area relationships. Conservation Biology, 35(2):711– 721, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Green:2024:CCC

[GCS^{+}24] Aalayna R. Green, Stotra Chakrabarti, Shweta Shivakumar, Courtney Hughes, Sayan Banerjee, Maureen W. Kinyanjui, Moreangels M. Mbizah, Omar Ohrens, and Abigail R. Thiemkey. Creating constellations of coexistence through connections between people in human–wildlife conflict areas. Conservation Biology, 38(6):e14402:1–e14402:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gilbert:2021:AEU

[GCSZ21] Neil A. Gilbert, John D. J. Clare, Jennifer L. Stenglein, and Benjamin Zuckerberg. Abundance estimation of unmarked animals based on camera-trap data. Conservation Biology, 35(1): 88–100, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gigliotti:2023:RRS

[GFF23] Franco N. Gigliotti, Thomas P. Franzem, and Paige F. B. Ferguson. Rapid, recurring, structured survey versus bioblitz for generating biodiversity data and analysis with a multispecies abundance model. Conservation Biology, 37(2):e13996:1– e13996:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaisberger:2022:TSA

[GFK^{+}22] Hannes Gaisberger, Tobias Fremout, Chris J. Kettle, Barbara Vinceti, Della Kemalasari, Tania Kanchanarak, Evert Thomas, Josep M. Serra-Diaz, Jens-Christian Svenning, Ferry Slik, Wichan Eiadthong, Kandasamy Palanisamy, Gudasalamani Ravikanth, Vilma Bodos, Julia Sang, Rekha R. Warrier, Alison K. S. Wee, Christian Elloran, Lawrence Tolentino Ramos, Matieu Henry, Md. Akhter Hossain, Ida Theilade, Simon Laegaard, K. M. A. Bandara, Dimantha Panduka Weerasinghe, Suchitra Changtragoon, Vivi Yuskianti, Peter Wilkie, Nguyen Hoang Nghia, Stephen Elliott, Greuk Pakkad, Pimonrat Tiansawat, Colin Maycock, Chaloun Bounithiphonh, Rozi Mohamed, M. Nazre, Baktiar Nur Siddiqui, Soon-Leong Lee, Chai-Ting Lee, Nurul Farhanah Zakaria, Ida Hartvig, Lutz Lehmann, Dzaeman B. Dzulkifli David, Jens-Peter Barnekow Lillesø, Chhang Phourin, Zheng Yongqi, Huang Ping, Hugo A. Volkaert, Lars Graudal, Arief Hamidi, So Thea, Sineath Sreng, David Boshier, Enrique Tolentino, Jr., Wickneswari Ratnam, Mu Mu Aung, Michael Galante, Siti Fatimah Md Isa, Nguyen Quoc Dung, Tran Thi Hoa, Tran Chan Le, Md. Danesh Miah, Abdul Lateef Mohd Zuhry, Deepani Alawathugoda, Amelia Azman, Gamini Pushpakumara, Nur Sumedi, Iskandar Z. Siregar, Hong Kyung Nak, Jean Linsky, Megan Barstow, Lian Pin Koh, and Riina Jalonen. Tropical and subtropical Asia’s valued tree species under threat. Conservation Biology, 36(3):e13873:1–e13873:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Galindo-Gonzalez:2024:ANU

[GG24] Jorge Galindo-González. Avoiding novel, unwanted interactions among species to decrease risk of zoonoses. Conservation Biology, 38(3):e14232:1–e14232:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Glentworth:2024:PPR

[GGA24] Joseph Glentworth, Anna Gilchrist, and Rowan Avery. The place for people in rewilding. Conservation Biology, 38(6): e14318:1–e14318:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Giannichi:2020:SDC

[GGB^{+}20a] Marta Lisli Giannichi, Yoni Gavish, Timothy R. Baker, Martin Dallimer, and Guy Ziv. Scale dependency of conservation outcomes in a forest-offsetting scheme. Conservation Biology, 34(1):148–157, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gnanalingam:2020:CSS

[GGB20b] Gaya Gnanalingam, Holly Gaff, and Mark J. Butler IV. Conserving spawning stocks through harvest slot limits and notake protected areas. Conservation Biology, 34(6):1492–1502, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goodchild:2020:ESA

[GGBD20] Christopher G. Goodchild, Kevin Grisham, Jason B. Belden, and Sarah E. DuRant. Effects of sublethal application of Deepwater Horizon oil to bird eggs on embryonic heart and metabolic rate. Conservation Biology, 34(5):1262–1270, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaget:2020:AEN

[GGJ^{+}20] Elie Gaget, Thomas Galewski, Frédéric Jiguet, Anis Guelmami, Christian Perennou, Coralie Beltrame, and Isabelle Le Viol. Antagonistic effect of natural habitat conversion on community adjustment to climate warming in nonbreeding waterbirds. Conservation Biology, 34(4):966–976, August 2020. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gardiner:2025:MPA

[GGLB25] Natasha Blaize Gardiner, Neil Gilbert, Daniela Liggett, and Michael Bode. Measuring the performance of Antarctic Treaty decision-making. Conservation Biology, 39(1):e14349:1– e14349:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gonzalez:2020:MAD

[GGSRF20] Alejandra V. González, Valeria Gómez-Silva, María José Ramírez, and Francisco E. Fontúrbel. Meta-analysis of the differential effects of habitat fragmentation and degradation on plant genetic diversity. Conservation Biology, 34(3):711–720, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Galea:2022:IMI

[GH22] Benjamin Galea and Tatyana Humle. Identifying and mitigating the impacts on primates of transportation and service corridors. Conservation Biology, 36(1):e13836:1–e13836:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gegear:2021:MSA

[GHR21] Robert J. Gegear, Kevin N. Heath, and Elizabeth F. Ryder. Modeling scale up of anthropogenic impacts from individual pollinator behavior to pollination systems. Conservation Biology, 35(5):1519–1529, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guilherme:2023:CBC

[GJC^{+}23] João L. Guilherme, Victoria R. Jones, Inês Catry, Martin Beal, Maria P. Dias, Steffen Oppel, Juliet A. Vickery, Chris M. Hewson, Stuart H. M. Butchart, and Ana S. L. Rodrigues. Connectivity between countries established by landbirds and raptors migrating along the African–Eurasian flyway. Conservation Biology, 37(1):e14002:1–e14002:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaget:2022:PAC

[GJPJ^{+}22] Elie Gaget, Alison Johnston, Diego Pavón-Jordán, Aleksi S. Lehikoinen, Brett K. Sandercock, Alaaeldin Soultan, Luka Božič, Preben Clausen, Koen Devos, Cristi Domsa, Vitor Encarnação, Sándor Faragó, Niamh Fitzgerald, Teresa Frost, Clemence Gaudard, Lívia Gosztonyi, Fredrik Haas, Menno Hornman, Tom Langendoen, Christina Ieronymidou, Leho Luigujõe, Włodzimierz Meissner, Tibor Mikuska, Blas Molina, Zuzana Musilová, Jean-Yves Paquet, Nicky Petkov, Danae Portolou, Jozef Ridzoň, Laimonas Sniauksta, Antra Stīpniece, Norbert Teufelbauer, Johannes Wahl, Marco Zenatello, and Jon E. Brommer. Protected area characteristics that help waterbirds respond to climate warming. Conservation Biology, 36 (4):e13877:1–e13877:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Godfery:2024:SPC

[GKH^{+}24] Tracey Godfery, John Kean, Daniel Hikuroa, Andrew Robinson, and Nari Williams. Shifting paradigms and creating space for Indigenous leadership in biosecurity management and decision-making. Conservation Biology, 38(6):e14399:1– e14399:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gao:2023:TCN

[GLL^{+}23] Yufang Gao, Yuntian Liu, Yu Luo, Duan Biggs, Weiyang Zhao, and Susan G. Clark. Tracking Chinese newspaper coverage of elephant ivory through topic modeling. Conservation Biology, 37(5):e14072:1–e14072:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gomides:2021:ASR

[GMEV^{+}21] Samuel C. Gomides, Talita M. Machado, Jôine C. Evangelista-Vale, Angele Tatiane Martins-Oliveira, João Carlos Pires-Oliveira, Angélica Muller, Lucas Barros da Rosa, Domingos Lucas Santos-Silva, and Pedro V. Eisenlohr. Assessing species reintroduction sites based on future climate suitability for food resources. Conservation Biology, 35(6):1821–1832, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grainger:2020:NTR

[GN20] Matthew J. Grainger and Erlend B. Nilsen. Need for transparent and repeatable conservation frameworks: reply to Child et al. 2019. Conservation Biology, 34(1):282–285, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gladstone:2022:SFG

[GNH^{+}22] Nicholas S. Gladstone, Matthew L. Niemiller, Benjamin Hutchins, Benjamin Schwartz, Alexander Czaja, Michael E. Slay, and Nathan V. Whelan. Subterranean freshwater gastropod biodiversity and conservation in the United States and Mexico. Conservation Biology, 36(1):e13722:1–e13722:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Goklany:2021:RGH

[Gok21] Indur M. Goklany. Reduction in global habitat loss from fossilfuel-dependent increases in cropland productivity. Conservation Biology, 35(3):766–774, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gonzalez:2020:RET

[Gon20] Ezequiel González. Rethinking the ecology of towns and villages. Conservation Biology, 34(5):1315–1316, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gonzalez:2021:CLC

[Gon21] Ezequiel González. Connecting landscapes for conservation in a changing world. Conservation Biology, 35(2):752–753, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gaget:2021:BPA

[GPJJ^{+}21] Elie Gaget, Diego Pavón-Jordán, Alison Johnston, Aleksi Lehikoinen, Wesley M. Hochachka, Brett K. Sandercock, Alaaeldin Soultan, Hichem Azafzaf, Nadjiba Bendjedda, Taulant Bino, Luka Božič, Preben Clausen, Mohamed Dakki, Koen Devos, Cristi Domsa, Vitor Encarnação, Kiraz Erciyas-Yavuz, Sándor Faragó, Teresa Frost, Clemence Gaudard, Lívia Gosztonyi, Fredrik Haas, Menno Hornman, Tom Langendoen, Christina Ieronymidou, Vasiliy A. Kostyushin, Lesley J. Lewis, Svein-Håkon Lorentsen, Leho Luigujõe, Włodzimierz Meissner, Tibor Mikuska, Blas Molina, Zuzana Musilová, Viktor Natykanets, Jean-Yves Paquet, Nicky Petkov, Danae Portolou, Jozef Ridzoň, Samir Sayoud, Marko Šćiban, Laimonas Sniauksta, Antra Stīpniece, Nicolas Strebel, Norbert Teufelbauer, Goran Topić, Danka Uzunova, Andrej Vizi, Johannes Wahl, Marco Zenatello, and Jon E. Brommer. Benefits of protected areas for nonbreeding waterbirds adjusting their distributions under climate warming. Conservation Biology, 35(3):834–845, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gorenflo:2021:LDC

[GR21] L. J. Gorenflo and Suzanne Romaine. Linguistic diversity and conservation opportunities at UNESCO World Heritage Sites in Africa. Conservation Biology, 35(5):1426–1436, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gallagher:2025:ICS

[GRP^{+}25] Rachael Gallagher, Erin Roger, Jasmin Packer, Cameron Slatyer, Jodi Rowley, Will Cornwell, Emilie Ens, Sarah Legge, Colin Simpfendorfer, Ruby Stephens, and Thomas Mesaglio. Incorporating citizen science into IUCN Red List assessments. Conservation Biology, 39(2):e14329:1–e14329:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbs:2023:ECC

[GRS^{+}23] Mitchell Gibbs, Pauline Ross, Elliot Scanes, James Gibbs, Raphaela Rotolo-Ross, and Laura Parker. Extending conservation of coastal and oyster reef restoration for First Nations cultural revitalization. Conservation Biology, 37(6):e14158:1– e14158:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gruber:2020:CPA

[Gru20] Thibaud Gruber. Conserving primates in the Anthropocene. Conservation Biology, 34(4):1057–1059, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gruber:2021:IC

[Gru21] Thibaud Gruber. The importance of context. Conservation Biology, 35(5):1706–1707, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gippet:2023:RSM

[GSB23] Jérôme M. W. Gippet, Zoé Sherpa, and Cleo Bertelsmeier. Reliability of social media data in monitoring the global pet trade in ants. Conservation Biology, 37(3):e13994:1–e13994:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Guerrero:2021:EIF

[GSMW21] A. M. Guerrero, I. Sporne, R. McKenna, and K. A. Wilson. Evaluating institutional fit for the conservation of threatened species. Conservation Biology, 35(5):1437–1450, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gibbons:2022:RIC

[GSS^{+}22] David W. Gibbons, Chris Sandbrook, William J. Sutherland, Rezvin Akter, Richard Bradbury, Steven Broad, Andy Clements, Humphrey Q. P. Crick, Joanna Elliott, Ngawang Gyeltshen, Melanie Heath, Jonathan Hughes, Richard K. B. Jenkins, Alastair H. Jones, Rocio Lopez de la Lama, Nicholas B. W. Macfarlane, Mike Maunder, Ravikash Prasad, Alfredo Romero-Muñoz, Noa Steiner, James Tremlett, Rosie Trevelyan, Savita Vijaykumar, Irushinie Wedage, and Nancy Ockendon. The relative importance of COVID-19 pandemic impacts on biodiversity conservation globally. Conservation Biology, 36(1): e13781:1–e13781:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gonzalez-Trujillo:2024:EPA

[GTAdPA24] Juan David González-Trujillo, Diogo Alagador, Pamela González del Pliego, and Miguel B. Araújo. Exposure of protected areas in Central America to extreme weather events. Conservation Biology, 38(4):e14251:1–e14251:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gomez:2021:CAF

[GTC21] Camila Gómez, Elkin A. Tenorio, and Carlos Daniel Cadena. Change in avian functional fingerprints of a neotropical montane forest over 100 years as an indicator of ecosystem integrity. Conservation Biology, 35(5):1552–1563, October 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gerling:2021:EEF

[GW21] Charlotte Gerling and Frank Wätzold. An economic evaluation framework for land-use-based conservation policy instruments in a changing climate. Conservation Biology, 35(3):824–833, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Griffiths:2024:DDA

[GWB^{+}24] Laura L. Griffiths, Joel Williams, Christina A. Buelow, Vivitskaia J. Tulloch, Mischa P. Turschwell, Max D. Campbell, David Harasti, Rod M. Connolly, and Christopher J. Brown. A data-driven approach to multiple-stressor impact assessment for a marine protected area. Conservation Biology, 38(2): e14177:1–e14177:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gatto:2023:MEC

[GWR23] Christopher R. Gatto, Sean A. Williamson, and Richard D. Reina. Mitigating the effects of climate change on the nests of sea turtles with artificial irrigation. Conservation Biology, 37(1):e14044:1–e14044:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gawel:2024:CPI

[GWR24] Ann Marie Gawel, Dara M. Wald, and Haldre S. Rogers. Community perceptions of invasive species and environmental management in a US island territory. Conservation Biology, 38(6): e14394:1–e14394:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Gallego-Zamorano:2020:CEL

[GZBLS^{+}20] Juan Gallego-Zamorano, Ana Benítez-López, Luca Santini, Jelle P. Hilbers, Mark A. J. Huijbregts, and Aafke M. Schipper. Combined effects of land use and hunting on distributions of tropical mammals. Conservation Biology, 34(5):1271–1280, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Grossmark:2024:IHW

[GZBP24] Yaara Grossmark, Barak Azriali Zohar, Adi Barash, and Michelle E. Portman. Incorporation of human–wildlife interactions in ecosystem-based management to enhance conservation of endangered guitarfish. Conservation Biology, 38(6): e14327:1–e14327:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holden:2024:IMP

[HAE^{+}24] Petra Holden, Molly Anderson, Frank Eckardt, Gina Ziervogel, Samuel L. Jack, Mark G. New, Julian Smit, Martine Visser, and M. Timm Hoffman. Importance of methodological pluralism in deriving counterfactuals for evidence-based conservation. Conservation Biology, 38(3):e14285:1–e14285:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Halley:2022:TVE

[Hal22] John M. Halley. Three visions of environmental apocalypse, or not. Conservation Biology, 36(4):e13923:1–e13923:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hawkins:2024:BGB

[HBB^{+}24] Frank Hawkins, Craig R. Beatty, Thomas M. Brooks, Rebekah Church, Wendy Elliott, Edit Kiss, Nicholas B. W. Macfarlane, Juliette Pugliesi, Aafke M. Schipper, and Maria Walsh. Bottom-up global biodiversity metrics needed for businesses to assess and manage their impact. Conservation Biology, 38 (2):e14183:1–e14183:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hatfield:2020:MAE

[HBJ^{+}20] Jack H. Hatfield, Jos Barlow, Carlos A. Joly, Alexander C. Lees, Celso Henrique de Freitas Parruco, Joseph A. Tobias, C. David L. Orme, and Cristina Banks-Leite. Mediation of area and edge effects in forest fragments by adjacent land use. Conservation Biology, 34(2):395–404, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrison:2020:EET

[HBL20] Michelle Lisa Kiri Harrison and Cristina Banks-Leite. Edge effects on trophic cascades in tropical rainforests. Conservation Biology, 34(4):977–987, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Healy:2023:EMS

[HBY^{+}23] Brian D. Healy, Phaedra Budy, Charles. B. Yackulic, Brendan P. Murphy, Robert C. Schelly, and Mark C. McKinstry. Exploring metapopulation-scale suppression alternatives for a global invader in a river network experiencing climate change. Conservation Biology, 37(1):e13993:1–e13993:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hilton:2022:DPT

[HC22] Mairi Hilton and Carly N. Cook. Defining performance thresholds for effective management of biodiversity within protected areas. Conservation Biology, 36(6):e13963:1–e13963:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hemming:2022:IDS

[HCA^{+}22] Victoria Hemming, Abbey E. Camaclang, Megan S. Adams, Mark Burgman, Katherine Carbeck, Josie Carwardine, Iadine Chadès, Lia Chalifour, Sarah J. Converse, Lindsay N. K. Davidson, Georgia E. Garrard, Riley Finn, Jesse R. Fleri, Jacqueline Huard, Helen J. Mayfield, Eve McDonald Madden, Ilona Naujokaitis-Lewis, Hugh P. Possingham, Libby Rumpff, Michael C. Runge, Daniel Stewart, Vivitskaia J. D. Tulloch, Terry Walshe, and Tara G. Martin. An introduction to decision science for conservation. Conservation Biology, 36(1): e13868:1–e13868:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hamilton:2021:AVB

[HCC^{+}21] Thalassa McMurdo Hamilton, Stefano Canessa, Katie Clark, Pani Gleeson, Fiona Mackenzie, Troy Makan, Gena Moses-Te Kani, Shona Oliver, Kevin A. Parker, and John G. Ewen. Applying a values-based decision process to facilitate comanagement of threatened species in Aotearoa, New Zealand. Conservation Biology, 35(4):1162–1173, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hart:2020:TPC

[HCD^{+}20] Adam G. Hart, Rosie Cooney, Amy Dickman, Darragh Hare, Charles Jonga, Paul K. Johnson, Maxi Pia Louis, Rodgers Lubilo, Dilys Roe, Catherine Semcer, and Keith Somerville. Threats posed to conservation by media misinformation. Conservation Biology, 34(6):1333–1334, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Holmes:2022:MHC

[HCJH^{+}22] George Holmes, Jonathan Carruthers-Jones, Graham Huggan, Eveline R. de Smalen, Katie Ritson, and Pavla Šimková. Mainstreaming the humanities in conservation. Conservation Biology, 36(4):e13824:1–e13824:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huang:2024:ARM

[HCZW24] Mingpan Huang, Yiting Chen, Wenliang Zhou, and Fuwen Wei. Assessing the response of marine fish communities to climate change and fishing. Conservation Biology, 38(6):e14291:1– e14291:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hennig:2022:CPV

[Hen22] Christian Hennig. The controversy over p-values as an illustration of the difficulty of statistics: response to Mayo (2022). Conservation Biology, 36(5):e13987:1–e13987:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [May22].

Hampton:2020:RH

[HFW20] Jordan O. Hampton, Penny M. Fisher, and Bruce Warburton. Reconsidering humaneness. Conservation Biology, 34(5): 1107–1113, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Heck:2024:GDM

[HGAB^{+}24] Nadine Heck, Liza Goldberg, Dominic A. Andradi-Brown, Anthony Campbell, Siddharth Narayan, Gabby N. Ahmadia, and David Lagomasino. Global drivers of mangrove loss in protected areas. Conservation Biology, 38(6):e14293:1–e14293:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hylander:2022:CAB

[HGCK22] Kristoffer Hylander, Caroline Greiser, Ditte M. Christiansen, and Irena A. Koelemeijer. Climate adaptation of biodiversity conservation in managed forest landscapes. Conservation Biology, 36(3):e13847:1–e13847:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hauser:2022:FEC

[HGM^{+}22] Cindy E. Hauser, Katherine M. Giljohann, Michael A. Mc-Carthy, Georgia E. Garrard, Andrew P. Robinson, Nicholas S. G. Williams, and Joslin L. Moore. A field experiment characterizing variable detection rates during plant surveys. Conservation Biology, 36(3):e13888:1–e13888:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hurtado:2023:CWH

[HHB23] Cindy Hurtado, Victoria Hemming, and Cole Burton. Comparing wildlife habitat suitability models based on expert opinion with camera trap detections. Conservation Biology, 37(5): e14113:1–e14113:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hatten:2024:ALI

[HHL^{+}24] Chloe E. R. Hatten, Yok Yok Hadiprakarsa, Jack Y. K. Lam, Janice Mak, Pavel Toropov, and Caroline Dingle. Assessing the legal, illegal, and gray ornamental trade of the critically endangered helmeted hornbill. Conservation Biology, 38(5): e14358:1–e14358:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hinsley:2024:CMI

[HHM24] Amy Hinsley, Alice Hughes, and Jared Margulies. Creating a more inclusive approach to wildlife trade management. Conservation Biology, 38(5):e14360:1–e14360:??, October 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hamm:2024:IEP

[HHMO24] Joseph Hamm, George Holmes, and Julia Martin-Ortega. The importance of equity in payments to encourage coexistence with large mammals. Conservation Biology, 38(3):e14207:1– e14207:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Harrould-Kolieb:2021:ESB

[HK21] Ellycia Harrould-Kolieb. Enhancing synergies between action on ocean acidification and the post-2020 global biodiversity framework. Conservation Biology, 35(2):548–558, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hopcraft:2020:MBL

[HL20] J. Grant C. Hopcraft and M. Karen Laurenson. Markus borner, a life at the conservation front line. Conservation Biology, 34 (3):769–770, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hu:2024:QSB

[HLL^{+}24] Sifan Hu, Zhijian Liang, Dan Liang, Yang Liu, Jia Zhong, Qian Wei, and Tien Ming Lee. Quantifying species biases among multidata sources on illegal wildlife trade and its implications for conservation. Conservation Biology, 38(5):e14351:1– e14351:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

He:2021:CEL

[HLZ^{+}21] Fengzhi He, Simone D. Langhans, Christiane Zarfl, Roland Wanke, Klement Tockner, and Sonja C. Jähnig. Combined effects of life-history traits and human impact on extinction risk of freshwater megafauna. Conservation Biology, 35(2):643–653, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hubschle:2024:NSH

[HM24] Annette Hübschle and Jared Margulies. The need for a socioecological harm reduction approach to reduce illegal wildlife trade. Conservation Biology, 38(5):e14335:1–e14335:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hardison:2023:ESR

[HMC23a] Sean B. Hardison, Karen J. McGlathery, and Max C. N. Castorani. Effects of seagrass restoration on coastal fish abundance and diversity. Conservation Biology, 37(6):e14147:1–e14147:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Howell:2023:MGB

[HMC^{+}23b] Lachlan G. Howell, Peter R. Mawson, Pierre Comizzoli, Ryan R. Witt, Richard Frankham, Simon Clulow, Justine K. O’Brien, John Clulow, Paul Marinari, and John C. Rodger. Modeling genetic benefits and financial costs of integrating biobanking into the conservation breeding of managed marsupials. Conservation Biology, 37(2):e14010:1–e14010:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hamede:2021:DDM

[HMM^{+}21] Rodrigo Hamede, Thomas Madsen, Hamish McCallum, Andrew Storfer, Paul A. Hohenlohe, Hannah Siddle, Jim Kaufman, Mathieu Giraudeau, Menna Jones, Frédéric Thomas, and Beata Ujvari. Darwin, the devil, and the management of transmissible cancers. Conservation Biology, 35(2):748–751, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hoppenreijs:2024:TMS

[HMM^{+}24] Jacqueline H. T. Hoppenreijs, Jeffery Marker, Ronald J. Maliao, Henry H. Hansen, Erika Juhász, Asko Lõhmus, Vassil Y. Altanov, Petra Horká, Annegret Larsen, Birgitta Malm-Renöfält, Kadri Runnel, John J. Piccolo, and Anne E. Magurran. Three major steps toward the conservation of freshwater and riparian biodiversity. Conservation Biology, 38(3): e14226:1–e14226:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hintsanen:2023:TNC

[HMS^{+}23] Leena Hintsanen, Emma-Liina Marjakangas, Andrea Santangeli, Alison Johnston, and Aleksi Lehikoinen. Temperature niche composition change inside and outside protected areas under climate warming. Conservation Biology, 37(5):e14134:1– e14134:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hagen:2024:UVI

[HNC^{+}24] Sarah Hagen, Christoph Nolte, Yuhe Chang, Seth Morgan, Giulio Boccaletti, and Sheila M. W. Reddy. Understanding variation in impacts from private protected areas across regions and protection mechanisms to inform organizational practices. Conservation Biology, 38(2):e14225:1–e14225:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hajjar:2025:NDC

[HOT^{+}25] Reem Hajjar, Johan A. Oldekop, Roberto Toto, Lucas Alencar, Samuel D. Bell, Katie Devenish, Duong T. Khuu, Mariana Hernandez-Montilla, Suhyun Jung, Sandy Nofyanza, and Lok Mani Sapkota. Navigating data challenges in socioeconomic impact assessments of conservation regimes. Conservation Biology, 39(2):e14457:1–e14457:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hunter:2024:EPO

[HOWM24] Sara Bronwen Hunter, Malik Oedin, Julie Weeds, and Fiona Mathews. Exploring the potential for online data sources to enhance species threat mapping through the case study of global bat exploitation. Conservation Biology, 38(4):e14242:1– e14242:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huang:2020:UMT

[HPG20] Ryan Huang, Stuart L. Pimm, and Chandra Giri. Using metapopulation theory for practical conservation of mangrove endemic birds. Conservation Biology, 34(1):266–275, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hintzen:2020:RBC

[HPM^{+}20] Rogier E. Hintzen, Marina Papadopoulou, Ross Mounce, Cristina Banks-Leite, Robert D. Holt, Morena Mills, Andrew T. Knight, Armand M. Leroi, and James Rosindell. Relationship between conservation biology and ecology shown through machine reading of 32,000 articles. Conservation Biology, 34(3):721–732, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huang:2024:GMK

[HQY^{+}24] Guangping Huang, Dunwu Qi, Zhisong Yang, Rong Hou, Wenyu Shi, Fangqing Zhao, Zitian Li, Li Yan, and Fuwen Wei. Gut microbiome as a key monitoring indicator for reintroductions of captive animals. Conservation Biology, 38(1):e14173:1– e14173:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hallman:2021:BBB

[HRCA21] Tyler A. Hallman, W. Douglas Robinson, Jenna R. Curtis, and Edward R. Alverson. Building a better baseline to estimate 160 years of avian population change and create historically informed conservation targets. Conservation Biology, 35(4): 1256–1267, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hanna:2020:ELU

[HRHB20] Dalal E. L. Hanna, Ciara Raudsepp-Hearne, and Elena M. Bennett. Effects of land use, cover, and protection on stream and riparian ecosystem services and biodiversity. Conservation Biology, 34(1):244–255, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hochkirch:2021:SND

[HSG^{+}21] Axel Hochkirch, Michael J. Samways, Justin Gerlach, Monika Böhm, Paul Williams, Pedro Cardoso, Neil Cumberlidge, P. J. Stephenson, Mary B. Seddon, Viola Clausnitzer, Paulo A. V. Borges, Gregory M. Mueller, Paul Pearce-Kelly, Domitilla C. Raimondo, Anja Danielczak, and Klaas-Douwe B. Dijkstra. A strategy for the next decade to address data deficiency in neglected biodiversity. Conservation Biology, 35(2):502–509, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hurst:2020:AIR

[HSHA20] Kristin Hurst, Marc J. Stern, R. Bruce Hull, and Danny Axsom. Addressing identity-related barriers to collaboration for conservation through self-affirmation theory and moral foundations theory. Conservation Biology, 34(3):572–580, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hagerman:2021:SCZ

[HSN^{+}21] Shannon Hagerman, Terre Satterfield, Sara Nawaz, Guillaume Peterson St-Laurent, Robert Kozak, and Robin Gregory. Social comfort zones for transformative conservation decisions in a changing climate. Conservation Biology, 35(6):1932–1943, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hanson:2025:SCP

[HSSM^{+}25] Jeffrey O. Hanson, Richard Schuster, Matthew Strimas-Mackey, Nina Morrell, Brandon P. M. Edwards, Peter Arcese, Joseph R. Bennett, and Hugh P. Possingham. Systematic conservation prioritization with the prioritizr R package. Conservation Biology, 39(1):e14376:1–e14376:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hollarsmith:2022:PIC

[HTC22] Jordan A. Hollarsmith, Thomas W. Therriault, and Isabelle M. Côté. Practical implementation of cumulative-effects management of marine ecosystems in western North America. Conservation Biology, 36(1):e13841:1–e13841:??, February 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Horigue:2025:DMC

[HTM25] Vera Horigue, Arthur Tuda, and Joseph Maina. Design and management considerations for the Kenya–Tanzania marine transboundary conservation area. Conservation Biology, 39(1): e14342:1–e14342:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hahlbeck:2022:CWH

[HTS^{+}22] Nick Hahlbeck, William R. Tinniswood, Matthew R. Sloat, Jordan D. Ortega, Matthew A. Wyatt, Mark E. Hereford, Ben S. Ramirez, David A. Crook, Kara J. Anlauf-Dunn, and Jonathan B. Armstrong. Contribution of warm habitat to cold-water fisheries. Conservation Biology, 36(3):e13857:1– e13857:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hiller:2025:SRI

[HtSR25] Christina Hiller and Michael ’t Sas-Rolfes. Systematic review of the impact of restrictive wildlife trade measures on conservation of iconic species in southern Africa. Conservation Biology, 39(1):e14262:1–e14262:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hulme:2022:IGI

[Hul22] Philip E. Hulme. Importance of greater interdisciplinarity and geographic scope when tackling the driving forces behind biological invasions. Conservation Biology, 36(2):e13817:1– e13817:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huynh:2023:UGT

[Huy23] Alex Van Huynh. Use of Google Trends categories in conservation culturomics. Conservation Biology, 37(4):e14103:1– e14103:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Horton:2021:NTE

[HVA^{+}21] Kyle G. Horton, Benjamin M. Van Doren, Heidi J. Albers, Andrew Farnsworth, and Daniel Sheldon. Near-term ecological forecasting for dynamic aeroconservation of migratory birds. Conservation Biology, 35(6):1777–1786, December 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hanson:2021:ESG

[HVVA^{+}21] Jeffrey O. Hanson, Ana Veríssimo, Guillermo Velo-Antón, Adam Marques, Miguel Camacho-Sanchez, Íñigo Martínez-Solano, Helena Gonçalves, Fernando Sequeira, Hugh P. Possingham, and Silvia B. Carvalho. Evaluating surrogates of genetic diversity for conservation planning. Conservation Biology, 35(2):634–642, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hinsley:2022:UWC

[HWG^{+}22] Amy Hinsley, Anita Kar Yan Wan, David Garshelis, Michael Hoffmann, Sifan Hu, Tien Ming Lee, Keila Meginnis, Brendan Moyle, Yingjie Qiu, Xiangdong Ruan, and E. J. Milner-Gulland. Understanding why consumers in China switch between wild, farmed, and synthetic bear bile products. Conservation Biology, 36(3):e13895:1–e13895:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hallam:2020:MIS

[HWK^{+}20] Christopher D. Hallam, Brendan A. Wintle, Heini Kujala, Amy L. Whitehead, and Emily Nicholson. Measuring impacts on species with models and metrics of varying ecological and computational complexity. Conservation Biology, 34(6):1512– 1524, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hilton:2022:LOD

[HWLC22] Mairi Hilton, Jessica C. Walsh, Erin Liddell, and Carly N. Cook. Lessons from other disciplines for setting management thresholds for biodiversity conservation. Conservation Biology, 36(1):e13865:1–e13865:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hill:2021:GPA

[HWP^{+}21] Narelle K. Hill, Bradley K. Woodworth, Stuart R. Phinn, Nicholas J. Murray, and Richard A. Fuller. Global protectedarea coverage and human pressure on tidal flats. Conservation Biology, 35(3):933–943, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huang:2023:PGT

[HXW^{+}23] Qiongyu Huang, Jin Xu, Jesse Pan Wong, Volker C. Radeloff, and Melissa Songer. Prioritizing global tall forests toward the 30×30 goals. Conservation Biology, 37(6):e14135:1–e14135:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Hyde:2024:WLE

[Hyd24] B. V. E. Hyde. Why the lack of environmental progress. Conservation Biology, 38(1):e14210:1–e14210:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Huang:2021:UHC

[HZW^{+}21] Kang Huang, He Zhang, Chengliang Wang, Rong Hou, Pei Zhang, Gang He, Songtao Guo, Shiyi Tang, Baoguo Li, Charles Oxnard, and Ruliang Pan. Use of historical and contemporary distribution of mammals in China to inform conservation. Conservation Biology, 35(6):1787–1796, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ibbett:2020:CPT

[IB20] Harriet Ibbett and Stephanie Brittain. Conservation publications and their provisions to protect research participants. Conservation Biology, 34(1):80–92, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Iglesias-Carrasco:2023:GEF

[ICMO23] Maider Iglesias-Carrasco, Iliana Medina, and Terry J. Ord. Global effects of forest modification on herpetofauna communities. Conservation Biology, 37(1):e13998:1–e13998:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ibbett:2022:EVS

[IDD^{+}22] Harriet Ibbett, Leejiah Dorward, Asri A. Dwiyahreni, Julia P. G. Jones, Joseph Kaduma, Edward M. Kohi, Jesca Mchomvu, Karlina Prayitno, Humairah Sabiladiyni, Stephen Sankeni, Andie Wijaya Saputra, Jatna Supriatna, and Freya A. V. St John. Experimental validation of specialized questioning techniques in conservation. Conservation Biology, 36(5): e13908:1–e13908:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ibbett:2025:ICA

[IDJ^{+}25] Harriet Ibbett, Leejiah Dorward, Julia P. G. Jones, Edward M. Kohi, Asri A. Dwiyahreni, Stephen Sankeni, Karlina Prayitno, Jesca Mchomvu, Joseph Kaduma, Andie Wijaya Saputra, Ika Yuni Agustin, Tyassanti Tryswidiarini, Rose Mawenya, Jatna Supriatna, and Freya A. V. St John. Improving compliance around protected areas through fair administration of rules. Conservation Biology, 39(1):e14332:1–e14332:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Itescu:2024:ACV

[IJ24] Yuval Itescu and Jonathan M. Jeschke. Assessing the conservation value of cemeteries to urban biota worldwide. Conservation Biology, 38(6):e14322:1–e14322:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Indigo:2021:ELA

[IJK^{+}21] Naomi L. Indigo, Chris J. Jolly, Ella Kelly, James Smith, Jonathan K. Webb, and Ben L. Phillips. Effects of learning and adaptation on population viability. Conservation Biology, 35(4):1245–1255, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Imboma:2022:GSA

[Imb22] Titus S. Imboma. A great service to African conservation science. Conservation Biology, 36(1):e13879:1–e13879:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ishiyama:2021:GDI

[IMI^{+}21] Nobuo Ishiyama, Kazuki Miura, Takahiro Inoue, Masanao Sueyoshi, and Futoshi Nakamura. Geology-dependent impacts of forest conversion on stream fish diversity. Conservation Biology, 35(3):884–896, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ioannidou:2023:BLS

[INO23] Christina T. Ioannidou, Thomas M. Neeson, and Jesse R. O’Hanley. Boosting large-scale river connectivity restoration by planning for the presence of unrecorded barriers. Conservation Biology, 37(3):e14093:1–e14093:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ionides:2022:SMP

[IR22] Edward Ionides and Ya’acov Ritov. The scientific method and p-values: response to Mayo (2022). Conservation Biology, 36(5):e13984:1–e13984:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [May22].

Izquierdo-Tort:2024:ELP

[ITAAG^{+}24] Santiago Izquierdo-Tort, Andrea Alatorre, Paulina Arroyo-Gerala, Elizabeth Shapiro-Garza, Julia Naime, and Jérôme Dupras. Exploring local perceptions and drivers of engagement in biodiversity monitoring among participants in payments for ecosystem services schemes in southeastern Mexico. Conservation Biology, 38(6):e14282:1–e14282:??, December 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jellesmark:2021:CAM

[JAB^{+}21a] Sean Jellesmark, Malcolm Ausden, Tim M. Blackburn, Richard D. Gregory, Mike Hoffmann, Dario Massimino, Louise McRae, and Piero Visconti. A counterfactual approach to measure the impact of wet grassland conservation on U.K. breeding bird populations. Conservation Biology, 35(5):1575–1585, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jucker:2021:SDP

[JAB^{+}21b] Tommaso Jucker, Tatsuya Amano, Alexandra Bell, Emma E. Garnett, Jan Laurens Geffert, Miriam K. Guth, Andrew Hacket-Pain, Sarah H. Luke, Hannah S. Mumby, Matheus Nunes, Tim Rademacher, David C. Rose, Judith Schleicher, Benno I. Simmons, Aiora Zabala, and Nibedita Mukherjee. Steps to diversify priority-setting research in conservation: reflections on de gracia 2021. Conservation Biology, 35(4):1324– 1326, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jachowski:2023:SCO

[JBB23] David S. Jachowski, Robert Baldwin, and Andy J. Boyce. Shifting the conservation Overton window. Conservation Biology, 37(2):e14080:1–e14080:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jaric:2021:ICI

[JBC^{+}21] Ivan Jarić, Céline Bellard, Ricardo A. Correia, Franck Courchamp, Karel Douda, Franz Essl, Jonathan M. Jeschke, Gregor Kalinkat, Lukáš Kalous, Robert J. Lennox, Ana Novoa, Raphaël Proulx, Petr Pyšek, Andrea Soriano-Redondo, Allan T. Souza, Reut Vardi, Diogo Veríssimo, and Uri Roll. Invasion culturomics and iEcology. Conservation Biology, 35 (2):447–451, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2022:CAB

[JBE^{+}22a] Gavin M. Jones, Berry Brosi, Jason M. Evans, Isabel G. W. Gottlieb, Xingwen Loy, Mauricio M. Núñez-Regueiro, Holly K. Ober, Elizabeth Pienaar, Rajeev Pillay, Kathryn Pisarello, Lora L. Smith, and Robert J. Fletcher, Jr. Conserving alpha and beta diversity in wood-production landscapes. Conservation Biology, 36(3):e13872:1–e13872:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2022:QUA

[JBE^{+}22b] Julia Patricia Gordon Jones, Megan Barnes, Johanna Eklund, Paul J. Ferraro, Jonas Geldmann, Johan A. Oldekop, and Judith Schleicher. Quantifying uncertainty about how interventions are assigned would improve impact evaluation in conservation: reply to Rasolofoson 2022. Conservation Biology, 36(6):e14007:1–e14007:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [Ras22].

Johnson:2023:ART

[JCD^{+}23] Thomas Frederick Johnson, Richard Cornford, Shawn Dove, Robin Freeman, and Joseph Millard. Achieving a real-time online monitoring system for conservation culturomics. Conservation Biology, 37(4):e14096:1–e14096:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Joshi:2021:CPC

[JDB^{+}21] Poorva Joshi, Neelesh Dahanukar, Shankar Bharade, Vijay Dethe, Smita Dethe, Neha Bhandare, and Milind Watve. Combining payment for crop damages and reward for productivity to address wildlife conflict. Conservation Biology, 35(6):1923– 1931, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jacobsen:2021:ITI

[JDM^{+}21] Kim S. Jacobsen, Amy J. Dickman, David W. Macdonald, Susana Mourato, Paul Johnson, Lovemore Sibanda, and Andrew Loveridge. The importance of tangible and intangible factors in human–carnivore coexistence. Conservation Biology, 35(4): 1233–1244, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jenkins:2022:PPC

[Jen22] Lekelia D. Jenkins. Power, politics, and culture of marine conservation technology in fisheries. Conservation Biology, 36(3): e13855:1–e13855:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2025:RNR

[JGAH25] Jacob Jones, Andrea S. Griffin, Frank W. Agbola, and Matt W. Hayward. Role of national regime ideology for predicting biodiversity outcomes. Conservation Biology, 39(1):e14314:1– e14314:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jorgensen:2021:CSE

[JJ21] Finn Arne Jørgensen and Dolly Jørgensen. Citizen science for environmental citizenship. Conservation Biology, 35(4):1344– 1347, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jerem:2021:TKG

[JM21] Paul Jerem and Fiona Mathews. Trends and knowledge gaps in field research investigating effects of anthropogenic noise. Conservation Biology, 35(1):115–129, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jarrad:2021:ITT

[JMB21] Frith Jarrad, Ellen Main, and Mark Burgman. Increasing transparency through open science badges. Conservation Biology, 35(3):764–765, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jolly:2021:REP

[JP21] Chris J. Jolly and Ben L. Phillips. Rapid evolution in predatorfree conservation havens and its effects on endangered species recovery. Conservation Biology, 35(1):383–385, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jackson:2022:GED

[JPAR^{+}22] Hazel A. Jackson, Lawrence Percival-Alwyn, Camilla Ryan, Mohammed F. Albeshr, Luca Venturi, Hernán E. Morales, Thomas C. Mathers, Jonathan Cocker, Samuel A. Speak, Gonzalo G. Accinelli, Tom Barker, Darren Heavens, Faye Willman, Deborah Dawson, Lauren Ward, Vikash Tatayah, Nicholas Zuël, Richard Young, Lianne Concannon, Harriet Whitford, Bernardo Clavijo, Nancy Bunbury, Kevin M. Tyler, Kevin Ruhomaun, Molly K. Grace, Michael W. Bruford, Carl G. Jones, Simon Tollington, Diana J. Bell, Jim J. Groombridge, Matt Clark, and Cock Van Oosterhout. Genomic erosion in a demographically recovered bird species during conservation rescue. Conservation Biology, 36(4):e13918:1–e13918:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jones:2021:BMT

[JPK^{+}21] Sorrel Jones, Sarah Papworth, Aidan M. Keane, Juliet Vickery, and Freya A. V. St John. The bean method as a tool to measure sensitive behavior. Conservation Biology, 35(2):722–732, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jiren:2021:PSP

[JRKF21] Tolera Senbeto Jiren, Maraja Riechers, Ruth Kansky, and Joern Fischer. Participatory scenario planning to facilitate human–wildlife coexistence. Conservation Biology, 35(6):1957– 1965, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jolly:2025:IHW

[JS25] Helina Jolly and Amanda Stronza. Insights on human–wildlife coexistence from social science and Indigenous and traditional knowledge. Conservation Biology, 39(2):e14460:1–e14460:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jordan:2020:AII

[JSA^{+}20] Neil R. Jordan, Bradley P. Smith, Robert G. Appleby, Lily M. van Eeden, and Hugh S. Webster. Addressing inequality and intolerance in human–wildlife coexistence. Conservation Biology, 34(4):803–810, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

James:2021:BGD

[JSGJ^{+}21] Tamora D. James, Roberto Salguero-Gómez, Owen R. Jones, Dylan Z. Childs, and Andrew P. Beckerman. Bridging gaps in demographic analysis with phylogenetic imputation. Conservation Biology, 35(4):1210–1221, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jolly:2022:IIH

[JSKR22] Helina Jolly, Terre Satterfield, Milind Kandlikar, and Suma T R. Indigenous insights on human–wildlife coexistence in southern India. Conservation Biology, 36(6):e13981:1–e13981:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jezequel:2020:FFD

[JTD^{+}20] Céline Jézéquel, Pablo A. Tedesco, William Darwall, Murilo S. Dias, Renata G. Frederico, Max Hidalgo, Bernard Hugueny, Javier Maldonado-Ocampo, Koen Martens, Hernan Ortega, Gislene Torrente-Vilara, Jansen Zuanon, and Thierry Oberdorff. Freshwater fish diversity hotspots for conservation priorities in the Amazon Basin. Conservation Biology, 34(4):956– 965, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Jiang:2023:ECC

[JZLP^{+}23] Dechun Jiang, Xumao Zhao, Jordi López-Pujol, Zhiqiang Wang, Yanhua Qu, Yanming Zhang, Tongzuo Zhang, Dayong Li, Ke Jiang, Bin Wang, Chaochao Yan, and Jia-Tang Li. Effects of climate change and anthropogenic activity on ranges of vertebrate species endemic to the Qinghai–Tibet Plateau over 40 years. Conservation Biology, 37(4):e14069:1–e14069:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kyle:2022:CSS

[KAD^{+}22] Kathleen E. Kyle, Michael C. Allen, Jeffrey Dragon, John F. Bunnell, Howard K. Reinert, Robert Zappalorti, Benjamin D. Jaffe, Jordan C. Angle, and Julie L. Lockwood. Combining surface and soil environmental DNA with artificial cover objects to improve terrestrial reptile survey detection. Conservation Biology, 36(6):e13939:1–e13939:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karlsson:2021:EBO

[KB21] Mikael Karlsson and Karin Edvardsson Björnberg. Ethics and biodiversity offsetting. Conservation Biology, 35(2):578–586, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Keith:2023:EIB

[KBB^{+}23] David A. Keith, Doug H. Benson, Ian R. C. Baird, Laura Watts, Christopher C. Simpson, Martin Krogh, Sarsha Gorissen, Jose R. Ferrer-Paris, and Tanya J. Mason. Effects of interactions between anthropogenic stressors and recurring perturbations on ecosystem resilience and collapse. Conservation Biology, 37(1):e13995:1–e13995:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kumschick:2024:UIE

[KBB^{+}24] Sabrina Kumschick, Sandro Bertolino, Tim M. Blackburn, Giuseppe Brundu, Katie E. Costello, Maarten de Groot, Thomas Evans, Belinda Gallardo, Piero Genovesi, Tanushri Govender, Jonathan M. Jeschke, Katharina Lapin, John Measey, Ana Novoa, Ana L. Nunes, Anna F. Probert, Petr Pyšek, Cristina Preda, Wolfgang Rabitsch, Helen E. Roy, Kevin G. Smith, Elena Tricarico, Montserrat Vilà, Giovanni Vimercati, and Sven Bacher. Using the IUCN Environmental Impact Classification for Alien Taxa to inform decision-making. Conservation Biology, 38(2):e14214:1–e14214:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kockel:2020:EAS

[KBCD20] Alessia Kockel, Natalie C. Ban, Maycira Costa, and Philip Dearden. Evaluating approaches for scaling-up communitybased marine-protected areas into socially equitable and ecologically representative networks. Conservation Biology, 34(1): 137–147, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kendall:2023:VCT

[KBL^{+}23] Corinne J. Kendall, Claire Bracebridge, Emily C. Lynch, Msafiri Mgumba, Ara Monadjem, Aaron Nicholas, and Adam Kane. Value of combining transect counts and telemetry data to determine short-term population trends in a globally threatened species. Conservation Biology, 37(6):e14146:1–e14146:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kadykalo:2021:BRP

[KBM^{+}21] Andrew N. Kadykalo, Rachel T. Buxton, Peter Morrison, Christine M. Anderson, Holly Bickerton, Charles M. Francis, Adam C. Smith, and Lenore Fahrig. Bridging research and practice in conservation. Conservation Biology, 35(6):1725– 1737, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelman:2023:FAP

[KBW^{+}23] Candice Carr Kelman, Chris J. Barton, Kyle Whitman, Simon Lhoest, Derrick M. Anderson, and Leah R. Gerber. Five approaches to producing actionable science in conservation. Conservation Biology, 37(2):e14039:1–e14039:??, April 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kemppinen:2020:GRB

[KCH^{+}20] Krista M. S. Kemppinen, Pamela M. Collins, David G. Hole, Christopher Wolf, William J. Ripple, and Leah R. Gerber.

Global reforestation and biodiversity conservation. Conservation Biology, 34(5):1221–1228, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kelly:2020:EII

[KCS20] Emily L. A. Kelly, Abigail L. Cannon, and Jennifer E. Smith. Environmental impacts and implications of tropical carrageenophyte seaweed farming. Conservation Biology, 34 (2):326–337, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kimmel:2022:IMS

[KCT22] Kaitlin Kimmel, Michael Clark, and David Tilman. Impact of multiple small and persistent threats on extinction risk. Conservation Biology, 36(5):e13901:1–e13901:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kusmanoff:2020:FLG

[KFG^{+}20] Alexander M. Kusmanoff, Fiona Fidler, Ascelin Gordon, Georgia E. Garrard, and Sarah A. Bekessy. Five lessons to guide more effective biodiversity conservation message framing. Conservation Biology, 34(5):1131–1141, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Knapp:2021:VPE

[KFN^{+}21] Wesley M. Knapp, Anne Frances, Reed Noss, Robert F. C. Naczi, Alan Weakley, George D. Gann, Bruce G. Baldwin, James Miller, Patrick McIntyre, Brent D. Mishler, Gerry Moore, Richard G. Olmstead, Anna Strong, Kathryn Kennedy, Bonnie Heidel, and Daniel Gluesenkamp. Vascular plant extinction in the continental United States and Canada. Conservation Biology, 35(1):360–368, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kaisin:2021:MAA

[KFP^{+}21] Olivier Kaisin, Lisieux Fuzessy, Pascal Poncin, Fany Brotcorne, and Laurence Culot. A meta-analysis of anthropogenic impacts on physiological stress in wild primates. Conservation Biology, 35(1):101–114, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Katayama:2024:EHD

[KFU^{+}24] Naoki Katayama, Taku Fujita, Mutsuyuki Ueta, Federico Morelli, and Tatsuya Amano. Effects of human depopulation and warming climate on bird populations in Japan. Conservation Biology, 38(2):e14175:1–e14175:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See erratum [Ano24f].

Keith:2024:CIR

[KGK^{+}24] David A. Keith, Somaya Magdy M. Ghoraba, Eric Kaly, Kendall R. Jones, Ané Oosthuizen, David Obura, Hugo M. Costa, Fahiema Daniels, Eleutério Duarte, Hedley Grantham, Mishal Gudka, Juliet Norman, Lynne J. Shannon, Andrew Skowno, and José R. Ferrer-Paris. Contributions of the IUCN Red List of Ecosystems to risk-based design and management of protected and conserved areas in Africa. Conservation Biology, 38(3):e14169:1–e14169:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Khorozyan:2022:DPR

[Kho22] Igor Khorozyan. Defining practical and robust study designs for interventions targeted at terrestrial mammalian predators. Conservation Biology, 36(2):e13805:1–e13805:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kophamel:2022:IHA

[KIA^{+}22] Sara Kophamel, Björn Illing, Ellen Ariel, Morgan Difalco, Lee F. Skerratt, Mark Hamann, Leigh C. Ward, Diana Méndez, and Suzanne L. Munns. Importance of health assessments for conservation in noncaptive wildlife. Conservation Biology, 36(1):e13724:1–e13724:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kinsley:2024:MRA

[KKE^{+}24] Amy C. Kinsley, Szu-Yu Zoe Kao, Eva A. Enns, Luis E. Escobar, Huijie Qiao, Nicholas Snellgrove, Ulirich Muellner, Petra Muellner, Ranjan Muthukrishnan, Meggan E. Craft, Daniel J. Larkin, and Nicholas B. D. Phelps. Modeling the risk of aquatic species invasion spread through boater movements and river connections. Conservation Biology, 38(6):e14260:1–e14260:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Konig:2020:HWC

[KKKS^{+}20] Hannes J. König, Christian Kiffner, Stephanie Kramer-Schadt, Christine Fürst, Oliver Keuling, and Adam T. Ford. Human– wildlife coexistence in a changing world. Conservation Biology, 34(4):786–794, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kyle:2020:TAP

[KLVW20] Gerard Kyle, Adam Landon, Jerry Vaske, and Kenneth Wallen. Tools for assessing the psychometric adequacy of latent variables in conservation research. Conservation Biology, 34(6): 1353–1363, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Karuno:2023:ICC

[KMC^{+}23] Alex Plimo Karuno, Xue Mi, Youhua Chen, Da-Hu Zou, Wei Gao, Bao-Lin Zhang, Wei Xu, Jie-Qiong Jin, Wen-Jing Shen, Song Huang, Wei-Wei Zhou, and Jing Che. Impacts of climate change on herpetofauna diversity in the Qinghai-Tibetan Plateau. Conservation Biology, 37(6):e14155:1–e14155:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Khalatbari:2021:RCS

[KOO^{+}21] Leili Khalatbari, Helen M. K. O’Neill, Stéphane Ostrowski, Gholam Hosein Yusefi, Arash Ghoddousi, Hamed Abolghasemi, Christine Breitenmoser-Würsten, Urs Breitenmoser, José Carlos Brito, and Sarah M. Durant. Risks to conservation of species in the wild from promoting ex situ management: response to Farhadinia et al. 2020. Conservation Biology, 35(4): 1327–1330, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [FJZ^{+}20, FJZ^{+}21].

Kuehne:2022:BPA

[KRB^{+}22] Lauren M. Kuehne, Robert J. Rolls, Kate J. Brandis, Kai Chen, Kevin M. Fraley, Lindsey K. Frost, Susie S. Ho, Erin H. Kunisch, Simone D. Langhans, Carri J. LeRoy, Gregory Mc-Donald, Paul J. McInerney, Katherine R. O’Brien, and Angela L. Strecker. Benefits of permanent adoption of virtual conferences for conservation science. Conservation Biology, 36 (4):e13884:1–e13884:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kristensen:2020:ERD

[KSC^{+}20] Nadiah P. Kristensen, Wei Wei Seah, Kwek Yan Chong, Yi Shuen Yeoh, Tak Fung, Laura M. Berman, Hui Zhen Tan, and Ryan A. Chisholm. Extinction rate of discovered and undiscovered plants in Singapore. Conservation Biology, 34(5): 1229–1240, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ke:2024:EAN

[KSF^{+}24] Alison Ke, Rahel Sollmann, Luke Frishkoff, Alejandra Echeverri, Jim Zook, and Daniel S. Karp. Effects of agriculture and nature reserves on avian behavior in northwestern Costa Rica. Conservation Biology, 38(4):e14241:1–e14241:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kinnebrew:2021:AIM

[KSFP^{+}21] Eva Kinnebrew, Elizabeth Shoffner, Aldo Farah-Pérez, Megan Mills-Novoa, and Katherine Siegel. Approaches to interdisciplinary mixed methods research in land-change science and environmental management. Conservation Biology, 35(1):130– 141, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kozak:2023:ICL

[KSZ^{+}23] Daniel Kozák, Marek Svitok, Veronika Zemlerová, Martin Mikoláš, Thibault Lachat, Laurent Larrieu, Yoan Paillet, Arne Buechling, Radek Bače, William S. Keeton, Lucie Vítková, Krešimir Begovič, Vojtěch Čada, Martin Dušátko, Matej Ferenčík, Michal Frankovič, Rhiannon Gloor, Jeňýk Hofmeister, Pavel Janda, Ondrej Kameniar, Tomáš Kníř, Linda Majdanová, Marek Mejstřík, Jakob Pavlin, Dheeraj Ralhan, Ruffy Rodrigo, Catalin-Constantin Roibu, Michal Synek, Ondřej Vostarek, and Miroslav Svoboda. Importance of conserving large and old trees to continuity of tree-related microhabitats. Conservation Biology, 37(3):e14066:1–e14066:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kuempel:2022:IMO

[KTG^{+}22] Caitlin D. Kuempel, Vivitskaia J. D. Tulloch, Alyssa L. Giffin, B. Alexander Simmons, Valerie Hagger, Carol Phua, and Ove Hoegh-Guldberg. Identifying management opportunities to combat climate, land, and marine threats across less climate exposed coral reefs. Conservation Biology, 36(3):e13856:1– e13856:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Katzner:2022:OSI

[KTH^{+}22] Todd Katzner, Eve Thomason, Karrin Huhmann, Tara Conkling, Camille Concepcion, Vince Slabe, and Sharon Poessel. Open-source intelligence for conservation biology. Conservation Biology, 36(6):e13988:1–e13988:??, December 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kusmanoff:2020:MFT

[Kus20] Alexander M. Kusmanoff. Message framing in the time of the precautionary principle: response to Martell and Rodewald 2019. Conservation Biology, 34(5):1061–1062, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kutszegi:2020:WGI

[Kut20] Gergely Kutszegi. Wheels for genus identification and presenting diversity in kingdom fungi. Conservation Biology, 34 (3):773–775, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Koot:2025:IEV

[KvZS^{+}25] Stasja Koot, Nowella Anyango van Zwieten, Sian Sullivan, Wolfram Dressler, Marja Spierenburg, Lisa Trogisch, Esther Marijnen, Robert Fletcher, Inaya Rakhmani, Suraya Abdulwahab Afiff, Tor A. Benjaminsen, Sarah Milne, Hanne Svarstad, Bram Büscher, Anwesha Dutta, Celia Lowe, and Nitin D. Rai. Intimidation as epistemological violence against social science conservation research. Conservation Biology, 39(2):e14454:1– e14454:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kong:2021:CUD

[KWY^{+}21] Fanhua Kong, Ding Wang, Haiwei Yin, Iryna Dronova, Fan Fei, Jiayu Chen, Yingxia Pu, and Manchun Li. Coupling urban 3-D information and circuit theory to advance the development of urban ecological networks. Conservation Biology, 35(4):1140– 1150, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ke:2025:PCP

[KxGG^{+}25] Can Ke, Li xin Gong, Yang Geng, Zhi qiang Wang, Wen jun Zhang, Jiang Feng, and Ting lei Jiang. Patterns and correlates of potential range shifts of bat species in China in the context of climate change. Conservation Biology, 39(1):e14310:1– e14310:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Kramer:2023:BHD

[KZM23] Jean M. Freitag Kramer, Victor P. Zwiener, and Sandra Cristina Müller. Biotic homogenization and differentiation of plant communities in tropical and subtropical forests. Conservation Biology, 37(1):e14025:1–e14025:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lau:2020:TLG

[Lau20] Jacqueline D. Lau. Three lessons for gender equity in biodiversity conservation. Conservation Biology, 34(6):1589–1591, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Laverty:2022:IEA

[LB22] Theresa M. Laverty and Joel Berger. Indirect effects of African megaherbivore conservation on bat diversity in the world’s oldest desert. Conservation Biology, 36(2):e13780:1–e13780:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lehnardt:2024:EWT

[LBBT24] Yael Lehnardt, Jesse R. Barber, and Oded Berger-Tal. Effects of wind turbine noise on songbird behavior during nonbreeding season. Conservation Biology, 38(2):e14188:1–e14188:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Larsen:2024:FPI

[LC24] Peter Bille Larsen and Chantaly Chanthavisouk. Free, prior, and informed consent, local officials, and changing biodiversity governance in Hin Nam No, Laos. Conservation Biology, 38(6): e14388:1–e14388:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lim:2024:LUC

[LCEM24] Felix K. S. Lim, Luis Roman Carrasco, David P. Edwards, and Jolian McHardy. Land-use change from market responses to oil palm intensification in Indonesia. Conservation Biology, 38(1): e14149:1–e14149:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lucas:2024:UCE

[LDC^{+}24] Pablo Miguel Lucas, Moreno Di Marco, Victor Cazalis, Jennifer Luedtke, Kelsey Neam, Mary H. Brown, Penny F. Langhammer, Giordano Mancini, and Luca Santini. Using comparative extinction risk analysis to prioritize the IUCN Red List reassessments of amphibians. Conservation Biology, 38(6): e14316:1–e14316:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lumbierres:2022:THC

[LDD^{+}22] Maria Lumbierres, Prabhat Raj Dahal, Moreno Di Marco, Stuart H. M. Butchart, Paul F. Donald, and Carlo Rondinini. Translating habitat class to land cover to map area of habitat of terrestrial vertebrates. Conservation Biology, 36(4):e13851:1– e13851:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lapola:2020:CCV

[LdSB^{+}20] David M. Lapola, José Maria C. da Silva, Diego R. Braga, Larissa Carpigiani, Fernanda Ogawa, Roger R. Torres, Luis C. F. Barbosa, Jean P. H. B. Ometto, and Carlos A. Joly. A climate-change vulnerability and adaptation assessment for Brazil’s protected areas. Conservation Biology, 34(2):427–437, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lear:2024:ECD

[LEF^{+}24] Karissa O. Lear, Brendan C. Ebner, Travis Fazeldean, Rebecca L. Bateman, and David L. Morgan. Effects of coastal development on sawfish movements and the need for marine animal crossing solutions. Conservation Biology, 38(4):e14263:1– e14263:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lengyel:2021:FPH

[Len21] Szabolcs Lengyel. Fire in the past and heat in the future. Conservation Biology, 35(4):1355–1357, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

LoParrino:2025:IAO

[LFD^{+}25] Elia Lo Parrino, Gentile Francesco Ficetola, Michel Devin, Raoul Manenti, and Mattia Falaschi. Integrating adult occurrence and reproduction data to identify conservation measures for amphibians. Conservation Biology, 39(1):e14343:1– e14343:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lawson:2021:DCN

[LFT^{+}21] Abigail J. Lawson, Brian Folt, Anna M. Tucker, Francesca Erickson, and Conor P. McGowan. Decision context as a necessary component of population viability analysis assessments: response to Chaudhary and Oli 2020. Conservation Biology, 35 (5):1683–1685, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [CO20].

Leong:2025:UAS

[LGD^{+}25] Kirsten Mya Leong, Ashley Rochelle Gramza, Jennifer N. Duberstein, Chelsey Bryson, and Angela Amlin. Using applied social science disciplines to implement creative outdoor cat management solutions and avoid the trap of one-size-fits-all policies. Conservation Biology, 39(1):e14321:1–e14321:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Leong:2020:UCC

[LGL20] Kirsten M. Leong, Ashley R. Gramza, and Christopher A. Lepczyk. Understanding conflicting cultural models of outdoor cats to overcome conservation impasse. Conservation Biology, 34 (5):1190–1199, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2021:UCS

[LH21] Juan Li and Qi Hu. Using culturomics and social media data to characterize wildlife consumption. Conservation Biology, 35 (2):452–459, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2022:FDL

[LHB^{+}22] Xueyou Li, Wenqiang Hu, William V. Bleisch, Quan Li, Hongjiao Wang, Wen Lu, Jun Sun, Fuyou Zhang, Bu Ti, and Xuelong Jiang. Functional diversity loss and change in nocturnal behavior of mammals under anthropogenic disturbance. Conservation Biology, 36(3):e13839:1–e13839:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2024:MES

[LHB^{+}24] Cheng Li, Rong Hou, Ziqiang Bao, Wei Wu, Jacob R. Owens, Wenlei Bi, Qiang Xu, XiaoDong Gu, Zuofu Xiang, and Dunwu Qi. Measuring ecosystem services and ecological sensitivity for comprehensive conservation in Giant Panda National Park. Conservation Biology, 38(2):e14215:1–e14215:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lavergne:2022:EFC

[LKA^{+}22] Edouard Lavergne, Manabu Kume, Hyojin Ahn, Yumi Henmi, Yuki Terashima, Feng Ye, Satoshi Kameyama, Yoshiaki Kai, Kohmei Kadowaki, Shiho Kobayashi, Yoh Yamashita, and Akihide Kasai. Effects of forest cover on richness of threatened fish species in Japan. Conservation Biology, 36(3):e13847:1– e13847:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lloyd:2023:PSC

[LKF^{+}23] Natasha A. Lloyd, Laura M. Keating, Alyssa J. Friesen, Dylan M. Cole, Jana M. McPherson, H. Resit Akçakaya, and Axel Moehrenschlager. Prioritizing species conservation programs based on IUCN Green Status and estimates of cost-sharing potential. Conservation Biology, 37(3):e14051:1–e14051:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lacher:2021:SRE

[LKL^{+}21] Thomas E. Lacher, Jr., Rosalind Kennerley, Barney Long, Shelby McCay, Nicolette S. Roach, Samuel T. Turvey, and Richard P. Young. Support for rodent ecology and conservation to advance zoonotic disease research. Conservation Biology, 35(4):1061–1062, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lin:2024:OER

[LLYK24] Li Lin, Yang Liu, Yang Yan, and Bin Kang. Optimizing efficiency and resilience of no-take marine protected areas for fish conservation under climate change along the coastlines of China Seas. Conservation Biology, 38(2):e14174:1–e14174:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Letessier:2023:CEI

[LMG^{+}23] Tom B. Letessier, Laura Mannocci, Brittney Goodwin, Clare Embling, Asha de Vos, R. Charles Anderson, Simon N. Ingram, Andy Rogan, and Samuel T. Turvey. Contrasting ecological information content in whaling archives with modern cetacean surveys for conservation planning and identification of historical distribution changes. Conservation Biology, 37 (3):e14043:1–e14043:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liang:2024:ASM

[LMY^{+}24] Dan Liang, Tong Mu, Ziyou Yang, Xingli Giam, Yudi Wang, Jing Li, Shangxiao Cai, Xuelian Zhang, Yixiao Wang, Yang Liu, and David S. Wilcove. Assessing shorebird mortalities due to razor clam aquaculture at key migratory stopover sites in southeastern China. Conservation Biology, 38(2):e14185:1– e14185:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lee:2021:ECT

[LNDM21] Calvin K. F. Lee, Emily Nicholson, Clare Duncan, and Nicholas J. Murray. Estimating changes and trends in ecosystem extent with dense time-series satellite remote sensing. Conservation Biology, 35(1):325–335, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Labao:2020:UDB

[LNYY20] Alfonso B. Labao, Prospero C. Naval, Jr., David Leonides T. Yap, and Helen T. Yap. Using deep-belief networks to understand propensity for livelihood change in a rural coastal community to further conservation. Conservation Biology, 34(4): 1008–1016, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lambertucci:2021:NGC

[LS21] Sergio A. Lambertucci and Karina L. Speziale. Need for global conservation assessments and frameworks to include airspace habitat. Conservation Biology, 35(4):1341–1343, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lynn:2020:MSE

[LSÁH^{+}20] William S. Lynn, Francisco J. Santiago-Ávila, John Hadidian, Arian D. Wallach, and Joann Lindenmayer. Misunderstandings of science and ethics in the moral panic over cats: reply to Crespin et al. 2020. Conservation Biology, 34(4):1038–1040, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [CMAS20].

Lenda:2020:EAM

[LSM^{+}20] Magdalena Lenda, Pjoter Skórka, Błażej Mazur, William Sutherland, Piotr Tryjanowski, Dawid Moroń, Erik Meijaard, Hugh P. Possingham, and Kerrie A. Wilson. Effects of amusing memes on concern for unappealing species. Conservation Biology, 34(5):1200–1209, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Linklater:2024:UAI

[LSM24] Wayne L. Linklater, Jamie Steer, and David Munro. An update on allegations of invasive species denialism. Conservation Biology, 38(1):e14223:1–e14223:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lum:2021:QRP

[LTH^{+}21] Deon Lum, Pablo A. Tedesco, Bernard Hugueny, Xingli Giam, and Ryan A. Chisholm. Quantifying the relative performance of two undetected-extinction models. Conservation Biology, 35(1):239–248, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lischka:2020:PDR

[LTJ^{+}20] Stacy A. Lischka, Tara L. Teel, Heather E. Johnson, Courtney Larson, Stewart Breck, and Kevin Crooks. Psychological drivers of risk-reducing behaviors to limit human–wildlife conflict. Conservation Biology, 34(6):1383–1392, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lennox:2020:SAM

[LVT^{+}20] Robert J. Lennox, Diogo Veríssimo, William M. Twardek, Colin R. Davis, and Ivan Jarić. Sentiment analysis as a measure of conservation culture in scientific literature. Conservation Biology, 34(2):462–471, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liang:2024:PCE

[LWC^{+}24] Jie Liang, Wanting Wang, Qing Cai, Xin Li, Ziqian Zhu, Yeqing Zhai, Xiaodong Li, Xiang Gao, and Yuru Yi. Prioritizing conservation efforts based on future habitat availability and accessibility under climate change. Conservation Biology, 38(3): e14204:1–e14204:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liao:2020:ENN

[LWD^{+}20] Huixuan Liao, Huijie Wang, Qiaohong Dong, Feihong Cheng, Ting Zhou, and Shaolin Peng. Estimating non-native plant richness with a species-accumulation model along roads. Conservation Biology, 34(2):472–481, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lewis:2021:UIA

[LWG21] Rebecca N. Lewis, Leah J. Williams, and R. Tucker Gilman. The uses and implications of avian vocalizations for conservation planning. Conservation Biology, 35(1):50–63, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lapp:2021:ADF

[LWRZ^{+}21] Sam Lapp, Tianhao Wu, Corinne Richards-Zawacki, Jamie Voyles, Keely Michelle Rodriguez, Hila Shamon, and Justin Kitzes. Automated detection of frog calls and choruses by pulse repetition rate. Conservation Biology, 35(5):1659–1668, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Lundgren:2024:PEA

[LWS^{+}24] Erick J. Lundgren, Arian D. Wallach, Jens-Christian Svenning, Martin A. Schlaepfer, Astrid L. A. Andersson, and Daniel Ramp. Preventing extinction in an age of species migration and planetary change. Conservation Biology, 38(6):e14270:1– e14270:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Liu:2022:ASD

[LXZ^{+}22] Jiajia Liu, Shangwen Xia, Di Zeng, Cong Liu, Yingjun Li, Wenjing Yang, Bao Yang, Jian Zhang, Ferry Slik, and David B. Lindenmayer. Age and spatial distribution of the world’s oldest trees. Conservation Biology, 36(4):e13907:1–e13907:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Li:2025:EEP

[LZG^{+}25] Yumei Li, Chao Zhang, Huayao Gao, Ming Lu, and Yonggang Nie. Effects of expansion of protected areas and habitat transformation on spatiotemporal variation in human–wildlife conflict. Conservation Biology, 39(2):e14393:1–e14393:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Majewska:2020:PGS

[MA20] Ania A. Majewska and Sonia Altizer. Planting gardens to support insect pollinators. Conservation Biology, 34(1):15–25, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mair:2023:QMS

[MAB^{+}23] Louise Mair, Eduardo Amorim, Monira Bicalho, Thomas M. Brooks, Vincente Calfo, Renata de T. Capellão, Colin Clubbe, Marianne Evju, Eduardo P. Fernandez, Gláucia C. Ferreira, Frank Hawkins, Randall R. Jiménez, Lucas S. B. Jordão, Magni Olsen Kyrkjeeide, Nicholas B. W. Macfarlane, Bianca C. Mattos, Pablo H. A. de Melo, Lara M. Monteiro, Eimear Nic Lughadha, Nina Pougy, Domitilla C. Raimondo, Trine Hay Setsaas, Xiaoli Shen, Marinez Ferreira de Siqueira, Bernardo B. N. Strassburg, and Philip J. K. McGowan. Quantifying and mapping species threat abatement opportunities to support national target setting. Conservation Biology, 37(1):e14046:1– e14046:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McAfee:2020:ESS

[MAC20] Dominic McAfee, Heidi K. Alleway, and Sean D. Connell. Environmental solutions sparked by environmental history. Conservation Biology, 34(2):386–394, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mayo:2022:SWI

[May22] Deborah G. Mayo. The statistics wars and intellectual conflicts of interest. Conservation Biology, 36(1):e13861:1–e13861:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See responses [IR22, Hen22, SP22, Sta22a].

Moon:2021:DSQ

[MB21] Katie Moon and Nicola K. Browne. Developing shared qualitative models for complex systems. Conservation Biology, 35 (3):1039–1050, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClure:2023:CEH

[MBB^{+}23] Christopher J. W. McClure, Igor Berkunsky, Evan R. Buechley, Leah Dunn, Jeff Johnson, Jennifer McCabe, Steffen Oppel, Brian W. Rolek, Luke J. Sutton, and Rikki Gumbs. Conserving the evolutionary history of birds. Conservation Biology, 37(6):e14141:1–e14141:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meyer:2021:BSC

[MBD^{+}21] Ninon F. V. Meyer, Niko Balkenhol, Trishna Dutta, Maarten Hofman, Jean-Yves Meyer, Euan G. Ritchie, Charlotte Alley, Chad Beranek, Cassandra K. Bugir, Alex Callen, Simon Clulow, Michael V. Cove, Kaya Klop-Toker, Omar R.

Lopez, Michael Mahony, Robert Scanlon, Sandeep Sharma, Elen Shute, Rose Upton, Emy Guilbault, Andrea S. Griffin, Edwin Hernández Pérez, Lachlan G. Howell, John-Paul King, Dean Lenga, Patrick O. Donoghue, and Matt W. Hayward. Beyond species counts for assessing, valuing, and conserving biodiversity: response to Wallach et al. 2019. Conservation Biology, 35(1):369–372, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marucco:2023:MAE

[MBD^{+}23] Francesca Marucco, Maria V. Boiani, Pierre Dupont, Cyril Milleret, Elisa Avanzinelli, Kristine Pilgrim, Michael K. Schwartz, Achaz von Hardenberg, Domenica Serena Perrone, Olivier P. Friard, Arianna Menzano, Francesco Bisi, Umberto Fattori, Michela Tomasella, Sonia Calderola, Sabrina Carolfi, Piero Ferrari, Christian Chioso, Fabrizio Truc, Giulia Bombieri, Luca Pedrotti, Davide Righetti, Pierluigi L. Acutis, Fabio Guglielmo, Heidi C. Hauffe, Chiara Rossi, Romolo Caniglia, Paola Aragno, Valentina La Morgia, Piero Genovesi, and Richard Bischof. A multidisciplinary approach to estimating wolf population size for long-term conservation. Conservation Biology, 37(6):e14132:1–e14132:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mentges:2021:ESS

[MBH^{+}21] Andrea Mentges, Shane A. Blowes, Dorothee Hodapp, Helmut Hillebrand, and Jonathan M. Chase. Effects of site-selection bias on estimates of biodiversity change. Conservation Biology, 35(2):688–698, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Murphy:2022:DFE

[MBI^{+}22] Erin L. Murphy, Miranda Bernard, Gwenllian Iacona, Stephanie B. Borrelle, Megan Barnes, Alexis McGivern, Jorge Emmanuel, and Leah R. Gerber. A decision framework for estimating the cost of marine plastic pollution interventions. Conservation Biology, 36(2):e13827:1–e13827:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Montgomery:2020:PHH

[MBK^{+}20] Robert A. Montgomery, Kendi Borona, Herbert Kasozi, Tutilo Mudumba, and Mordecai Ogada. Positioning human heritage at the center of conservation practice. Conservation Biology, 34 (5):1122–1130, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mikolajczak:2023:EIN

[MBL^{+}23] Katarzyna M. Mikołajczak, Jos Barlow, Alexander C. Lees, Christopher D. Ives, Micha Strack, Oriana Trindade de Almeida, Agnis C. Souza, Frazer Sinclair, and Luke Parry. Evaluating the influence of nature connection and values on conservation attitudes at a tropical deforestation frontier. Conservation Biology, 37(4):e14067:1–e14067:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mathon:2025:TDC

[MBLD^{+}25] Laetitia Mathon, Florian Baletaud, Anne Lebourges-Dhaussy, Gaël Lecellier, Christophe Menkes, Céline Bachelier, Claire Bonneville, Tony Dejean, Mahé Dumas, Sylvie Fiat, Jacques Grelet, Jérémie Habasque, Stéphanie Manel, Laura Mannocci, David Mouillot, Maëlis Peran, Gildas Roudaut, Christine Sidobre, David Varillon, and Laurent Vigliola. Three-dimensional conservation planning of fish biodiversity metrics to achieve the deep-sea 30 × 30 conservation target. Conservation Biology, 39 (2):e14368:1–e14368:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moussy:2022:QGR

[MBS^{+}22] Caroline Moussy, Ian J. Burfield, P. J. Stephenson, Arabella F. E. Newton, Stuart H. M. Butchart, William J. Sutherland, Richard D. Gregory, Louise McRae, Philip Bubb, Ignacio Roesler, Cynthia Ursino, Yanqing Wu, Ernst F. Retief, Jihad S. Udin, Ruslan Urazaliyev, Lina M. Sánchez-Clavijo, Eric Lartey, and Paul F. Donald. A quantitative global review of species population monitoring. Conservation Biology, 36(1): e13721:1–e13721:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moor:2024:BPA

[MBV^{+}24] Helen Moor, Ariel Bergamini, Christoph Vorburger, Rolf Holderegger, Christoph Bühler, Nicolas Bircher, and Benedikt R. Schmidt. Building pondscapes for amphibian metapopulations. Conservation Biology, 38(6):e14165:1–e14165:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marshall:2022:LMN

[MBW^{+}22] Imogen R. Marshall, Chris J. Brauer, Scotte D. Wedderburn, Nick S. Whiterod, Michael P. Hammer, Thomas C. Barnes, Catherine R. M. Attard, Luciana M. Möller, and Luciano B. Beheregaray. Longitudinal monitoring of neutral and adaptive genomic diversity in a reintroduction. Conservation Biology, 36 (4):e13889:1–e13889:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Montgomery:2020:CMT

[MCB^{+}20] Robert A. Montgomery, Madeline Carr, Charlie R. Booher, Abigail M. Pointer, Brendan M. Mitchell, Natalie Smith, Keegan Calnan, Georgina M. Montgomery, Mordecai Ogada, and Daniel B. Kramer. Characteristics that make trophy hunting of giant pandas inconceivable. Conservation Biology, 34(4):915– 924, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Muenzel:2023:CSC

[MCC^{+}23a] Dominic Muenzel, Kay Critchell, Courtney Cox, Stuart J. Campbell, Raymond Jakub, Iliana Chollett, Nils Krueck, Daniel Holstein, Eric A. Treml, and Maria Beger. Comparing spatial conservation prioritization methods with site- versus spatial dependency-based connectivity. Conservation Biology, 37(2):e14008:1–e14008:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Muenzel:2023:ILC

[MCC^{+}23b] Dominic Muenzel, Kay Critchell, Courtney Cox, Stuart J. Campbell, Raymond Jakub, Wahid Suherfian, La Sara, Iliana Chollett, Eric A. Treml, and Maria Beger. Integrating larval connectivity into the marine conservation decision-making process across spatial scales. Conservation Biology, 37(3): e14038:1–e14038:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:2024:UCG

[McC24] Tim R. McClanahan. Usage and coordination of governance principles to address proximate and distal drivers of conflicts in fisheries commons. Conservation Biology, 38(2):e14178:1– e14178:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marques:2021:UED

[MCF^{+}21] Virginie Marques, Paul Castagné, Andréa Polanco Fernández, Giomar Helena Borrero-Pérez, Régis Hocdé, Pierre-Édouard Guérin, Jean-Baptiste Juhel, Laure Velez, Nicolas Loiseau, Tom Bech Letessier, Sandra Bessudo, Alice Valentini, Tony Dejean, David Mouillot, Loïc Pellissier, and Sébastien Villéger. Use of environmental DNA in assessment of fish functional and phylogenetic diversity. Conservation Biology, 35(6):1944–1956, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McCarthy:2024:CGB

[MCF^{+}24] Orion S. McCarthy, Kanisha Contractor, Will F. Figueira, Arthur C. R. Gleason, T. Shay Viehman, Clinton B. Edwards, and Stuart A. Sandin. Closing the gap between existing largearea imaging research and marine conservation needs. Conservation Biology, 38(1):e14145:1–e14145:??, February 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mittermeier:2021:UWM

[MCG^{+}21] John C. Mittermeier, Ricardo Correia, Rich Grenyer, Tuuli Toivonen, and Uri Roll. Using Wikipedia to measure public interest in biodiversity and conservation. Conservation Biology, 35(2):412–423, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mothes:2020:USM

[MCH^{+}20] Caitlin C. Mothes, Stephanie L. Clements, Dishane K. Hewavithana, Hunter J. Howell, Aaron S. David, Nicole D. Leventhal, and Christopher A. Searcy. Use of standardized methods to improve extinction-risk classification. Conservation Biology, 34(3):754–761, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meierhofer:2022:PPE

[MCLM22] Melissa B. Meierhofer, Pedro Cardoso, Thomas Lilley, and Stefano Mammola. The promise and perils of engineering cave climates: response to Turner et al. Conservation Biology, 36 (4):e13927:1–e13927:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moore:2021:FAC

[MCM^{+}21] Joslin L. Moore, Abbey E. Camaclang, Alana L. Moore, Cindy E. Hauser, Michael C. Runge, Victor Picheny, and Libby Rumpff. A framework for allocating conservation resources among multiple threats and actions. Conservation Biology, 35 (5):1639–1649, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maurenza:2025:EDM

[MCP^{+}25] Daniel Maurenza, Renato Crouzeilles, Jayme Augusto Prevedello, Mauricio Almeida-Gomes, Marina Schmoeler, Renata Pardini, Cristina Banks-Leite, Marcus Vinicius Vieira, Jean Paul Metzger, Carlos Roberto Fonseca, Marina Zanin, Alex Fernando Mendes, Andrea Larissa Boesing, Andréia Alves Rezende, Bruno Karol Cordeiro Filgueiras, Camila dos Santos de Barros, Candelaria Estavillo, Carlos A. Peres, Carolina Franco Esteves, Dary Rigueira, Deborah Faria, Eduardo Mariano-Neto, Eliana Cazetta, Elivane Salete Capellesso, Emerson Monteiro Vieira, Erica Hasui, Euvaldo Marciano Santos Silva Júnior, Flavio Nunes Ramos, Francisco Sanches Gomes, Gabriela Paise, Inara Roberta Leal, José Carlos Morante-Filho, Juliano André Bogoni, Katia Maria Paschoaletto Micchi de Barros Ferraz, Larissa Rocha-Santos, Luciana Carvalho dos Reis, Luciano Carramaschi de Alagão Querido, Luiz Fernando Silva Magnago, Luiz Gustavo Rodrigues Oliveira Santos, Marcelo Passamani, Marcelo Tabarelli, Marcia Cristina Mendes Marques, Marilia Mascarenhas Lima, Marlla Alves Matos, Maurício Eduardo Graipel, Maxwell Souza Silveira, Michaele de Souza Pessoa, Nathália Vieira Hissa Safar, Pedro Henrique Santin Brancalion, Tiago Jordão Porto, and Thomas Püttker. Effects of deforestation on multitaxa community similarity in the Brazilian Atlantic Forest. Conservation Biology, 39(2):e14419:1–e14419:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moore:2024:QBV

[MCW^{+}24] Alex C. Moore, Kendall Calhoun, Christine E. Wilkinson, Elan Alford, Alycia Ellington, Cesar O. Estien, Gabriela S. Fleury, Nepsis Garcia, Akiebia Hicks, Camille Mosley, and Neshima Vitale-Penniman. Queer black voices in conservation. Conservation Biology, 38(6):e14385:1–e14385:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:2024:DRT

[MDB^{+}24] Tim R. McClanahan, Emily S. Darling, Maria Beger, Helen E. Fox, Hedley S. Grantham, Stacy D. Jupiter, Cheryl A. Logan, Elizabeth Mcleod, Lisa C. McManus, Remy M. Oddenyo, Gautam S. Surya, Amelia. S. Wenger, Jens Zinke, and Joseph M. Maina. Diversification of refugia types needed to secure the future of coral reefs subject to climate change. Conservation Biology, 38(1):e14108:1–e14108:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martinet:2021:GEE

[MDG^{+}21] Baptiste Martinet, Simon Dellicour, Guillaume Ghisbain, Kimberly Przybyla, Ella Zambra, Thomas Lecocq, Mira Boustani, Ruslan Baghirov, Denis Michez, and Pierre Rasmont. Global effects of extreme temperatures on wild bumblebees. Conservation Biology, 35(5):1507–1518, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moreto:2025:ICC

[ME25] William D. Moreto and Richard L. Elligson. Integrating criminology, criminal justice, and crime science in conservation science and practice. Conservation Biology, 39(2):e70006:1– e70006:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mychajliw:2022:LCB

[MEA^{+}22] Alexis M. Mychajliw, Elizabeth R. Ellwood, Peter S. Alagona, R. Scott Anderson, Mairin A. Balisi, Eric Biber, Justin L. Brown, Jessie George, Austin J. W. Hendy, Lila Higgins, Courtney A. Hofman, Ashley Leger, Miguel A. Ordeñana, Gregory B. Pauly, Breanna J. Putman, John M. Randall, Seth P. D. Riley, Allison J. Shultz, M. Allison Stegner, Thomas A. Wake, and Emily L. Lindsey. Lessons for conservation from beneath the pavement. Conservation Biology, 36(6):e13983:1– e13983:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Muscatello:2021:HDA

[MEK21] Angela Muscatello, Jane Elith, and Heini Kujala. How decisions about fitting species distribution models affect conservation outcomes. Conservation Biology, 35(4):1309–1320, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mayor:2022:WMT

[MEM^{+}22] Pedro Mayor, Hani R. El Bizri, Thais Q. Morcatty, Kelly Moya, Nora Bendayán, Samantha Solis, Carlos F. A. Vasconcelos Neto, Maire Kirkland, Omar Arevalo, Tula G. Fang, Pedro E. Pérez-Peña, and Richard E. Bodmer. Wild meat trade over the last 45 years in the Peruvian Amazon. Conservation Biology, 36(2):e13801:1–e13801:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:2024:TCB

[MFW^{+}24] Tim R. McClanahan, Alan M. Friedlander, Julien Wickel, Nicholas A. J. Graham, J. Henrich Bruggemann, Mireille M. M. Guillaume, P. Chabanet, Sean Porter, Michael H. Schleyer, M. Kodia Azali, and N. A. Muthiga. Testing for concordance between predicted species richness, past prioritization, and marine protected area designations in the western Indian Ocean. Conservation Biology, 38(4):e14256:1–e14256:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Millard:2021:SAI

[MGJF21] Joseph W. Millard, Richard D. Gregory, Kate E. Jones, and Robin Freeman. The species awareness index as a conservation culturomics metric for public biodiversity awareness. Conservation Biology, 35(2):472–482, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2023:CCS

[MGS^{+}23] David M. Martin, Joshua G. Goldstein, David R. Smith, Jessica Musengezi, Jessie G. Rountree, G. A. Pabodha Galgamuwa, Aileen Craig, Michelle Dietz, and Caitlin Kerr. Creating conservation strategies with value-focused thinking. Conservation Biology, 37(5):e14109:1–e14109:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marsh:2022:PSU

[MHB^{+}22] Sophie M. E. Marsh, Michael Hoffmann, Neil D. Burgess, Thomas M. Brooks, Daniel W. S. Challender, Patricia J. Cremona, Craig Hilton-Taylor, Flore Lafaye de Micheaux, Gabriela Lichtenstein, Dilys Roe, and Monika Böhm. Prevalence of sustainable and unsustainable use of wild species inferred from the IUCN red list of threatened species. Conservation Biology, 36(2):e13844:1–e13844:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mankad:2021:EKE

[MHC21] Aditi Mankad, Elizabeth V. Hobman, and Lucy Carter. Effects of knowledge and emotion on support for novel synthetic biology applications. Conservation Biology, 35(2):623–633, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marquina:2022:LEV

[MHD^{+}22] Tatiana Marquina, Diana Hackenburg, Hannah Duray, Brendan Fisher, and Rachelle K. Gould. Lessons from an experiment with values-based messaging to support watershed conservation. Conservation Biology, 36(5):e13910:1–e13910:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mbaru:2021:EOC

[MHGC21] Emmanuel K. Mbaru, Christina C. Hicks, Georgina G. Gurney, and Joshua E. Cinner. Evaluating outcomes of conservation with multidimensional indicators of well-being. Conservation Biology, 35(5):1417–1425, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marraffini:2024:EIM

[MHJ^{+}24] M. L. Marraffini, S. L. Hamilton, J. R. Marin Jarrin, M. Ladd, G. Koval, J. R. Madden, I. Mangino, L. M. Parker, K. A. Emery, K. Terhaar, D. M. Hubbard, R. J. Miller, and J. E. Dugan. Evaluating the influence of marine protected areas on surf zone fish. Conservation Biology, 38(6):e14296:1–e14296:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ma:2023:DEC

[MHJD23] Liang Ma, Chao Hou, Zhong-Wen Jiang, and Wei-Guo Du. Divergent effects of climate change on the egg-laying opportunity of species in cold and warm regions. Conservation Biology, 37 (3):e14056:1–e14056:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mi:2022:ECH

[MHL^{+}22] Chunrong Mi, Falk Huettmann, Xinhai Li, Zhongwen Jiang, Weiguo Du, and Baojun Sun. Effects of climate and human activity on the current distribution of amphibians in China. Conservation Biology, 36(6):e13964:1–e13964:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Munstermann:2022:GES

[MHM^{+}22] Maya J. Munstermann, Noel A. Heim, Douglas J. McCauley, Jonathan L. Payne, Nathan S. Upham, Steve C. Wang, and Matthew L. Knope. A global ecological signal of extinction risk in terrestrial vertebrates. Conservation Biology, 36(3): e13852:1–e13852:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McEachran:2024:RWD

[MHM^{+}24] Margaret C. McEachran, Johanna A. Harvey, Riley O. Mummah, Molly C. Bletz, Claire S. Teitelbaum, Elias Rosenblatt, F. Javiera Rudolph, Fernando Arce, Shenglai Yin, Diann J. Prosser, Brittany A. Mosher, Jennifer M. Mullinax, Graziella V. DiRenzo, Jannelle Couret, Michael C. Runge, Evan H. Campbell Grant, and Jonathan D. Cook. Reframing wildlife disease management problems with decision analysis. Conservation Biology, 38(4):e14284:1–e14284:??, August 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Medina:2023:RWR

[MHP23] Wilderson Medina, Ryan M. Huang, and Stuart L. Pimm. Region-wide retreats from lower elevations of range-restricted birds across the Northern Andes. Conservation Biology, 37(5): e14127:1–e14127:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Munno:2022:MCG

[MHR^{+}22] Keenan Munno, Paul A. Helm, Chelsea Rochman, Tara George, and Donald A. Jackson. Microplastic contamination in Great Lakes fish. Conservation Biology, 36(1):e13794:1– e13794:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mainali:2020:MER

[MHRF20] Kumar Mainali, Trevor Hefley, Leslie Ries, and William F. Fagan. Matching expert range maps with species distribution model predictions. Conservation Biology, 34(5):1292–1304, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClure:2021:RIE

[MHS^{+}21] Eva C. McClure, Andrew S. Hoey, Katie T. Sievers, Rene A. Abesamis, and Garry R. Russ. Relative influence of environmental factors and fishing on coral reef fish assemblages. Conservation Biology, 35(3):976–990, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Melanidis:2023:EET

[MHSL^{+}23] Marina Stavroula Melanidis, Shannon Hagerman, Guillaume Peterson St-Laurent, Lauren E. Oakes, and Molly S. Cross. Exploring the emergence of a tipping point for conservation with increased recognition of social considerations. Conservation Biology, 37(4):e14086:1–e14086:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martinez-Jauregui:2020:ASA

[MJDMAS20] María Martínez-Jauregui, Miguel Delibes-Mateos, Beatriz Arroyo, and Mario Soliño. Addressing social attitudes toward lethal control of wildlife in national parks. Conservation Biology, 34(4):868–878, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meierhofer:2024:ECS

[MJPJ^{+}24] Melissa B. Meierhofer, Joseph S. Johnson, Janette Perez-Jimenez, Fernanda Ito, Paul W. Webela, Sigit Wiantoro, Enrico Bernard, Krizler C. Tanalgo, Alice Hughes, Pedro Cardoso, Thomas Lilley, and Stefano Mammola. Effective conservation of subterranean-roosting bats. Conservation Biology, 38(1): e14157:1–e14157:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moilanen:2021:TWD

[MK21] Atte Moilanen and Janne S. Kotiaho. Three ways to deliver a net positive impact with biodiversity offsets. Conservation Biology, 35(1):197–205, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mitani:2022:MAL

[ML22] Yohei Mitani and Henrik Lindhjem. Meta-analysis of landowner participation in voluntary incentive programs for provision of forest ecosystem services. Conservation Biology, 36(1):e13729:1–e13729:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Milson:2025:NCF

[MLIE25] Caroline E. Milson, Jun Ying Lim, Daniel J. Ingram, and David P. Edwards. The need for carbon finance schemes to tackle overexploitation of tropical forest wildlife. Conservation Biology, 39(1):e14406:1–e14406:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Manlik:2022:SME

[MLS^{+}22] Oliver Manlik, Robert C. Lacy, William B. Sherwin, Hugh Finn, Neil R. Loneragan, and Simon J. Allen. A stochastic model for estimating sustainable limits to wildlife mortality in a changing world. Conservation Biology, 36(4):e13897:1– e13897:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McAfee:2022:TLR

[MMA^{+}22] Dominic McAfee, Ian M. McLeod, Heidi K. Alleway, Melanie J. Bishop, Simon Branigan, Sean D. Connell, Craig Copeland, Christine M. Crawford, Ben K. Diggles, James A. Fitzsimons, Ben L. Gilby, Paul Hamer, Boze Hancock, Robert Pearce, Kylie Russell, and Chris L. Gillies. Turning a lost reef ecosystem into a national restoration program. Conservation Biology, 36(6): e13958:1–e13958:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Margulies:2023:PPI

[MMG^{+}23] Jared D. Margulies, Francesca R. Moorman, Bárbara Goettsch, Jan C. Axmacher, and Amy Hinsley. Prevalence and perspectives of illegal trade in cacti and succulent plants in the collector community. Conservation Biology, 37(3):e14030:1– e14030:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moseby:2020:EIE

[MMHR20] Katherine E. Moseby, Hugh McGregor, Brydie M. Hill, and John L. Read. Exploring the internal and external wildlife gradients created by conservation fences. Conservation Biology, 34(1):220–231, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morcatty:2020:ITW

[MMN^{+}20] Thais Q. Morcatty, Jonathan C. Bausch Macedo, K. Anne-Isola Nekaris, Qingyong Ni, Carlos C. Durigan, Magdalena S. Svensson, and Vincent Nijman. Illegal trade in wild cats and its link to Chinese-led development in Central and South America. Conservation Biology, 34(6):1525–1535, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Meyer:2020:EPI

[MMS^{+}20] Ninon F. V. Meyer, Ricardo Moreno, Christopher Sutherland, J. Antonio de la Torre, Helen J. Esser, Christopher A. Jordan, Melva Olmos, Josué Ortega, Rafael Reyna-Hurtado, Samuel Valdes, and Patrick A. Jansen. Effectiveness of Panama as an intercontinental land bridge for large mammals. Conservation Biology, 34(1):207–219, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martinez-Nunez:2023:PAC

[MNMPGN23] Carlos Martínez-Núñez, Ricardo Martínez-Prentice, and Vicente García-Navas. Protected area coverage of vulnerable regions to conserve functional diversity of birds. Conservation Biology, 37(6):e14131:1–e14131:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McClanahan:2025:SCG

[MO25] T. R. McClanahan and R. M. Oddenyo. Stimulating the capacity to govern the commons. Conservation Biology, 39(1): e14333:1–e14333:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Molnar:2022:BMI

[Mol22] Zsolt Molnár. From biodiversity management to interdisciplinary governance. Conservation Biology, 36(4):e13937:1– e13937:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moorhouse:2021:CAC

[Moo21] Ron Moorhouse. Cats among the conservationists. Conservation Biology, 35(3):1054–1055, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mahajan:2024:IEM

[MOO^{+}24] Shauna L. Mahajan, Samson Obiene, Lenice Ojwang, Nasser Olwero, Abel Valdivia, Adaoma Wosu, Emily Adrid, Dominic A. Andradi-Brown, Gildas Andriamalala, Natalie C. Ban, Nathan J. Bennett, Jessica Blythe, Samantha H. Cheng, Emily Darling, Matheus De Nardo, Elizabeth Drury O’Neill, Graham Epstein, Robert Y. Fidler, Kim Fisher, Jonas Geldmann, David A. Gill, Rachel Golden Kroner, Georgina Gurney, Arundhati Jagadish, Harry D. Jonas, Muhammad Erdi Lazuardi, Samantha Petersen, Valencia V. Ranarivelo, Lilia Rasoloformanana, Tojo M. Rasolozaka, Daniel J. Read, Elia Sabula Mwaiteleke, and Gabby Ahmadia. Introducing Elinor for monitoring the governance and management of areabased conservation. Conservation Biology, 38(2):e14213:1– e14213:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Maroso:2023:FCA

[MPM^{+}23] Francesco Maroso, Giada Padovani, Víctor Hugo Muñoz Mora, Francesco Giannelli, Emiliano Trucchi, and Giorgio Bertorelle. Fitness consequences and ancestry loss in the Apennine brown bear after a simulated genetic rescue intervention. Conservation Biology, 37(6):e14133:1–e14133:??, December 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martell:2020:FFS

[MR20a] Josephine E. M. Martell and Amanda D. Rodewald. Frames, facts, and the science of communicating environmental crises. Conservation Biology, 34(3):766–768, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martell:2020:MTE

[MR20b] Josephine E. M. Martell and Amanda D. Rodewald. Media transparency and evidence-based framing: reply to Kusmanoff. Conservation Biology, 34(5):1063–1064, October 2020. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martin:2022:ULU

[MR22] Dominic A. Martin and Estelle Raveloaritiana. Using landuse history and multiple baselines to determine bird responses to cocoa agroforestry. Conservation Biology, 36(4):e13920:1– e13920:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mozelewski:2022:FIC

[MRS22] Tina G. Mozelewski, Zachary J. Robbins, and Robert M. Scheller. Forecasting the influence of conservation strategies on landscape connectivity. Conservation Biology, 36(5):e13904:1– e13904:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moseby:2024:IIE

[MRTV24] Katherine E. Moseby, John L. Read, Katherine Tuft, and Leanne K. Van der Weyde. Influence of interactive effects on long-term population trajectories in multispecies reintroductions. Conservation Biology, 38(3):e14209:1–e14209:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

MacKinlay:2023:SRA

[MS23] Regan D. MacKinlay and Rachael C. Shaw. A systematic review of animal personality in conservation science. Conservation Biology, 37(1):e13935:1–e13935:??, February 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marino:2025:CPF

[MSB25] Clara Marino, Filipa Coutinho Soares, and Céline Bellard. Conservation priorities for functionally unique and specialized terrestrial vertebrates threatened by biological invasions. Conservation Biology, 39(2):e14401:1–e14401:??, April 2025. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mancini:2024:SAI

[MSC^{+}24] Giordano Mancini, Luca Santini, Victor Cazalis, H. Reşit Akçakaya, Pablo M. Lucas, Thomas M. Brooks, Wendy Foden, and Moreno Di Marco. A standard approach for including climate change responses in IUCN Red List assessments. Conservation Biology, 38(3):e14227:1–e14227:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morton:2023:ART

[MSHE23] Oscar Morton, Brett R. Scheffers, Torbjørn Haugaasen, and David P. Edwards. Association of reproductive traits with captive- versus wild-sourced birds in trade. Conservation Biology, 37(4):e14076:1–e14076:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Muller:2020:EEE

[MSJ20] Anke Müller, Uwe A. Schneider, and Kerstin Jantke. Evaluating and expanding the European Union’s protected-area network toward potential post-2020 coverage targets. Conservation Biology, 34(3):654–665, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Munoz-Saez:2021:BFW

[MSKM21] Andrés Muñoz-Sáez, Justin Kitzes, and Adina M. Merenlender. Bird-friendly wine country through diversified vineyards. Conservation Biology, 35(1):274–284, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

March-Salas:2023:CPE

[MSLBS23] Martí March-Salas, Juan Lorite, Oliver Bossdorf, and J. F. Scheepens. Cliffs as priority ecosystems. Conservation Biology, 37(5):e14166:1–e14166:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Martens:2022:CIF

[MSMH22] Carola Martens, Simon Scheiter, Guy F. Midgley, and Thomas Hickler. Combined impacts of future climate-driven vegetation changes and socioeconomic pressures on protected areas in Africa. Conservation Biology, 36(6):e13968:1–e13968:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Michel:2021:ESP

[MSMW21] Nicole L. Michel, Sarah P. Saunders, Timothy D. Meehan, and Chad B. Wilsey. Effects of stewardship on protected area effectiveness for coastal birds. Conservation Biology, 35(5): 1484–1495, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Munteanu:2022:UHS

[MSN^{+}22] Catalina Munteanu, Cornelius Senf, Mihai D. Nita, Francesco Maria Sabatini, Julian Oeser, Rupert Seidl, and Tobias Kuemmerle. Using historical spy satellite photographs and recent remote sensing data to identify high-conservation-value forests. Conservation Biology, 36(2):e13820:1–e13820:??, April 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Molder:2020:HTP

[MSPM20] Andreas Mölder, Marcus Schmidt, Tobias Plieninger, and Peter Meyer. Habitat-tree protection concepts over 200 years. Conservation Biology, 34(6):1444–1451, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mumby:2021:MRF

[MSRP21] Peter J. Mumby, Robert S. Steneck, George Roff, and Valerie J. Paul. Marine reserves, fisheries ban, and 20 years of positive change in a coral reef ecosystem. Conservation Biology, 35(5): 1473–1483, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Morcatty:2024:NEC

[MSS^{+}24] Thais Q. Morcatty, Shan Su, Penthai Siriwat, Astrid Alex Andersson, Sadek Atoussi, Kim Feddema, Sergio Henriques, Jordi Janssen, Anushri Karve, Jennifer Pytka, Ruth M. Thompson, Vincent Nijman, Joss Wright, and David L. Roberts. Navigating ethical challenges in online wildlife trade research. Conservation Biology, 38(5):e14341:1–e14341:??, October 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Manfredo:2020:CSC

[MTC^{+}20] Michael J. Manfredo, Tara L. Teel, Andrew W. Don Carlos, Leeann Sullivan, Alan D. Bright, Alia M. Dietsch, Jeremy Bruskotter, and David Fulton. The changing sociocultural context of wildlife conservation. Conservation Biology, 34(6): 1549–1559, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Margulies:2024:EAI

[MTH^{+}24] Jared D. Margulies, Benjamin Trost, Laura Hamon, Natalie Z. Kerr, Michael Kunz, John L. Randall, Roger D. Shew, Dale M. Shew, Lesley Starke, Dale Suiter, and Zachary West. Expert assessment of illegal collecting impacts on Venus flytraps and priorities for research on illegal trade. Conservation Biology, 38(5):e14320:1–e14320:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moranta:2022:TCG

[MTM^{+}22] Joan Moranta, Cati Torres, Ivan Murray, Manuel Hidalgo, Hilmar Hinz, and Adam Gouraguine. Transcending capitalism growth strategies for biodiversity conservation. Conservation Biology, 36(2):e13821:1–e13821:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marselle:2021:ABP

[MTS^{+}21] Melissa R. Marselle, Anne Turbe, Assaf Shwartz, Aletta Bonn, and Agathe Colléony. Addressing behavior in pollinator conservation policies to combat the implementation gap. Conservation Biology, 35(2):610–622, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Marshall:2021:QIV

[MVC^{+}21] Erica Marshall, Roozbeh Valavi, Louise O’ Connor, Natasha Cadenhead, Darren Southwell, Brendan A. Wintle, and Heini Kujala. Quantifying the impact of vegetation-based metrics on species persistence when choosing offsets for habitat destruction. Conservation Biology, 35(2):567–577, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mannocci:2022:LSM

[MVC^{+}22] Laura Mannocci, Sébastien Villon, Marc Chaumont, Nacim Guellati, Nicolas Mouquet, Corina Iovan, Laurent Vigliola, and David Mouillot. Leveraging social media and deep learning to detect rare megafauna in video surveys. Conservation Biology, 36(1):e13798:1–e13798:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Moreau:2025:ITC

[MW25] Marie-Annick Moreau and Emily Woodhouse. Improving transparency in conservation social science research to enhance quality, equity, and collaboration. Conservation Biology, 39 (2):e70003:1–e70003:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McGowan:2020:PDC

[MWC^{+}20] Jennifer McGowan, Rob Weary, Leah Carriere, Edward T. Game, Joanna L. Smith, Melissa Garvey, and Hugh P. Possingham. Prioritizing debt conversion opportunities for marine conservation. Conservation Biology, 34(5):1065–1075, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mokany:2022:HBB

[MWH^{+}22] Karel Mokany, Chris Ware, Thomas D. Harwood, Rebecca K. Schmidt, and Simon Ferrier. Habitat-based biodiversity assessment for ecosystem accounting in the Murray–Darling Basin. Conservation Biology, 36(5):e13915:1–e13915:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McNally:2025:MAE

[MWS^{+}25] Xavier McNally, Thomas L. Webb, Charlotte Smith, Andrew Moss, and Jilly Gibson-Miller. A meta-analysis of the effect of visiting zoos and aquariums on visitors’ conservation knowledge, beliefs, and behavior. Conservation Biology, 39(1): e14237:1–e14237:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

McDonald:2020:CSF

[MWV^{+}20] Gavin McDonald, Molly Wilson, Diogo Veríssimo, Rebecca Twohey, Michaela Clemence, Dean Apistar, Stephen Box, Paul Butler, Fel Cesar Cadiz, Stuart J. Campbell, Courtney Cox, Micah Effron, Steve Gaines, Raymond Jakub, Roquelito H. Mancao, Pablo T. Rojas, Rocky Sanchez Tirona, and Gabriel Vianna. Catalyzing sustainable fisheries management through behavior change interventions. Conservation Biology, 34(5): 1176–1189, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ma:2024:CER

[MWW^{+}24] Liang Ma, Dan-Yang Wu, Yang Wang, Joshua M. Hall, Chun-Rong Mi, Hong-Xin Xie, Wei-Jie Tao, Chao Hou, Kun-Ming Cheng, Yong-Pu Zhang, Ji-Chao Wang, Hong-Liang Lu, Wei-Guo Du, and Bao-Jun Sun. Collective effects of rising average temperatures and heat events on oviparous embryos. Conservation Biology, 38(4):e14266:1–e14266:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Mathwin:2021:MWA

[MWY^{+}21] Rupert Mathwin, Skye Wassens, Jeanne Young, Qifeng Ye, and Corey J. A. Bradshaw. Manipulating water for amphibian conservation. Conservation Biology, 35(1):24–34, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niemiec:2020:USE

[NAG^{+}20] Rebecca M. Niemiec, Greg P. Asner, Julie A. Gaertner, Philip G. Brodrick, Nick Vaughn, Joseph Heckler, Flint Hughes, Lisa Keith, and Tracie Matsumoto. Using spatially explicit, time-dependent analysis to understand how social factors influence conservation outcomes. Conservation Biology, 34 (2):505–514, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Natusch:2021:PFS

[NAS21a] Daniel J. D. Natusch, Patrick W. Aust, and Richard Shine. The perils of flawed science in wildlife trade literature. Conservation Biology, 35(5):1396–1404, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [SP21] and replies [EDA^{+}21, NAS21b].

Natusch:2021:PES

[NAS21b] Daniel J. D. Natusch, Patrick W. Aust, and Richard Shine. Pitfalls in evaluating the sustainability of wildlife trade: reply to Sosnowski and Petrossian and Edwards et al. Conservation Biology, 35(5):1695–1697, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SP21].

Nelson:2021:CIC

[NBB^{+}21] Michael Paul Nelson, Chelsea Batavia, Kate J. Brandis, Scott P. Carroll, Danielle Celermajer, Wayne Linklater, Erick Lundgren, Daniel Ramp, Jamie Steer, Esty Yanco, and Arian D. Wallach. Challenges at the intersection of conservation and ethics: Reply to Meyer et al. 2021. Conservation Biology, 35(1):373–377, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nowak:2024:TCN

[NBD^{+}24] Katarzyna Nowak, Dinah Bear, Anwesha Dutta, Myles Traphagen, Michał Żmihorski, and Bogdan Jaroszewicz. Threats to conservation from national security interests. Conservation Biology, 38(1):e14193:1–e14193:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naito:2024:ASA

[NCdlLZ24] Rumi Naito, Kai M. A. Chan, Rocío López de la Lama, and Jiaying Zhao. Audience segmentation approach to conservation messaging for transforming the exotic pet trade. Conservation Biology, 38(6):e14267:1–e14267:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noss:2021:ISS

[NCE^{+}21] Reed F. Noss, Jennifer M. Cartwright, Dwayne Estes, Theo Witsell, Gregg Elliott, Daniel Adams, Matthew Albrecht, Ryan Boyles, Patrick Comer, Chris Doffitt, Don Faber-Langendoen, JoVonn Hill, William C. Hunter, Wesley M. Knapp, Michael E. Marshall, Jason Singhurst, Christopher Tracey, Jeffrey Walck, and Alan Weakley. Improving species status assessments under the U.S. Endangered Species Act and implications for multispecies conservation challenges worldwide. Conservation Biology, 35(6):1715–1724, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Negret:2020:ESA

[NDS^{+}20] Pablo Jose Negret, Moreno Di Marco, Laura J. Sonter, Jonathan Rhodes, Hugh P. Possingham, and Martine Maron.

Effects of spatial autocorrelation and sampling design on estimates of protected area effectiveness. Conservation Biology, 34(6):1452–1462, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nilsson:2020:ACI

[NFD20] Danielle Nilsson, Kelly Fielding, and Angela J. Dean. Achieving conservation impact by shifting focus from human attitudes to behaviors. Conservation Biology, 34(1):93–102, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Noonan:2020:EBS

[NFT^{+}20] Michael J. Noonan, Christen H. Fleming, Marlee A. Tucker, Roland Kays, Autumn-Lynn Harrison, Margaret C. Crofoot, Briana Abrahms, Susan C. Alberts, Abdullahi H. Ali, Jeanne Altmann, Pamela Castro Antunes, Nina Attias, Jerrold L. Belant, Dean E. Beyer, Jr., Laura R. Bidner, Niels Blaum, Randall B. Boone, Damien Caillaud, Rogerio Cunha de Paula, J. Antonio de la Torre, Jasja Dekker, Christopher S. DePerno, Mohammad Farhadinia, Julian Fennessy, Claudia Fichtel, Christina Fischer, Adam Ford, Jacob R. Goheen, Rasmus W. Havmøller, Ben T. Hirsch, Cindy Hurtado, Lynne A. Isbell, René Janssen, Florian Jeltsch, Petra Kaczensky, Yayoi Kaneko, Peter Kappeler, Anjan Katna, Matthew Kauffman, Flavia Koch, Abhijeet Kulkarni, Scott LaPoint, Peter Leimgruber, David W. Macdonald, A. Catherine Markham, Laura McMahon, Katherine Mertes, Christopher E. Moorman, Ronaldo G. Morato, Alexander M. Moßbrucker, Guilherme Mourão, David O’Connor, Luiz Gustavo R. Oliveira-Santos, Jennifer Pastorini, Bruce D. Patterson, Janet Rachlow, Dustin H. Ranglack, Neil Reid, David M. Scantlebury, Dawn M. Scott, Nuria Selva, Agnieszka Sergiel, Melissa Songer, Nucharin Songsasen, Jared A. Stabach, Jenna Stacy-Dawes, Morgan B. Swingen, Jeffrey J. Thompson, Wiebke Ullmann, Abi Tamim Vanak, Maria Thaker, John W. Wilson, Koji Yamazaki, Richard W. Yarnell, Filip Zieba, Tomasz Zwijacz-Kozica, William F. Fagan, Thomas Mueller, and Justin M. Calabrese. Effects of body size on estimation of mammalian area requirements. Conservation Biology, 34(4):1017–1028, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Namkhan:2021:LVL

[NGST21] Maliwan Namkhan, George A. Gale, Tommaso Savini, and Naruemon Tantipisanuh. Loss and vulnerability of lowland forests in mainland Southeast Asia. Conservation Biology, 35 (1):206–215, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Navarro-Herrero:2024:UBB

[NHSM^{+}24] Leia Navarro-Herrero, Sarah Saldanha, Teresa Militão, Diego Vicente-Sastre, David March, and Jacob González-Solís. Use of bird-borne radar to examine shearwater interactions with legal and illegal fisheries. Conservation Biology, 38(3):e14224:1– e14224:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nielsen:2020:EDR

[NJ20] Martin Reinhardt Nielsen and Jette Bredahl Jacobsen. Effect of decision rules in choice experiments on hunting and bushmeat trade. Conservation Biology, 34(6):1393–1403, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Niemiec:2021:EBN

[NJLC21] Rebecca Niemiec, Megan S. Jones, Stacy Lischka, and Veronica Champine. Efficacy-based and normative interventions for facilitating the diffusion of conservation behavior through social networks. Conservation Biology, 35(4):1073–1085, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Neville:2020:AMP

[NLD^{+}20] Helen M. Neville, Douglas R. Leasure, Daniel C. Dauwalter, Jason B. Dunham, Robin Bjork, Kurt A. Fesenmyer, Nathan D. Chelgren, Mary M. Peacock, Charles H. Luce, Daniel J. Isaak, Lee Ann Carranza, Jon Sjoberg, and Seth J. Wenger. Application of multiple-population viability analysis to evaluate species recovery alternatives. Conservation Biology, 34(2):482– 493, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nori:2020:PAC

[NLV20] Javier Nori, Rafael Loyola, and Fabricio Villalobos. Priority areas for conservation of and research focused on terrestrial vertebrates. Conservation Biology, 34(5):1281–1291, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Novotny:2023:TDC

[Nov23] Vojtech Novotny. Two doses of carbon budget realism. Conservation Biology, 37(1):e14049:1–e14049:??, February 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nelson:2024:SNA

[NRN^{+}24] Scott B. Maresh Nelson, Christine A. Ribic, Neal D. Niemuth, Jacy Bernath-Plaisted, and Benjamin Zuckerberg. Sensitivity of North American grassland birds to weather and climate variability. Conservation Biology, 38(1):e14143:1–e14143:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nunez-Regueiro:2021:EBB

[NRSF21] Mauricio M. Núñez-Regueiro, Sharmin F. Siddiqui, and Robert J. Fletcher, Jr. Effects of bioenergy on biodiversity arising from land-use change and crop type. Conservation Biology, 35(1):77–87, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nogueira:2021:ADF

[NSB^{+}21] Joana Garrido Nogueira, Ronaldo Sousa, Hassan Benaissa, Geert De Knijf, Sónia Ferreira, Mohamed Ghamizi, Duarte V. Gonçalves, Richard Lansdown, Catherine Numa, Vincent Prié, Nicoletta Riccardi, Mary Seddon, Maria Urbańska, Alice Valentini, Ilya Vikhrev, Simone Varandas, Amílcar Teixeira, and Manuel Lopes-Lima. Alarming decline of freshwater trigger species in western Mediterranean key biodiversity areas. Conservation Biology, 35(5):1367–1379, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Narain:2023:GAB

[NSL^{+}23] Divya Narain, Laura J. Sonter, Alex Mark Lechner, James E. M. Watson, Jeremy S. Simmonds, and Martine Maron.

Global assessment of the biodiversity safeguards of development banks that finance infrastructure. Conservation Biology, 37(4):e14095:1–e14095:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Naia:2021:PNE

[NTS^{+}21] Marisa Naia, Pedro Tarroso, Andack Saad Sow, André Vicente Liz, Duarte Vasconcelos Gonçalves, Fernando Martínez-Freiría, Frederico Santarém, Gholam Hosein Yusefi, Guillermo Velo-Antón, Ignazio Avella, Jeffrey O. Hanson, Leili Khalatbari, Maria Joana Ferreira da Silva, Miguel Camacho-Sanchez, Zbyszek Boratyński, Sílvia B. Carvalho, and José Carlos Brito. Potential negative effects of the Green Wall on Sahel’s biodiversity. Conservation Biology, 35(6):1966–1968, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nuttman:2025:SNM

[Nut25] Clive Nuttman. The state of nature in Madagascar. Conservation Biology, 39(1):e14378:1–e14378:??, February 2025. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nabutanyi:2022:MMV

[NW22] Peter Nabutanyi and Meike J. Wittmann. Modeling minimum viable population size with multiple genetic problems of small populations. Conservation Biology, 36(5):e13940:1–e13940:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Nyssa:2024:FPD

[NWSL^{+}24] Zoe Nyssa, Sophia Winkler-Schor, Diele Lobo, Harold N. Eyster, and Andrew J. Wright. A framework for promoting disciplinary diversity and inclusion through epistemic justice. Conservation Biology, 38(6):e14409:1–e14409:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oyanedel:2024:API

[OABDCM24] Rodrigo Oyanedel, Erendira Aceves-Bueno, Lucia Davids, and Miguel Ángel Cisneros-Mata. An assessment of potential interventions to reduce the totoaba illegal trade market. Conservation Biology, 38(5):e14356:1–e14356:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Osborne:2024:GEE

[OAY^{+}24] Megan J. Osborne, Thomas P. Archdeacon, Charles B. Yackulic, Robert K. Dudley, Guilherme Caeiro-Dias, and Thomas F. Turner. Genetic erosion in an endangered desert fish during a megadrought despite long-term supportive breeding. Conservation Biology, 38(1):e14154:1–e14154:??, February 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olden:2024:GIB

[OC24] Julian D. Olden and Francisco A. C. Carvalho. Global invasion and biosecurity risk from the online trade in ornamental crayfish. Conservation Biology, 38(5):e14359:1–e14359:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ovando:2021:APL

[OCC^{+}21] Daniel Ovando, Jennifer E. Caselle, Christopher Costello, Olivier Deschenes, Steven D. Gaines, Ray Hilborn, and Owen Liu. Assessing the population-level conservation effects of marine protected areas. Conservation Biology, 35(6):1861–1870, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olimpi:2022:IEM

[ODG^{+}22] Elissa M. Olimpi, Hallie Daly, Karina Garcia, Victoria M. Glynn, David J. Gonthier, Claire Kremen, Leithen K. M’Gonigle, and Daniel S. Karp. Interactive effects of multiscale diversification practices on farmland bird stress. Conservation Biology, 36(4):e13902:1–e13902:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oficialdegui:2020:RLI

[ODMG^{+}20] Francisco J. Oficialdegui, Miguel Delibes-Mateos, Andy J. Green, Marta I. Sánchez, Luz Boyero, and Miguel Clavero. Rigid laws and invasive species management. Conservation Biology, 34(4):1047–1050, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Overton:2024:ACR

[ODS^{+}24] Kathy Overton, Tim Dempster, Stephen E. Swearer, Rebecca L. Morris, and Luke T. Barrett. Achieving conservation and restoration outcomes through ecologically beneficial aquaculture. Conservation Biology, 38(1):e14065:1–e14065:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

OBryan:2021:IIP

[OGF^{+}21] Christopher J. O’Bryan, Stephen T. Garnett, Julia E. Fa, Ian Leiper, Jose A. Rehbein, Álvaro Fernández-Llamazares, Micha V. Jackson, Harry D. Jonas, Eduardo S. Brondizio, Neil D. Burgess, Catherine J. Robinson, Kerstin K. Zander, Zsolt Molnár, Oscar Venter, and James E. M. Watson. The importance of indigenous peoples’ lands for the conservation of terrestrial mammals. Conservation Biology, 35(3):1002–1008, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ogg:2020:TFP

[Ogg20] Clayton Ogg. Transforming farm-program incentives to preserve tropical forests. Conservation Biology, 34(3):762–765, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oyanedel:2022:DSM

[OGMMG22] Rodrigo Oyanedel, Stefan Gelcich, Emile Mathieu, and E. J. Milner-Gulland. A dynamic simulation model to support reduction in illegal trade within legal wildlife markets. Conservation Biology, 36(2):e13814:1–e13814:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Oyanedel:2024:CIF

[OLF24] Rodrigo Oyanedel, Marios Levi, and Giuliana Furci. A call to include fungi in wildlife trade research and policy. Conservation Biology, 38(5):e14340:1–e14340:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Otto:2024:PIR

[OPF24] Mathias Otto, Phillip Papastefanou, and Lorenz Fahse. Pressure from insect-resistant maize on protected butterflies and moths. Conservation Biology, 38(2):e14222:1–e14222:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Olmedo:2024:DCE

[OVC^{+}24] Alegría Olmedo, Diogo Veríssimo, Daniel W. S. Challender, Huong Thi Thu Dao, Elena C. Rubino, and E. J. Milner-Gulland. Designing celebrity-endorsed behavioral interventions in conservation. Conservation Biology, 38(3):e14273:1– e14273:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Parks:2022:FEB

[PAFB^{+}22] Danni Parks, Nida Al-Fulaij, Clare Brook, Stuart H. M. Butchart, Jean-Gaël Collomb, David Cope, Simon Dowell, Becky Falkingham, Winifred F. Frick, Douglas Gibbs, Emily E. Gray, Nicolas Heard, Anastasios (Tasso) Leventis, Kate Mastro, Helen Meredith, Simon Mickleburgh, Florence Miller, Matthew Muir, Rascha J. M. Nuijten, Nancy Ockendon, Nisha R. Owen, Jacob R. Owens, Jon Paul Rodríguez, Elizabeth Tully, and Jean-Christophe Vié. Funding evidence-based conservation. Conservation Biology, 36(6):e13991:1–e13991:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peterson:2021:UEM

[PB21] Katie Peterson and Michael Bode. Using ensemble modeling to predict the impacts of assisted migration on recipient ecosystems. Conservation Biology, 35(2):678–687, April 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Prieto:2022:PLS

[PBB^{+}22] Pablo V. Prieto, Jacob J. Bukoski, Felipe S. M. Barros, Hawthorne L. Beyer, Alvaro Iribarrem, Pedro H. S. Brancalion, Robin L. Chazdon, David B. Lindenmayer, Bernardo B. N. Strassburg, Manuel R. Guariguata, and Renato Crouzeilles. Predicting landscape-scale biodiversity recovery by natural tropical forest regrowth. Conservation Biology, 36(3):e13842:1– e13842:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pooley:2021:RSH

[PBV21] Simon Pooley, Saloni Bhatia, and Anirudhkumar Vasava. Rethinking the study of human–wildlife coexistence. Conservation Biology, 35(3):784–793, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pearson:2022:EUC

[PCH22] Dean E. Pearson, Tyler J. Clark, and Philip G. Hahn. Evaluating unintended consequences of intentional species introductions and eradications for improved conservation management. Conservation Biology, 36(1):e13734:1–e13734:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pereyra:2020:RNR

[Per20] Patricio Javier Pereyra. Rethinking the native range concept. Conservation Biology, 34(2):373–377, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pereyra:2020:DIM

[PG20] Patricio Javier Pereyra and Radu Cornel Guiaşu. Debate over the importance and meaning of native range in invasion biology: reply to Courchamp et al. Conservation Biology, 34(4): 1044–1046, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [CHP20].

Perez-Gomez:2023:PCJ

[PGRDOGG23] Álvaro Pérez-Gómez, Irene Repeto-Deudero, Fernando Ojeda, and Susana Gómez-González. Poor communication jeopardizes biodiversity. Conservation Biology, 37(6):e14181:1–e14181:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pavlacky:2021:SPL

[PHB^{+}21] David C. Pavlacky, Jr., Christian A. Hagen, Anne M. Bartuszevige, Rich Iovanna, Thomas Luke George, and David E. Naugle. Scaling up private land conservation to meet recovery goals for grassland birds. Conservation Biology, 35(5): 1564–1574, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Plumptre:2025:SCS

[PHB^{+}25] Andy Plumptre, Jack Hayes, Daniele Baisero, Rob Rose, S. Holness, Lize von Staden, and Robert J. Smith. Strengths and complementarity of systematic conservation planning and Key Biodiversity Area approaches for spatial planning. Conservation Biology, 39(2):e14400:1–e14400:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pausch:2025:REV

[PHK^{+}25] Rachel E. Pausch, Jessica R. Hale, Peter Kiffney, Beth Sanderson, Sara Azat, Katie Barnas, W. Bryant Chesney, Natalie Cosentino-Manning, Stephanie Ehinger, Dayv Lowry, and Steve Marx. Review of ecological valuation and equivalency analysis methods for assessing temperate nearshore submerged aquatic vegetation. Conservation Biology, 39(1):e14380:1– e14380:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perez-Hammerle:2022:WFT

[PHMVL^{+}22] Katharina-Victoria Pérez-Hämmerle, Katie Moon, Rubén Venegas-Li, Sean Maxwell, Jeremy S. Simmonds, Oscar Venter, Stephen T. Garnett, Hugh P. Possingham, and James E. M. Watson. Wilderness forms and their implications for global environmental policy and conservation. Conservation Biology, 36 (4):e13875:1–e13875:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Piccolo:2020:CAL

[Pic20] John J. Piccolo. Celebrating aldo Leopold’s land ethic at 70. Conservation Biology, 34(6):1586–1588, December 2020. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pittock:2021:URC

[Pit21] Jamie Pittock. Use of the Ramsar Convention to protect springs and other wetlands. Conservation Biology, 35(6):1969– 1971, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pienkowski:2022:BMD

[PKC^{+}22] Thomas Pienkowski, Aidan Keane, Sofia Castelló y Tickell, Mirjam Hazenbosch, William N. S. Arlidge, Gergő Baranyi, Stephanie Brittain, Emiel de Lange, Munib Khanyari, Sarah Papworth, and E. J. Milner-Gulland. Balancing making a difference with making a living in the conservation sector. Conservation Biology, 36(4):e1346:1–e1346:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pienkowski:2023:SCM

[PKC^{+}23a] Thomas Pienkowski, Aidan Keane, Sofia Castelló y Tickell, Emiel de Lange, Mirjam Hazenbosch, Munib Khanyari, William N. S. Arlidge, Gergő Baranyi, Stephanie Brittain, Vena Kapoor, Vik Mohan, Sarah Papworth, Roshni Ravi, Izak P. J. Smit, and E. J. Milner-Gulland. Supporting conservationists’ mental health through better working conditions. Conservation Biology, 37(5):e14097:1–e14097:??, October 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pienkowski:2023:RRA

[PKC^{+}23b] Thomas Pienkowski, Laur Kiik, Allison Catalano, Mirjam Hazenbosch, Santiago Izquierdo-Tort, Munib Khanyari, Roshni Kutty, Claudia Martins, Fleur Nash, Omar Saif, and Chris Sandbrook. Recognizing reflexivity among conservation practitioners. Conservation Biology, 37(2):e14022:1–e14022:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palacio:2020:BEC

[PKP20] Ruben D. Palacio, Gustavo H. Kattan, and Stuart L. Pimm. Bird extirpations and community dynamics in an Andean cloud forest over 100 years of land-use change. Conservation Biology, 34(3):677–687, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phillips:2022:ISB

[PLC22] Laura M. Phillips, Rachel I. Leihy, and Steven L. Chown. Improving species-based area protection in Antarctica. Conservation Biology, 36(4):e13885:1–e13885:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pozo:2021:MAW

[PLD^{+}21] Rocío A. Pozo, Eric G. LeFlore, A. Bradley Duthie, Nils Bunnefeld, Isabel L. Jones, Jeroen Minderman, O. Sarobidy Rakotonarivo, and Jeremy J. Cusack. A multispecies assessment of wildlife impacts on local community livelihoods. Conservation Biology, 35(1):297–306, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pratt:2024:IIE

[PLKP24] Elizabeth N. Pratt, Julie L. Lockwood, Elizabeth G. King, and Elizabeth F. Pienaar. Identifying inconsistencies in exotic pet regulations that perpetuate trade in risky species. Conservation Biology, 38(2):e14189:1–e14189:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pocknee:2023:MMR

[PLMF23] Christopher A. Pocknee, Sarah M. Legge, Jane McDonald, and Diana O. Fisher. Modeling mammal response to fire based on species’ traits. Conservation Biology, 37(4):e14062:1– e14062:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Philippe-Lesaffre:2023:RIS

[PLTC^{+}23] Martin Philippe-Lesaffre, Martin Thibault, Stephane Caut, Karen Bourgeois, Tristan Berr, Andreas Ravache, Eric Vidal, Franck Courchamp, and Elsa Bonnaud. Recovery of insular seabird populations years after rodent eradication. Conservation Biology, 37(3):e14042:1–e14042:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Perry:2020:RPD

[PMLM20] Laura R. Perry, Tom P. Moorhouse, Andrew J. Loveridge, and David W. Macdonald. The role of psychology in determining human–predator conflict across southern Kenya. Conservation Biology, 34(4):879–890, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pillay:2022:UIS

[PMR^{+}22] Rajeev Pillay, David A. W. Miller, R. Raghunath, Atul A. Joshi, Charudutt Mishra, A. J. T. Johnsingh, and M. D. Madhusudan. Using interview surveys and multispecies occupancy models to inform vertebrate conservation. Conservation Biology, 36(2):e13832:1–e13832:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Peters:2020:IDE

[PMS^{+}20] Andrew Peters, Anna Meredith, Lee Skerratt, Scott Carver, and Shane Raidal. Infectious disease and emergency conservation interventions. Conservation Biology, 34(4):784–785, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pettersen:2022:IMP

[PMSC22] Amanda K. Pettersen, Ezequiel M. Marzinelli, Peter D. Steinberg, and Melinda A. Coleman. Impact of marine protected areas on temporal stability of fish species diversity. Conservation Biology, 36(2):e13815:1–e13815:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pagani-Nunez:2022:TOB

[PNXY^{+}22] Emilio Pagani-Núñez, Yang Xu, Mingxiao Yan, Jiekun He, Zifei Jiang, and Haisheng Jiang. Trade-offs between economic development and biodiversity conservation on a tropical island. Conservation Biology, 36(5):e139121:1–e139121:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Palfrey:2021:CSO

[POH21] Rachel Palfrey, Johan Oldekop, and George Holmes. Conservation and social outcomes of private protected areas. Conservation Biology, 35(4):1098–1110, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Plein:2022:MTP

[POH^{+}22] Michaela Plein, Katherine R. O’Brien, Matthew H. Holden, Matthew P. Adams, Christopher M. Baker, Nigel G. Bean, Scott A. Sisson, Michael Bode, Kerrie L. Mengersen, and Eve McDonald-Madden. Modeling total predation to avoid perverse outcomes from cat control in a data-poor island ecosystem. Conservation Biology, 36(5):e13916:1–e13916:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pooley:2025:CCC

[Poo25] Simon Pooley. Conservation and coexistence at a crossroads. Conservation Biology, 39(1):e14433:1–e14433:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pozsgai:2020:LOT

[Poz20] Gabor Pozsgai. The land of orangutans through the lens of science. Conservation Biology, 34(5):1316–1317, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Park:2025:LEP

[PP25] Tae Jung Park and Sung-Pil Park. Legal and economic perspectives on fair and equitable benefit sharing in the Nagoya Protocol. Conservation Biology, 39(2):e14410:1–e14410:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pergl:2020:NRT

[PPE^{+}20] Jan Pergl, Petr Pyšek, Franz Essl, Jonathan M. Jeschke, Franck Courchamp, Juergen Geist, Martin Hejda, Ingo Kowarik, Aileen Mill, Camille Musseau, Pavel Pipek, Wolf-Christian Saul, Menja von Schmalensee, and David Strayer. Need for routine tracking of biological invasions. Conservation Biology, 34(5):1311–1314, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Panisi:2024:TSI

[PPN^{+}24] Martina Panisi, Ana R. Pereira, Bruzinay C. Neto, Paulo J. de Sousa, Gabriel Oquiongo, Jorge M. Palmeirim, Ricardo F. de Lima, and Ana Nuno. Trade and socioeconomic importance of an invasive giant snail in the endemic-rich island of São Tomé, Central Africa. Conservation Biology, 38(5):e14357:1– e14357:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Price:2021:PSL

[Pri21a] Trevor Price. Pakistan’s snow leopards. Conservation Biology, 35(6):1972–1973, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Price:2021:HS

[Pri21b] Trevor Price. Hunting for a solution. Conservation Biology, 35 (4):1353–1355, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Price:2025:PTP

[Pri25] Trevor Price. Pray for the tiger or prey for the tiger? Conservation Biology, 39(2):e14444:1–e14444:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petriello:2025:SEC

[PRS^{+}25] Michael A. Petriello, Lauren Redmore, Aby L. Sène, Dhananjaya Katju, Lilian Barraclough, Sara Boyd, Carly Madge, Anastasia Papadopoulos, and Reddi S. Yalamala. The scope of empowerment for conservation and communities. Conservation Biology, 39(1):e14249:1–e14249:??, February 2025. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Petriello:2020:CHC

[PS20] Michael A. Petriello and Amanda L. Stronza. Campesino hunting and conservation in Latin America. Conservation Biology, 34(2):338–353, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Patrick:2024:BCH

[PS24] Patricia G. Patrick and Abdul Rahman Singkam. Biodiversity conservation, human–animal interactions, and zootherapy in ecological knowledge of Indonesian healers. Conservation Biology, 38(6):e14278:1–e14278:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Powlen:2023:ISB

[PSJG23] Kathryn A. Powlen, Jonathan Salerno, Kelly W. Jones, and Michael C. Gavin. Identifying socioeconomic and biophysical factors driving forest loss in protected areas. Conservation Biology, 37(4):e14058:1–e14058:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pritchard:2022:DJB

[PSO^{+}22] Rose Pritchard, Laura Aileen Sauls, Johan A. Oldekop, Wilhelm Andrew Kiwango, and Dan Brockington. Data justice and biodiversity conservation. Conservation Biology, 36(5): e13919:1–e13919:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pellissier:2020:ENN

[PSP^{+}20] V. Pellissier, R. Schmucki, G. Pe’er, A. Aunins, T. M. Brereton, L. Brotons, J. Carnicer, T. Chodkiewicz, P. Chylarecki, J. C. del Moral, V. Escandell, D. Evans, R. Foppen, A. Harpke, J. Heliölä, S. Herrando, M. Kuussaari, E. Kühn, A. Lehikoinen, Å. Lindström, C. M. Moshøj, M. Musche, D. Noble, T. H. Oliver, J. Reif, D. Richard, D. B. Roy, O. Schweiger, J. Settele, C. Stefanescu, N. Teufelbauer, J. Touroult, S. Trautmann, A. J. van Strien, C. A. M. van Swaay, C. van Turnhout, Z. Vermouzek, P. Voříšek, F. Jiguet, and R. Julliard. Effects of Natura 2000 on nontarget bird and butterfly species based on citizen science data. Conservation Biology, 34(3):666–676, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Phelps:2020:TEC

[PSS20] Michael P. Phelps, Lisa W. Seeb, and James E. Seeb. Transforming ecology and conservation biology through genome editing. Conservation Biology, 34(1):54–65, February 2020. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Proverbs:2021:DEH

[PSV^{+}21] Tracey A. Proverbs, Abraham R. Stewart, Alice Vittrekwa, Ernest Vittrekwa, Rachel A. Hovel, and Emma E. Hodgson. Disrupted ecosystem and human phenology at the climate frontline in Gwich’in First Nation territory. Conservation Biology, 35(4):1348–1352, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Pascual:2021:BGT

[PWB^{+}21] Maria Pascual, James Wingard, Naila Bhatri, Alyona Rydannykh, and Jacob Phelps. Building a global taxonomy of wildlife offenses. Conservation Biology, 35(6):1903–1912, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Post:2022:ACS

[PWW^{+}22] John R. Post, Hillary G. M. Ward, Kyle L. Wilson, George L. Sterling, Ariane Cantin, and Eric B. Taylor. Assessing conservation status with extensive but low-resolution data: Application of frequentist and Bayesian models to endangered Athabasca River rainbow trout. Conservation Biology, 36(3): e13783:1–e13783:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quartuch:2025:SSI

[QGC^{+}25] Michael R. Quartuch, Ashley Gramza, Chelsey Crandall, Emily Pomeranz, Rene Valdez, Natalie Sexton, Ann B. Forstchen, and Coren Jagnow. Social science integration at state and federal fish and wildlife organizations in the United States. Conservation Biology, 39(2):e70004:1–e70004:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quigley:2022:TGR

[QHS22] Kate M. Quigley, Margaux Hein, and David J. Suggett. Translating the 10 golden rules of reforestation for coral reef restoration. Conservation Biology, 36(4):e13890:1–e13890:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Quinlan:2021:UFC

[QOF^{+}21] Jessica R. Quinlan, Shannon J. O’Leary, Andrew T. Fields, Martin Benavides, Emily Stumpf, Ramon Carcamo, Joel Cruz, Derrick Lewis, Beverly Wade, George Amato, Sergios-Orestis Kolokotronis, Gina M. Clementi, and Demian D. Chapman. Using fisher-contributed secondary fins to fill critical sharkfisheries data gaps. Conservation Biology, 35(3):991–1001, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Qin:2024:ADD

[QZX^{+}24] Sheng-Yuan Qin, Zheng-Yu Zuo, Shuang-Xiu Xu, Jie Liu, Feng-Mao Yang, Ya-Huang Luo, Jun-Wei Ye, Yao Zhao, Jun Rong, Bing Liu, Peng-Fei Ma, and De-Zhu Li. Anthropogenic disturbance driving population decline of a dominant tree in East Asia evergreen broadleaved forests over the last 11,000 years. Conservation Biology, 38(1):e14180:1–e14180:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rothenberger:2020:BID

[RAG^{+}20] Megan Rothenberger, Andrea Armstrong, Trent Gaugler, Sarah Massaro, William Pfadenhauer, and Juliana Ventresca. Bridging information domains to improve ecological understanding of biological invasions in a marine ecosystem. Conservation Biology, 34(6):1560–1570, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rapson:2021:TAP

[Rap21] Gillian L. Rapson. Testing the acid in plant succession. Conservation Biology, 35(5):1704–1705, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rasolofoson:2022:SMC

[Ras22] Ranaivo A. Rasolofoson. Statistical matching for conservation science revisited: response to Schleicher et al. 2020. Conservation Biology, 36(6):e14006:1–e14006:??, December 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [SEB^{+}20] and reply [JBE^{+}22b].

Ratto:2022:RAI

[RBC^{+}22] Fabrizia Ratto, Tom D. Breeze, Lorna J. Cole, Michael P. D. Garratt, David Kleijn, Bill Kunin, Denis Michez, Rory O’Connor, Jeff Ollerton, Robert J. Paxton, Guy M. Poppy, Simon G. Potts, Deepa Senapathi, Rosalind Shaw, Lynn V. Dicks, and Kelvin S.-H. Peh. Rapid assessment of insect pollination services to inform decision-making. Conservation Biology, 36(4):e13886:1–e13886:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2021:EOG

[RBD^{+}21] Zachary L. Robinson, Donovan A. Bell, Tashi Dhendup, Gordon Luikart, Andrew R. Whiteley, and Marty Kardos. Evaluating the outcomes of genetic rescue attempts. Conservation Biology, 35(2):666–677, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rowland:2021:GRV

[RBJN21] Jessica A. Rowland, Lucie M. Bland, Simon James, and Emily Nicholson. A guide to representing variability and uncertainty in biodiversity indicators. Conservation Biology, 35(5):1669– 1682, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reed:2021:IGE

[RBLN21] Graeme Reed, Nicolas D. Brunet, Sheri Longboat, and David C. Natcher. Indigenous guardians as an emerging approach to indigenous environmental governance. Conservation Biology, 35(1):179–189, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:2021:ACS

[RCBT21] Kelsey E. Roberts, Carly N. Cook, Jutta Beher, and Eric A. Treml. Assessing the current state of ecological connectivity in a large marine protected area system. Conservation Biology, 35(2):699–710, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ramsey:2022:QPD

[RCL^{+}22] David S. L. Ramsey, Karl J. Campbell, Christian Lavoie, Norm Macdonald, and Scott A. Morrison. Quantifying the probability of detection of wild ungulates with the Judas technique.

Conservation Biology, 36(4):e13898:1–e13898:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rees:2020:RAC

[RCRF20] Matthew W. Rees, Josie Carwardine, Andrew Reeson, and Jennifer Firn. Rapidly assessing cobenefits to advance threatmanagement alliances. Conservation Biology, 34(4):843–853, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rojas-Castillo:2024:MAC

[RCRJ^{+}24] Oscar Alberto Rojas-Castillo, Sebastian Kepfer Rojas, Leandro Juen, Luciano Fogaça de Assis Montag, Fernando Geraldo Carvalho, Thiago Pereira Mendes, Kenny Wei Jie Chua, Clare L. Wilkinson, Mohammad Noor Azmai Amal, Muhammad Fahmi-Ahmad, and Dean Jacobsen. Meta-analysis contrasting freshwater biodiversity in forests and oil palm plantations with and without riparian buffers. Conservation Biology, 38(1):e14172:1–e14172:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ricci:2024:MAN

[RDS^{+}24] Lorenzo Ricci, Michele Di Musciano, Francesco Maria Sabatini, Alessandro Chiarucci, Piero Zannini, Roberto Cazzolla Gatti, Carl Beierkuhnlein, Anna Walentowitz, Alexandra Lawrence, Anna Rita Frattaroli, and Samuel Hoffmann. A multitaxonomic assessment of Natura 2000 effectiveness across European biogeographic regions. Conservation Biology, 38(3):e14212:1– e14212:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Redford:2023:ECI

[Red23] Kent H. Redford. Extending conservation to include Earth’s microbiome. Conservation Biology, 37(3):e14088:1–e14088:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Riva:2023:OBC

[RF23] Federico Riva and Lenore Fahrig. Obstruction of biodiversity conservation by minimum patch size criteria. Conservation Biology, 37(5):e14092:1–e14092:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rhodes:2020:FIW

[RGBC20] Jonathan R. Rhodes, Angela M. Guerrero, Örjan Bodin, and Iadine Chadès. Fundamental insights on when social network data are most critical for conservation planning. Conservation Biology, 34(6):1463–1472, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Romer:2025:ESF

[RGD^{+}25] Alexander S. Romer, Matthew Grisnik, Jason W. Dallas, William Sutton, Christopher M. Murray, Rebecca H. Hardman, Tom Blanchard, Ryan J. Hanscom, Rulon W. Clark, Cody Godwin, N. Reed Alexander, Kylie C. Moe, Vincent A. Cobb, Jesse Eaker, Rob Colvin, Dustin Thames, Chris Ogle, Josh Campbell, Carlin Frost, Rachel L. Brubaker, Shawn D. Snyder, Alexander J. Rurik, Chloe E. Cummins, David W. Ludwig, Joshua L. Phillips, and Donald M. Walker. Effects of snake fungal disease (ophidiomycosis) on the skin microbiome across two major experimental scales. Conservation Biology, 39 (2):e14411:1–e14411:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reyes-Galindo:2023:EPR

[RGJCN^{+}23] Verónica Reyes-Galindo, Juan Pablo Jaramillo-Correa, Karina Carrasco Nava, Alejandra Elizabeth De la Rosa-González, David Flores Flores, Mauricio Martínez, Luis Alberto Monroy-De la Rosa, Miguel Ángel Morelos Zamora, Billy Emmanuel Ramírez Morales, Oliver Tanui Ramírez Morales, María del Pilar Rodríguez, Maurilio Salazar Zamora, Claudio Zamora Callejas, Rafael Zamora Callejas, César Zamora, Tonatiuh Zamora, Victor Alejandro González-Camacho, Erick Rebollo, Ricardo Torres-Jardón, Ana Wegier, and Alicia Mastretta-Yanes. Evaluating pollution-related damage and restoration success in urban forests with participatory monitoring and digital tools. Conservation Biology, 37(5):e14112:1– e14112:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roberts:2023:UDE

[RHFV23] David L. Roberts, Amy Hinsley, Sicily Fiennes, and Diogo Veríssimo. Understanding the drivers of expert opinion when classifying species as extinct. Conservation Biology, 37(1): e13968:1–e13968:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rivera-Hechem:2021:EEA

[RHGRSG21] María I. Rivera-Hechem, Ricardo A. Guzmán, Carlos Rodríguez-Sickert, and Stefan Gelcich. Effects of experience with access regimes on stewardship behaviors of small-scale fishers. Conservation Biology, 35(6):1913–1922, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ruegg:2020:GNM

[RHS^{+}20] Kristen C. Ruegg, Ryan J. Harrigan, James F. Saracco, Thomas B. Smith, and Caz M. Taylor. A genoscape-network model for conservation prioritization in a migratory bird. Conservation Biology, 34(6):1482–1491, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Richards:2021:ASC

[Ric21] Jeannine H. Richards. Assessing the strength of climate and land-use influences on montane epiphyte communities. Conservation Biology, 35(5):1496–1506, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rice:2022:IDE

[Ric22] Wayne Stanley Rice. Identifying and developing effective post-2020 conservation bridging leaders. Conservation Biology, 36 (6):e13980:1–e13980:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rayden:2023:ICB

[RJAR23] Tim Rayden, Kendall R. Jones, Kemen Austin, and Jeremy Radachowsky. Improving climate and biodiversity outcomes through restoration of forest integrity. Conservation Biology, 37(6):e14163:1–e14163:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rojas:2022:LSF

[RJC^{+}22] Isabel M. Rojas, Megan K. Jennings, Erin Conlisk, Alexandra D. Syphard, Jack Mikesell, Alicia M. Kinoshita, Krista West, Doug Stow, Emanuel Storey, Mark E. De Guzman, Diane Foote, Alexandria Warneke, Amber Pairis, Sherry Ryan, Lorraine E. Flint, Alan L. Flint, and Rebecca L. Lewison.

A landscape-scale framework to identify refugia from multiple stressors. Conservation Biology, 36(1):e13834:1–e13834:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rivera:2024:ARO

[RJV^{+}24] Shannon Noelle Rivera, Laure Joanny, Isabel Vique, Rebekah Middleton, and Diogo Veríssimo. Assessing the risk of overexploitation to a tarantula species in the pet trade. Conservation Biology, 38(5):e14362:1–e14362:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roth:2021:NEN

[RKR^{+}21] Tobias Roth, Lukas Kohli, Beat Rihm, Reto Meier, and Valentin Amrhein. Negative effects of nitrogen deposition on Swiss butterflies. Conservation Biology, 35(6):1766–1776, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ricci:2024:EYB

[RLSH24] Kyra Ricci, Kathleen Lu, Grascen Shidemantle, and Jessica Hua. Engaging youth in biodiversity education through visual narrative. Conservation Biology, 38(6):e14386:1–e14386:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ryan:2020:EBC

[RMD^{+}20] Jillian Ryan, Sarah Mellish, Jillian Dorrian, Tony Winefield, and Carla Litchfield. Effectiveness of biodiversity-conservation marketing. Conservation Biology, 34(2):354–367, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roast:2023:HDI

[RMFP^{+}23] Michael J. Roast, Samir Martins, Lourdes Fernández-Peralta, José Carlos Báez, Ahmed Diame, David March, Jazel Ouled-Cheikh, Adolfo Marco, Jacob González-Solís, and Luis Cardona. Hidden demographic impacts of fishing and environmental drivers of fecundity in a sea turtle population. Conservation Biology, 37(5):e14110:1–e14110:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Reum:2020:RAM

[RML^{+}20] Jonathan C. P. Reum, P. Sean McDonald, W. Christopher Long, Kirstin K. Holsman, Lauren Divine, David Armstrong, and Jan Armstrong. Rapid assessment of management options for promoting stock rebuilding in data-poor species under climate change. Conservation Biology, 34(3):611–621, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Romanelli:2021:RER

[RMN^{+}21] João P. Romanelli, Paula Meli, Rafaela P. Naves, Marcelo C. Alves, and Ricardo R. Rodrigues. Reliability of evidence-review methods in restoration ecology. Conservation Biology, 35(1): 142–154, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rowland:2023:HEI

[RMW23] Jessica A. Rowland, Joslin L. Moore, and Jessica C. Walsh. How expert insight into alpine peatland conservation complements global scientific evidence. Conservation Biology, 37(3): e14081:1–e14081:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robert:2024:BRN

[Rob24] Alexandre Robert. Building references for nature conservation. Conservation Biology, 38(2):e14202:1–e14202:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Roche:2022:CKA

[ROK^{+}22] Dominique G. Roche, Rose E. O’Dea, Kecia A. Kerr, Trina Rytwinski, Richard Schuster, Vivian M. Nguyen, Nathan Young, Joseph R. Bennett, and Steven J. Cooke. Closing the knowledge-action gap in conservation with open science. Conservation Biology, 36(3):e13835:1–e13835:??, June 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Robinson:2024:DLN

[RR24] Janine E. Robinson and David L. Roberts. Determining the legality of newly described CITES-listed species in the horticulture trade of tropical pitcher plants (Nepenthes). Conservation Biology, 38(5):e14361:1–e14361:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ridley:2024:GMB

[RRH^{+}24] Francesca A. Ridley, Stephen P. Rushton, Emily J. Hickinbotham, Andrew J. Suggitt, Philip J. K. McGowan, and Louise Mair. Global mismatches between threat mapping research effort and the potential of threat abatement actions to reduce extinction risk. Conservation Biology, 38(4):e14271:1–e14271:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rasambainarivo:2024:AIP

[RRR^{+}24] Fidisoa T. Rasambainarivo, Santatra Randrianarisoa, Olivier A. Rasolofoniaina, Benjamin L. Rice, and C. Jessica E. Metcalf. Assessing the impact of preventative measures to limit the spread of Toxoplasma gondii in wild carnivores of Madagascar. Conservation Biology, 38(6):e14300:1–e14300:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Riisager-Simonsen:2020:UES

[RSRG^{+}20] Christian Riisager-Simonsen, Olivia Rendon, Anders Galatius, Morten Tange Olsen, and Nicola Beaumont. Using ecosystemservices assessments to determine trade-offs in ecosystem-based management of marine mammals. Conservation Biology, 34(5): 1152–1164, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Regos:2020:IIG

[RVLD20] Adrián Regos, María Vidal, Miguel Lorenzo, and Jesús Domínguez. Integrating intraseasonal grassland dynamics in cross-scale distribution modeling to support waterbird recovery plans. Conservation Biology, 34(2):494–504, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Radinger:2022:ECD

[RvTW22] Johannes Radinger, Ruben van Treeck, and Christian Wolter. Evident but context-dependent mortality of fish passing hydroelectric turbines. Conservation Biology, 36(3):e13870:1– e13870:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Read:2023:AIE

[RW23] Daniel J. Read and Lisa Wainger. Assessing intervention effectiveness at promoting voluntary conservation practice adoption in agrienvironments. Conservation Biology, 37(1):e14009:1– e14009:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ronsted:2022:ERE

[RWC^{+}22] Nina Rønsted, Seana K. Walsh, Margaret Clark, Merlin Edmonds, Tim Flynn, Scott Heintzman, Alexander Loomis, David Lorence, Uma Nagendra, Ben Nyberg, Michael Opgenorth, Lauren Weisenberger, Adam Williams, Dustin Wolkis, Kenneth R. Wood, and Matthew Keir. Extinction risk of the endemic vascular flora of Kauai, Hawaii, based on IUCN assessments. Conservation Biology, 36(4):e13896:1–e13896:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Rampling:2024:ABN

[RzEHB24] Emily E. Rampling, Sophus O. S. E. zu Ermgassen, Isobel Hawkins, and Joseph W. Bull. Achieving biodiversity net gain by addressing governance gaps underpinning ecological compensation policies. Conservation Biology, 38(2):e14198:1– e14198:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stark:2022:IEC

[SAB^{+}22] Keila Stark, Vanessa Adams, Christopher J. Brown, Alienor Chauvenet, Katrina Davis, Edward T. Game, Benjamin S. Halpern, John Lynham, Bonnie Mappin, Kim Selkoe, James E. M. Watson, Hugh P. Possingham, and Carissa J. Klein. Importance of equitable cost sharing in the Convention on Biological Diversity’s protected area agenda. Conservation Biology, 36(2):e13812:1–e13812:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sutherland:2020:ETC

[SACA^{+}20] William J. Sutherland, Sergio Ticul Alvarez-Castañeda, Tatsuya Amano, Roberto Ambrosini, Philip Atkinson, John M. Baxter, Alexander L. Bond, Philip J. Boon, Katherine L. Buchanan, Jos Barlow, Giuseppe Bogliani, Olivia M. Bragg, Mark Burgman, Marc W. Cadotte, Michael Calver, Steven J.

Cooke, Richard T. Corlett, Vincent Devictor, John G. Ewen, Martin Fisher, Guy Freeman, Edward Game, Brendan J. Godley, Christian Gortázar, Ian R. Hartley, David L. Hawksworth, Keith A. Hobson, Ming-Lun Lu, Berta Martín-López, Keping Ma, Antonio Machado, Dirk Maes, Marco Mangiacotti, Dominic J. McCafferty, Victoria Melfi, Sanjay Molur, Allen J. Moore, Stephen D. Murphy, Darren Norris, Alexander P. E. van Oudenhoven, Jennifer Powers, Eileen C. Rees, Mark W. Schwartz, Ilse Storch, and Claire Wordley. Ensuring tests of conservation interventions build on existing literature. Conservation Biology, 34(4):781–783, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sagoff:2020:FVIa

[Sag20a] Mark Sagoff. Fact and value in invasion biology. Conservation Biology, 34(3):581–588, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [CBB^{+}20].

Sagoff:2020:FVIb

[Sag20b] Mark Sagoff. Fact and value in invasion biology: reply to Cuthbert et al. 2020. Conservation Biology, 34(6):1583–1585, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [CBB^{+}20].

Salvidio:2024:EHW

[Sal24] Sebastiano Salvidio. Expanding the human–wildlife interaction framework of Galindo–González 2023. Conservation Biology, 38(6):e14374:1–e14374:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sandler:2020:EGE

[San20] Ronald Sandler. The ethics of genetic engineering and gene drives in conservation. Conservation Biology, 34(2):378–385, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sarospataki:2020:PTI

[Sár20] Miklós Sárospataki. Pollinators and their importance. Conservation Biology, 34(2):530, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sarospataki:2023:BRT

[Sár23] Miklós Sárospataki. Book reviews: Taxonomy, behavior, and intelligence of bees Bumblebees of Europe and Neighbouring Regions By Rasmont P., Ghisbain G. and Terzo M. 2021. NAP Editions, Verrieres le Buisson, France. 632 pp. EUR 87. (hardback). ISBN 978-2-913688-38-4. The Mind of a Bee, By Chittka L. 2022. Princeton University Press, Princeton, NJ, U.S.A. 272 pp. US$29.95 (hardcover). ISBN 978-0-691-18047-2. Conservation Biology, 37(2):e14085:1–e14085:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Santiago-Avila:2021:TMJ

[SÁT21] Francisco J. Santiago-Ávila and Adrian Treves. Toward multispecies justice in human–wildlife coexistence: reply to Clark et al. Conservation Biology, 35(4):1337–1340, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stirling:2021:SCS

[SB21] Andy Stirling and Mark A. Burgman. Strengthening conservation science as a crisis discipline by addressing challenges of precaution, privilege, and individualism. Conservation Biology, 35(6):1738–1746, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simons:2025:IES

[SBE25] D.-L. Simons, R. B. Bradbury, and K. L. Evans. Impacts of ecosystem service message framing and dynamic social norms on public support for tropical forest restoration. Conservation Biology, 39(2):e14373:1–e14373:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salerno:2020:WIV

[SBG^{+}20] Jonathan Salerno, Karen Bailey, Andrea E. Gaughan, Forrest R. Stevens, Tom Hilton, Lin Cassidy, Michael D. Drake, Narcisa G. Pricope, and Joel Hartter. Wildlife impacts and vulnerable livelihoods in a transfrontier conservation landscape. Conservation Biology, 34(4):891–902, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Selinske:2022:PBB

[SBG^{+}22] Matthew J. Selinske, Sarah A. Bekessy, William L. Geary, Richard Faulkner, Fern Hames, Charlotte Fletcher, Zoe E. Squires, and Georgia E. Garrard. Projecting biodiversity benefits of conservation behavior-change programs. Conservation Biology, 36(3):e13845:1–e13845:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Santostasi:2025:SEW

[SBG^{+}25] Nina Luisa Santostasi, Sarah Bauduin, Oksana Grente, Olivier Gimenez, and Paolo Ciucci. Simulating the efficacy of wolf– dog hybridization management with individual-based modeling. Conservation Biology, 39(1):e14312:1–e14312:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schuster:2023:PAP

[SBH^{+}23] Richard Schuster, Rachel Buxton, Jeffrey O. Hanson, Allison D. Binley, Jeremy Pittman, Vivitskaia Tulloch, Frank A. La Sorte, Patrick R. Roehrdanz, Peter H. Verburg, Amanda D. Rodewald, Scott Wilson, Hugh P. Possingham, and Joseph R. Bennett. Protected area planning to conserve biodiversity in an uncertain future. Conservation Biology, 37(3):e14048:1– e14048:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwartz:2023:ECR

[SBP^{+}23] Timothée Schwartz, Aurélien Besnard, Christophe Pin, Olivier Scher, Thomas Blanchon, Arnaud Béchet, and Nicolas Sadoul. Efficacy of created and restored nesting sites for the conservation of colonial Laridae in the South of France. Conservation Biology, 37(2):e14005:1–e14005:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soto:2022:HDS

[SBW^{+}22] Juan S. Vargas Soto, Christopher Beirne, Andrew Whitworth, Juan Carlos Cruz Diaz, Eleanor Flatt, Ruthmery Pillco-Huarcaya, Erik R. Olson, Alejandro Azofeifa, Guido Saborío-R, Roberto Salom-Pérez, Deiver Espinoza-Muñoz, Leslie Hay, Lawrence Whittaker, Carmen Roldán, Ricardo Bedoya-Arrieta, Eben North Broadbent, and Péter K. Molnár. Human disturbance and shifts in vertebrate community composition in a biodiversity hotspot. Conservation Biology, 36(2):e13813:1– e13813:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saltz:2023:NPP

[SC23] David Saltz and Shlomo Cohen. Naturalness and principle pluralism in conservation. Conservation Biology, 37(6):e14137:1– e14137:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Skroblin:2021:IIK

[SCB^{+}21] Anja Skroblin, Tracy Carboon, Gladys Bidu, Nganjapayi Chapman, Minyawu Miller, Karnu Taylor, Waka Taylor, Edward T. Game, and Brendan A. Wintle. Including indigenous knowledge in species distribution modeling for increased ecological insights. Conservation Biology, 35(2):587–597, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Santos:2025:RRF

[SCB^{+}25] Xavier Santos, Brahim Chergui, Josabel Belliure, Francisco Moreira, and Juli G. Pausas. Reptile responses to fire across the western Mediterranean Basin. Conservation Biology, 39(1): e14326:1–e14326:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2023:MAE

[SCC23] Rachel S. Smith, Selina L. Cheng, and Max C. N. Castorani. Meta-analysis of ecosystem services associated with oyster restoration. Conservation Biology, 37(1):e13966:1–e13966:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Salafsky:2024:SLT

[SCdM^{+}24] Nick Salafsky, Paola Mejía Cortez, Kalli de Meyer, Nigel Dudley, Helen Klimmek, Alanah Lewis, Duncan MacRae, Brent A. Mitchell, Kent H. Redford, and Mitali Sharma. A standard lexicon of terms for area-based conservation version 1.0. Conservation Biology, 38(4):e14269:1–e14269:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwartz:2021:CLT

[Sch21] Mark W. Schwartz. Conservation lessons from taboos and trolley problems. Conservation Biology, 35(3):794–803, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:2020:BOW

[Sco20] Michael Scott. Birds of the oceans and waterways. Conservation Biology, 34(3):771–773, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:2021:PVF

[Sco21] Michael Scott. A place for volunteers in field sciences. Conservation Biology, 35(1):386–388, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:2022:SIC

[Sco22] Michael Scott. Solving identification challenges in cetaceans and seabirds. Conservation Biology, 36(2):e13899:1–e13899:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scott:2024:CM

[Sco24] M. Scott. A conservation masterclass. Conservation Biology, 38(6):e14377:1–e14377:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sopniewski:2024:NBA

[SCW^{+}24] Jarrod Sopniewski, Renee Catullo, Michelle Ward, Nicola Mitchell, and Ben C. Scheele. Niche-based approach to explore the impacts of environmental disturbances on biodiversity. Conservation Biology, 38(6):e14277:1–e14277:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sitters:2020:IFC

[SD20] Holly Sitters and Julian Di Stefano. Integrating functional connectivity and fire management for better conservation outcomes. Conservation Biology, 34(3):550–560, June 2020. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smyth:2022:CFI

[SD22] Eric R. B. Smyth and D. Andrew R. Drake. A classification framework for interspecific trade-offs in aquatic ecology. Conservation Biology, 36(1):e13762:1–e13762:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simkins:2024:RTC

[SDB^{+}24] Ashley Thomas Simkins, Paul F. Donald, Alison E. Beresford, Stuart H. M. Butchart, Julia E. Fa, Alvaro Onrubia Fernández-Llamazares, Stephen T. Garnett, and Graeme M. Buchanan. Rates of tree cover loss in key biodiversity areas on Indigenous Peoples’ lands. Conservation Biology, 38(3):e14195:1– e14195:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sibanda:2025:AIC

[SDH^{+}25] Lovemore Sibanda, Amy Dickman, Courtney Hughes, Jessica Tacey, Emily Madsen, Lessah Mandoloma, Moreangels M. Mbizah, Yolanda Mutinhima, Betty Rono, Salum Kulunge, David Kimaili, Trisha Bhujle, David W. Macdonald, and Darragh Hare. Avoiding an impending collision in international conservation. Conservation Biology, 39(2):e14450:1–e14450:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simpson:2022:EEI

[SdVD^{+}22] Katherine Hannah Simpson, Frans P. de Vries, Martin Dallimer, Paul R. Armsworth, and Nick Hanley. Ecological and economic implications of alternative metrics in biodiversity offset markets. Conservation Biology, 36(5):e13906:1–e13906:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schleicher:2020:SMC

[SEB^{+}20] Judith Schleicher, Johanna Eklund, Megan D. Barnes, Jonas Geldmann, Johan A. Oldekop, and Julia P. G. Jones. Statistical matching for conservation science. Conservation Biology, 34(3):538–549, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [May22].

Stanley:2024:APC

[SENTI24] Annabelle E. Stanley, Rebecca Epanchin-Niell, Tyler Treakle, and Gwenllian D. Iacona. Attributes of preemptive conservation efforts for species precluded from listing under the U.S. Endangered Species Act. Conservation Biology, 38(2):e14200:1– e14200:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Singeo:2023:LPE

[SF23] Ann Singeo and Caroline E. Ferguson. Lessons from Palau to end parachute science in international conservation research. Conservation Biology, 37(1):e13971:1–e13971:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stahl:2020:MLA

[SFC20] Amanda T. Stahl, Alexander K. Fremier, and Barbara A. Cosens. Mapping legal authority for terrestrial conservation corridors along streams. Conservation Biology, 34(4):943–955, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soofi:2024:AIT

[SGA^{+}24] Mahmood Soofi, Benjamin Ghasemi, Mohsen Ahmadpour, Mobin Soufi, Iman Islami, Alaina Eckert, Mohammad Hossein Gorjian Arabi, Ali T. Qashqaei, Javad Selyari, Kamran Nasirahmadi, Johannes Kamp, Matthias Waltert, Jane Addison, and Chris R. Pavey. Application of the integrated threat theory to conservation law enforcement. Conservation Biology, 38(6):e14248:1–e14248:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Santangeli:2020:ICA

[SGB^{+}20] Andrea Santangeli, Marco Girardello, Evan R. Buechley, Andre Botha, Enrico Di Minin, and Atte Moilanen. Importance of complementary approaches for efficient vulture conservation: reply to Efrat et al. Conservation Biology, 34(5):1308–1310, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schutter:2021:SAB

[SH21] Marleen S. Schutter and Christina C. Hicks. Speaking across boundaries to explore the potential for interdisciplinarity in ecosystem services knowledge production. Conservation Biology, 35(4):1198–1209, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Soares:2023:RLC

[SHA^{+}23] António O. Soares, Danny Haelewaters, Olga M. C. C. Ameixa, Isabel Borges, Peter M. J. Brown, Pedro Cardoso, Michiel D. de Groot, Edward W. Evans, Audrey A. Grez, Axel Hochkirch, Milada Holecová, Alois Honěk, Ján Kulfan, Ana I. Lillebø, Zdenka Martinková, J. P. Michaud, Oldřich Nedvěd, Omkar, Helen E. Roy, Swati Saxena, Apoorva Shandilya, Arnaud Sentis, Jiri Skuhrovec, Sandra Viglášová, Peter Zach, Tania Zaviezo, and John E. Losey. A roadmap for ladybird conservation and recovery. Conservation Biology, 37(1):e13965:1–e13965:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schmidt:2023:GDI

[SHH^{+}23] Chloé Schmidt, Sean Hoban, Margaret Hunter, Ivan Paz-Vinas, and Colin J. Garroway. Genetic diversity and IUCN Red List status. Conservation Biology, 37(4):e14064:1–e14064:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Strampelli:2022:HUT

[SHS^{+}22] Paolo Strampelli, Philipp Henschel, Charlotte E. Searle, David W. Macdonald, and Amy J. Dickman. Habitat use of and threats to African large carnivores in a mixed-use landscape. Conservation Biology, 36(6):e13943:1–e13943:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sharma:2023:EIP

[SJRG23] Roshan Sharma, Simon Jones, Doug Robinson, and Ascelin Gordon. Evaluating the impact of private land conservation with synthetic control design. Conservation Biology, 37(6): e14150:1–e14150:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Saif:2022:MCC

[SKS22] Omar Saif, Aidan Keane, and Sam Staddon. Making a case for the consideration of trust, justice, and power in conservation relationships. Conservation Biology, 36(4):e13903:1–e13903:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shen:2020:EMZ

[SLM^{+}20] Xiaoli Shen, Sheng Li, William J. McShea, Dajun Wang, Jianping Yu, Xiaogang Shi, Wei Dong, Xiangcheng Mi, and Keping Ma. Effectiveness of management zoning designed for flagship species in protecting sympatric species. Conservation Biology, 34(1):158–167, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanders:2020:PIM

[SMB^{+}20] Michele Jeanette Sanders, Laura Miller, Shonil A. Bhagwat, Jesse Marije Anne van der Grient, and Alex David Rogers. Practitioner insights as a means of setting a context for conservation. Conservation Biology, 34(1):113–124, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Siegel:2024:GMA

[SMLL24] Tovah Siegel, Ainhoa Magrach, William F. Laurance, and David Luther. A global meta-analysis of the impacts of forest fragmentation on biotic mutualisms and antagonisms. Conservation Biology, 38(3):e14206:1–e14206:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sullivan:2021:CID

[SMSS^{+}21] Lauren L. Sullivan, Matthew J. Michalska-Smith, Katie P. Sperry, David A. Moeller, and Allison K. Shaw. Consequences of ignoring dispersal variation in network models for landscape connectivity. Conservation Biology, 35(3):944–954, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwacke:2022:MPE

[SMT^{+}22] Lori H. Schwacke, Tiago A. Marques, Len Thomas, Cormac G. Booth, Brian C. Balmer, Ashley Barratclough, Kathleen Colegrove, Sylvain De Guise, Lance P. Garrison, Forrest M. Gomez, Jeanine S. Morey, Keith D. Mullin, Brian M. Quigley, Patricia E. Rosel, Teresa K. Rowles, Ryan Takeshita, Forrest I. Townsend, Todd R. Speakman, Randall S. Wells, Eric S. Zolman, and Cynthia R. Smith. Modeling population effects of the Deepwater Horizon oil spill on a long-lived species. Conservation Biology, 36(4):e13878:1–e13878:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stevens:2022:MME

[SN22] Madison Stevens and D. Ryan Norris. A mixed methodology for evaluating use of evidence in conservation planning. Conservation Biology, 36(4):e13876:1–e13876:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schoen:2022:SHC

[SNC^{+}22] Jay M. Schoen, Amrita Neelakantan, Samuel A. Cushman, Trishna Dutta, Bilal Habib, Yadvendradev V. Jhala, Indranil Mondal, Uma Ramakrishnan, P. Anuradha Reddy, Swati Saini, Sandeep Sharma, Prachi Thatte, Bibek Yumnam, and Ruth De Fries. Synthesizing habitat connectivity analyses of a globally important human-dominated tiger-conservation landscape. Conservation Biology, 36(4):e13909:1–e13909:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Speaker:2022:GCS

[SOW^{+}22] Talia Speaker, Stephanie O’Donnell, George Wittemyer, Brett Bruyere, Colby Loucks, Anthony Dancer, Marianne Carter, Eric Fegraus, Jonathan Palmer, Ellie Warren, and Jennifer Solomon. A global community-sourced assessment of the state of conservation technology. Conservation Biology, 36(3): e13871:1–e13871:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sosnowski:2021:ILW

[SP21] Monique C. Sosnowski and Gohar A. Petrossian. Illegal and legal wildlife trade analysis discourse: response to Natusch et al. 2021. Conservation Biology, 35(5):1689–1691, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [NAS21a, NAS21b].

Staley:2022:SSC

[SP22] Kent Staley and John Park. Statistical significance and conflicts of interest: Response to Mayo (2022). Conservation Biology, 36(5):e13985:1–e13985:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [May22].

Sales:2023:ICC

[SP23a] Lilian P. Sales and Mathias M. Pires. Identifying climate change refugia for South American biodiversity. Conservation Biology, 37(4):e14087:1–e14087:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sandvik:2023:MQH

[SP23b] Hanno Sandvik and Bård Pedersen. Metrics for quantifying how much different threats contribute to red lists of species and ecosystems. Conservation Biology, 37(5):e14105:1–e14105:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Scholte:2022:COL

[SPA^{+}22] Paul Scholte, Olivier Pays, Saleh Adam, Bertrand Chardonnet, Hervé Fritz, Jean-Baptiste Mamang, Herbert H. T. Prins, Pierre-Cyril Renaud, Patrick Tadjo, and Mark Moritz. Conservation overstretch and long-term decline of wildlife and tourism in the Central African savannas. Conservation Biology, 36(2): e13860:1–e13860:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Siegel:2022:IQQ

[SPK^{+}22] Katherine Siegel, Aldo Farah Perez, Eva Kinnebrew, Megan Mills-Novoa, José Ochoa, and Elizabeth Shoffner. Integration of qualitative and quantitative methods for land-use-change modeling in a deforestation frontier. Conservation Biology, 36 (6):e13924:1–e13924:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smith:2022:WAD

[SQJ^{+}22] William J. Smith, Claudio S. Quilodrán, Michał T. Jezierski, Ashley T. Sendell-Price, and Sonya M. Clegg. The wild ancestors of domestic animals as a neglected and threatened component of biodiversity. Conservation Biology, 36(3):e13867:1– e13867:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Snyder:2020:RAS

[SR20] Kristen Denninger Snyder and Dennis Rentsch. Rethinking assessment of success of mitigation strategies for elephantinduced crop damage. Conservation Biology, 34(4):829–842, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sramko:2023:DDE

[Sra23] Gábor Sramkó. A deep dive into evolutionary genomics and its use in conservation science. Conservation Biology, 37(4): e14111:1–e14111:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Souza:2023:EPB

[SRCS23] Barbara Almeida Souza, Josianne Claudia Sales Rosa, Pedro Bueno R. Campos, and Luis Enrique Sánchez. Evaluating the potential of biodiversity offsets to achieve net gain. Conservation Biology, 37(4):e14094:1–e14094:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schilbert:2023:LOM

[SS23] Jana Schilbert and Annette Scheersoi. Learning outcomes measured in zoo and aquarium conservation education. Conservation Biology, 37(1):e13891:1–e13891:??, February 2023. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Srednick:2024:EPT

[SS24] Griffin Srednick and Stephen E. Swearer. Effects of protection and temperature variation on temporal stability in a marine reserve network. Conservation Biology, 38(3):e14220:1– e14220:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Simmonds:2023:RNV

[SSCR^{+}23] Jeremy S. Simmonds, Andres Felipe Suarez-Castro, April E. Reside, James E. M. Watson, James R. Allan, Scott C. Atkinson, Pasquale Borrelli, Nigel Dudley, Stephen Edwards, Richard A. Fuller, Edward T. Game, Simon Linke, Sean L. Maxwell, Panos Panagos, Philippe Puydarrieux, Fabien Quétier, Rebecca K. Runting, Talitha Santini, Laura J.

Sonter, and Martine Maron. Retaining natural vegetation to safeguard biodiversity and humanity. Conservation Biology, 37 (3):e14040:1–e14040:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sykes:2020:ERF

[SSEN20] Louise Sykes, Luca Santini, Adrienne Etard, and Tim Newbold. Effects of rarity form on species’ responses to land use. Conservation Biology, 34(3):688–696, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shaw:2023:LLH

[SSG^{+}23] Robyn E. Shaw, Peter B. Spencer, Lesley A. Gibson, Judy A. Dunlop, Janine E. Kinloch, Karel Mokany, Margaret Byrne, Craig Moritz, Harriet Davie, Kenny J. Travouillon, and Kym M. Ottewell. Linking life history to landscape for threatened species conservation in a multiuse region. Conservation Biology, 37(1):e13989:1–e13989:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stewart-Sinclair:2021:SCB

[SSKBL21] Phoebe J. Stewart-Sinclair, Carissa J. Klein, Ian J. Bateman, and Catherine E. Lovelock. Spatial cost–benefit analysis of blue restoration and factors driving net benefits globally. Conservation Biology, 35(6):1850–1860, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shumway:2023:ERB

[SSN^{+}23] Nicole Shumway, Megan I. Saunders, Sam Nicol, Richard A. Fuller, Noam Ben-Moshe, Takuya Iwamura, Sun W. Kim, Nicholas J. Murray, James E. M. Watson, and Martine Maron. Exploring the risks and benefits of flexibility in biodiversity offset location in a case study of migratory shorebirds. Conservation Biology, 37(2):e14031:1–e14031:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stark:2022:RPV

[Sta22a] Philip B. Stark. Reproducibility, p-values, and type III errors: response to Mayo (2022). Conservation Biology, 36(5): e13986:1–e13986:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [May22].

Strzelecka:2022:ECN

[STA^{+}22b] Marianna Strzelecka, Joanna Tusznio, Arash Akhshik, Marcin Rechcinski, and Malgorzata Grodzinska-Jurczak. Effects of connection to nature on residents’ perceptions of conservation policy justice of Natura 2000. Conservation Biology, 36(5): e13944:1–e13944:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sakurai:2023:PAI

[STES23] Ryo Sakurai, Hiroshi Tsunoda, Hiroto Enari, and Richard C. Stedman. Public attitudes and intentions toward engaging in reintroduction of wolves to Japan. Conservation Biology, 37(6): e14130:1–e14130:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sigsgaard:2020:UVE

[STF^{+}20] Eva Egelyng Sigsgaard, Felipe Torquato, Tobias Guldberg Frøslev, Alec B. M. Moore, Johan Mølgård Sørensen, Pedro Range, Radhouane Ben-Hamadou, Steffen Sanvig Bach, Peter Rask Møller, and Philip Francis Thomsen. Using vertebrate environmental DNA from seawater in biomonitoring of marine habitats. Conservation Biology, 34(3):697–710, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Spanowicz:2020:APP

[STJ20] Ariel G. Spanowicz, Fernanda Zimmermann Teixeira, and Jochen A. G. Jaeger. An adaptive plan for prioritizing road sections for fencing to reduce animal mortality. Conservation Biology, 34(5):1210–1220, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stringham:2021:GUI

[STK^{+}21] Oliver C. Stringham, Adam Toomes, Aurelie M. Kanishka, Lewis Mitchell, Sarah Heinrich, Joshua V. Ross, and Phillip Cassey. A guide to using the internet to monitor and quantify the wildlife trade. Conservation Biology, 35(4):1130–1139, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sewall:2022:CSE

[STSJ22] Brent J. Sewall, Gregory G. Turner, Michael R. Scafini, and Joseph S. Johnson. Cooling subterranean environments for climate adaptation and disease management: reply to Meierhofer et al. Conservation Biology, 36(4):e13928:1–e13928:??, June 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shreedhar:2022:EEI

[STW22] Ganga Shreedhar and Laura Thomas-Walters. Experimental evidence of the impact of framing of actors and victims in conservation narratives. Conservation Biology, 36(6):e14015:1– e14015:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Schwacke:2024:EBS

[STW^{+}24] Lori H. Schwacke, Len Thomas, Randall S. Wells, Teresa K. Rowles, Gregory D. Bossart, Forrest Townsend, Jr., Marilyn Mazzoil, Jason B. Allen, Brian C. Balmer, Aaron A. Barleycorn, Ashley Barratclough, Louise Burt, Sylvain De Guise, Deborah Fauquier, Forrest M. Gomez, Nicholas M. Kellar, John H. Schwacke, Todd R. Speakman, Eric D. Stolen, Brian M. Quigley, Eric S. Zolman, and Cynthia R. Smith. An expert-based system to predict population survival rate from health data. Conservation Biology, 38(1):e14073:1–e14073:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Slatyer:2022:ERV

[SUA22] Rachel A. Slatyer, Kate D. L. Umbers, and Pieter A. Arnold. Ecological responses to variation in seasonal snow cover. Conservation Biology, 36(1):e13727:1–e13727:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sullivan:2024:HC

[Sul24] S. Sullivan. Hunting for conservation. Conservation Biology, 38 (6):e14258:1–e14258:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Shackleton:2022:CCD

[SVP^{+}22] Ross T. Shackleton, Giovanni Vimercati, Anna F. Probert, Sven Bacher, Christian A. Kull, and Ana Novoa. Consensus and controversy in the discipline of invasion science. Conservation Biology, 36(5):e13931:1–e13931:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stoffels:2025:QRL

[SW25] Rick J. Stoffels and Richard S. A. White. Quantifying regulatory limits for multiple stressors in an open and transparent way. Conservation Biology, 39(2):e14375:1–e14375:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stevenson:2020:TPM

[SWBD20] Simone L. Stevenson, Skipton N. C. Woolley, Jon Barnett, and Piers Dunstan. Testing the presence of marine protected areas against their ability to reduce pressures on biodiversity. Conservation Biology, 34(3):622–631, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Stevenson:2021:MBI

[SWC^{+}21] Simone L. Stevenson, Kate Watermeyer, Giovanni Caggiano, Elizabeth A. Fulton, Simon Ferrier, and Emily Nicholson. Matching biodiversity indicators to policy needs. Conservation Biology, 35(2):522–532, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Smallhorn-West:2022:LSS

[SWCP^{+}22] Patrick Smallhorn-West, Philippa J. Cohen, Michael Phillips, Stacy D. Jupiter, Hugh Govan, and Robert L. Pressey. Linking small-scale fisheries co-management to U.N. Sustainable Development Goals. Conservation Biology, 36(6):e13977:1– e13977:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sanders:2024:ERW

[SWM^{+}24] Emmalie Sanders, Skye Wassens, Damian R. Michael, Dale G. Nimmo, and James M. Turner. Extinction risk of the world’s freshwater mammals. Conservation Biology, 38(1):e14168:1– e14168:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sithole:2024:YEG

[SWMM24] Samantha S. Sithole, Gretchen M. Walters, Philile Mbatha, and Frank Matose. Youth engagement in global conservation governance. Conservation Biology, 38(6):e14387:1–e14387:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Sousa:2023:RCF

[SZH^{+}23] Ronaldo Sousa, Tadeusz Zajac, Dariusz Halabowski, Olga V. Aksenova, Yulia V. Bespalaya, Francisco Carvalho, Paulo Castro, Karel Douda, Janine P. da Silva, Noé Ferreira-Rodríguez, Juergen Geist, Clemens Gumpinger, Anna M. Labecka, Jasna Lajtner, Iga Lewin, Manuel Lopes-Lima, Alexandra Meira, Keiko Nakamura, Joana Garrido Nogueira, Paz Ondina, Małgorzata Ożgo, Joaquim Reis, Nicoletta Riccardi, Spase Shumka, Mikhail O. Son, Amílcar Teixeira, Frankie Thielen, Maria Urbańska, Simone Varandas, Niklas Wengström, Katarzyna Zajac, Alexandra Zieritz, and David C. Aldridge. A roadmap for the conservation of freshwater mussels in Europe. Conservation Biology, 37(2):e13994:1–e13994:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tims:2022:PBC

[TA22] Amy R. Tims and John Alroy. Phylogeny-based conservation priorities for Australian freshwater fishes. Conservation Biology, 36(2):e13811:1–e13811:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thumsova:2024:RBT

[TAA^{+}24] Barbora Thumsová, Gonzalo Alarcos, Cesar Ayres, Gonçalo M. Rosa, and Jaime Bosch. Relationship between two pathogens in an amphibian community that experienced mass mortalities. Conservation Biology, 38(3):e14196:1–e14196:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2020:MSS

[Tay20] Bron Taylor. Michael Soulé (1936–2020) on spirituality, ethics, and conservation biology. Conservation Biology, 34(6):1426– 1432, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tukiainen:2023:EGN

[TB23] Helena Tukiainen and Joseph J. Bailey. Enhancing global nature conservation by integrating geodiversity in policy and practice. Conservation Biology, 37(3):e14024:1–e14024:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Trimbach:2025:HWB

[TB25] David J. Trimbach and Kelly Biedenweg. Human well-being indicators as a boundary object for social science integration into conservation. Conservation Biology, 39(2):e14459:1–e14459:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Temple:2024:LER

[TBJ^{+}24] Andrew J. Temple, Per Berggren, Narriman Jiddawi, Nina Wambiji, Chris N. S. Poonian, Yussuf N. Salmin, Michael L. Berumen, and Selina M. Stead. Linking extinction risk to the economic and nutritional value of sharks in small-scale fisheries. Conservation Biology, 38(6):e14292:1–e14292:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:2022:AER

[TBP^{+}22] Elin A. Thomas, Monika Böhm, Caroline Pollock, Chong Chen, Mary Seddon, and Julia D. Sigwart. Assessing the extinction risk of insular, understudied marine species. Conservation Biology, 36(2):e13854:1–e13854:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thiault:2021:CSE

[TCG^{+}21] Lauric Thiault, Matthew I. Curnock, Georgina G. Gurney, Scott F. Heron, Nadine A. Marshall, Erin Bohensky, Nao Nakamura, Petina L. Pert, and Joachim Claudet. Convergence of stakeholders’ environmental threat perceptions following mass coral bleaching of the great barrier reef. Conservation Biology, 35(2):598–609, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2020:NEB

[TCK^{+}20] Bron Taylor, Guillaume Chapron, Helen Kopnina, Ewa Orlikowska, Joe Gray, and John J. Piccolo. The need for ecocentrism in biodiversity conservation. Conservation Biology, 34 (5):1089–1096, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tao:2023:BFF

[TDC^{+}23] Juan Tao, Chengzhi Ding, Jinnan Chen, Liuyong Ding, Sébastien Brosse, Jani Heino, Virgilio Hermoso, Ruidong Wu, Ziwang Wang, Jiaxin Hu, Rongxiao Che, Xiaowei Jin, Songhao Ji, and Dekui He. Boosting freshwater fish conservation with high-resolution distribution mapping across a large territory. Conservation Biology, 37(3):e14036:1–e14036:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Travers:2023:IES

[TDS^{+}23] Samantha K. Travers, Josh Dorrough, Ian Shannon, James Val, Mitchell L. Scott, Claudine J. Moutou, and Ian Oliver. The importance of expert selection when identifying threatened ecosystems. Conservation Biology, 37(6):e14151:1–e14151:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tsunoda:2020:SWM

[TE20] Hiroshi Tsunoda and Hiroto Enari. A strategy for wildlife management in depopulating rural areas of Japan. Conservation Biology, 34(4):819–828, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Treakle:2023:FAP

[TENI23] Tyler Treakle, Rebecca Epanchin-Niell, and Gwenllian D. Iacona. Factors associated with preemptive conservation under the U.S. Endangered Species Act. Conservation Biology, 37(5): e14104:1–e14104:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tensen:2024:EHS

[TF24] Laura Tensen and Klaus Fischer. Evaluating hybrid speciation and swamping in wild carnivores with a decision-tree approach. Conservation Biology, 38(1):e14197:1–e14197:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turner:2022:ENG

[TFFM22] Rachel A. Turner, Johanna Forster, Clare Fitzsimmons, and Robin Mahon. Expanding narratives of governance constraints to improve coral reef conservation. Conservation Biology, 36(5):e13933:1–e13933:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thornton:2025:COI

[TFJ^{+}25] Sara A. Thornton, Will Freeman, Lucy L. Jupe, Julia L. Newth, Jonathan P. Reeves, Kevin A. Wood, and Stephen Woroniecki. Challenges and opportunities for integrating social science in a conservation nongovernmental organization. Conservation Biology, 39(2):e70002:1–e70002:??, April 2025. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tan:2024:DBA

[TFT^{+}24] David J. X. Tan, Nicholas A. Freymueller, Kah Ming Teo, William S. Symes, Shawn K. Y. Lum, and Frank E. Rheindt. Disentangling the biotic and abiotic drivers of bird–building collisions in a tropical Asian city with ecological niche modeling. Conservation Biology, 38(4):e14255:1–e14255:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turo:2021:EUG

[TG21] Katherine J. Turo and Mary M. Gardiner. Effects of urban greenspace configuration and native vegetation on bee and wasp reproduction. Conservation Biology, 35(6):1755–1765, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tulloch:2020:CEM

[TGJ^{+}20] Vivitskaia Tulloch, Alana Grech, Ian Jonsen, Vanessa Pirotta, and Rob Harcourt. Cost-effective mitigation strategies to reduce bycatch threats to cetaceans identified using returnon-investment analysis. Conservation Biology, 34(1):168–179, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thurman:2022:AAA

[TGM^{+}22] Lindsey L. Thurman, John E. Gross, Claudia Mengelt, Erik A. Beever, Laura M. Thompson, Gregor W. Schuurman, Christopher L. Hoving, and Julian D. Olden. Applying assessments of adaptive capacity to inform natural-resource management in a changing climate. Conservation Biology, 36(2):e13838:1– e13838:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Theis:2024:SAT

[The24] Sebastian Theis. Sentiment and attitudes toward offsetting and the biodiversity market in online media articles. Conservation Biology, 38(6):e14274:1–e14274:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thompson:2021:EGI

[THM^{+}21] Ruth M. Thompson, Jordan Hall, Chris Morrison, Nicole R. Palmer, and David L. Roberts. Ethics and governance for Internet-based conservation science research. Conservation Biology, 35(6):1747–1754, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Taylor:2024:ICP

[THSA24] Liam U. Taylor, Wriley Hodge, Katherine R. Shlepr, and John Anderson. Interspecies conflict, precarious reasoning, and the gull problem in the Gulf of Maine. Conservation Biology, 38(6): e14299:1–e14299:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tulloch:2023:EDM

[TJB^{+}23] Ayesha I. T. Tulloch, Micha V. Jackson, Elisa Bayraktarov, Alexander R. Carey, Diego F. Correa-Gomez, Michael Driessen, Ian C. Gynther, Mel Hardie, Katherine Moseby, Liana Joseph, Harriet Preece, Andrés Felipe Suarez-Castro, Stephanie Stuart, John C. Z. Woinarski, and Hugh P. Possingham. Effects of different management strategies on long-term trends of Australian threatened and near-threatened mammals. Conservation Biology, 37(2):e14032:1–e14032:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Turnbull:2021:ESE

[TJC21] John W. Turnbull, Emma L. Johnston, and Graeme F. Clark. Evaluating the social and ecological effectiveness of partially protected marine areas. Conservation Biology, 35(3):921–932, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Theobald:2024:EEP

[TJE^{+}24] David M. Theobald, Aerin L. Jacob, Paul R. Elsen, Erik A. Beever, Libby Ehlers, and Jodi Hilty. Evaluating ecosystem protection and fragmentation of the world’s major mountain regions. Conservation Biology, 38(3):e14240:1–e14240:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Tonos:2025:ESP

[TPP^{+}25] Jadelys Tonos, Bastien Papinot, Daniel S. Park, Mihajatiana Raelison, Herilantonirina Ramaroson, Jessica Stubbs, and Onja H. Razafindratsima. Examining the structure of plant– lemur interactions in the face of imperfect knowledge. Conservation Biology, 39(1):e14323:1–e14323:??, February 2025. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Theis:2020:CEF

[TRR^{+}20] Sebastian Theis, Jonathan L. W. Ruppert, Karling N. Roberts, Charles K. Minns, Marten Koops, and Mark S. Poesch. Compliance with and ecosystem function of biodiversity offsets in North American and European freshwaters. Conservation Biology, 34(1):41–53, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Treves:2020:MAA

[TSÁ20] Adrian Treves and Francisco J. Santiago-Ávila. Myths and assumptions about human-wildlife conflict and coexistence. Conservation Biology, 34(4):811–818, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [CBD^{+}21].

Turner:2022:CBH

[TSS^{+}22] Gregory G. Turner, Brent J. Sewall, Michael R. Scafini, Thomas M. Lilley, Daniel Bitz, and Joseph S. Johnson. Cooling of bat hibernacula to mitigate white-nose syndrome. Conservation Biology, 36(2):e13803:1–e13803:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Teixeira:2021:LHE

[TTDdLM^{+}21] Lucas Teixeira, Karina Campos Tisovec-Dufner, Gabriela de Lima Marin, Silvio Marchini, Ine Dorresteijn, and Renata Pardini. Linking human and ecological components to understand human–wildlife conflicts across landscapes and species. Conservation Biology, 35(1):285–296, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thorne:2024:EDR

[TW24] L. H. Thorne and D. N. Wiley. Evaluating drivers of recent large whale strandings on the East Coast of the United States. Conservation Biology, 38(6):e14302:1–e14302:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas-Walters:2025:EST

[TWGGB^{+}25] Laura Thomas-Walters, Francisco Gelves-Gomez, Stephanie Brittain, Lily M. van Eeden, Nick Harvey Sky, Amit Kaushik, Kaylan Kemink, Patricia Manzano-Fischer, Kyle Plotsky, and Matthew Selinske. Essential skills for the training of conservation social scientists. Conservation Biology, 39(2):e14456:1– e14456:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas-Walters:2021:MUC

[TWHB^{+}21] Laura Thomas-Walters, Amy Hinsley, Daniel Bergin, Gayle Burgess, Hunter Doughty, Sara Eppel, Douglas MacFarlane, Wander Meijer, Tien Ming Lee, Jacob Phelps, Robert J. Smith, Anita K. Y. Wan, and Diogo Veríssimo. Motivations for the use and consumption of wildlife products. Conservation Biology, 35(2):483–491, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas-Walters:2023:SRC

[TWMM^{+}23] Laura Thomas-Walters, Jamie McCallum, Ryan Montgomery, Claire Petros, Anita K. Y. Wan, and Diogo Veríssimo. Systematic review of conservation interventions to promote voluntary behavior change. Conservation Biology, 37(1):e14000:1– e14000:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Thomas:2025:PCE

[TWS^{+}25] Hannah Thomas, Michelle S. Ward, Jeremy S. Simmonds, Martin F. J. Taylor, and Martine Maron. Poor compliance and exemptions facilitate ongoing deforestation. Conservation Biology, 39(1):e14354:1–e14354:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Torres:2024:MTH

[TzEN^{+}24] Aurora Torres, Sophus O. S. E. zu Ermgassen, Laetitia M. Navarro, Francisco Ferri-Yanez, Fernanda Z. Teixeira, Constanze Wittkopp, Isabel M. D. Rosa, and Jianguo Liu. Mining threats in high-level biodiversity conservation policies. Conservation Biology, 38(4):e14261:1–e14261:??, August 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Uddin:2023:LPR

[UEH^{+}23] Nasir Uddin, Sam Enoch, Abishek Harihar, Rob S. A. Pickles, Tasnim Ara, and Alice C. Hughes. Learning from perpetrator replacement to remove crime opportunities and prevent poaching of the Sundarbans tiger. Conservation Biology, 37 (2):e13997:1–e13997:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Utset:2024:CPS

[Uts24] Emily Utset. Combating parachute science in Latin America through indigenous agency. Conservation Biology, 38(2): e14217:1–e14217:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valko:2020:DFC

[Val20] Orsolya Valkó. Driving forces for conserving dynamic landscapes. Conservation Biology, 34(1):293–294, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanHuynh:2023:EIR

[Van23] Alex Van Huynh. Effect of IUCN Red List category on public attention to mammals. Conservation Biology, 37(3):e14050:1– e14050:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vercammen:2025:TFD

[VBC^{+}25] Ans Vercammen, Sayan Banerjee, Kyle Clifton, Matthew Selinske, and Chris Sandbrook. Trends and future directions in the conservation social sciences. Conservation Biology, 39 (2):e70011:1–e70011:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

vanBergen:2020:ESD

[vBDD^{+}20] Erik van Bergen, Tad Dallas, Michelle F. DiLeo, Aapo Kahilainen, Anniina L. K. Mattila, Miska Luoto, and Marjo Saastamoinen. The effect of summer drought on the predictability of local extinctions in a butterfly metapopulation. Conservation Biology, 34(6):1503–1511, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanBuskirk:2023:DSC

[VBK^{+}23] Amanda N. Van Buskirk, Bertram Bynum Boley, Charlie H. Killmaster, Kristina L. Johannsen, and Gino J. D’Angelo. Drivers of support for citizen science across state wildlife management agencies in the United States. Conservation Biology, 37(3):e14028:1–e14028:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Volenec:2020:CVS

[VD20] Zoe M. Volenec and Andrew P. Dobson. Conservation value of small reserves. Conservation Biology, 34(1):66–79, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valdez:2024:SAI

[VDO^{+}24] Jose Valdez, Gabriella Damasceno, Rachel R. Y. Oh, Laura Catalina Quintero Uribe, Martha Paola Barajas Barbosa, Talita Ferreira Amado, Chloé Schmidt, Miguel Fernandez, and Sandeep Sharma. Strategies for advancing inclusive biodiversity research through equitable practices and collective responsibility. Conservation Biology, 38(6):e14325:1–e14325:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vegvari:2024:EIT

[Veg24] Zsolt Vegvári. Ecological impacts of traffication. Conservation Biology, 38(3):e14275:1–e14275:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Verissimo:2023:TLA

[Ver23] Diogo Veríssimo. Think local, act local. Conservation Biology, 37(3):e13894:1–e13894:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

vanEeden:2020:ENS

[vESN^{+}20] Lily M. van Eeden, Kristina Slagle, Thomas M. Newsome, Mathew S. Crowther, Christopher R. Dickman, and Jeremy T. Bruskotter. Exploring nationality and social identity to explain attitudes toward conservation actions in the United States and Australia. Conservation Biology, 34(5):1165–1175, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vave:2024:CES

[VFKT24] Ron Vave, Alan M. Friedlander, John N. Kittinger, and Tamara Ticktin. Cultural ecosystem services and the conservation challenges for an Indigenous people’s aquatic protected area practice. Conservation Biology, 38(6):e14403:1–e14403:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vaidyanathan:2024:PAM

[VFRV24] Tanvi Vaidyanathan, Sarah J. Foster, B. Ramkumar, and Amanda C. J. Vincent. A practical approach to meeting national obligations for sustainable trade under CITES. Conservation Biology, 38(5):e14337:1–e14337:??, October 2024. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vasudev:2020:BHA

[VG20] Divya Vasudev and Varun R. Goswami. A Bayesian hierarchical approach to quantifying stakeholder attitudes toward conservation in the presence of reporting error. Conservation Biology, 34(2):515–526, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Verniest:2024:EWI

[VGB^{+}24] Fabien Verniest, Thomas Galewski, Olivier Boutron, Laura Dami, Pierre Defos du Rau, Anis Guelmami, Romain Julliard, Nadège Popoff, Marie Suet, Loïc Willm, Wed Abdou, Hichem Azafzaf, Nadjiba Bendjedda, Taulant Bino, John J. Borg, Luka Božič, Mohamed Dakki, Rhimou El Hamoumi, Vitor Encarnação, Kiraz Erciyas-Yavuz, Khaled Etayeb, Valeri Georgiev, Ayman Hamada, Ohad Hatzofe, Christina Ieronymidou, Tom Langendoen, Tibor Mikuska, Blas Molina, Filipe Moniz, Caroline Moussy, Asmaâ Ouassou, Nicky Petkov, Danae Portolou, Tareq Qaneer, Samir Sayoud, Marko Šćiban, Goran Topić, Danka Uzunova, Gal Vine, Andrej Vizi, Erald Xeka, Marco Zenatello, Elie Gaget, and Isabelle Le Viol. Exposure of wetlands important for nonbreeding waterbirds to sea-level rise in the Mediterranean. Conservation Biology, 38(6):e14288:1– e14288:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Visintin:2025:DCE

[VGW^{+}25] Casey Visintin, Georgia E. Garrard, Wolfgang W. Weisser, Mauro Baracco, Richard J. Hobbs, and Sarah A. Bekessy. Designing cities for everyday nature. Conservation Biology, 39(1): e14328:1–e14328:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanRossum:2022:GGM

[VH22] Fabienne Van Rossum and Olivier J. Hardy. Guidelines for genetic monitoring of translocated plant populations. Conservation Biology, 36(1):e13670:1–e13670:??, February 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vaisanen:2021:EHN

[VHHT21] Tuomas Väisänen, Vuokko Heikinheimo, Tuomo Hiippala, and Tuuli Toivonen. Exploring human–nature interactions in national parks with social media photographs and computer vision. Conservation Biology, 35(2):424–436, April 2021. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vincze:2022:BCH

[Vin22] Ernő Vincze. Bird conservation in human-made habitats. Conservation Biology, 36(5):e13999:1–e13999:??, October 2022.

CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Villalobo-Lopez:2024:ETP

[VLPVM^{+}24] Angelica Villalobo-Lopez, Carol M. Peña, Antonio Varas-Myrik, Michiel Pillet, Paulina Jahnsen, Patricio Pliscoff, Bárbara Goettsch, and Pablo C. Guerrero. Effects of trade and poaching pressure on extinction risk for cacti in the Atacama Desert. Conservation Biology, 38(5):e14353:1–e14353:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vallet:2023:KCI

[VLVD^{+}23] Améline Vallet, Bruno Locatelli, Merelyn Valdivia-Díaz, Yésica Quispe Conde, Gerardina Matencio García, Alejandrina Ramos Criales, Francisca Valverde Huamanñahui, Santusa Ramos Criales, David Makowski, and Sandra Lavorel. Knowledge coproduction to improve assessments of nature’s contributions to people. Conservation Biology, 37(6):e14182:1– e14182:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vimal:2020:UKM

[VM20] Ruppert Vimal and Courtney Morgans. Using knowledge mapping to rethink the gap between science and action. Conservation Biology, 34(6):1433–1443, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Velez:2024:ISD

[VMP^{+}24] Juliana Vélez, William McShea, Budhan Pukazhenthi, Pablo Stevenson, and John Fieberg. Implications of the scale of detection for inferring co-occurrence patterns from paired camera traps and acoustic recorders. Conservation Biology, 38(3): e14218:1–e14218:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vardi:2021:CCS

[VMR21] Reut Vardi, John C. Mittermeier, and Uri Roll. Combining culturomic sources to uncover trends in popularity and seasonal interest in plants. Conservation Biology, 35(2):460–471, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

VanderZanden:2024:GEB

[VNC^{+}24] Hannah B. Vander Zanden, David M. Nelson, Tara J. Conkling, Taber D. Allison, Jay E. Diffendorfer, Thomas V. Dietsch, Amy L. Fesnock, Scott R. Loss, Patricia A. Ortiz, Robin Paulman, Krysta H. Rogers, Peter M. Sanzenbacher, and Todd E. Katzner. The geographic extent of bird populations affected by renewable-energy development. Conservation Biology, 38 (2):e14191:1–e14191:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vogel:2022:JFT

[VPS22] Susanne Marieke Vogel, Maya Pasgaard, and Jens-Christian Svenning. Joining forces toward proactive elephant and rhinoceros conservation. Conservation Biology, 36(1):e13726:1– e13726:??, February 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Vallejos:2020:ECD

[VPVdAMF20] Marcelo Alejandro Villegas Vallejos, André Andrian Padial, Jean Ricardo Simões Vitule, and Emygdio Leite de Araujo Monteiro-Filho. Effects of crowding due to habitat loss on species assemblage patterns. Conservation Biology, 34(2):405– 415, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Villalva:2025:EUT

[VPZ25] Pablo Villalva, Francisco Palomares, and Marina Zanin. Effect of uneven tolerance to human disturbance on dominance interactions of top predators. Conservation Biology, 39(2): e14364:1–e14364:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Villarreal-Rosas:2024:ISE

[VRBAB^{+}24] Jaramar Villarreal-Rosas, Christopher J. Brown, Dominic A. Andradi-Brown, Ricardo Domínguez, Pilar Jacobo, Anuar Martínez, Coral Mascote, Eduardo Najera, Yves Paiz, Víctor Hugo Vázquez Moran, Jaime Villarreal, and María F. Adame. Integrating socioeconomic and ecological data into restoration practice. Conservation Biology, 38(4):e14286:1–e14286:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Valente:2023:FEE

[VRY^{+}23] Jonathon J. Valente, James W. Rivers, Zhiqiang Yang, S. Kim Nelson, Joseph M. Northrup, Daniel D. Roby, Carolyn B. Meyer, and Matthew G. Betts. Fragmentation effects on an endangered species across a gradient from the interior to edge of its range. Conservation Biology, 37(5):e14091:1–e14091:??, October 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiedenfeld:2021:CRA

[WAA^{+}21] David A. Wiedenfeld, Allison C. Alberts, Ariadne Angulo, Elizabeth L. Bennett, Onnie Byers, Topiltzin Contreras-MacBeath, Gláucia Drummond, Gustavo A. B. da Fonseca, Claude Gascon, Ian Harrison, Nicolas Heard, Axel Hochkirch, William Konstant, Penny F. Langhammer, Olivier Langrand, Frederic Launay, Daniel J. Lebbin, Susan Lieberman, Barney Long, Zhi Lu, Michael Maunder, Russell A. Mittermeier, Sanjay Molur, Razan Khalifa al Mubarak, Michael J. Parr, Jonah Ratsimbazafy, Anders G. J. Rhodin, Anthony B. Rylands, Jim Sanderson, Wes Sechrest, Pritpal Soorae, Jatna Supriatna, Amy Upgren, Jean-Christophe Vié, and Li Zhang. Conservation resource allocation, small population resiliency, and the fallacy of conservation triage. Conservation Biology, 35(5):1388–1395, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See response [WSW^{+}22].

White:2024:MIM

[WAB^{+}24] Elroy White, Kyle A. Artelle, Ed Brown, Kelly Brown, Diana E. Chan, and William Housty. Mn.úxv?it model for centering Indigenous knowledge and governance. Conservation Biology, 38(6):e14398:1–e14398:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallen:2024:SMT

[Wal24] Kenneth E. Wallen. Show me the theory: Response to Birdsong et al. 2024. Conservation Biology, 38(6):e14379:1–e14379:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [BMM^{+}24].

Williams:2023:PHD

[WB23] Peter Jeffrey Williams and Jedediah F. Brodie. Predicting how defaunation-induced changes in seed predation and dispersal will affect tropical tree populations. Conservation Biology, 37 (2):e14014:1–e14014:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallach:2020:RAP

[WBB^{+}20] Arian D. Wallach, Chelsea Batavia, Marc Bekoff, Shelley Alexander, Liv Baker, Dror Ben-Ami, Louise Boronyak, Adam P. A. Cardilin, Yohay Carmel, Danielle Celermajer, Simon Coghlan, Yara Dahdal, Jonatan J. Gomez, Gisela Kaplan, Oded Keynan, Anton Khalilieh, Helen Kopnina, William S. Lynn, Yamini Narayanan, Sophie Riley, Francisco J. Santiago-Ávila, Esty Yanco, Miriam A. Zemanova, and Daniel Ramp. Recognizing animal personhood in compassionate conservation. Conservation Biology, 34(5):1097–1106, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wood:2020:AZM

[WBC^{+}20] Jordan Wood, Jonathan D. Ballou, Taylor Callicrate, Jeremie B. Fant, M. Patrick Griffith, Andrea T. Kramer, Robert C. Lacy, Abby Meyer, Sara Sullivan, Kathy Traylor-Holzer, Seana K. Walsh, and Kayri Havens. Applying the zoo model to conservation of threatened exceptional plant species. Conservation Biology, 34(6):1416–1425, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Welch:2020:DST

[WBJ^{+}20] Heather Welch, Stephanie Brodie, Michael G. Jacox, Steven J. Bograd, and Elliott L. Hazen. Decision-support tools for dynamic management. Conservation Biology, 34(3):589–599, June 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wen:2024:RTT

[WBR^{+}24] Xin Wen, Joseph R. Bennett, Trina Rytwinski, Sahebeh Karimi, Matthew Spetka, Jessica J. Taylor, and Paul A. Smith. Review of terrestrial temporarily conserved areas in Canada, the United States, and Mexico. Conservation Biology, 38(1): e14160:1–e14160:??, February 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ward:2022:HPS

[WCW^{+}22] Michelle Ward, Josie Carwardine, James E. M. Watson, Anna Pintor, Stephanie Stuart, Hugh P. Possingham, Jonathan R. Rhodes, Alexander R. Carey, Nancy Auerbach, April Reside, Chuan Ji Yong, and Ayesha I. T. Tulloch. How to prioritize species recovery after a megafire. Conservation Biology, 36(5): e13936:1–e13936:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2025:EPA

[WCY^{+}25] Zhining Wang, Tao Chen, Li Yang, Colin A. Chapman, and Pengfei Fan. Effects of protected area coverage and research on conservation status of primates globally. Conservation Biology, 39(1):e14311:1–e14311:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wardropper:2021:CCS

[WDGM21] Chloe B. Wardropper, Ashley A. Dayer, Madeline S. Goebel, and Victoria Y. Martin. Conducting conservation social science surveys online. Conservation Biology, 35(5):1650–1658, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williamson:2021:IIA

[WDH^{+}21] Matthew A. Williamson, Brett G. Dickson, Mevin B. Hooten, Rose A. Graves, Mark N. Lubell, and Mark W. Schwartz. Improving inferences about private land conservation by accounting for incomplete reporting. Conservation Biology, 35 (4):1174–1185, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wu:2021:RBA

[WDHM21] Chung-Huey Wu, Aaron J. Dodd, Cindy E. Hauser, and Michael A. McCarthy. Reallocating budgets among ongoing and emerging conservation projects. Conservation Biology, 35 (3):955–966, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Webster:2022:KHB

[WDHR22] Hugh Webster, Amy Dickman, Adam Hart, and Dilys Roe. Keeping hunting bans on target. Conservation Biology, 36(4): e13932:1–e13932:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whiting-Fawcett:2025:PVB

[WFBT^{+}25] F. Whiting-Fawcett, A. S. Blomberg, T. Troitsky, M. B. Meierhofer, K. A. Field, S. J. Puechmaille, and T. M. Lilley. A Palearctic view of a bat fungal disease. Conservation Biology, 39(1):e14265:1–e14265:??, February 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woodworth:2021:TSB

[WFH^{+}21] Bradley K. Woodworth, Richard A. Fuller, Graham Hemson, Andrew McDougall, Bradley C. Congdon, and Matthew Low. Trends in seabird breeding populations across the great barrier reef. Conservation Biology, 35(3):846–858, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watermeyer:2021:UDS

[WGAB^{+}21] Kate E. Watermeyer, Gurutzeta Guillera-Arroita, Payal Bal, Michael J. Burgass, Lucie M. Bland, Ben Collen, Chris Hallam, Luke T. Kelly, Michael A. McCarthy, Tracey J. Regan, Simone Stevenson, Brendan A. Wintle, and Emily Nicholson. Using decision science to evaluate global biodiversity indices. Conservation Biology, 35(2):492–501, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Woinarski:2024:SVB

[WGZ24] John C. Z. Woinarski, Stephen T. Garnett, and Kerstin K. Zander. Social valuation of biodiversity relative to other types of assets at risk in wildfire. Conservation Biology, 38(3):e14230:1– e14230:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watters:2022:CSS

[WH22] George M. Watters and Jefferson T. Hinke. Conservation in the Scotia Sea in light of expiring regulations and disrupted negotiations. Conservation Biology, 36(5):e13925:1–e13925:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wei:2021:MAG

[WJ21] Xinzeng Wei and Mingxi Jiang. Meta-analysis of genetic representativeness of plant populations under ex situ conservation in contrast to wild source populations. Conservation Biology, 35 (1):12–23, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wiik:2020:MII

[WJP^{+}20] Emma Wiik, Julia P. G. Jones, Edwin Pynegar, Patrick Bottazzi, Nigel Asquith, James Gibbons, and Andreas Kontoleon. Mechanisms and impacts of an incentive-based conservation program with evidence from a randomized control trial. Conservation Biology, 34(5):1076–1088, October 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wood:2021:UES

[WKG^{+}21] Connor M. Wood, Holger Klinck, Michaela Gustafson, John J. Keane, Sarah C. Sawyer, R. J. Gutiérrez, and M. Zachariah Peery. Using the ecological significance of animal vocalizations to improve inference in acoustic monitoring programs. Conservation Biology, 35(1):336–345, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watts:2021:SPC

[WKT^{+}21] Matthew Watts, Carissa J. Klein, Vivitskaia J. D. Tulloch, Silvia B. Carvalho, and Hugh P. Possingham. Software for prioritizing conservation actions based on probabilistic information. Conservation Biology, 35(4):1299–1308, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallen:2020:SMC

[WL20] Kenneth E. Wallen and Adam C. Landon. Systematic map of conservation psychology. Conservation Biology, 34(6):1339– 1352, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitburn:2020:MAH

[WLA20] Julie Whitburn, Wayne Linklater, and Wokje Abrahamse. Meta-analysis of human connection to nature and proenvironmental behavior. Conservation Biology, 34(1):180–193, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallach:2020:WAL

[WLB^{+}20] Arian D. Wallach, Erick Lundgren, Chelsea Batavia, Michael Paul Nelson, Esty Yanco, Wayne L. Linklater, Scott P. Carroll, Danielle Celermajer, Kate J. Brandis, Jamie Steer, and Daniel Ramp. When all life counts in conservation. Conservation Biology, 34(4):997–1007, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Walker:2023:EBG

[WLB^{+}23] Barnaby E. Walker, Tarciso C. C. Leão, Steven P. Bachman, Eve Lucas, and Eimear Nic Lughadha. Evidencebased guidelines for automated conservation assessments of plant species. Conservation Biology, 37(1):e13992:1–e13992:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2022:EAL

[WLR^{+}22] Xiaodan Wang, Xinhai Li, Xiaotong Ren, Micha V. Jackson, Richard A. Fuller, David S. Melville, Tatsuya Amano, and Zhijun Ma. Effects of anthropogenic landscapes on population maintenance of waterbirds. Conservation Biology, 36(2): e13808:1–e13808:??, April 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Welch:2024:SPU

[WLR^{+}24] Heather Welch, Owen R. Liu, Leena Riekkola, Briana Abrahms, Elliott L. Hazen, and Jameal F. Samhouri. Selection of planning unit size in dynamic management strategies to reduce human–wildlife conflict. Conservation Biology, 38 (3):e14201:1–e14201:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wyborn:2021:ARA

[WMK^{+}21] C. Wyborn, J. Montana, N. Kalas, S. Clement, F. Davila, N. Knowles, E. Louder, M. Balan, J. Chambers, L. Christel, T. Forsyth, G. Henderson, S. Izquierdo Tort, M. Lim, M. J. Martinez-Harms, J. Merçon, E. Nuesiri, L. Pereira, V. Pilbeam, E. Turnhout, S. Wood, and M. Ryan. An agenda for research and action toward diverse and just futures for life on Earth. Conservation Biology, 35(4):1086–1097, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wilkinson:2020:EFC

[WMM^{+}20] Christine E. Wilkinson, Alex McInturff, Jennifer R. B. Miller, Veronica Yovovich, Kaitlyn M. Gaynor, Kendall Calhoun, Harshad Karandikar, Jeff Vance Martin, Phoebe Parker-Shames, Avery Shawler, Amy Van Scoyoc, and Justin S. Brashares. An ecological framework for contextualizing carnivore–livestock conflict. Conservation Biology, 34(4):854–867, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

White:2020:TRI

[WOF^{+}20] Timothy D. White, Tiffany Ong, Francesco Ferretti, Barbara A. Block, Douglas J. McCauley, Fiorenza Micheli, and Giulio A. De Leo. Tracking the response of industrial fishing fleets to large marine protected areas in the Pacific Ocean. Conservation Biology, 34(6):1571–1578, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wong:2024:MIS

[Won24] Mark K. L. Wong. Misrepresentation of invasive species in the mass media with images of unrelated organisms. Conservation Biology, 38(6):e14382:1–e14382:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wallach:2023:SPI

[WRBL^{+}23] Arian D. Wallach, Daniel Ramp, Ana Benítez-López, Eamonn I. F. Wooster, Scott Carroll, Alexandra J. R. Carthey, Erin I. E. Rogers, Owen Middleton, Kyle J. A. Zawada, Jens-Christian Svenning, Ella Avidor, and Erick Lundgren. Savviness of prey to introduced predators. Conservation Biology, 37(2): e14012:1–e14012:??, April 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Ward:2020:USS

[WRW^{+}20] Michelle Ward, Jonathan R. Rhodes, James E. M. Watson, James Lefevre, Scott Atkinson, and Hugh P. Possingham. Use of surrogate species to cost-effectively prioritize conservation actions. Conservation Biology, 34(3):600–610, June 2020. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:2020:ICC

[WSB^{+}20] Sara H. Williams, Sarah A. Scriven, David F. R. P. Burslem, Jane K. Hill, Glen Reynolds, Agnes L. Agama, Frederick Kugan, Colin R. Maycock, Eyen Khoo, Alexander Y. L. Hastie, John B. Sugau, Reuben Nilus, Joan T. Pereira, Sandy L. T.

Tsen, Leung Y. Lee, Suzika Juiling, Jenny A. Hodgson, Lydia E. S. Cole, Gregory P. Asner, Luke J. Evans, and Jedediah F. Brodie. Incorporating connectivity into conservation planning for the optimal representation of multiple species and ecosystem services. Conservation Biology, 34(4):934–942, August 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

White:2023:FBC

[WSC^{+}23] Rehema M. White, Birgit Schmook, Sophie Calmé, Anthony J. Giordano, Yves Hausser, Lynn Kimmel, Lou Lecuyer, Mauro Lucherini, Crisol Méndez-Medina, and Juan L. Peña-Mondragón. Facilitating biodiversity conservation through partnerships to achieve transformative outcomes. Conservation Biology, 37(3):e14057:1–e14057:??, June 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winkler-Schor:2024:EDD

[WSEL^{+}24] Sophia Winkler-Schor, Harold N. Eyster, Diele Lobo, Lauren Redmore, Andrew J. Wright, Victoria M. Lukasik, Wendy Chávez-Páez, Brooke Tully, Sarah Beard, Kwan-Lamar Blount-Hill, Catherine Christen, and Zoe Nyssa. Enhancing disciplinary diversity and inclusion in conservation science and practice based on a case study of the Society for Conservation Biology. Conservation Biology, 38(6):e14395:1–e14395:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:2022:RUL

[WSK^{+}22] Nicholas E. Williams, Seeta A. Sistla, Daniel B. Kramer, Kara J. Stevens, and Adam B. Roddy. Resource users as landsea links in coastal and marine socioecological systems. Conservation Biology, 36(1):e13784:1–e13784:??, February 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watters:2022:UMI

[WSSC22] Freyja Watters, Oliver Stringham, Chris R. Shepherd, and Phillip Cassey. The U.S. market for imported wildlife not listed in the CITES multilateral treaty. Conservation Biology, 36(6): e13978:1–e13978:??, December 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Winkler-Schor:2020:DRE

[WSvRLK20] Sophia Winkler-Schor, Carena J. van Riper, Adam Landon, and Rose Keller. Determining the role of eudaimonic values in conservation behavior. Conservation Biology, 34(6):1404–1415, December 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Watson:2022:CTC

[WSW^{+}22] James E. M. Watson, Jeremy S. Simmonds, Michelle Ward, Chuan J. Yong, April E. Reside, Hugh P. Possingham, Andrew Rogers, and Josie Carwardine. Communicating the true challenges of saving species: response to Wiedenfeld et al. Conservation Biology, 36(4):e13961:1–e13961:??, August 2022. CO-DEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic). See [WAA^{+}21].

White:2023:ATB

[WSW23] Richard S. A. White, Rick J. Stoffels, and Amy L. Whitehead. Assigning trend-based conservation status despite high uncertainty. Conservation Biology, 37(4):e14084:1–e14084:??, August 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Whitehouse-Tedd:2021:EUP

[WTAD21] K. Whitehouse-Tedd, J. Abell, and A. K. Dunn. Evaluation of the use of psychometric scales in human–wildlife interaction research to determine attitudes and tolerance toward wildlife. Conservation Biology, 35(2):533–547, April 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Williams:2022:GRI

[WWB^{+}22] Brooke A. Williams, James E. M. Watson, Hawthorne L. Beyer, Carissa J. Klein, Jamie Montgomery, Rebecca K. Runting, Leslie A. Roberson, Benjamin S. Halpern, Hedley S. Grantham, Caitlin D. Kuempel, Melanie Frazier, Oscar Venter, and Amelia Wenger. Global rarity of intact coastal regions. Conservation Biology, 36(4):e13874:1–e13874:??, August 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wu:2025:VSH

[WWVC^{+}25] Nicholas C. Wu, Justin A. Welbergen, Tomás Villada-Cadavid, Lindy F. Lumsden, and Christopher Turbill. Vulnerability of Southern Hemisphere bats to white-nose syndrome based on global analysis of fungal host specificity and cave temperatures. Conservation Biology, 39(2):e14390:1–e14390:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Wang:2021:CPC

[WYB^{+}21] Lin Wang, Bin Yang, Yang Bai, Xiaoqiang Lu, Richard T. Corlett, Yunhong Tan, Xiao-Yong Chen, Jianguo Zhu, Yan Liu, and Rui-Chang Quan. Conservation planning on China’s borders with Myanmar, Laos, and Vietnam. Conservation Biology, 35(6):1797–1808, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Xuan:2021:VHS

[XAAY^{+}21] Bui Bich Xuan, Claire W. Armstrong, Isaac Ankamah-Yeboah, Stephen Hynes, and Katherine Needham. Valuing high-seas ecosystem conservation. Conservation Biology, 35(5):1405– 1416, October 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Xuereb:2021:IPN

[XDA^{+}21] Amanda Xuereb, Cassidy C. D’Aloia, Marco Andrello, Louis Bernatchez, and Marie-Josée Fortin. Incorporating putatively neutral and adaptive genomic data into marine conservation planning. Conservation Biology, 35(3):909–920, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Xu:2020:NAP

[XST^{+}20] Yanjie Xu, Yali Si, John Takekawa, Qiang Liu, Herbert H. T. Prins, Shenglai Yin, Diann J. Prosser, Peng Gong, and Willem F. de Boer. A network approach to prioritize conservation efforts for migratory birds. Conservation Biology, 34 (2):416–426, April 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yeo:2024:UMA

[YCH^{+}24] Darren Yeo, Amy H. J. Chan, Kum Chew Hiong, Jasmine Ong, Jun Yuan Ng, Jie Min Lim, Wendy Zhang, Sara R. Lim, Charlene J. Fernandez, Anna M-S. Wong, Benjamin P. Y-H. Lee, Max D. Y. Khoo, Thaddaeus X. W. Cheng, Bryan T. M. Lim, Hazelina H. T. Yeo, Maxine M. Q. Tan, Wendy B. G. Sng, Shaun S. Adam, Wee Foong Ang, Choon Beng How, Renhui Xie, Samuel K. Wasser, Kristen N. Finch, Adrian H. B. Loo, Him Hoo Yap, Chee Chiew Leong, and Kenneth B. H. Er. Uncovering the magnitude of African pangolin poaching with extensive nanopore DNA genotyping of seized scales. Conservation Biology, 38(2):e14162:1–e14162:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yang:2023:ECL

[YCS^{+}23] Li Yang, Tao Chen, Kai-Chong Shi, Lu Zhang, Ngwe Lwin, and Peng-Fei Fan. Effects of climate and land-cover change on the conservation status of gibbons. Conservation Biology, 37(1):e14045:1–e14045:??, February 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yoh:2025:UAC

[YHW^{+}25] Natalie Yoh, Mukhlish Jamal Musa Holle, Jasmin Willis, Lauren F. Rudd, Iain M. Fraser, and Diogo Veríssimo. Understanding author choices in the current conservation publishing landscape. Conservation Biology, 39(2):e14369:1–e14369:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yamaura:2020:SEE

[YLY^{+}20] Yuichi Yamaura, David Lindenmayer, Yusuke Yamada, Hao Gong, Toshiya Matsuura, Yasushi Mitsuda, and Takashi Masaki. A spatially explicit empirical model of structural development processes in natural forests based on climate and topography. Conservation Biology, 34(1):194–206, February 2020. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yeiser:2021:OCS

[YMB^{+}21] John M. Yeiser, John J. Morgan, Danna L. Baxley, Richard B. Chandler, and James A. Martin. Optimizing conservation in species-specific agricultural landscapes. Conservation Biology, 35(6):1871–1881, December 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yan:2025:ANP

[YTW^{+}25] Yanyun Yan, Song Ling Tan, Edward L. Webb, James E. M. Watson, and L. Roman Carrasco. Ability of new protected areas to counteract losses from downgrading, downsizing, and degazettement. Conservation Biology, 39(2):e14381:1– e14381:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yoon:2022:ASH

[YVA22] Hyun Seok Yoon, Varsha Vijay, and Paul R. Armsworth. Accounting for spatial heterogeneity in the added conservation value of land protection when prioritizing protected areas. Conservation Biology, 36(5):e13960:1–e13960:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Yasue:2024:EEI

[YWH^{+}24] Maï Yasué, Netta Weinstein, Sara E. Harris, I-Chant A. Chiang, Nicole Legate, Ashley J. Moore, and Nadia Joe. Embedding equity and inclusion in universities through motivational theory and community-based conservation approaches. Conservation Biology, 38(6):e14384:1–e14384:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zimmerman:2024:PIG

[ZAS^{+}24] Shawna J. Zimmerman, Cameron L. Aldridge, Michael A. Schroeder, Jennifer A. Fike, Robert Scott Cornman, and Sara J. Oyler-McCance. The potential influence of genomewide adaptive divergence on conservation translocation outcome in an isolated greater sage-grouse population. Conservation Biology, 38(4):e14254:1–e14254:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zeiss:2024:ECD

[ZBM^{+}24] Romy Zeiss, Maria J. I. Briones, Jérome Mathieu, Angela Lomba, Jessica Dahlke, Laura-Fiona Heptner, Gabriel Salako, Nico Eisenhauer, and Carlos A. Guerra. Effects of climate on the distribution and conservation of commonly observed European earthworms. Conservation Biology, 38(2):e14187:1– e14187:??, April 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zeiss:2022:COP

[ZEO^{+}22] Romy Zeiss, Nico Eisenhauer, Alberto Orgiazzi, Matthias Rillig, François Buscot, Arwyn Jones, Anika Lehmann, Thomas Reitz, Linnea Smith, and Carlos A. Guerra. Challenges of and opportunities for protecting European soil biodiversity. Conservation Biology, 36(5):e13930:1–e13930:??, October 2022. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhao:2021:EAP

[ZLZ^{+}21] Peng Zhao, Shuming Liu, Yi Zhou, Tim Lynch, Wenhu Lu, Tao Zhang, and Hongsheng Yang. Estimating animal population size with very high-resolution satellite imagery. Conservation Biology, 35(1):316–324, February 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhang:2024:EPT

[ZSL^{+}24a] Anning Zhang, Hongxian Song, Ziyang Liu, Hanwen Cui, Haitao Ding, Shuyan Chen, Sa Xiao, Lizhe An, and Pedro Cardoso. Effects of plant taxonomic position on soil nematode communities in Antarctica. Conservation Biology, 38(4): e14264:1–e14264:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zitong:2024:ISD

[ZSL24b] Zuo Zitong, Wang Shuyi, and Zhong Le. Identification, spatial distribution, and associated factors of urban protected areas in China. Conservation Biology, 38(6):e14298:1–e14298:??, December 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zizka:2021:ACA

[ZSVK21] Alexander Zizka, Daniele Silvestro, Pati Vitt, and Tiffany M. Knight. Automated conservation assessment of the orchid family with deep learning. Conservation Biology, 35(3):897–908, June 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhang:2021:EPA

[ZTCF21] Lu Zhang, Samuel T. Turvey, Colin Chapman, and Pengfei Fan. Effects of protected areas on survival of threatened gibbons in China. Conservation Biology, 35(4):1288–1298, August 2021. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhang:2024:GPH

[ZWC^{+}24a] Chengcheng Zhang, Yihong Wang, Jiang Chang, Junsheng Li, Shengkai Pan, Biao Yang, Xiangjiang Zhan, and Qiang Dai. Global patterns of human–wildlife spatial associations and implications for differentiating conservation strategies. Conservation Biology, 38(6):e14279:1–e14279:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhang:2024:EIS

[ZWC^{+}24b] Eden Zhang, Sin Yin Wong, Paul Czechowski, Aleks Terauds, Angelique E. Ray, Nicole Benaud, Devan S. Chelliah, Daniel Wilkins, Kate Montgomery, and Belinda C. Ferrari. Effects of increasing soil moisture on Antarctic desert microbial ecosystems. Conservation Biology, 38(4):e14268:1–e14268:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zheng:2024:UCI

[ZWC^{+}24c] Wuji Zheng, Anita Kar Yan Wan, Zhan Chen, Alexander Clark, Caitlin Court, Yuanlu Gu, Toby Park, Jake Reynolds, Xiaoxi Zhang, Lishu Li, and Tien Ming Lee. Use of consumer insights to inform behavior change interventions aimed at illegal pet turtle trade in China. Conservation Biology, 38(5): e14352:1–e14352:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhang:2024:QEG

[ZXR24] Taxing Zhang, Yu Xu, and Jianghong Ran. Quantitative evaluation of the global impacts of human land modification on raptors. Conservation Biology, 38(3):e14228:1–e14228:??, June 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhu:2023:CPE

[ZXXL23] Yingying Zhu, Xiaoting Xu, Zhenxiang Xi, and Jianquan Liu. Conservation priorities for endangered trees facing multiple threats around the world. Conservation Biology, 37(6): e14142:1–e14142:??, December 2023. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhou:2024:CCP

[ZYM^{+}24] Xuehong Zhou, Ami Yang, Zhen Miao, Wei Zhang, Qiang Wang, and Douglas C. MacMillan. Consumer characteristics and preferences for mobulid gill plates in China. Conservation Biology, 38(6):e14244:1–e14244:??, August 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhao:2025:THQ

[ZYNC25] Jianqiao Zhao, Le Yu, Tim Newbold, and Xin Chen. Trends in habitat quality and habitat degradation in terrestrial protected areas. Conservation Biology, 39(2):e14348:1–e14348:??, April 2025. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).

Zhu:2024:FSM

[ZZ24] Annah Lake Zhu and George Zhu. Financial speculation meets cultural heritage in China’s wildlife markets. Conservation Biology, 38(5):e14339:1–e14339:??, October 2024. CODEN CBIOEF. ISSN 0888-8892 (print), 1523-1739 (electronic).